A Complete Bibliography of Publications in Journal of Parallel and Distributed Computing: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

10 June 2022 Version 2.00

Title word cross-reference

(f, g) [CDD^{+}15]. 0 [dADC18]. 1 [dADC18, JM14]. 1 − m [SJG19]. 2 [CHCG18, DJDK19, KRKS11, LXLS12, TSFZ14, Tur12, YA11]. 2.5 [MPG17b]. 2 log N − 1 [CC14]. 3 [AA14, AA16, BDRB14, GOH^{+}13, GRS19, LLFJ18, NM17, OGRV^{+}12, PYP^{+}10, SLV19, YA11, ZHW19, ZLS17, Zsa16]. 4 [KMC16]. 45 [HRF^{+}11]. d [DTK11b]. G [BFKW13]. GF(2^{m}) [SKH15]. h [KLP10]. hp [PPTV^{+}10]. K [DBCF13, WL11, Amm16, CDDL10, RDA18, SSKS11, SDG17, WCH^{+}17, YTZ19, ZHT16]. K_{1,3} [LLFJ18]. L [ZBW^{+}17]. LT Q_{n} [XHZZ16]. LU [FHL^{+}15, SLV19]. m [ABBD14, Kar19, SJG19]. n [LLFJ18]. O(1) [Can18]. O(log_{2}(min(m, n))) [XL11]. O(log m̄, log N) [CC14]. O(n) [DLV11]. p [YBX^{+}13]. P_{N} [OGM^{+}19]. P_{N−2} [OGM^{+}19]. π [EHKSS19]. QR [BDG^{+}15, FHL^{+}15]. t [CRHC19].

-alliances [CDD^{+}15]. -ary [LLFJ18, SJG19]. -aware [MSRB19]. -Cipher [EHKSS19]. -clustering [CDDL10]. -connectivity [ZHW19]. -coverage [Amm16]. -covered [CHCG18]. -cubes [LLFJ18]. -D [LXLS12]. -dimensional [DTK11b]. -disjoint [KMC16]. -distributed [CRHC19]. -Gaussian [WL11]. -hop [JM14]. -Means [DBCF13]. -MSA [BFKW13]. -mutual [RDA18]. -nearest [SDG17]. -NN [ZHT16]. -packing [TSFZ14]. -PIC [YBX^{+}13]. -plex [WCH^{+}17]. -relations [KLP10]. -stage [CC14]. -structure [LLFJ18]. -systems [ZBW^{+}17]. -time [DLV11]. -wave [Kar19].

/many [KSG13].

0/1 [BW18]. 0/1-Knapsack [BW18].

1-Knapsack [BW18]. 1-type [GA18]. 10 [LB12]. 113 [KN18b].

2010 [Phi13]. 2011 [Mue13]. 2014 [Ben15]. 2016 [PGKV18, SV19]. 26th [OY13].

3 [KMC16].

5G [DAPR18].

71 [LSS^{+}11a].

802.11 [ZBR11]. 802.11n [GZY14a].

A-GHSOM [IZ12]. A-Star [SRT^{+}18]. AAIA [TFV^{+}15]. abstract [RJKL11]. Abstraction [DDO^{+}18, IRRS16, LSZJ15, HCR12]. ACAS [MBR19]. accelerate [SJVRVVS19, SDG17]. Accelerated [AB13, BMS19, CRHC19, DGNW13, DCA^{+}15, Eme13, GOH^{+}13, KDO^{+}13, LMSK18, SHA17, WLL16, Zsa16]. Accelerating [AVAH18, DFST13, GK19, GAOHG17, RCG18, SKH15, WD13, YL12, YZG18, ZXB14, ZCS^{+}18, AM12a, VBDRC13]. acceleration [BAT^{+}19, KCP19, LLY15, NMS^{+}18, OGM^{+}19, UGG^{+}11]. accelerator [CNLGRL18, ICQO^{+}12, PP13, PG19]. Accelerators [DF12, MLK12, RBN11]. Access [ALLM11, AM13, BC11, ETS14, FLC14, KKK11a, KGN11, LZ11, LWZZ12, LC11, LS19, MLZY17, MYYY17, NSDZ18, NKK16, TODQ18, WBRT13]. access-aware [MYYY17]. AccessAuth [TODQ18]. accessing [CJYC19]. Accident [CCW14]. accrual [CRJ10b]. accuracy [CRWX12, CDY^{+}19]. accurate [BFKW13, CGL^{+}14, GJ12, HZDP12, NV19]. Accurately [LC13]. achieve [LCB16]. Achieving [XLC^{+}18]. Acknowledgment [Gra10a]. ACOR [BFVB19]. Acoustic [LPLFMC^{+}12]. across [MB19, SGdSS13]. Action [Sie16, CDY^{+}19]. active [KV10, PSGS17, SI13]. activity [CWZ^{+}18, HES11, SAR^{+}18, Udd19]. actor [YpGyLlC13]. actuator [KKKP12, SCN12]. Ad [LC14b, RBP^{+}11, TM10, AH11, AH12, BM11, FCW11, FGL^{+}11, JLWX11, Kim11, KSK15, LHW14, NMN^{+}14, OM10, SNCP12, SJS11, TC13, WHS^{+}18, YSS11, YWW12]. ad-hoc [KSK15, LHW14, NMN^{+}14]. Ada [WSX^{+}19]. Ada-Things [WSX^{+}19]. Adaboost [CLZ19]. adaptable [LLLC15, LFGM17]. adaptation [KGN11, NZY^{+}11, WMC^{+}18, WWY^{+}18, YHWY18a]. adapting [SW18, WRW13]. Adaptive [AA16, ACPT15, BFVB19, BOT13, CGM14, KG10, LPK^{+}10, LC11, UBES10, VMMB10, WL10, ZHLQ12, APK18, AF17, BM17a, BMT12, BBS13, BEN12, CMMN10, CKC19, CAF^{+}11, DLW^{+}12, DAB^{+}14, GA16, GNZ18, HHK15, IZ12, KK17, KBK^{+}19, LST17, LHX^{+}16, LWW18, MPG17a, MPN17, MSEM^{+}19, NKK16, PPTV^{+}10, SMO14, SB12, SMB10, SHC14, TLY12, WSX^{+}19, ZXYO11, ZLCZ18]. Address [YQTV12, WZ13, YGZ^{+}10, YC12]. addressable [KRM14]. adjacent [CFJW13]. adjusted [TDBL13]. administration [LB17]. Admission [MBO11, AAA^{+}10, MCZ14, YJKD10]. advance [CRH11]. Advanced [HDCM11, MCP^{+}18, PH18, PSGS17, DMK19, ZXMR18]. advancement [LZ11]. Advances [DDE19, GA16, HAC^{+}19]. adversarial [WLK^{+}19]. adversary [dOCS14]. advice [DP12, GGY19]. Advisor [uRIL^{+}18]. aerial [SRB^{+}19]. AES [ABO^{+}17]. affected [LdPLC^{+}19]. affinity [HD10]. Agate [CZPP16]. agent [Rao16, YZS15, YHWY18a]. agent-based [Rao16, YHWY18a]. agents [FP17, SGAC14, SMO^{+}18, BJ18]. aggregate [AMT13]. aggregated [WE13]. Aggregation [MBMC19, BCO^{+}12, HBSASA19, JBA15, JBS14, JHPL13, SSKS11, XHZ^{+}10, ZSCX18, Zsa16]. Aging [BM17a, LC14a]. Aging-aware [BM17a]. agreement [AP16, HC11, LLW12, REK10a, REK10b]. ahead [mH14, SHL^{+}13, TLL^{+}18]. AHMW [BMT12]. aid [CVK^{+}18b]. aided [SV18]. air [YBM13]. algebra [Eme13, FHL^{+}15, ICQO^{+}12, LKD14]. Algorithm [Ren11, AA10, ALM^{+}16, AA14, AA16, ALLM11, ADDB18, ARDQ18, BDL^{+}19, BCH15, BFKW13, BBD18, CDDL10, CC14, CK13, CLOL17, CPLY18, COF^{+}17, CSW^{+}17, CDW^{+}19, CCC^{+}19, DFHH13, DK11, DLV11, FZWL12, GLW14, GG19, GL12, GAOHG17, HES10, HSS10, HES11, HSY10, HVW16, HWY^{+}10, JXZ^{+}19, KG19, KR10b, KHW13, Kar19, Kim17, KIH15, LHWJ19, LASS15, LLCZ19, LLT12, LÜ14, LW16b, LYIP19, LFEP19, MSEM^{+}19, NST19, NHO^{+}13, OGM^{+}19, PDP17, PB15, RT18, RBG17, RBOH^{+}18, RDA18, SLV19, SMP17, SWW^{+}17, Tam18, TLQS12, Tát11, Ter16, TYA16, TSFZ14, WLL16, WCL^{+}13, WWW17a, WJ12, gWW18, XL11, XYZW14, XSYG18, YWJ^{+}18, YÖ11, YSS11, ZQMM11, dOBG^{+}15, CMR10, KM17, LY12]. algorithm/implementation [HVW16]. algorithmic [BBH^{+}17, CG11, CLZ19, JF12]. Algorithms [Ben15, FTM^{+}14, TFV^{+}15, ANEA13, ASC^{+}18, Ara13, ACCP12, AAC10, AF17, ARVZ14, BKC^{+}15, BBBC12, Ben19, BMT12, BKCM17, CP10a, CRH11, CRSB13, CRD17, Cuz11, Cuz13, DJ16, DCA^{+}15, DKU15, FHL^{+}15, GMMP12, GZY14a, GM14a, GK10, GN15, ICQO^{+}12, JST12, KR10a, KHT^{+}14, KSSG14, KK10, KMS10, KR11, LNW^{+}12, LLS^{+}16, MCAS12, MRS^{+}14, MPR19, MKM16, MSAZ10a, MSAZ10b, NTN12, OA10, PKN10, PH18, PH16, PPSV15, PS14, QGZP19, SPH13, SHRM19, SSVC10, SBRM19, Tal19, TWQS12, Tur12, VAF19, VS16, YZG18, ZGJ^{+}18, ZXMR18]. alignment [BFKW13, SPRG^{+}12, SRT^{+}18]. All-Pairs [DCA^{+}15]. all-to-all [PW16, ZTFK16]. alliances [CDD^{+}15]. allocating [AS19a]. allocation [AAA^{+}10, ADD17, ACCP12, BSS^{+}13, CCA18, CDS10, CPLY18, DW12, HWY^{+}10, HLL^{+}19, HB11, HGX^{+}19, JL11, KR10a, KR10b, KHW13, KS18, LL10, LL12a, LL12b, LDP^{+}14, MLK^{+}16, NVK^{+}11, PKN10, SSM^{+}16, SNCP12, SCW^{+}18, SCMS12, SHL^{+}13, SSVC10, SZB16, TFMS15, YYWZ19, ZG13]. almost [EB13]. almost-optimal [EB13]. alternating [HWY^{+}10]. alternative [Can18]. always [AD10]. always-on [AD10]. ambiguities [RK18]. ambiguity [LDS16]. Amdahl [CN14, NZ17, SC10]. among [ST12, ZWY^{+}15, ZCW19]. AMR [RV13]. AMTE [HCM11]. Analysis [CSMML10, LDS16, LE19, SWHB17, WCF14, AFK14, AK18, Ben19, BBH^{+}17, BFL^{+}13, BC11, BF13, CSL15, CLZ19, CKT11, CLXX19, CWCW18, CJA^{+}19, GZG^{+}17, GHC^{+}17, HSH10, HB11, JF12, KBK^{+}19, KK10, KKK^{+}11b, LMSK18, LdSB^{+}18, LpJS^{+}18, LZC11, MAKWZ13, MBO11, MEMEMH17, NV19, NSKN17, PLK^{+}18, SMW18, SWLP19, SPPA19, TLY12, VLW18, Wu11, XLW^{+}18, ZKZF18, ZPK^{+}14, DFLO17]. analytic [Ale19b, LWC^{+}18]. analytical [AHZ11, KyLPC17]. Analytics [AS13, AS15, CJ17, Eck18, KKKG14, PS14, PAG^{+}18, TH19, VLGV^{+}18, YLB^{+}15]. analyze [LZN19]. Analyzing [LB12, PB19, RB12, RPS19]. anchors [MKM16]. Android [TY17]. animation [JdSJC^{+}15]. ANMR [BM17a]. annealing [dADC18]. annuli [Li14]. anomaly [AKK^{+}19, IZ12, KKTZ13, MBR19, RLP14]. anomaly-based [MBR19]. anonymity [DKJG19]. anonymous [KS13, LWK^{+}19]. answer [BYG^{+}18]. Ant [COV13, CGN^{+}13, CLA^{+}18, DDGK13, CCK11, Ski16]. Anti [GSASA19]. Anti-spoofing [GSASA19]. Apache [KKH17]. API [HLS12]. applicability [Can18]. Application [DKK18, OGRV^{+}12, WLST16, dKG^{+}10, AHA^{+}16, AAI^{+}15, BM16, BMT12, CKMP17, DWYB10, FCP^{+}15, HSS17, KG19, KAA^{+}19a, Kub17, LGRV19, LW16a, Li17, LAS^{+}19, MLZY17, MCM^{+}11, MRJ^{+}19, PVP18, PS14, PVRS17, WW18b, WJ14, YÖ11]. Applications [PGRP17, PJ18, SFC17, TFV^{+}15, ALM^{+}16, AC16, ASSS19, BCD^{+}15, BHLT14, CCA18, CLW^{+}19, CGL^{+}14, CGM14, CSMML10, CP10b, CMC^{+}19, CDAN14, FCML13, FPF14, FRM15, GQZ18, GLC14, GYAB11, GVBB13, GRR13, HZZ^{+}19, Kol19, KBC^{+}10, LWCC15, LFGM17, MSRB19, MA19, MLK12, NV19, NLB^{+}18, NPS^{+}19, NMS^{+}18, NVK^{+}11, NC13, OTKT12, PCMM^{+}17, PH18, PB19, PMAL11, PA15, PCLP16, RCG18, RJKL11, SWW^{+}17, SR16, SSGZ13, TP18, TPLY18, TOR^{+}14, TKX^{+}13, WSX^{+}19, WBS19, WIR^{+}18, XCC^{+}19, ZVL11, ZZJ^{+}18, ZSW14, ZXMR18, dSS11, FTM^{+}14]. apply [NZ17]. applying [CCK11, Kol19]. approach [AS19a, AFD^{+}11, AJG18, AS18, BM11, BCK^{+}13, CTS17, CHX^{+}17, CZZ^{+}17, CDPS18, DKKV15, dADC18, GDL^{+}11, GXYZ13, ICQO^{+}12, KYS13, KBC19, KSJC17, KZ11, KCFP18, LL19, LXW^{+}11, LZN19, MHLZ16, MLCFH^{+}18, MBMC19, MGRRK14, NTN12, NHO^{+}13, OPR18, Ozt11, PD19, PST^{+}19, PQ19, RK18, RMGM19, SW18, SGVRP19, SDG17, SK11, TP18, TXLL14, TY17, TM10, VLW18, WML^{+}18, WWY^{+}18, WZQ^{+}13, WLZ^{+}18, XRB12, XLH18, YHWY18a, YAA10, YDTZ18, YWG15, ZHH15, ZS13, ZTGL17]. Approaches [FMIF18, CB11, DBA^{+}18, KWZ19, KBK^{+}19, LE19, RPN19, Upa13]. Approximate [LHW14, LRS18, ST12, BDL^{+}19, CLOL17, HN19, JHL^{+}18]. Approximation [GM14a, JST12, LLCZ19, DKU15, FZWL12, GG19, PSRS12]. arbitrary [BCF14, Kar19]. arbitration [HRG^{+}11]. Architectural [TGPUC16, KWZ19, NXTK17]. Architecture [DB18, MEMEMH17, PSGS17, YPCW16, AA10, AA16, ABO^{+}17, BJS18, BGA12, BBCQ13, CLMRL15, CS17, FSP18, GRZ^{+}18, KK17, KNHH18, KH12, LLKY13, LHHH11, LLY15, MCM^{+}11, MYD^{+}11, NPS^{+}19, NPVG^{+}19, NKV14, PPP14, PCMM^{+}17, PYP^{+}10, PGP^{+}12, PTK^{+}13, SAR^{+}18, SLKK12, WTWZ16, YFBY17, ZMZJ17, ZPK^{+}14, KCSS18, VRGS17]. Architectures [FTM^{+}14, FPS12, AA14, AG12, BKC^{+}15, BYG^{+}18, CGC16, CJ17, DKRC^{+}15, DKU15, FPS11, FTM^{+}19, GMS^{+}13, GMSS^{+}11, HDCM11, JJ12, KHT^{+}14, LGRV19, Lla17, MSGS^{+}13, MP10, PD19, SPPA19, SGdSS13, TRS^{+}12, WQZ^{+}13, TFV^{+}15]. Archive [FTK14, JKIE13]. Area [CDR12, ABO^{+}17, CHCG18, JKV15, LdSB^{+}18, LMJC11]. Area-maximizing [CDR12]. Ariadne [MM15]. arithmetic [Dav17]. ARM [AG12]. arrangement [Ten16]. Array [NKV14, CS10, Man13]. array-based [CS10]. arrays [Can18, EAB^{+}19, JWSG14, WCF14, XS11]. art [PSC^{+}16]. arts [NDW17]. ary [LLFJ18, SJG19]. Aspen [UMM^{+}18]. Aspen-based [UMM^{+}18]. Assessing [BCD^{+}15]. assessment [CG17, FGL^{+}11, LC14a, SJVRVVS19]. assigned [HMR15]. Assigning [CCK11]. Assignment [Cza13, SSZ10, ABBD14, CS10, DLM19, GQZ18, GDL^{+}11, GZY14a, JTZZ11, Kim11, LZLX11, NDP13, PLY15, SLKK13, WW18b, YZX11]. assisted [GM13, HMY^{+}18, KO12, MBBD13, NS12, SRT^{+}18]. Associate [Ano16k]. Associations [GPJA10]. Associative [TJCB10, HDCM11, YBM13]. assumption [Pen11]. assumptions [MS15]. assured [AKK^{+}19]. asymmetric [CMC^{+}19, KNHH18, SPC^{+}17]. asymmetrically [ATKT19]. Asymptotically [Li10]. Async [ARP18]. Asynchronous [KVNV17, TP18, ATDH13, CPA^{+}11, DBCF13, DPBNT12, GLGLBG12, IRRS16, Kak15, KMS10, KS13, MEMEMH17, MSEM^{+}19, RV13, SMO^{+}18, WPC19]. ATAPE [PW17]. atmosphere [KVNV17]. atomic [BOT13]. atomicity [RHH12]. attack [BK18, KCFP18]. Attacking [ZWY^{+}15]. attacks [LLWC17, TAM^{+}19, UGG^{+}11, YXX13]. attention [PLSM18]. attribute [LSS^{+}11a, LSS^{+}11b]. attributed [LKB^{+}15]. auction [GVBB13, RA11, ZG13]. auction-based [ZG13]. auction-inspired [GVBB13]. audiences [LMB^{+}17]. audio [TYD^{+}19, WIR^{+}18]. Audit [HLS12]. auditing [XLC^{+}18]. augmentation [BCH15]. augmented [KM17, KAA^{+}19b, MKW18]. Authentic [GPJA10, SZMK13]. authenticated [YTH^{+}19]. Authentication [ZBR11, BDM18, LMJC11, LWK^{+}19, PRN^{+}19, TODQ18, YTZ19]. Author [Ano10a, Ano11j, Ano12m, Ano14f]. authority [ZCMY12]. Auto [PSB^{+}19, CXX^{+}18, KKR14, KGN11, MBR19, TLL^{+}18, VD18]. auto-adaptation [KGN11]. auto-clean [CXX^{+}18]. auto-encoders [TLL^{+}18]. auto-scaler [VD18]. auto-scaling [MBR19]. Auto-tuned [PSB^{+}19]. auto-tuning [KKR14]. autoencoder [WMC^{+}18]. automata [EM11, GKS15, MBO11, PD19, RT18, TM10, ZBW^{+}17]. automata-based [EM11, RT18]. automated [CV16]. Automatic [AD12, CLXX19, HZZ^{+}19, LZZ^{+}11, NCB^{+}17, AM17, DXS^{+}19, GLC14, GFPC14, MLCFH^{+}18, NVK^{+}11]. automatically [DSEP17]. automation [HKK^{+}18, PD19]. automaton [LSZZ15, Pet18]. automaton-based [LSZZ15]. automotive [RAN^{+}17]. autonomic [AZC13, LS10, RDA18, XRB12]. autonomous [CKT11, CKMP17, WZZ^{+}17, ZWW17]. availability [AGMS16, Fu10, KVA18, LKM12, LAC18, PMMMA15]. average [WWW17a]. avoid [DP16]. Avoiding [SI13, SLV19]. awards [OY13]. Aware [DR18, DKK18, SDS^{+}18, AH12, AYB^{+}15, BM17a, BGO19, CWZ^{+}18, CLW^{+}19, CCW14, CWP12, CHCG18, CKML12, EHL^{+}15, FCW11, FCJG^{+}18, Fu10, GQZ18, GPSH19, HMR15, HWL18, HV13, JAB12, JHF^{+}17, KKK11a, KK11, KCR14, KBC^{+}10, LL19, LBMG15, LFS16, LR14, LDZ^{+}14, LZI^{+}11, LW16a, LNAL17, LLCZ19, LY13, LHLM14, MBBD13, MHLZ16, MYYY17, MSRB19, MTL^{+}18a, MLK^{+}16, MMK^{+}11, MBR19, MA19, MSEM^{+}19, NHX^{+}19, NL19, ORWT^{+}18, OMT^{+}17, OJP^{+}18, PD19, RBN11, RCG18, SNMB16, SJB12, SA19, SKK14, SCW^{+}18, SP13, STK11, SZL10, TLLV10, TODQ18, TVT^{+}17, UM17, VMMB10, WQL14, WMY^{+}17, WHC^{+}18, YZX11, YJKD10, ZLKK19, ZVL15, ZXYO11, ZTFK16, ZLJ^{+}19, ZWQ^{+}16, Sie16]. awareness [HRH18, LWZZ12, ZXGD18]. Axiom [ABLP17]. Axiom-based [ABLP17].

B&B [BMT12, DBA^{+}18]. back [LKD14, WMES12]. backends [IEWK17]. backfilling [GMVRGS16]. backtracking [AKDMN15]. bad [Sch14]. bag [BHLT14, dSS11]. bag-of-tasks [BHLT14, dSS11]. balance [ZW11, ZWY^{+}15]. balanced [ASES15, GHY10, SB15, YMLP14]. balancing [AES11, BFMT^{+}18, CKLW19, DB11, DLW^{+}12, GLC14, HLL^{+}19, JL11, KAA^{+}19a, MPV12, NHO^{+}13, PC11, PRN^{+}19, SBÇ12a, TVT^{+}17, YJL16, YAA10, ZV14, ZSW14, ZLMC14]. Balls [BBFN12, BBFN14]. band [WIR^{+}18]. Bandwidth [KK17, BHK17, FFYH19, HWY^{+}10, HB11, MSK^{+}16, SSGG18, YYWZ19]. bandwidth-efficient [BHK17]. barriers [HS12]. based [APRA18, ASA18, AA10, ASKTZ13, AVAH18, ALLM11, AHG12, AMSÅ19, AYB^{+}15, AP16, AK18, ABLP17, ABF^{+}14, AJG18, AS18, BDM18, BNBR16, BOY10, BCMV15, BCH15, BDRB14, BFKW13, BYG^{+}18, BK18, BAT^{+}19, BEN12, BYH^{+}17, BBB11, CWZ^{+}18, CLW^{+}19, CG12, CLMRL15, CK13, CVK^{+}18b, CP10b, CS10, CHX^{+}17, CLOL17, CXQ^{+}18, CRJ10a, CJ17, CTT16, CAF^{+}11, CKMP17, CRD12, CDW^{+}19, DBA^{+}18, DKKV15, DB11, DR19, DBW^{+}18, DKC14, DWYB10, DXS^{+}19, ESGQ^{+}14, ESGQ^{+}18, EM11, ECP^{+}18, FLL14, FCML13, FLCB10, FGL^{+}11, GOH^{+}13, GMMP12, GPJA10, GTGLSA12, GL12, GSASA19, GA16, GNZ18, GRZ^{+}18, GXYZ13, HBS17, HRH18, HWY^{+}10, HZL18, HMY^{+}18, HGX^{+}19, IIH16, IIH^{+}17, JTC^{+}18, JM14, KKR14, KyLPC17, KAA^{+}19a, KKS^{+}12, KKLJ14, KCP19, KKTZ13, LK15, LC14a, LdSB^{+}18, LSH^{+}13, LZI^{+}11, LMJC11, LW16a, LLWC17, LNW^{+}12, LÜ14, LZC11]. based [LSZZ15, LZY^{+}18, LCJ^{+}18, LLDL15, LPLFMC^{+}12, Lop18, LACJ18, LAC18, LFEP19, LLFJ18, MS19, MAGL13, MM15, MP10, MAKWZ13, MBR19, MBO11, MH18, MSAZ10a, MSAZ10b, MRJ^{+}19, MZZC12, MCZ14, NV19, NSKN17, NST19, NCA^{+}12, NTN12, NHO^{+}13, NC13, No12, NPVG^{+}19, OM10, OJP^{+}18, Ozt11, PARB14, PD19, PLSM18, PDP17, PMAL11, RLP14, RT18, Rao16, RA11, RTZ11, RDA18, RDCQ17, SMW18, SSM^{+}16, SMPMLVLS11, SHSH17, SCG10, SPH13, SLW10, ST12, Ski16, SQQL19, Suk18, SK11, SLHS19, TBG^{+}17, TFMS15, TCMB^{+}19, TW15, TKKH17, TC13, TJCB10, TWQS12, Udd19, URK^{+}19, UMM^{+}18, UM17, VETT18, VMMB10, VS18, WW12, WCL^{+}13, WRW13, WYW15, WWW17b, WML^{+}18, WMC^{+}18, WWY^{+}18, WXZ^{+}18, WXMZ19, WIR^{+}18, WLYS19, WMG13, WD18, WD13, WCCH18, WWA^{+}18, WHW^{+}19, XLHT13, YWJ^{+}18, YZC^{+}19, YL12, YHWY18a, YHWY18b, YXW^{+}18, YAA10, ZG13]. based [ZGJ^{+}18, ZAAB17, ZFT^{+}18, ZW13, ZPK^{+}14, ZLL14, ZLT^{+}19, ZV12, ZGG^{+}14, ZXGD18, dSAJ15, dAAD^{+}19, HRF^{+}11, TOR^{+}14, ZBR11]. bat [CCC^{+}19]. batch [CMR19]. batched [HSH10]. Bayesian [DKC14, LWC^{+}18, NZA13, RFP^{+}19, SHK19, YWAT13]. be [HBS17, KSSK16, STKW12, BGA12]. beacons [DWX10]. beeps [CD19]. Before [HCR12]. behavior [ACD^{+}18, CL14, LWXX19, dAMFdS13, RA11]. behaviour [CMMN10]. belief [HMY^{+}18]. benchmark [DMS^{+}16, GN15]. Benchmarking [BBR13, YYLC11]. benchmarks [HZZ^{+}19, KC17]. Benefit [BHK17]. Best [Mue13, OY13, Phi13, QGZP17]. best-effort [QGZP17]. better [STKW12]. between [KR17, LCB16, MP15, NM17, XS11, ZXGD18]. Beyond [LdPLC^{+}19, CC14]. BFS [BCMV15, DJ16, RGAN18]. BGP [LXW^{+}11]. bi [AM11, MMK^{+}11]. bi-modal [AM11]. bi-objective [MMK^{+}11]. Biased [AS19b, RM11]. bichromatic [NMN^{+}14]. bidirectional [PD19]. BiELL [ZGG^{+}14]. Big [AS13, AS15, APK18, LWWQ18, MNR^{+}19, SFC17, ACPT15, CF19, CJYC19, CCC^{+}19, Eck18, FRM15, KKKG14, NXTK17, NST19, PST^{+}19, RBA^{+}18, SMW18, TH19, WWW17b, YBX^{+}13, ACB^{+}15]. Bimodal [UM17]. binary [ASC^{+}18, Can18, CJDC10, LFZ^{+}17, WD18]. bins [BBFN12, BBFN14]. Bio [Hua17]. Bio-Grid [Hua17]. biological [SMB10]. biologically [FTM^{+}19]. biomedical [WZY^{+}19]. biometric [CNLGRL18, GSASA19]. bipartite [DKU15]. bipartitioning [PB15]. Bisection [AK17, ZGG^{+}14]. bit [KIH15]. bit-parallel [KIH15]. Bitonic [BM14, TW15]. BitTorrent [ARD14, CTC11, LXZ13]. BitTorrent-like [CTC11]. Black [PSC^{+}16, BE13, SGAC14]. BlackOut [ZCF^{+}17]. BlobCR [NC13]. BlobSeer [NAB^{+}11]. Block [ADV14, ATDH13, DAB^{+}14, FLCB10, LYIP19, MRJ^{+}19, MRT18, PP13, SPH13, WZZ^{+}17, XLW^{+}18]. block-asynchronous [ATDH13]. block-level [FLCB10]. block-structured [DAB^{+}14]. Blockchain [ZLT^{+}19, DKJG19]. blocking [ASES15, CASD18, DBA^{+}18, ESGQ^{+}11, KR17, MPN17]. blocks [RJKL11]. Bloom [SMPMLVLS11]. BMB [WD18]. BMMI [SJG19]. BMMI-tree [SJG19]. Board [Ano18v, Ano18w, Ano18x, Ano18q, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m]. Board [Ano18n, Ano18o, Ano18p, Ano18r, Ano18s, Ano18t, Ano18u, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x]. body [CHCG18, YJL16]. Boltzmann [WZY^{+}19]. Bone [AFK14]. Boolean [ESCV15, OT19]. boosting [AC16]. both [WTY^{+}18]. bottom [LXZ13]. bottom-up [LXZ13]. bound [AMM^{+}18, Kub17, Li19b]. bound-consistency [Kub17]. boundary [SMP17, TRS^{+}12, ZQMM11]. bounded [JM14, Sta17]. bounds [LXLS12, LYW^{+}16, NDP13]. brain [RBOH^{+}18]. branch [AMM^{+}18]. brawny [LNC13]. breadth [MB13, ZCS^{+}18]. breadth-first [ZCS^{+}18]. breast [HES11, XTN12]. bridging [KLJ^{+}11]. broad [LMB^{+}17]. Broadband [XP10, XTN12]. broadcast [AA10, CB15, KYS13, KG10, LDZ^{+}14, LDZ^{+}17, LSWC14, LSZZ15, MT14, MPS16, WWW17a, WIR^{+}18, dAAD^{+}19]. broadcast-based [AA10]. broadcasting [LSWC14, OWK14]. Brokering [CPJ^{+}19]. Brown [DTK11a]. Brunotte [Tát11]. BTS [BKK^{+}11]. bubble [GNZ18]. bubble-type [GNZ18]. budget [ZVL15]. budget-aware [ZVL15]. budgeted [Sta17]. buffer [CHX^{+}17, IIH16, WCWO17, WYW15]. bufferless [LLT12]. buffers [EKNS17, ZCF^{+}17]. build [ZHH15]. Building [RJKL11, Suk18, ZW13, HSS10]. burst [WCWO17]. bus [PTN^{+}19, TJCB10, YGZ^{+}10]. bus-based [TJCB10]. buses [BG16]. Business [MBS^{+}12]. Business-driven [MBS^{+}12]. Butterfly [VAF19]. bypass [dOBG^{+}15]. Bypassing [DKK18]. Byzantine [DPBNT12, HC11, IRRS16, LHW14, MT14].

C [EFG^{+}14, HCM11]. C-AMTE [HCM11]. C2FPGA [CSJ^{+}13]. Cache [DKK18, CK13, CDAN14, FABG^{+}19, HCM11, KK11, LC11, LZLX11, MYYY17, MPG17a, MA11, SS17, VRGS17, WLZ^{+}18]. caches [ATKT19, DMI^{+}19, EHL^{+}15, NSAS10, SS17]. Caching [CS17, FCW11, FCML13, LAK10, MA11, TC13, ZVL11]. CAFES [MCM^{+}11]. calculations [KVNV17]. calibration [SDG17]. call [LGK^{+}12]. campus [CJYC19]. Can [KSSK16, HBS17]. cancer [XTN12]. CANE [LAS^{+}19]. capabilities [FEH^{+}14, GRS19, RBN11]. capabilities-aware [RBN11]. Capability [JLWX11, BJ15, CKC19, dOBG^{+}15]. capable [SMP17]. Capacity [TODQ18, ACCP12, BKK^{+}11, JHPL13]. Capacity-aware [TODQ18]. capture [BOT13]. Carlo [PAS15, ZS13]. carrier [BC11]. Cartesian [GOH^{+}13, ISAZ10, MSAZ11]. Case [PP13, AGMS16, AES11, BJ18, CHLL18, FRM15, HdR13, PCMM^{+}17, RÖE^{+}18, SJVRVVS19, TdAR18, WLCZ15, WMG13, ZKZF18, ZLJ^{+}19]. CASS [FPS11]. Cassandra [PMMMA15]. Cataloging [RFP^{+}19]. causal [EDH^{+}17, dAAD^{+}19, HCR12]. Causality [MCS14]. cause [LXW^{+}11, MBR19]. caused [Zha11]. CBase [ZLZ^{+}19]. ccNUMA [MTM10]. CEA [LY12]. cell [LWCC15, MFT^{+}19, BGA12, XP10, XTN12]. cell-centered [LWCC15]. Cell/BE [BGA12]. cells [FTM^{+}19, ZPK^{+}14]. cellular [ANEA13, EM11, GKS15, LMSK18, PD19, PSRS12, Pet18, ZBW^{+}17]. center [BFH^{+}17, CGC16, SCW^{+}18, ZLZ^{+}19]. centered [LWCC15]. centers [AG12, AK18, GYAB11, HTB19, MB19, MLK^{+}16, OJP^{+}18, RT18, TVT^{+}17, YAK15, ZV14, ZV12]. centrality [JL11, SSKÇ15, WBRT13]. centric [KTP17, LAS^{+}19, VS18, XYZW14]. CFPA [MSEM^{+}19]. chain [ASKO16, LWW19]. chained [BM14, CMR^{+}18]. chained-cubic [BM14]. chains [LBMG15]. Challenges [NKSA17, PJ18, AFH^{+}19, PSC^{+}16]. Channel [LWLD12, SSZ10, GDL^{+}11, GZY14a, GZY14b, KKK11a, Kim11, ZMG^{+}16]. channels [LSWC14]. chaos [DZC17]. chaos-oriented [DZC17]. Characteristics [DKK18, BCD^{+}15, RGAN18, SWHB17]. Characteristics-Aware [DKK18]. characterization [DWYB10]. Characterizing [HRF^{+}11, ZSW14]. chart [LZN19]. Check [MC17, LXW^{+}11]. checking [BBBC12, CAK13, MMN^{+}18, XYZW14]. Checkpoint [LACJ18, NC13, WCWO17]. Checkpoint-Restart [LACJ18, NC13]. Checkpointing [ARVZ14, BCC^{+}18]. checkpoints [AD10]. chemical [CP10b, XLHT13]. chess [WW18b]. Chief [Pra16]. Chinese [XLW^{+}18]. Chip [BJS18, MSEM^{+}19, BYG^{+}18, DJDK19, DR19, DMS^{+}16, GJ12, HCM11, HRG^{+}11, KK11, KH12, KKK^{+}11b, LNA12, LLKY13, LSXX14, LLT12, LY13, LHLM14, LWCG14, MYD^{+}11, PMCC18, SAJ13, TCHC12, UM17, AA14, ALLM11, LK11, MEMEMH17, ORWT^{+}18, PR13, ZCF^{+}17]. chip-multiprocessors [LWCG14]. Chips [LK10, RGAN18]. choices [BBFN12]. Cholesky [GLW14]. chordal [BCH15]. chromatin [MFT^{+}19]. Chronos [LHNBB19]. Cipher [EHKSS19]. circuit [LC14a, LWCG14]. circuit-level [LC14a]. circuit-switched [LWCG14]. circuits [EB13, HBS17, MH18, OOSGVG^{+}16]. circulation [Nes10]. Cities [NCRK19, AKK^{+}19, AMSÅ19, DFLO17, NML^{+}19, SLZ^{+}19, ZCW19]. city [AKSZ19, HRH18, KDSS18, NPS^{+}19, ZLJ^{+}19]. clairvoyant [Li19a]. CLAP [KK17]. CLAP-NET [KK17]. class [EB13]. classes [FP17]. CLASSIC [PG19]. classification [COV13, CK13, PDP17, TPLY18]. classifier [BOKS19, SDG17, UGG^{+}11]. classifying [Luc18]. clean [CXX^{+}18]. clients [GZY14a]. Climate [AFH^{+}19, DXS^{+}19]. clinical [KDO^{+}13]. clique [WCH^{+}17]. cliques [SMT15]. clocks [TPLY18]. cloning [RMHR17]. Cloud [CDJL11, FEH^{+}14, LAC18, PR13, VS18, ASKO16, ASHO19, AKK^{+}19, Ale19b, Ale19a, AZC13, AM12a, AMSÅ19, ACCP12, BYH^{+}17, CL14, CCA18, CPJ^{+}19, CXY14, CTKA17, DKRC^{+}15, FRM15, FCJG^{+}18, FMIF18, GQZ18, GYAB11, GSASA19, GPSH19, HTB19, HRM17, HMY^{+}18, JAB12, KVA18, KBC19, KS18, KBK^{+}19, KSSK16, LWZZ12, LQM^{+}12, LLB^{+}18, LGM18, LZWZ19, LWH^{+}19, MHLZ16, MYYY17, MXSL12, MMK^{+}11, MA19, PLSM18, PH18, RT18, SMW18, SWW^{+}17, SFHS19, SLZ^{+}19, TCMB^{+}19, TKR^{+}19, TLW18, TKX^{+}13, VD18, WXMZ19, WCCH18, gWW18, XLC^{+}18, XRB12, XSYG18, YTH^{+}19, YYWZ19, YYLC11, ZV14, ZLL14, ZHT16, HAC^{+}19, NLB^{+}18]. cloud-based [AMSÅ19, GSASA19, TCMB^{+}19, WCCH18]. Cloud-centric [VS18]. cloud-of-things [TKR^{+}19, HAC^{+}19]. cloud-oriented [GYAB11, HRM17, MXSL12]. cloudlets [TPS^{+}18]. clouds [ACPT15, ACB^{+}15, CKMP17, KM17, KKLJ14, LYJ^{+}19, LTWW12, LWWQ18, MBR19, NC13, NKK16, PVP18, ZG13, ZVL15]. Cluster [JKV15, CCA18, CDS10, CDY^{+}19, FLCB10, GRR13, IEWK17, JGMY17, LAK10, LML^{+}10, LÜ14, LZC11, LB17, LB18, NDP13, NVK^{+}11, PKW^{+}10, RLP14, SBÇ12b, SHL^{+}13, SMH^{+}14, WLL16]. cluster-based [FLCB10, LÜ14]. Cluster-to-cluster [JKV15]. clustered [GZY14b, Lop18, NS12, SFT^{+}13]. Clustering [TM10, ASKTZ13, AYB^{+}15, AS18, BM16, BM17b, BDL^{+}19, BF13, CDDL10, CLC^{+}17, DBCF13, DKM10, GYP13, KG19, KKH17, LK15, SAL10, TLW18, YBX^{+}13, YÖ11, YWW12, ZMCP11]. clusters [ARP18, DCA^{+}15, Fu10, GRS19, GYY^{+}14, HV13, JM14, KG19, KKH17, KCR14, MMVL11, SS11, VBDRC13, WQL14, uRIL^{+}18, WH17, YJKD10, ZMCP11, ZHLQ12]. CMOS [KRM14]. CMPs [AFA13, APRA18, FABG^{+}19, FLC14, HRF^{+}11, OOSGVG^{+}16]. CNN [CXX^{+}18]. CNNs [CDW^{+}19]. Co [AHA^{+}16, KN18b, KN18a, RBG17, BBH^{+}17, HVW16, HD10, NVK^{+}11, OJP^{+}18, PSB^{+}19]. co-allocation [NVK^{+}11]. Co-Design [RBG17, BBH^{+}17]. co-evolutionary [HD10]. co-execution [PSB^{+}19]. co-location [OJP^{+}18]. co-optimization [HVW16]. Co-optimizing [AHA^{+}16]. Co-processing [KN18b, KN18a]. coalition [YZS15]. Cobb [HGX^{+}19]. Code [TAM^{+}19, CDW^{+}19, MMN^{+}18]. Code-driven [TAM^{+}19]. codes [AM13, CP10a]. coding [DFHH13, ZY12]. CODISC [MA11]. coevolutionary [ADDB18]. cognitive [FCZ^{+}12, GDCC18, SA19]. cognizant [LK13]. coherence [APRA18, CDAN14, CRD12, MPG17a, SPPA19]. cohort [AKBD10]. coin [AAC10]. cold [MB19, GPSH19]. cold-spot [MB19]. collaboration [LR14]. Collaborative [MA11, SLHS19, CJDC10, DBLB^{+}12, LLWC17, LHCC19, NKK16, RJKL11, VLW18, ZLJ^{+}19]. collection [Amm16, ZWW17]. collective [MBBD13, NKK16]. collectives [Zah12]. collectors [VRM10]. college [NDW17]. Collision [LDZ^{+}17, CXQ^{+}18, JBS14, MBMC19]. collision-free [JBS14, MBMC19]. Collision-tolerant [LDZ^{+}17]. collusion [AFD^{+}11]. Colony [CGN^{+}13, CLA^{+}18, DDGK13, Ski16, CCK11]. coloring [GK10]. combination [DKC14, YFBY17]. Combinatorial [Ben15, DDE19, ZG13, CMMT13, Men18, PPSV15, WMG13]. Combined [GDCC18, VAS^{+}13]. Combining [AAC10, CMMT13, SR16, UBES10, WMY^{+}17, BCC^{+}18]. committee [BDP16]. commodity [ZXB14]. Communicating [CD19, DRR13]. Communication [FPS11, FLM^{+}19, MLK^{+}16, RFS^{+}12, AFA13, ARP18, ALTV13, AM12a, BM17b, BBR13, DB11, DKUÇ15, DAPR18, EDH^{+}17, GM13, JJ12, JZZ^{+}17, LAK10, LWCG14, LLW12, dAMFdS13, MTL^{+}18a, MCM^{+}11, MPG17b, REK10a, REK10b, SAL10, SLV19, SRI14, SLKK12, SZB16, SSGZ13, Tam18, TW15, WPC19, YCH^{+}10, YQTV12, FPS12]. communication-avoiding [SLV19]. Communication-free [FLM^{+}19]. communication-intensive [MLK^{+}16]. communication-optimal [MPG17b]. communications [MBBD13, PGP^{+}12, TP18, TKG^{+}17]. community [CTC^{+}10, LpJS^{+}18, WLYS19, ZLL14]. community-based [ZLL14]. compact [NKV14]. Compaction [SLHS19]. Comparative [SJVRVVS19, KBK^{+}19]. Comparison [GRS19, AG12, Ben19, GHC^{+}17, JKIE13, MP10, NSKN17, SMB10, ZTFK16]. Comparisons [YBM13]. compass [AKBD10]. compass-free [AKBD10]. competition [WSLC11]. Competitive [SHC14, LHHH11]. competitiveness [GK15]. compilation [MH18, PAG^{+}18, WQZ^{+}13]. Compiler [NS12]. Compiler-assisted [NS12]. complementary [ZPK^{+}14]. complete [LFZ^{+}17]. completely [SPC^{+}17]. Complex [DDO^{+}18, HASB16, CM12, DF17, HHA14, JKD^{+}15, RBP^{+}11, SW12, SJG19]. Complexity [KCP19, AEF11, FWM^{+}10, Sol13, XL11]. complexity-effective [FWM^{+}10]. Component [AHG12, HdR13, KRKS11, VLW18]. Component-based [AHG12]. component-oriented [HdR13]. components [GHIJ19]. composition [CJ17, WMY^{+}17]. compositions [FZ14]. comprehensive [CJA^{+}19, GM14b, MFT^{+}19, TH19, Upa13, uRIL^{+}18, ZAB18]. compression [CW15, HBSASA19, JKV15, KP17, AHG12]. Comput [KN18b, LSS^{+}11a, MSAZ10a, PCX^{+}14, REK10a]. computation [Ale19a, CXY14, DB11, LMB^{+}17, LGM18, MCS14, RMU14, SS11, XLH18, YJL16]. Computational [APV18, AAH17, CCE^{+}17, CLZ19, KHT^{+}14, LGRV19, MS19, Pen11, QGZP19, RBN11, SMO14, SNCP12, ZLKK19]. computational-power [ZLKK19]. computations [BD11, CG10, IEWK17, RV13, SSKÇ15, SBÇ12a, SWLP19, SMH^{+}14, WJ14]. computed [KDO^{+}13]. Computer [DB18, BDRB14, Eme13, FSP18, LMB^{+}17, LB17, PSC^{+}16, PR13]. computers [AM13, Li16]. Computing [BS11, CDJL11, CGFH19, DDO^{+}18, DB18, Eme13, FTM^{+}14, HAC^{+}19, KRS13, LK11, MSGS^{+}13, MNK12, MBG^{+}17, Nee17, PBB^{+}17, SLL18, STS19, SFC17, TPS^{+}18, TJCB10, TPJ^{+}19, ALM^{+}16, AAK^{+}13, AMU^{+}19, AZC13, AM12a, AMT13, ASC^{+}18, ACB^{+}15, ABLP17, BFL^{+}13, BHS13, BLZ^{+}18, BYH^{+}17, CFI^{+}18, CMMT13, CVK^{+}18b, CPJ^{+}19, CTKA17, CDR12, DDG^{+}17, DF12, EFG^{+}14, ES12, FPF14, FKR^{+}17, FP17, Fu10, FX10, GQZ18, GMSS^{+}11, GAC^{+}17, GRZ^{+}18, HES10, HRH18, HZL18, HLL^{+}19, mH14, JHL^{+}18, JTC^{+}18, JdSJC^{+}15, KHW13, KDO^{+}13, KBC19, KV10, KCR14, Kol19, KCFP18, KDSS18, Las12, Las13, LHWJ19, LZY11, LLCZ19, LBT19, LWH^{+}19, LYJ^{+}19, LS10, LML^{+}10, LE19, LB18, MYYY17, MSRB19, MNR^{+}19, MZC18, MFT^{+}19, MMK^{+}11, MA19, MKN14, NXTK17, NML^{+}19, NDW17, NSDZ18, PLD14, PH18, PGKV18, RBN11, Ren11, SMW18, SJB12, SSM^{+}16, SAR^{+}18]. computing [Sch14, SFT^{+}13, Sie16, SZL10, Suk18, TZI11, TLLL10, TLLV10, TFMS15, TXLL14, Udd19, WZX^{+}19, WZH^{+}19, WG11, gWW18, XLHT13, YLL17, YWJ^{+}18, YTH^{+}19, YLZW18, YBM13, ZAB18, ZKZF18, ZGW^{+}19, ZLL14, AS13, Cuz11, FPS11, GMSS^{+}11, KRS13, KRS14, Las12, MMVL11, TH11]. Concept [DFLO17]. concepts [MAGL13, NKSA17]. Concurrency [ADD17, SRI14, UBES10]. Concurrent [ACHY18, CMN12, DBLB^{+}12, BE13, CTS17, PVP18, SW18]. Conditional [CSS11]. conditions [Ste17]. Confident [YDZ^{+}18]. confidentiality [ZHT16]. Configurable [TCMB^{+}19, ZMZJ17]. configuration [LB17, VAS^{+}13, WZ13, WLST16]. conformation [MFT^{+}19]. conforming [LGM18]. Congestion [MSEM^{+}19, AA10, BM11, BFVB19, ESGQ^{+}14, ESGQ^{+}18, YJKD10]. congestion-oblivious [BFVB19]. conjugate [GLW14, LR14]. Connected [KRKS11, BJ18, JPD17, KO12, YSS11, YWW12, ZAAB17]. connection [LXLS12]. connectivity [BCMV15, Ten16, ZHW19]. conquer [GDL^{+}11, PV19, TP18]. conscious [GYAB11]. consensus [AAI^{+}15, LHW14, WWW17a]. consequences [YBM13]. conservation [SHRM19, XS11]. considerations [RSK19]. considering [MLMSMG12]. consistency [KBK^{+}19, Kub17, LC11, LHZ^{+}18, RHH12]. consolidation [MA19, RT18, ZLCZ18]. Constrained [BSH15, CHX^{+}17, JHF^{+}17, JZZ^{+}17, KSK15, LFS16, LL10, Li16, MSK^{+}16, VMMB10, XLL15, YAK15, ZWWX16]. constraint [Ozt11]. constraint-based [Ozt11]. constraints [BMS19, Pet19, SZB16, VRM10, WMY^{+}17, WHS^{+}18, YA11]. Constructing [CFL^{+}19, HS12, Lai15, MKW18, YWW12, WJ12, YSS11]. construction [AK19, CJYC19, CFJW13, HN19, JPD17, JM14, Lai14, Lai17, OOSGVG^{+}16, SB12, WIB12]. consumer [GLGLBG12, KK11]. consumption [AH12, GHY10, GMRRG19, LM16, RTZ11, TKX^{+}13, XLPL19, ZW11]. container [AZW13]. Containers [LACJ18, LAC18, MFT^{+}19, Str12]. contamination [SWLP19]. contended [AFA13]. Content [SLW10, DMI^{+}19, KTP17, KRM14, NKK16, ST12, SK11, WLZ^{+}18, ZW13]. content-based [ST12, SK11, ZW13]. Contention [FCW11, STK11, AEY12, HHS12, KH12, LW16a, SW18, Zah12]. Contention-aware [FCW11, STK11, LW16a]. contention-free [KH12]. Contents [PSGS17]. Context [AHG12, CWZ^{+}18, ORWT^{+}18, Sie16]. Context-aware [CWZ^{+}18, ORWT^{+}18, Sie16]. contexts [KHT^{+}14]. contextual [Ana14]. continued [Ano18v, Ano18w, Ano18x]. Continuous [JHPL13, Luc18, NST19, TCS^{+}10]. contraction [LGK^{+}12, SMH^{+}14]. contractions [IEWK17]. Contribution [AS19b]. contributory [SA19]. control [AA10, AAA^{+}10, BCO^{+}12, BWP^{+}11, BJ18, CG17, CWP12, ESGQ^{+}18, GL12, GAOHG17, HMY^{+}18, JTZZ11, Kim11, KGN11, LWLD12, LL12a, MLZY17, MBO11, MCZ14, RCG^{+}11, SRI14, WRW13, WHS^{+}18, XLW^{+}18, YBM13, YJKD10, ZMZJ17, ZBW^{+}17]. controllable [ZHT16]. Controlled [WXMZ19, BYT19]. convergecast [PLY15]. convergence [CDD^{+}15, PH18]. converging [BHK17]. conversion [FC14]. convolution [CLXX19, XLW^{+}18]. convolutional [ZLS17]. cool [LFS16]. Cooled [SWHB17]. cooling [MLK^{+}16, SWHB17]. cooperation [YQTV12]. Cooperative [LTWW12, SRB^{+}19, SZL10, ADDB18, DDG^{+}17, FCML13, FZ14, GZY14b, KK10, LGM18, TC13, TVT^{+}17, WLL16, WHC^{+}18, XHZ^{+}10, YpGyLlC13]. Coordinated [DDG^{+}17, MCZ14]. Coordinating [LZI^{+}11]. Coordination [FCZ^{+}12, SCN12, SZB16, BDP16, Wu11]. copies [RS19]. Coping [BGBC^{+}16, BCC^{+}18]. coprocessor [KVNV17, SA11, ZMZJ17]. CoQoS [LZI^{+}11]. Core [DDO^{+}18, AFA13, APRA18, AA16, AVAH18, ARI17, ABLP17, BBBC12, BLMB13, CDD^{+}19, CMMT13, CHLL18, CKK^{+}13, CMC^{+}19, DBA^{+}18, DWYB10, FTM^{+}19, GZG^{+}17, GS18, GKS15, Hus17, JHF^{+}17, KSG13, KR11, LWC^{+}18, LKS14, LNAL17, LSC^{+}15, LLS^{+}16, MBBD13, MNR^{+}19, MZC18, MAHKZ12, MGRRK14, NPVG^{+}19, PCMM^{+}17, PGP^{+}12, PTK^{+}13, PR13, RLA^{+}16, RLA^{+}17, SNMB16, SFT^{+}13, Sol13, SAJ13, SHRM19, TCHC12, WQZ^{+}13, WH17, ZXB14, Zha11]. core-periphery [ABLP17]. cores [CVK^{+}18a, LNC13, LTG14, TGPUC16, ZLS17]. correction [LSH^{+}13]. correlations [FX10, WZQ^{+}13]. Corrigendum [KN18b, LSS^{+}11a, MSAZ10a, REK10a]. corrupted [DP16, XSYG18]. cortex [PG19]. cortex-inspired [PG19]. cortical [NFHL13]. cosine [AK19]. Cost [DJDK19, AMU^{+}19, AM12a, AD12, CPLY18, DKUÇ15, ESGQ^{+}11, HS12, JLWX11, KSK15, Li17, MP15, NV19, NML^{+}19, PTN^{+}19, TCMB^{+}19, YGZ^{+}10, YYLC11]. cost-effective [AMU^{+}19, AM12a, JLWX11, NML^{+}19]. cost-efficient [ESGQ^{+}11]. cost-quality [TCMB^{+}19]. CoT [HMY^{+}18]. Count [MPS16]. Counting [AP16]. Counting-based [AP16]. coupled [DXS^{+}19, FPM^{+}14, IEWK17, SMH^{+}14]. coupled-cluster [SMH^{+}14]. course [Bog17, Eck18, LB17, LB18, PSGS17]. courses [FSP18, Kum17]. cover [KO12]. Coverage [ZCW19, Amm16, DGBN14, GM14a, PSRS12, PCX^{+}11, PCX^{+}14, REZN17, YDZ^{+}18]. covered [CHCG18]. covering [KCR14, ST12]. CPS [CHX^{+}17]. CPU [ASSS19, DV13, DBA^{+}18, GKS15, KAA^{+}19a, KLJ^{+}11, LR14, LLKY13, PV19, Ren11, TRS^{+}12, TYA16, UFF19, VLW18, WLL16, WTWZ16, YLL17]. CPU-GPU [DV13]. CPU/GPU [LR14]. CPUs [AVAH18]. CR [LACJ18]. crash [DDG^{+}17, DGDF10, PMHM19]. crash-faults [PMHM19]. crash-prone [DDG^{+}17]. crashes [GK15]. Cray [YQTV12]. credible [YGWJ19]. criteria [ZWWX16]. Critical [LC14a]. criticality [ZZJ^{+}18]. Cross [IEWK17, PQ19, SJS11, WXZ^{+}18, LST^{+}13, WCL^{+}13, YFBY17]. cross-architecture [YFBY17]. cross-layer [WCL^{+}13]. Cross-scale [IEWK17]. Cross-Site [WXZ^{+}18]. Cross-state [PQ19]. crossbar [KK17, ZPK^{+}14]. crossed [CFJW13]. crossing [JD12]. crowd [KDSS18]. crowdsourced [VLGV^{+}18]. crypto [SA11]. cryptographic [ABO^{+}17]. Cryptoprocessor [EHKSS19]. cryptosystems [AVAH18]. CSD [KHT^{+}14]. CTS [ASSS19]. cubes [CFJW13, LFZ^{+}17, LLFJ18, MKW18, WFZJ12, XHZZ16]. cubic [BM14]. cuckoo [CSW^{+}17]. CUDA [BSH15, CB11, Cza13, KRKS11, LYIP19, dlAMCFN12]. CUIRRE [ZSW14]. currency [HBF12]. current [MMCL^{+}17]. curriculum [NDW17]. Curve [LZY^{+}18, SKH15]. curves [ST12]. cut [CRD17]. cuts [LÜ14]. Cyber [HRM17, QGB^{+}17, CWCW18, CSW^{+}17, DZC17, GQZ18, JWH^{+}17, LLWC17, LMXJ18, MMN^{+}18, PST^{+}19, SLG^{+}18, ZXMR18]. cyber-enabled [GQZ18, LMXJ18, ZXMR18]. Cyber-Physical [QGB^{+}17, HRM17, CSW^{+}17, JWH^{+}17, LLWC17]. cyberthreat [KAA^{+}19b]. cycle [LLFJ18]. cycled [LDZ^{+}17, LDZ^{+}14]. cycles [HBAD15, JD12, LdSB^{+}18]. cyclic [SPH13, LY12].

D [AA14, GOH^{+}13, AA16, BDRB14, DJDK19, GRS19, KRKS11, LXLS12, MPG17b, NM17, OGRV^{+}12, PYP^{+}10, SLV19, YA11, ZLS17, Zsa16]. D-NoC [AA16]. DADTA [ZLCZ18]. daemon [BBD18]. DAG [Tam18, XLHT13, ZS13]. DAGs [BSS^{+}13, CDR12, MSV19]. Dandelion [CP10a]. Dandelion-like [CP10a]. Dark [SDS^{+}18]. Dark-Silicon [SDS^{+}18]. Data [AS13, AS15, BS11, CGN^{+}13, LW19, LSWC14, LWWQ18, MD13, Ozt11, SPPA19, Str12, SFC17, SLHS19, XMMD17, ZMCP11, ZTFK16, AAA^{+}15, AMU^{+}19, ASB18, Amm16, AH12, AGWY11, ACPT15, AG12, APK18, AYB^{+}15, AEY12, AK18, ARDQ18, AS18, BFH^{+}17, BCO^{+}12, BEN12, BMLLC^{+}19, CF19, CMR^{+}18, CJYC19, CGC16, CLC^{+}17, CPLY18, CW15, CTT16, CCC^{+}19, Cuz11, Cuz13, DF17, DMG18, DTK11a, Eck18, ECP^{+}18, FCW11, FRM15, GYAB11, GLGLBG12, GM14b, GB11, HTB19, JLY12, JBS14, JHPL13, JHL^{+}18, JWH^{+}17, JdSJC^{+}15, JKV15, KKKG14, KCR14, KSB11, KKTZ13, LWC^{+}18, LL19, LWZZ12, LC11, LY12, LLWC17, LZWZ19, LAS^{+}19, LBT19, LJQ^{+}19, LSZZ15, LWW18, LZY^{+}18, LGK^{+}12, LSZJ15, MB19, MLK^{+}16, MBMC19, NLB^{+}18, NCA^{+}12, NCB^{+}17, NAB^{+}11, NKK16, OWK14, OM10, OJP^{+}18, PST^{+}19, PS14, RBN11, RT18, RB12, Ren11, RMU14, RBA^{+}18, RAN^{+}17]. data [RJKL11, SMW18, SHK19, SCW^{+}18, SCMH13, SWW^{+}17, TH19, TVT^{+}17, TLW18, VETT18, VLGV^{+}18, VMMB10, VRM10, WCWO17, WZZ^{+}17, WWW17b, WCH^{+}17, WW18a, WXMZ19, WG11, WLZ^{+}18, XHZ^{+}10, XSYG18, YBX^{+}13, YZC^{+}19, YAK15, ZV14, ZKZF18, ZLZ^{+}19, ZV12, ZWW17, ZSCX18, ZHT16, ACB^{+}15, LSZJ15, PJ18]. Data-aware [ZTFK16, AYB^{+}15, VMMB10]. data-centric [LAS^{+}19]. data-driven [NCB^{+}17, WLZ^{+}18]. Data-flow [SPPA19]. Data-intensive [ZMCP11, RBN11, WZZ^{+}17, WG11]. Data-oriented [LWWQ18]. database [LLB^{+}18, TR16, XLC^{+}18]. databases [Ale19b, GPSH19, PLK^{+}18]. datacenter [CPLY18, YYWZ19]. datacenters [PRN^{+}19]. dataflow [ESCV15, TBG^{+}17]. DATALET [PST^{+}19]. dataset [YYLC11]. datasets [CLOL17, KSJC17, KN18b, KN18a, YÖ11, YLB^{+}15]. day [TLL^{+}18]. day-ahead [TLL^{+}18]. DCell [WFLJ16]. DCT [VAF19]. DDS [SMPMLVLS11]. Deadline [RCG^{+}11, SCW^{+}18, LFS16, MGSG12]. Deadline-aware [SCW^{+}18]. Deadline-sensitive [RCG^{+}11]. deadlines [WMG13]. Deadlock [BYT19, AA14, XL11]. deadlock-free [AA14]. deal [ESGQ^{+}14]. debug [MH18]. debugging [CV16, LZZ^{+}11]. Decentralised [YZS15, DBCF13]. Decentralized [AM11, DW12, GHK^{+}12, AS18, BHK17, LWH^{+}19, MAPF14, WZQ^{+}13]. Decidability [FP17]. decision [AKK^{+}19, KDSS18, SV18]. decision-making [AKK^{+}19]. declustering [WZZ^{+}17]. decoder [MC17]. decoding [CP10a]. decomposition [CZZ^{+}17, CKLW19, NST19, RSK19, WD18]. dedicated [HLL^{+}19]. deep [CXQ^{+}18, FFYH19, HMY^{+}18, HKK^{+}18, SRB^{+}19, TLL^{+}18, WW18b, WBS19, WDS^{+}18, XCC^{+}19, ZWW17, ZHG^{+}19, MLCFH^{+}18]. defense [WHW^{+}19]. deformable [SQQL19]. degree [BCF14, BPBR11, KSK15, Sta17]. Deister [WZZ^{+}17]. Delay [AH11, GZG^{+}17, Hu11, GL12, Li19b, NLB^{+}18, WW12, WYW15, WHC^{+}18, WHS^{+}18, YA11, YWG15, ZWW17, KSSK16]. delay-aware [WHC^{+}18]. Delay-sensitive [Hu11, NLB^{+}18]. Delay-tolerant [AH11, WYW15]. delays [RWB^{+}13, WPC19]. deliveries [WE13]. Delivery [THGY15, AH11, WLZ^{+}18, ZLT^{+}19]. Dellat [THGY15]. Delta [ASB18, WBS19]. demand [HZDP12, KyLPC17, LSZZ15, NKK16, YYLC11]. demands [SLW10]. Denial [BK18]. Denial-of-Service [BK18]. denoising [TLL^{+}18]. dense [FHL^{+}15, ICQO^{+}12, LKD14, RM10]. densities [LHWJ19]. Density [MC17, BAT^{+}19, WCXL11]. dependability [CJA^{+}19]. dependability-driven [CJA^{+}19]. dependable [NPGV10]. dependence [NCA^{+}12]. dependencies [ZLKK19]. dependency [CSJ^{+}13]. dependency-timing [CSJ^{+}13]. dependent [LSWC14]. deployable [YC12]. deployment [EM11, SMO^{+}18, TWQS12]. depth [YWJ^{+}18]. describe [JWH^{+}17]. description [MRS^{+}14]. descriptor [HN19]. descriptors [LNW^{+}12]. Design [AFA13, AM17, AC16, CTKA17, CKK^{+}13, DR19, JZZ^{+}17, LLKY13, MP10, PD19, PMCC18, RBG17, SDS^{+}18, Sol13, TTH12, ZPK^{+}14, Ada17, ABLP17, BBH^{+}17, BMS19, CG11, CSJ^{+}13, CK13, CHX^{+}17, DFHH13, EFG^{+}14, FHL^{+}15, FCG^{+}14, HKK^{+}18, KMC16, LÜ14, MCM^{+}11, NV19, ORWT^{+}18, OMT^{+}17, RA11, SDS10, VRGS17, VLL^{+}14, Wu11, ZMZJ17, ZY12]. designed [BSH15]. Designing [BBBC12, FSP18, GMS^{+}13, YCH^{+}10, YFBY17]. designs [MC17, Man13, PGRP17]. Desktop [LSH^{+}13, AAD10]. detect [UGG^{+}11]. Detecting [CL14, SKK14, YXX13, TAM^{+}19]. Detection [SQQL19, SJS11, YHWY18b, AKK^{+}19, AFD^{+}11, CDW^{+}19, DKKV15, ECWV19, GDCC18, GHIJ19, HMY^{+}18, IZ12, KCFP18, Ksh12, KKTZ13, LLLC15, LLWC17, LHLM14, NHO^{+}13, PMHM19, PH16, RLP14, ST12, SMP17, TRS^{+}12, TY17, TCS^{+}10, WL11, WML^{+}18, WXZ^{+}18, XL11, XTN12, XSYG18, YDZ^{+}18]. detector [DMI^{+}19]. detectors [AAI^{+}15, BGBC^{+}16]. deterministic [GTGLSA12, MPR19, WZZ^{+}17, ZLWL12]. Development [KHT^{+}14, AM17]. device [PVP18, VFAD17]. devices [Kim17, MXSL12, WCF14]. DGIN [KMC16]. DGIN-3 [KMC16]. DHT [CTT16, HASB16]. DHT-based [CTT16]. DHTs [GTGLSA12, SAL10]. diagnosis [CJA^{+}19, CAF^{+}11, VS18]. Diagnostics [DMG18]. diameter [BBD18]. difference [CLXX19, SS11]. different [CG17, LCB16, MFT^{+}19]. Differential [CDD^{+}19, URK^{+}19, WLYS19, WRW13]. differentiation [MCZ14]. Diffusion [DM17, HES11, ZXGD18, Zsa16]. diffusion-drift [HES11]. diffusion-limited [Zsa16]. dilation [LST17]. Dimension [CFJW13]. Dimension-adjacent [CFJW13]. dimensional [ANEA13, CMR19, DTK11b, GB11, HN19, HS17, KKN13, KN18b, KN18a, LZY11, LZWZ19, LDS16, WRW13]. dimensionality [BV13]. dining [AFNT17]. DINO [RMHR17]. Direct [FLC14, MRJ^{+}19, SWHB17, PPTV^{+}10, Tam18]. directed [MTM10, TDP15]. Direction [BEN12, MSAZ10a, MSAZ10b]. Direction-based [BEN12, MSAZ10a, MSAZ10b]. directions [ACB^{+}15, PSC^{+}16]. Directive [MM15]. Directive-based [MM15]. directory [SB15, VRGS17]. disaster [SZB16]. Disconnected [GHIJ19, MCS14]. Discovering [TFV19]. discovery [CKC19, FZ14, REZN17, RSL12, SMPMLVLS11, SK11, ZAB18, ZMG^{+}16]. discrete [AZC13, IIH16, Li16, PQ19, SS17]. discretization [SWLZ17]. disease [VS18, ZXGD18]. disjoint [HBAD15, KMC16, Lai14, Lai15, Lai17, MT14, SMP17, WFLJ16]. disk [GGY19, JPD17, KR12, NC13, NZY^{+}11, SRT^{+}18, XS11]. disk-assisted [SRT^{+}18]. Disks [KR11, RWB^{+}13]. dispatch [YZS15]. disruptive [SI13]. dissemination [AHZ11, DF17, LWW19, MSF^{+}13, WW18a]. Distance [NM17, Tur12]. distance- [Tur12]. distinguishability [ZCW19]. Distrib [KN18b, LSS^{+}11a, MSAZ10a, PCX^{+}14, REK10a]. Distributed [AAA^{+}15, AK17, AS13, ABLP17, BWP^{+}11, CG11, CFI^{+}18, CWCW18, CWD11, Cuz11, DFLO17, DDE19, FTM^{+}14, FPS11, GHY10, GMSS^{+}11, GZY14a, HWLR14, HWY^{+}10, JPD17, JKD^{+}15, KSSL16, KDO^{+}13, KKH17, KSK15, KRS13, KKTZ13, Las12, LLWC17, LACJ18, LK11, MSGS^{+}13, MNK12, OY13, PSRS12, PBB^{+}17, PRS14, SHSH17, SLG^{+}18, TH11, TT10, TSFZ14, Ale19b, ACCP12, AAI^{+}15, AM11, BBCQ13, BDL^{+}19, Ben19, BCMV15, BOKS19, BHLT14, BFL^{+}13, BLZ^{+}18, BJ18, CG12, CGL^{+}14, CRHC19, CVK^{+}18b, CKWT17, CKLW19, CTT16, Cuz13, DB11, DLM19, DKM10, DTK11a, Eij18, EHL^{+}15, ES12, FFYH19, FPF14, FKR^{+}17, Fu10, FLC14, FLM^{+}19, GYAB11, GL12, GN15, HD10, HLL^{+}19, HHK15, JF12, JKIE13, JZZ^{+}17, Kak15, KHW13, KSG13, Kim11, KKS^{+}12, KCFP18, KS13, LL19, Las13, LASS15, LdPLC^{+}19, LL18, LWW19, LYIP19, Lop13]. distributed [Lop18, LS19, LSZJ15, MLZY17, MAPF14, MA11, MBMC19, MPR19, NSAS10, NL19, NML^{+}19, NTN12, NDW17, NSDZ18, OPR18, PKN10, PC11, Pet19, PH16, PMdO11, PGKV18, PRN^{+}19, RLP14, RAN^{+}17, RDA18, SSKS11, SW12, SMP15, SB15, SCMS12, SS18, SCMH13, ST14, SLKK13, Suk18, TLLV10, TXLL14, TM10, TVT^{+}17, TWQS12, WL11, WML^{+}18, WHC^{+}18, WD13, WSLC11, WZQ^{+}13, YZS15, YHWY18a, YHWY18b, YLB^{+}15, YZG18, YWG15, ZAB18, ZLKK19, ZZJ^{+}18, ZCMY12, ZTFK16, ZBW^{+}17, ZHG^{+}19, DLLL11]. Distribution [SLW10, KG19, Li17, NM17, SFHS19, WZZ^{+}17, gWW18, YJL16]. distributions [BKMT14, PCX^{+}11, PCX^{+}14]. DITVA [KCSS18]. Divergent [RMHR17]. diversity [SSFP11]. divide [GDL^{+}11, PV19, TP18]. divide-and-conquer [GDL^{+}11, PV19]. divisible [BD11, CG12, HV13, KVA18, LML^{+}10, MLDG12]. division [Dav17, HRG^{+}11]. do [LTG14, KMS10]. Do-All [KMS10]. Docker [MFT^{+}19]. document [SGVRP19, UGG^{+}11, ZMCP11]. document-similarity [UGG^{+}11]. doing [MBG^{+}17]. DOM [WXZ^{+}18]. Domain [CZZ^{+}17, KRS13, KRS14, MRS^{+}14, RMGM19, SS11, WMC^{+}18]. Domain-Specific [KRS13, KRS14, MRS^{+}14, RMGM19]. dominated [AM12b]. dominating [JPD17, YSS11, YWW12]. domination [GK10]. dOpenCL [KSG13]. Double [GVBB13, XLHT13]. Douglas [HGX^{+}19]. DPI [HVW16]. dragonfly [BFVB19]. DRAM [ZLH^{+}18]. DRAM/NVM [ZLH^{+}18]. Drawing [DP12]. drawings [JD12]. drift [HES11]. drive [LTG14]. driven [ASES15, CJA^{+}19, CTT16, LWZZ12, LS10, LJQ^{+}19, LGK^{+}12, MBS^{+}12, NCB^{+}17, SS18, TLQS12, TAM^{+}19, WLZ^{+}18, XLL15]. drives [GFPC14]. DSL [MRJ^{+}19]. DSL-based [MRJ^{+}19]. DSP [DSEP17]. DSPONE48 [DSEP17]. DTNs [MPS16]. Dual [ACCP12, LSXX14]. duplication [STK11, TLLL10, WCEA10]. duty [LDZ^{+}17, LDZ^{+}14]. duty-cycled [LDZ^{+}17, LDZ^{+}14]. DV [CSW^{+}17]. DV-Hop [CSW^{+}17]. DVFS [CG17, ECLV12, LSC^{+}15, RTZ11]. DVFS-based [RTZ11]. DVS [ZHLQ12]. DVS-enabled [ZHLQ12]. Dwarf [DTK11a]. Dynamic [AAD10, ANEA13, CDAN14, DB11, GM14b, KCSS18, KR10a, KVA18, MD13, MB19, NPVG^{+}19, OOSGVG^{+}16, SHSH17, SS17, TT10, TDP15, AKSZ19, BBCQ13, BCF14, CSMML10, CPLY18, CKML12, CWD11, DLW^{+}12, GÖÖ16, GFPC14, JBA15, KZ11, LGZ^{+}10, Li10, LLY15, LFEP19, LLW12, MYYY17, MCS14, NDP13, NLB^{+}18, NHO^{+}13, PKN10, PW17, RK18, RCG18, SNMB16, SSM^{+}16, SSdlB^{+}10, SZB16, TW15, TMK^{+}17, WW12, WXZ^{+}18, XLC^{+}18, YS11, ZXYO11, ZCS^{+}18]. dynamic-warp [NHO^{+}13]. Dynamically [PPP14, GK15, MSV19, ORWT^{+}18]. dynamics [CDPS18, DAG^{+}17, LHWJ19, LGRV19, PARB14, PTK^{+}13, WYTX13].

e-infrastructure [HPB^{+}10]. e-payments [CSS11]. e-voting [LWH^{+}19]. Eagle [KS18]. Early [GRJ^{+}15, AMT13]. early-stopping [AMT13]. ECG [ZAAB17]. ECHO [HASB16, SAL10]. ecosystem [LZN19]. EDF [dOCS14]. Edge [AMU^{+}19, HAC^{+}19, HRH18, HBAD15, TPS^{+}18, TPJ^{+}19, WZX^{+}19, Ale19b, Ale19a, JTC^{+}18, KCFP18, LBT19, LYJ^{+}19, LWWQ18, MS19, MA19, PRN^{+}19, Udd19, WZH^{+}19, YWJ^{+}18, ZCS^{+}18, ZGW^{+}19]. Edge-of-things [AMU^{+}19]. edge/cloud [Ale19b, MA19]. edges [BKCM17]. editing [LHCC19]. Editor [Pra16]. Editor-in-Chief [Pra16]. Editorial [AS15, Ano18v, Ano18w, Ano18x, Ano18q, CGFH19, Hol17, SLL18, STS19, DF12, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h]. Editorial [Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18r, Ano18s, Ano18t, Ano18u, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x]. Editors [Ano16k, TFV^{+}15]. Edu [PGKV18]. Edu-2016 [PGKV18]. educating [LMB^{+}17]. education [APV18, BLZ^{+}18, CVK^{+}18b, Hua17, MBG^{+}17, Nee17, NKSA17, NSDZ18]. Effective [KO11, AMU^{+}19, AM12a, BV13, BCK^{+}13, Cza13, DJDK19, FZWL12, FWM^{+}10, JLWX11, KHW13, LJQ^{+}19, NML^{+}19, SNCP12, WMY^{+}17, YCH^{+}10]. effectiveness [MA11]. Efficiency [LdSB^{+}18, AHG12, AG12, BC11, BYH^{+}17, ESCV15, FRM15, FCP^{+}15, HRM17, HJLR12, LB12, LHW^{+}19, PB19, Ren11, SWHB17, SHC14, VETT18]. Efficient [AZC13, BCO^{+}12, BM16, CWP12, HASB16, IR12, JBS14, KK10, KS13, KR11, LHHH11, LDP^{+}14, LJZ^{+}19, LZWZ19, LW19, MLDG12, MB13, ND12, PP13, Pen11, RV13, Rao16, RMU14, SMP15, Sch13, SSGG18, SMP17, SWLZ17, SCLL10, Tur12, UFF19, WCCH18, XLH18, YTH^{+}19, ZHG^{+}19, dSAJ15, AAH17, AFA13, ARI17, Ale19a, ACA^{+}19, Ara13, AS19b, BFH^{+}17, BM11, BKC^{+}15, BK13, BDL^{+}19, BOY10, BBD18, BCK^{+}13, BHK17, CWZ^{+}18, CMR^{+}18, CP10b, CMN12, DKM10, ESGQ^{+}11, EDH^{+}17, FTM^{+}19, GDCC18, GKS15, GYP13, HSS10, HBSASA19, HZHS18, IEWK17, KTP17, KVA18, KyLPC17, KHK18, Kol19, KSSK16, LK13, Lai14, LW16a, LLB^{+}18, LSC^{+}15, LZY^{+}18, LL18, LCJ^{+}18, LLDL15, LFEP19, MGSG12, MNR^{+}19, MSF^{+}13, MPS16, MPN17, MAHKZ12, MCP^{+}18, NMS^{+}18, NF16, PPSV15, RM11, RLA^{+}16, RLA^{+}17, RFS^{+}12, RT18, RGAN18, SB12, SZMK13, SJG19]. efficient [Tam18, TLY12, TGPUC16, TMK^{+}17, TLL^{+}18, UBES10, VRGS17, VAF19, WLST16, WTWZ16, WHC^{+}18, WIB12, WH17, gWW18, XLC^{+}18, XCC^{+}19, XHZ^{+}10, YSS11, YLB^{+}15, ZCMY12, ZLL14, ZSCX18, ZWWX16, ZLCZ18, ZHLQ12, ZTGL17]. effort [QGZP17]. EFS [MSK^{+}16]. egress [MCAS12]. eigenanalysis [TYA16]. eigensolver [ABGV11]. Eisenstein [HBAD15, HS17]. Elastic [FGG17]. elasticity [LHNBB19, MMVL11]. elderly [HRM17]. election [DLV11, DGDF10]. electric [AK18, AKSZ19]. electricity [TLL^{+}18]. electron [DAG^{+}17]. electrophysiological [HES11]. Element [PPTV^{+}10, FC14, Ren11]. elimination [ESGQ^{+}11]. elliptic [SKH15]. ELLPACK [ZGG^{+}14]. ELLPACK-based [ZGG^{+}14]. ELM [CLOL17]. EM-KDE [EHL^{+}15]. embedded [BM17a, CNLGRL18, CRJ10b, FWM^{+}10, GZG^{+}17, HLL^{+}19, LLLC15, LCB16, MGRRK14, NL19, XLL15, XLPL19, YZX11, FWM^{+}10]. Embedded-TM [FWM^{+}10]. Embedding [PW16, CRHC19, LFZ^{+}17, PW17, YLZW18]. embeddings [LLFJ18]. emergency [HPB^{+}10]. emerging [BKC^{+}15, KHT^{+}14]. Emitter [FPM^{+}14]. Emitter-coupled [FPM^{+}14]. Empirical [LGK^{+}12, NXTK17]. Employing [PKW^{+}10]. Emulating [KMS10]. emulation [LST17]. enabled [CSL15, GQZ18, GRJ^{+}15, HTB19, LMXJ18, NML^{+}19, SLZ^{+}19, TODQ18, ZXMR18, ZHLQ12]. Enabling [ETS14, FCG^{+}14, JKIE13, SA19, TT10, ZCF^{+}17, DKKV15, HRH18]. encoders [TLL^{+}18]. encoding [CP10a, WLCZ15, ZWQ^{+}16]. encrypted [LZWZ19, SWW^{+}17, ZHT16]. encryption [WXMZ19, WCCH18, ZAAB17]. End [Ano10a, Ano10b, Ano11j, Ano11k, Ano12m, Ano12n, Ano14f, Ano14g, Ano15k, ZLCJ12, CF19, CXQ^{+}18, ORWT^{+}18, RSVW19, WG11, XLL15]. End-to-end [ZLCJ12, RSVW19, WG11, XLL15]. enDebug [CV16]. endurance [WCWO17]. Energy [ACA^{+}19, AKSZ19, BOY10, BYH^{+}17, DKM10, FWM^{+}10, GQZ18, GYP13, KR12, LK13, LBMG15, LL10, LW16a, Li16, LNAL17, LSC^{+}15, LY13, MGSG12, MTL^{+}18a, NMS^{+}18, RM11, SP13, SSGZ13, WHC^{+}18, WH17, XHZ^{+}10, ZZJ^{+}18, AS19a, AHG12, AK18, CV16, ECLV12, FRM15, FCJG^{+}18, FCP^{+}15, GHY10, GDCC18, GMRRG19, GL12, GPSH19, HRM17, JZZ^{+}17, JZF^{+}15, KR10a, KyLPC17, KCR14, KSSK16, LR14, LL12b, LLCZ19, Li19b, LZC11, LLDL15, LCB16, LFEP19, MNR^{+}19, MMK^{+}11, NS12, OMT^{+}17, PCMM^{+}17, PB19, RWB^{+}13, RLA^{+}16, RLA^{+}17, RFS^{+}12, RT18, RTZ11, TLY12, URK^{+}19, UMM^{+}18, VRGS17, WLST16, gWW18, XS11, XLPL19, YL12, YZS15, YAK15, ZW11, ZWY^{+}15, ZWWX16, ZLCZ18, ZHLQ12, MSK^{+}16]. Energy-aware [GQZ18, LBMG15, LNAL17, LY13, FCJG^{+}18, LR14, LLCZ19, MMK^{+}11]. energy-constrained [JZZ^{+}17]. Energy-efficient [DKM10, GYP13, LK13, LW16a, LSC^{+}15, MGSG12, NMS^{+}18, WHC^{+}18, WH17, XHZ^{+}10, GDCC18, KyLPC17, KSSK16, LLDL15, MNR^{+}19, TLY12, VRGS17, WLST16, ZHLQ12]. Energy-Friendly [MSK^{+}16]. energy-performance [ECLV12]. energy/power [OMT^{+}17]. energy/power-aware [OMT^{+}17]. enforcing [Kub17]. engine [HVW16, XTN12, XP10]. engineer [GS18]. engineering [BCD^{+}15, CCE^{+}17, Nee17]. enhance [DZC17]. Enhanced [MD13, RK18, LLDL15, dOBG^{+}15]. EnhancedBit [ARD14]. enhancement [KS18, NGQM12, TBG^{+}17]. enhancements [ESGQ^{+}18, LÜ14]. Enhancing [CGN^{+}13, GRR13, HWLR14, dAMFdS13, MH18, OM10, QGZP17, VETT18]. ensemble [KBC19, SV18]. entangled [EAB^{+}19]. enterprise [GSASA19, LSH^{+}13]. entity [MPN17]. entropy [DFHH13]. enumeration [WCH^{+}17]. Environment [CSMML10, GYAB11, AKSZ19, Ben19, BLZ^{+}18, JLWX11, KK10, LL18, MYYY17, MLK12, PST^{+}19, SSKS11, VD18, WD13]. Environment-conscious [GYAB11]. environments [CF19, CPJ^{+}19, CTKA17, DWX10, ECP^{+}18, ECLV12, FRM15, FCJG^{+}18, FMIF18, KV10, KLJ^{+}11, KCFP18, Ksh12, LSH^{+}13, LML^{+}10, LSWC14, MSRB19, NPS^{+}19, SJB12, SZB16, SZL10, SJS11, TZI11, WMES12, WG11, YWG15, ZLWZ18]. Ephemeral [AGMS16]. epidemic [AHZ11, LpJS^{+}18, MSF^{+}13]. epidemiological [Rao16]. EPLS [CLC^{+}17]. equation [JGMY17, WJ14]. equations [SPH13, Ter16]. equivalencing [ES12]. era [MBG^{+}17, SC10]. Erlang [CLG^{+}16]. error [BGBC^{+}16, DFHH13, OWK14]. error-prone [OWK14]. error-resilient [DFHH13]. errors [BCC^{+}18]. establishing [GPJA10]. establishment [SZMK13]. estimate [BKK^{+}11]. estimates [TDBL13]. Estimating [LGL13]. Estimation [GMRRG19, KC17, gWW18, ZRN^{+}14, DLLL11]. estimator [SIY14]. evaluate [dOCS14]. Evaluating [AFNT17, Ale19b, MMN^{+}18]. evaluation [ASHO19, AMSÅ19, AB13, CTKA17, CB11, dADC18, DR19, DMS^{+}16, HBS17, HTB19, LZY11, LNW^{+}12, MGRRK14, PMCC18, SA11, Sol13, SE15, WLZ^{+}18, YL12]. evasion [YpGyLlC13]. Event [ECP^{+}18, AZC13, BM17b, CM12, FX10, JKD^{+}15, PQ19, SW12, WZQ^{+}13]. Event-based [ECP^{+}18]. events [PQ19]. everybody [KSSK16]. Everything [NPS^{+}19]. EvoDeep [MLCFH^{+}18]. Evolution [CDD^{+}19, RBB17, HWY^{+}10, Li10, SV18, WRW13]. Evolutionary [YLZW18, AS19a, COF^{+}17, HD10, MLCFH^{+}18, RPN19, Tal19]. evolvable [KKKP12]. exact [GA18, PB15, XP10]. exascale [APV18, CCAAS19, RPS19]. exchange [CMR^{+}18, ECP^{+}18, NKK16, PW16]. exclusion [Ara13, LASS15, NTN12, RDA18]. exclusive [DMI^{+}19, WW18a]. executing [Sol13]. execution [CWCW18, ESCV15, GYY^{+}14, LFS16, LR14, LPK^{+}10, Li19b, MTL^{+}18b, PP13, PSB^{+}19, WLST16, uRIL^{+}18, dKG^{+}10]. executions [FCP^{+}15, KVNV17, RV13]. exercises [Suk18]. expanding [RM10]. Expansion [LY12]. Expectation [YZG18]. Experience [FTK14, MSRB19, LBT19, NGQM12]. experiences [GRJ^{+}15]. Experimental [CKT11, dADC18, GHC^{+}17]. experiments [LYW^{+}16]. explicit [RCG^{+}11, Rao16]. exploit [YCH^{+}10]. exploitation [SWLP19, VFAD17]. Exploiting [CDD^{+}19, CB15, FTM^{+}19, LKS14, MFT^{+}19, RGAN18, VBF13, WYTX13, ZLWL12, CDAN14]. Exploration [SDS^{+}18, BKC^{+}15, CKK^{+}13, LLKY13, NV19, OT19, TKKH17]. Exploring [ARP18, NXTK17, PCMM^{+}17, ROB^{+}18]. express [APRA18]. expression [CLXX19]. expressions [LGK^{+}12]. expressiveness [HdR13]. extended [LWWQ18, YWW12]. Extending [CMR10, EMC19]. Extensible [FLCB10, HGFF10]. extracting [BCH15]. extraction [CLC^{+}17, LLS^{+}16, MM15, WZY^{+}19, dAT17]. extractive [SGVRP19]. Extrapolated [DM17]. extremal [FSV14]. Extreme [SFT^{+}13, YZW^{+}15].

fabric [PD19]. fabrics [ZRN^{+}14]. face [CMN12, NHO^{+}13]. Facial [CLXX19]. facilitate [Udd19]. factorization [CLW^{+}19, CASD18, FHL^{+}15, OT19, SLV19]. fail [BCC^{+}18]. fail-stop [BCC^{+}18]. failovers [SI13]. Failure [AAI^{+}15, Fu10, JAB12, BKMT14, FX10, JKIE13, KV10, LGZ^{+}10, PCLP16, YHWY18b, JKIE13]. Failure-aware [Fu10, JAB12]. failures [DGDF10, LY10, SCMS12]. Fair [BHLT14, KNHH18, Tau16, AS19b, LASS15, SPC^{+}17, SCG10, ZLL14, ZQMM11]. Fair-share [KNHH18]. fairness [Ara13, SHC14, ZLCJ12]. false [LLWC17]. families [FSV17]. farms [JTZZ11, MCP^{+}18]. Fast [BV13, CXX^{+}18, CDY^{+}19, DMK19, HZA^{+}15, KSSG14, LYIP19, SWLP19, TPLY18, ZLZ^{+}19, AGMS16, BBBC12, BFKW13, BHK17, Can18, NV19, PH16, WWW17a, WZY^{+}19, WJ12, XLH18, CVK^{+}18a]. Fat [Zah12, ESGQ^{+}11, ESGQ^{+}14, YMLP14]. Fat-tree [Zah12]. fat-trees [ESGQ^{+}11, ESGQ^{+}14, YMLP14]. Fattened [GMVRGS16]. Fault [ABBD14, BKMT14, CLMRL15, DBCF13, JBA15, LFZ^{+}17, MPG17b, XHZZ16, AA14, AA16, ANEA13, ARVZ14, BJ15, CAF^{+}11, DTK11a, FABG^{+}19, JBS14, KG10, LHLM14, LFGM17, LAC18, MSEM^{+}19, PD19, PAS15, TCHC12]. fault-tolerance [BJ15]. Fault-tolerant [ABBD14, BKMT14, JBA15, LFZ^{+}17, XHZZ16, AA14, AA16, ANEA13, DTK11a, FABG^{+}19, JBS14, KG10, TCHC12]. faults [CP17, LLFJ18, PMHM19]. FCFS [Ara13]. FDTD [SS11]. feasibility [MAKWZ13, RB12]. Feasible [ESGQ^{+}18]. feature [CLC^{+}17, DKC14, LLS^{+}16, PLSM18]. features [CGC16, LMXJ18, dAT17]. federated [SJB12, TODQ18]. federated-IoT-enabled [TODQ18]. federation [CTC^{+}10]. Feedback [MTM10, HWL18]. Feedback-directed [MTM10]. few [Sch14]. FFT [DMK19, JGMY17, VAF19, WJ14]. field [EAB^{+}19]. fields [LdPLC^{+}19]. File [HWLR14, ZLH^{+}18, CTC11, DT11, DLW^{+}12, KYS13, LHZ^{+}18, MXSL12, No12, SSX14, WZZ^{+}17]. files [ARDQ18, WJ12]. filling [ST12]. Filter [VRGS17, SMPMLVLS11]. filter-based [SMPMLVLS11]. filtered [LKB^{+}15]. filtering [BW18, ZLJ^{+}19]. financial [CLZ19, PVRS17]. Finding [FSV14]. Fine [Man13, MPV12, CHLL18, ZCF^{+}17]. Fine-grained [Man13, CHLL18, ZCF^{+}17]. finite [FC14, LWCC15, SS11, PPTV^{+}10]. finite-difference [SS11]. FireGrid [HPB^{+}10]. first [DAG^{+}17, MB13, MAKWZ13, SWHB17, ZCS^{+}18]. first-principles [DAG^{+}17]. fission [GÖÖ16]. fitness [YXW^{+}18]. fitting [LZY^{+}18].

FIWARE [AMSÅ19]. Fixed [KP17, MT14]. flabellate [LSS^{+}11a, LSS^{+}11b]. flags [TdAR18]. FLAME [ICQO^{+}12]. flash [No12]. Flexible [JWSG14, CS17, HCM11, LL12a, LFEP19, SDS10]. flip [LDS16]. Floating [CNLGRL18, Can18, Dav17]. Floating-point [CNLGRL18]. Flocking [TWQS12]. Flooding [BCF14]. Flow [JBA15, CWP12, JTZZ11, KM17, LFEP19, NPE^{+}19, SPPA19, ZCW19]. flow-level [NPE^{+}19]. flows [VBDRC13]. flowshop [CB11]. fluid [CVK^{+}18a, LGRV19]. fluids [JdSJC^{+}15]. FluteDB [LLB^{+}18]. FMM [LPLFMC^{+}12]. FOCAN [NPS^{+}19]. focus [DSEP17]. focusing [FSP18]. Fog [NPS^{+}19, NML^{+}19, JHL^{+}18, WML^{+}18, SAR^{+}18, MSRB19]. fog-based [WML^{+}18]. Fog-supported [NPS^{+}19]. folded [Lai14, Lai17]. food [CXX^{+}18]. footprint [MSV19]. foraged [PST^{+}19]. forces [LHWJ19]. forecasting [TLL^{+}18]. forest [CFL^{+}19]. ForestLayer [ZHG^{+}19]. forests [ZHG^{+}19]. form [NCB^{+}17]. formal [Ben19, SHSH17]. formalism [MBO11]. Formalization [BFL^{+}13]. format [ZGG^{+}14]. formation [KSK15, YZS15]. Forward [Lla17, dOBG^{+}15]. Forwarding [AD10, Ana14, HDCM11, KHK18, LWW18, STMZ18]. fossil [SWLP19]. Foundations [BFL^{+}13]. Fourier [CVK^{+}18a]. FPGA [CNLGRL18, CS17, CGFH19, HBS17, IIH^{+}17, KG19, MH18, NV19, NSKN17, PD19, Pet18, SA11, TYA16, TOR^{+}14, WLCZ15, WIR^{+}18]. FPGA-based [HBS17, IIH^{+}17, NSKN17, WIR^{+}18]. FPGAs [AD12, LdSB^{+}18, MC17, NMS^{+}18, WD18]. Fractal [ASKTZ13, NST19]. frame [SCG10]. Frames [LNA12]. framework [AAA^{+}15, AMU^{+}19, Amm16, AM12a, AC16, BK13, BMT12, BJ18, CCAAS19, CCA18, CF19, CV16, CHX^{+}17, CDPS18, DV13, GM13, GFPC14, HSH10, HRM17, HRH18, KTP17, KKS^{+}12, KBC^{+}10, LV15, MCM^{+}11, Men18, MBR19, NLB^{+}18, NPE^{+}19, PMAL11, PAG^{+}18, RBN11, ROB^{+}18, SAL10, SMH^{+}14, SGdSS13, TLW18, VS18, WTWZ16, WHW^{+}17, WXZ^{+}18, WMG13, YLB^{+}15, ZGW^{+}19, dAT17]. Frameworks [KRS13, KRS14, DAB^{+}14, LHNBB19, UMM^{+}18, uRIL^{+}18, ZKZF18]. free [AA14, AKBD10, ACH18, Dav17, FKKR16, FLM^{+}19, HSY10, JBS14, KH12, LASS15, LWW18, MYM10, MBMC19, MKM16, Pen11, SSdlB^{+}10, VBDRC13, Zah12, dOBG^{+}15]. free-surface [VBDRC13]. FREP [KR12]. frequencies [LdSB^{+}18]. frequency [MYD^{+}11, RTZ11]. frequent [LT10, YZG18, BMLLC^{+}19]. Friendly [MSK^{+}16]. Frog [KM17]. front [ORWT^{+}18]. front-end [ORWT^{+}18]. FSI [KHT^{+}14]. fuel [SWLP19]. Full [Ano18y, Ano18z, SWW^{+}17]. full-text [SWW^{+}17]. Fully [WZY^{+}19, DTK11a]. fully-distributed [DTK11a]. function [ABO^{+}17, BNBR16, LRS18]. functional [WMY^{+}17, WD18]. functions [AMT13, CMR^{+}18, MM15, RMU14, SJVRVVS19, WD18]. fusion [ECP^{+}18]. Future [HAC^{+}19, MNK12, PJ18, ACB^{+}15, ECLV12, LY13, MKN14, PSC^{+}16, RPS19]. Fuzzy [DFLO17, TZI11, KKTZ13, SMO14, ESCV15].

G [GDL^{+}11, GA18]. G-PaMeLA [GDL^{+}11]. G/M/1 [GA18]. G/M/1-type [GA18]. Game [KK10, PC11, JTC^{+}18, RPN19, YpGyLlC13]. Game-theoretic [PC11]. gamma [KMC16]. gap [KLJ^{+}11, KR17]. gas [OGRV^{+}12]. gate [EAB^{+}19, NKV14, WCF14]. gateway [KKKP12]. Gathering [PMHM19, JLY12]. gating [CZPP16, ZCF^{+}17]. Gauss [Dav17]. Gaussian [HAC17, WL11]. GbE [LB12]. GCHAR [CWZ^{+}18]. GEMM [JM15]. gene [WCEA10]. general [AZW13, CW15, GRS19, GFPC14, LV15, LCB16, MSAZ10a, MSAZ10b, Pet19]. general-purpose [GRS19, LCB16]. generalized [HSH10, WRW13, ZHW19, ZLMC14]. generated [MTM10]. Generating [AAK^{+}13, SCMH13]. generation [BCK^{+}13, FLM^{+}19, GNZ18, HPB^{+}10, HZZ^{+}19, LK13, NAB^{+}11, ORWT^{+}18, Zsa16]. generator [KCP19, Pet18]. generic [Ben19, GM13]. Genetic [OA10, ALM^{+}16, ANEA13, AB13, DK11, LFEP19, PKN10]. genomes [SPRG^{+}12]. genre [WIR^{+}18]. Geographic [AD10, SJS11]. geometric [CMN12, LZWZ19, MRS^{+}14, MPR19, TSFZ14]. geometry [QGZP19, WZY^{+}19]. GET [HLS12]. GET/PUT [HLS12]. GHSOM [IZ12]. global [GVBB13, MS15, WWW17a, Zah12, ZLWZ18, dOCS14, YQTV12]. globally [CWP12, NZA13, LNA12]. globally-aware [CWP12]. Goal [CJ17, XLPL19]. Goal-based [CJ17]. goals [TdAR18]. Godson [PTK^{+}13]. Godson-T [PTK^{+}13]. Golgi [FTM^{+}19]. GOM [YLB^{+}15]. GOM-Hadoop [YLB^{+}15]. Google [DKC14]. Gossip [FCML13, AS18, LT10, WWW17a]. Gossip-based [FCML13]. GPGPU [DFST13, KWZ19, OGRV^{+}12, SJVRVVS19, WMG13, YPCW16]. GPGPUs [AFK14, DKK18]. GPS [AKBD10, LWW18]. GPS-free [AKBD10, LWW18]. GPU [YJL16, ARP18, BCMV15, BDRB14, BFKW13, BMS19, BHS13, CDD^{+}19, CSL15, CMMT13, CRHC19, CMR19, CW15, DV13, DBA^{+}18, DFHH13, DCA^{+}15, Eme13, FSV14, FSV17, GMMP12, GLW14, GKS15, GMS^{+}13, GRS19, GK19, HVW16, IIH16, JGMY17, JdSJC^{+}15, KP17, KAA^{+}19a, KKN13, KCP19, KC17, LR14, LLKY13, LST^{+}13, LW19, LPLFMC^{+}12, LFEP19, MB13, MRT18, NFHL13, PV19, PDP17, PDB13, RV13, RS19, Ren11, RMU14, ROB^{+}18, Sch13, SS11, SCMH13, SDG17, Ski16, TH11, TRS^{+}12, TYA16, VBDRC13, VLW18, WLL16, WD13, WH17, XLH18, YLL17, ZMCP11, ZHH15, ZWQ^{+}16, dSAJ15, dMS18]. GPU-accelerated [DCA^{+}15, Eme13]. GPU-based [BCMV15, BDRB14, BFKW13, DBA^{+}18, GMMP12, KCP19, PDP17, Ski16]. GPU-Investigations [Sch13]. GPUDirect [ARP18]. GPUs [AVAH18, ASES15, BBBC12, BBR13, BCK^{+}13, COV13, CGN^{+}13, CDY^{+}19, DP16, GOH^{+}13, JM15, LMGLGLG17, LJZ^{+}19, LW16b, LV15, MLW^{+}19, MBW16, NSKN17, NHO^{+}13, PVRS17, SBRM19, TH13, ZSW14, ZGG^{+}14]. Graceful [AA14]. gradient [GLW14, LR14]. Gradual [ADDP19]. graduate [APV18]. grained [CHLL18, Man13, MPV12, ZCF^{+}17]. Grammatical [RBB17]. grand [SIY14]. granularity [WCL^{+}13]. GRAP [FGL^{+}11]. graph [AK19, BKC^{+}15, BCK^{+}13, CSJ^{+}13, CFL^{+}19, DCA^{+}15, FLM^{+}19, GHC^{+}17, KSSG14, LK15, NXTK17, PV19, PS14, RGAN18, SSKÇ15, SMT15, WCH^{+}17, YFBY17, ZCS^{+}18]. graph-partitioning [GHC^{+}17]. graphene [KRM14]. graphene-CMOS [KRM14]. graphic [SKH15]. Graphics [BHS13, DDGK13, ATDH13, BK13, CLA^{+}18, PYP^{+}10, SIY14, ZMCP11, Eme13, GLGLBG12, YL12, YJL16]. graphs [AAK^{+}13, BCF14, BKCM17, CDDL10, CDS10, DM17, GK10, JPD17, LKB^{+}15, MPR19, NCA^{+}12, STMZ18, TBG^{+}17, Ten16, TSFZ14, WWW17a]. Grasping [KR17]. greedy [KHW13, STMZ18, dOBG^{+}15]. Green [DAPR18, AG12, BFH^{+}17, SWLP19, WCL^{+}13]. Grex [BK13]. Grey [FGL^{+}11]. Grid [Hua17, MD13, MFT^{+}19, AAH17, Ben19, CP10b, JdSJC^{+}15, KRKS11, KV10, LL12a, LLWC17, SMB10, SZL10, TLQS12, WH17, dKG^{+}10, DW12, LSH^{+}13, LS10, SJB12]. grid-based [CP10b]. gridding [GOH^{+}13]. gridding-accelerated [GOH^{+}13]. grids [ARDQ18, BMT12, DJH11, GVBB13, GM14b, LKS14, LL10, LWK^{+}19, SMO14, YZS15, AAD10, SNCP12]. grooming [WCL^{+}13]. Grøstl [ABO^{+}17]. groundwater [SWLP19]. Group [CWZ^{+}18, KKLJ14, LLW12, CJDC10, EDH^{+}17, LC14b, dAMFdS13, TC13]. Group-based [CWZ^{+}18, KKLJ14, TC13]. groups [ARDQ18, LKM12, WLZ^{+}18]. growing [IZ12, OGRV^{+}12]. guarantee [JM14, MZZC12]. guaranteed [LNA12, LNAL17, NGQM12, WCWO17]. guarantees [ESCV15]. Guest [CGFH19, TFV^{+}15].

Hadoop [FRM15, GYY^{+}14, HWL18, HWLR14, MNR^{+}19, YLB^{+}15]. Hamiltonian [HBAD15, LLFJ18]. Hamiltonicity [Ste17]. handle [RK18]. handling [KVA18, KV10, LNW^{+}12]. handoff [FCZ^{+}12, ZBR11]. Happened [HCR12]. Happened-Before [HCR12]. happy [KSSK16]. hard [GFPC14]. Hardware [BK18, CRB19, DGNW13, MCAS12, SHA17, Zsa16, ABO^{+}17, BDM18, CDD^{+}19, CV16, CGC16, CP17, CM12, EAB^{+}19, FWM^{+}10, GKS15, Hus17, JJ12, KDO^{+}13, KC17, LMSK18, MTM10, MCS^{+}19, PAG^{+}18, PG19, QGZP17, QGZP19, SV18]. Hardware-accelerated [DGNW13, Zsa16]. hardware-entangled [EAB^{+}19]. hardware-generated [MTM10]. hardware-software [CV16]. hardwares [SKH15]. harmony [ES12]. Harnessing [MTL^{+}18b]. harvesting [RB12]. Hash [LACJ18, TT10, ABO^{+}17, SRT^{+}18]. hashing [BMLLC^{+}19, HDCM11]. Haste [JXZ^{+}19]. HBS [CK13]. HD [GB11]. HDL [DSEP17]. Head [ESGQ^{+}11]. Head-of-Line [ESGQ^{+}11]. health [ZAAB17]. healthcare [AMU^{+}19, SMW18, Udd19, VS18]. heat [LE19]. Help [IR12]. helping [ACH18]. herd [KS18]. Heterogeneity [Las12, Las13, XLL15, KWZ19, LpJS^{+}18]. Heterogeneity-driven [XLL15]. Heterogeneous [EKNS17, KRM14, WMC^{+}18, ALM^{+}16, AAD10, Amm16, BKC^{+}15, BKCM17, BEN12, BSS^{+}13, CCK11, CJ17, DK11, GQZ18, GZY14a, GAOHG17, HLL^{+}19, Hus17, JST12, KG19, KHN17, KyLPC17, KSG13, KN18b, KN18a, LK13, LWC^{+}18, LLKY13, LGRV19, LL12b, LDP^{+}14, LLY15, LNAL17, LLCZ19, LV15, LE19, LFGM17, LXZ13, MGSG12, MCS^{+}19, NDP13, NFHL13, ND12, OPR18, OJP^{+}18, PKN10, PP13, PSB^{+}19, REK10a, REK10b, RGAN18, SSFP11, SSM^{+}16, SLV19, SS11, SCMS12, SZMK13, SHL^{+}13, SPPA19, TLLL10, TLLV10, TFMS15, VLW18, VBF13, WQL14, WTWZ16, WJ12, WG11, WYTX13, WJ14, XLHT13, XLPL19, YLL17, ZMG^{+}16, ZTFK16, ZLWZ18, ZGW^{+}19, ZSCX18, ZHLQ12, VAF19, VBF13, VFAD17]. heuristic [DK11, KKS^{+}12, PKN10, YFBY17]. heuristic-genetic [DK11]. heuristics [JST12]. HEVC [Lla17]. HiCOO [YQTV12]. hidden [HB11]. Hierarchical [MD13, TTH12, YQTV12, AAH17, BJS18, BMT12, DR19, EDH^{+}17, GHY10, IZ12, LK13, TLQS12, WCWO17, dSS11]. hierarchies [DTK11b]. hierarchy [Ale19b, WYTX13]. High [BYG^{+}18, CLA^{+}18, DYL^{+}12, DB18, FC14, HES10, HN19, KRS13, KRS14, MBG^{+}17, Nee17, PBB^{+}17, SWHB17, TPJ^{+}19, VFAD17, XMMD17, AM13, ARI17, AGWY11, BAT^{+}19, CVK^{+}18b, CDY^{+}19, Cuz11, Cuz13, DKK18, DF12, DAB^{+}14, DMS^{+}16, FHL^{+}15, Fu10, GOH^{+}13, GMSS^{+}11, HRG^{+}11, HVW16, ICQO^{+}12, KVNV17, KSB11, LWC^{+}18, LMSK18, LSXX14, LJZ^{+}19, LB18, LAC18, MSGS^{+}13, MZC18, MLK12, No12, PGKV18, SPRG^{+}12, SA11, SQQL19, SGdSS13, TYD^{+}19, VAS^{+}13, WRW13, ZW13, ZWQ^{+}16, dAT17, MMVL11]. high-availability [Fu10]. High-dimensional [HN19, WRW13]. High-Level [KRS13, KRS14, BYG^{+}18, DMS^{+}16, SGdSS13]. High-Performance [KRS13, KRS14, PBB^{+}17, TPJ^{+}19, Cuz11, Cuz13, DF12, FHL^{+}15, GMSS^{+}11, HRG^{+}11, ICQO^{+}12, LSXX14, LJZ^{+}19, LB18, MSGS^{+}13, PGKV18, ZW13, ZWQ^{+}16]. high-radix [VAS^{+}13]. high-resolution [GOH^{+}13, SQQL19]. High-Temperature [SWHB17]. High-throughput [CLA^{+}18, HVW16]. Higher [HS17]. Highly [DF17, KHT^{+}14, AFA13, HN19, SMT15, Ter16]. Hint [CK13]. Hint-based [CK13]. Hints [GLC14]. histograms [CL14]. HLA [DB11]. HLA-based [DB11]. HLS [MH18]. HLS-based [MH18]. HMFS [LHZ^{+}18]. HMIPv6 [CKML12]. HMVFS [ZLH^{+}18]. Hoc [RBP^{+}11, TM10, AH11, AH12, BM11, FCW11, FGL^{+}11, JLWX11, Kim11, KSK15, LHW14, LC14b, NMN^{+}14, OM10, SNCP12, SJS11, TC13, WHS^{+}18, YSS11, YWW12]. HOG [RBG17]. hole [LZC11, PSC^{+}16, SGAC14, YDZ^{+}18, dOBG^{+}15]. holistic [WL10, ZHH15]. home [HRM17]. homogeneous [BM17a, CRJ10a, OOSGVG^{+}16]. homology [DKKV15]. homonymous [AAI^{+}15]. hop [FCW11, FCZ^{+}12, JLWX11, JM14, KHK18, MPV12, RFS^{+}12, RB12, ZMG^{+}16, CSW^{+}17]. host [LLWC17]. host-based [LLWC17]. hosting [SSVC10]. hostload [DKC14]. hot [EMC19, MB19, GPSH19]. Hot-N-Cold [GPSH19]. hot-spot [MB19]. Householder [BDG^{+}15]. HPC [APV18, CVK^{+}18b, CJA^{+}19, ECLV12, GYAB11, NV19, NKSA17, NC13, PCLP16, RBA^{+}18, RMHR17, RÖE^{+}18, WMES12, uRIL^{+}18, YFS^{+}15]. HTM [PB19]. human [CWZ^{+}18, WDS^{+}18]. Hungarian [LYIP19]. hunt [MP15]. HW [RBG17]. HW/SW [RBG17]. Hybrid [BJL18, DBA^{+}18, DR18, LWCG14, PA15, VD18, YS11, ZLH^{+}18, ALM^{+}16, CB15, CJ17, DK11, GLC14, HZL18, HGX^{+}19, JAB12, KS18, KSJC17, LY13, LHZ^{+}18, MBS^{+}12, MMK^{+}11, No12, PARB14, PV19, TY17, WLL16, WHW^{+}17, YLL17, ZFT^{+}18, MMCL^{+}17]. hydrodynamics [VBDRC13]. HyPar [PV19]. hyperbolic [SHRM19]. hypercube [CS10, CFL^{+}19, TMK^{+}17]. hypercubes [Lai14, Lai17, WIB12]. Hypergraph [DKUÇ15]. hypergraphs [STA12]. Hyperspherical [RLP14].

I/O [DLW^{+}12, GFPC14, HZZ^{+}19, No12, WHW^{+}17]. I/O-intensive [HZZ^{+}19]. IaaS [LQM^{+}12, NC13, NKK16]. IC [CMR10]. IC-scheduling [CMR10]. IceCube [AAA^{+}15]. IceProd [AAA^{+}15]. ICS [HMY^{+}18]. ICT [CTS17]. ideas [Sch14]. identical [GG19]. identification [GSASA19, MMN^{+}18, SRB^{+}19, XCC^{+}19, ZAAB17]. Identifying [LT10]. identity [WXMZ19]. idling [CFI^{+}18]. IEEE [HB11, ZBR11]. II [KHT^{+}14, RLA^{+}17, SMO14]. III [CP10b]. image [HLBM16, HN19, LMSK18, LLS^{+}16, dAT17, FC14]. Image-to-Mesh [FC14]. imagery [SQQL19]. images [CDD^{+}19]. imaging [KDO^{+}13]. Immediate [Ksh12]. immersive [TYD^{+}19]. immune [HD10]. Impact [YAA10, HHS12, HRF^{+}11, RBP^{+}11, SFT^{+}13, WCF14]. Impacts [PCX^{+}11, PCX^{+}14]. IMPATIENT [GOH^{+}13]. Implementation [ABGV11, SDS10, SE15, ADV14, CP10b, CWP12, GKS15, HES11, HVW16, JM15, KHT^{+}14, LYIP19, MCAS12, MP10, MRT18, OGRV^{+}12, SA11, Sol13, TdAR18, YÖ11, dlAMCFN12]. implementations [ICQO^{+}12, Tát11, TYA16, YBM13]. implementing [TR16, YFBY17]. Implications [HKK^{+}18]. implicit [HWL18]. importance [MLMSMG12]. impossibility [AP16]. improve [CSW^{+}17, GLC14, VRM10]. Improved [CP10b, HSH10, TC13, Ara13, CLZ19, GMVRGS16, dlAMCFN12]. improvement [CZZ^{+}17]. Improving [AM13, AHG12, CLG^{+}16, CRWX12, CKWT17, CAF^{+}11, GYY^{+}14, HHK15, KZ11, QGZP19, RMGM19, SLKK12, AA10, HBSASA19, KWZ19, LBT19, SAL10, SK11, MMCL^{+}17]. IMSuite [GN15]. In-Memory [SLL18, LLB^{+}18, LHZ^{+}18, VETT18, ZKZF18]. in-network [BCO^{+}12, JF12]. incentive [CG12, YAA10, ZCMY12]. incentive-based [CG12, YAA10]. inclusion [Kak15, dMS18]. incomplete [CASD18, GLW14]. Incorporating [WTY^{+}18]. Incremental [ESCV15, LRS18]. incrementally [YC12]. independence [GK10]. Independent [WFZJ12, AFD^{+}11, CFJW13, CFL^{+}19, EB13, HAC17, Li19a, MPR19, PDB13, SSM^{+}16, SBÇ12b, WCF14, WIB12]. independent-gate [WCF14]. Index [Ano10a, Ano10b, Ano11j, Ano11k, Ano12m, Ano12n, Ano14f, Ano14g, Ano15k, AS19b, KN18b, KN18a, LSZZ15, PCLP16]. indexing [WIR^{+}18]. Indian [Nee17]. indirect [HBF12]. individuality [TYD^{+}19]. industrial [HMY^{+}18, KKTZ13]. Inexact [Pla13]. inference [RFP^{+}19, SHK19, XP10, YWAT13]. Inferencing [TFV19]. InfiniBand [ARP18, ESGQ^{+}14, ESGQ^{+}18, GRJ^{+}15]. InfiniBand-based [ESGQ^{+}14, ESGQ^{+}18]. influence [MCS14]. Info [NTN12]. Info-based [NTN12]. information [AHZ11, AH11, Ana14, JLWX11, KTP17, LSWC14, MP15, SHK19, TKG^{+}17, WHW^{+}19, YDZ^{+}18]. infrastructure [AFA13, HPB^{+}10, JAB12, KKKP12, MBS^{+}12, SW12, SWHB17, ZCMY12]. infrastructures [FPF14, NAB^{+}11]. inherent [CB15]. inhomogeneous [LWW19]. initial [YS11]. Initiatives [Hua17]. injected [GK15]. injection [CP17, LLWC17]. inland [SQQL19]. innovative [MCS^{+}19]. INRFlow [NPE^{+}19]. insertion [SS17]. inspired [CMMN10, GVBB13, HD10, PG19]. instances [PDB13, ZG13]. Instantly [TOR^{+}14]. institute [Nee17]. instruction [NHX^{+}19, PYP^{+}10]. instruction-systolic [PYP^{+}10]. instructions [Sol13]. instruments [CKK^{+}13]. InteGrade [dKG^{+}10]. integrated [LMXJ18, WCL^{+}13, YWG15]. Integrating [DT11, KKKP12]. Integration [YJKD10, Kum17, ZMZJ17]. integrity [BCO^{+}12]. Intel [CHLL18, LTG14, RPN19, Zha11]. intelligence [KAA^{+}19b, LdPLC^{+}19, ZGJ^{+}18]. intelligence-based [ZGJ^{+}18]. Intelligent [LHW^{+}19, KBC19, KDSS18, PLSM18, YXW^{+}18]. Intel(R) [KVNV17]. Intended [CTC11]. Intensive [BS11, DF17, HZZ^{+}19, HWLR14, MLK^{+}16, RBN11, Ren11, WZZ^{+}17, WG11, ZMCP11]. Inter [KCSS18, GZG^{+}17, RS19]. inter-core [GZG^{+}17]. inter-node [RS19]. Inter-Thread [KCSS18]. interaction [HWLR14, YJL16]. interaction-intensive [HWLR14]. interactions [PARB14]. interactive [CTS17, HSS17]. interconnect [AHA^{+}16, UM17]. interconnected [CPJ^{+}19, SGAC14]. interconnection [ARI17, BM14, BJ15, CCAAS19, GJ12, KMC16, LLKY13, NPE^{+}19, PW16, PW17, PMCC18]. interdependent [SNCP12]. Interdisciplinary [NKSA17, CCE^{+}17, Hua17]. interest [Ano16l, REZN17, CTC11]. Interest-Intended [CTC11]. interface [JM15, NSDZ18]. interfaces [NGQM12]. interference [GZG^{+}17, WHS^{+}18]. intermediate [YYLC11]. internal [TAM^{+}19]. International [OY13]. Internet [BJ18, CXQ^{+}18, CCC^{+}19, CDPS18, DAPR18, ECP^{+}18, HMY^{+}18, LAS^{+}19, LJQ^{+}19, MS19, MXSL12, MZZC12, NPS^{+}19, PJ18, RSVW19, WSX^{+}19, WHC^{+}18, WCCH18, YWJ^{+}18, ZLT^{+}19]. interoperability [AZW13]. Interplay [ZXGD18]. Interrupting [AST12]. Intersecting [FSV17]. interval [BBCQ13, MHLZ16]. intra [GM13]. intra-node [GM13]. intrachip [MCM^{+}11]. Intrinsic [PAS15]. Introducing [CCE^{+}17, Ada17, BLZ^{+}18]. Introduction [KRS14, LK11, MKN14, PRS14, TFV^{+}15]. introductory [Bog17]. Intrusion [WL11, WML^{+}18]. inventory [GAOHG17]. Inverse [Lla17, GRS19]. inverted [WJ12]. Investigating [LCB16]. investigation [GKS15, PHW^{+}13]. Investigations [Sch13]. invocations [BVGV14]. IoT [Ale19b, AMSÅ19, CLW^{+}19, DBW^{+}18, DKJG19, GRZ^{+}18, HRH18, LAS^{+}19, LWWQ18, MA19, PH18, SCW^{+}18, SLZ^{+}19, TODQ18, TAM^{+}19, VS18, YXW^{+}18, ZGJ^{+}18, ZXMR18]. IoT-based [YXW^{+}18, ZGJ^{+}18]. IoT-CANE [LAS^{+}19]. IoT-enabled [SLZ^{+}19]. IoTDeM [LWWQ18]. IOV [DYL^{+}12, GRJ^{+}15]. iPACS [KCR14]. IPDPS [OY13, Ben15, Mue13, Phi13]. IPDPS’18 [BÇ19]. IPv6 [WZ13]. Irregular [FTM^{+}14, PGRP17, TFV^{+}15, AM13, AC16, FCP^{+}15, KG19, LWCC15, MSAZ10a, MSAZ10b, PCMM^{+}17, PA15, TP18, ZSW14]. ISA [KNHH18, SSFP11, SPC^{+}17]. isogeometric [SWLP19]. isomorphism [Pla13]. Issue [AS13, CDJL11, CGFH19, DDE19, FTM^{+}14, GMSS^{+}11, KRS13, KRS14, LK10, MNK12, QGB^{+}17, SLL18, STS19, SFC17, TFV^{+}15, TFV19, XMMD17, BÇ19, BS11, Cuz11, DF12, DB18, FPS11, FPS12, Gra10a, LZ11, Las12, LK11, MSGS^{+}13, MKN14, PRS14, RLA^{+}16, RLA^{+}17, TH11]. Issues [NCRK19]. items [LT10, ST14]. Itemsets [BMLLC^{+}19]. iterated [KHW13]. iteration [YBX^{+}13]. iterations [CASD18, YS11]. Iterative [CG10, BDRB14, NVK^{+}11].

J [KN18b, LSS^{+}11a, MSAZ10a, PCX^{+}14, REK10a]. Jacobi [CASD18, HBAD15, HS17, TYA16, WPC19]. Janus [DMG18]. Java [AST12, BVGV14]. JDPC [MSGS^{+}13]. Job [TDBL13, Ben19, DBA^{+}18, EHL^{+}15, GMVRGS16, GYY^{+}14, MLK^{+}16]. job-scheduling [Ben19]. jobs [HSH10, LYW^{+}16, SHC14]. join [GK19, LL18]. joins [CTKA17, CKWT17, CKLW19]. Joint [AAA^{+}10, ABF^{+}14, LBT19, LYW^{+}16, LZLX11, AS19a, CCA18, GDL^{+}11, ZY12]. Jones [NHO^{+}13]. Jordan [Dav17]. Journal [AS13, Cuz11, FTM^{+}14, FPS11, GMSS^{+}11, KRS13, Las12, LK11, MSGS^{+}13, MNK12, TH11]. JPDC [LK11, KRS14, MKN14, PRS14]. JPEG [WLCZ15]. Julia [RFP^{+}19]. JVM [AC16].

KD [HN19]. KD-tree [HN19]. KDE [EHL^{+}15]. Keep [LFS16]. Keeping [PBB^{+}17, PGKV18]. Kernel [MBBD13, GM13, KC17, PSB^{+}19, dSAJ15]. Kernel-assisted [MBBD13, GM13]. kernelized [PDP17]. Key [SLHS19, AVAH18, BCD^{+}15, GTGLSA12, HZHS18, JXZ^{+}19, LAK10, LLW12, REK10a, REK10b, SZMK13, ZWQ^{+}16, ZHT16]. key-based [GTGLSA12]. Key-Value [SLHS19]. keyword [HWL18]. keyword-aware [HWL18]. kinetic [LMB^{+}17]. Knapsack [BW18, dADC18, WYW15]. knapsack-based [WYW15]. KNEM [GM13]. knowledge [KKS^{+}12, LAS^{+}19, MS15, SLG^{+}18, YL12]. knowledge-based [YL12]. Kokkos [ETS14]. krill [KS18]. Kronecker [JD12, LNW^{+}12].

L [CRJ10a]. L2 [KK11, Zha11]. L2-prefetch-caused [Zha11]. lab [FSP18]. labeling [KRKS11, TMK^{+}17]. LABS [LDZ^{+}14]. LaDAR [YWAT13]. Lagrangian [BHLT14]. Lamport [TPLY18]. language [JWH^{+}17, LMY^{+}11, MRS^{+}14, RK18]. Languages [KRS13, KRS14]. Large [AFG^{+}19, MRJ^{+}19, WBRT13, XMMD17, AM13, BKC^{+}15, BDL^{+}19, CC16, CLOL17, CTKA17, DV13, DB11, DBCF13, DLW^{+}12, KSSL16, KSJC17, KBC^{+}10, LGZ^{+}10, LZY11, LHW^{+}19, Luc18, LWCG14, MYM10, MBMC19, MVP17, NAB^{+}11, PP13, PB19, PDB13, PLK^{+}18, SS17, SMT15, WCWO17, YÖ11, ZVL11]. Large-scale [AFG^{+}19, MRJ^{+}19, WBRT13, CC16, CLOL17, DB11, DBCF13, DLW^{+}12, KSSL16, KBC^{+}10, LGZ^{+}10, LZY11, Luc18, LWCG14, MBMC19, PB19, WCWO17, ZVL11]. large/irregular [AM13]. last [DMI^{+}19, FABG^{+}19, SS17]. last-level [DMI^{+}19]. Latency [LDZ^{+}14, THGY15, ASA18, ASSS19, CRD12, CM12, Dav17, LW19, MS19, NCB^{+}17, RM11, SLKK12, SFHS19, SLZ^{+}19, TVT^{+}17]. latency-sensitive [ASSS19]. latency-tail-tolerance [SLZ^{+}19]. latency-tolerant [NCB^{+}17]. lattice [AVAH18, WZY^{+}19]. law [NZ17, SC10, CN14]. laws [SHRM19]. Layer [DDO^{+}18, PKW^{+}10, WCL^{+}13, YYWZ19, dlAMCFN12]. layered [LL12a]. layout [KMC16, LGK^{+}12, Str12]. LDA [BOKS19]. LEACH [CCC^{+}19, NSA11]. leader [DLV11, DGDF10]. leakage [KK11, NKV14]. leakage-aware [KK11]. Leaping [KM17]. Learning [BM11, MBG^{+}17, TFV19, CXQ^{+}18, EM11, FFYH19, GMRRG19, HSS17, HKK^{+}18, HHK15, KAA^{+}19a, KCFP18, LGZ^{+}10, LHHH11, LCJ^{+}18, MCZ14, NZA13, OPR18, PSGS17, RT18, RDCQ17, TXLL14, TM10, Upa13, WRW13, WBS19, WLK^{+}19, WDS^{+}18, XCC^{+}19, XRB12, ZGW^{+}19]. learning-based [MCZ14, RDCQ17]. Learning-TCP [BM11]. length [KP17]. lengths [KIH15]. LEON3 [TdAR18]. Let [CVK^{+}18b]. Level [AC16, KRS13, KRS14, BBH^{+}17, BYG^{+}18, CLMRL15, DMI^{+}19, DAB^{+}14, DMS^{+}16, FABG^{+}19, FLCB10, GAC^{+}17, HES10, LC14a, LWH^{+}19, LPLFMC^{+}12, MEMEMH17, NPE^{+}19, OWK14, OMT^{+}17, Ren11, SS17, SGdSS13, WMES12]. level-set [HES10]. levels [Kum17, Li16]. Leveraging [SSFP11, CFI^{+}18]. LeWI [GLC14]. lexicographic [BMLLC^{+}19]. LFRic [AFH^{+}19]. liberal [NDW17]. library [HZHS18, LGK^{+}12, ZSW14, VAF19, VBF13, VFAD17]. lifetime [EMC19, LL12b, Li14, LZC11, VRM10, WBS19]. lifting [IIH16]. lifting-based [IIH16]. light [PR12, WZZ^{+}17, ZFT^{+}18]. light-trails [PR12]. light-weight [WZZ^{+}17, ZFT^{+}18]. Lightweight [MSF^{+}13, KP17, Kim17, DKJG19, MP10]. like [CP10a, CTC11]. limited [SSGG18, VS16, Zsa16]. Limiting [MSV19]. limits [dSS11]. Line [DGBN14, CFL^{+}19, ESGQ^{+}11]. linear [CMR19, CP10b, FHL^{+}15, ICQO^{+}12, LMXJ18, LWXX19, LKD14, MRT18, TFMS15, Ter16, XYZW14, YÖ11, KCP19]. linearizability [KKW17]. lines [DJDK19]. Link [SJS11, FCZ^{+}12, LST17, MCAS12, MVP17]. links [AGMS16, ACA^{+}19, SHK19]. LinuX [LACJ18, LAC18]. Liquid [SWHB17, SWLP19]. List [TLLL10, FPF14, WLL16]. little [MNR^{+}19]. live [GRJ^{+}15, HTB19, WMES12]. Load [CKLW19, SBÇ12a, AES11, ACCP12, ASES15, BFMT^{+}18, DB11, DLW^{+}12, GLC14, HLL^{+}19, JL11, KAA^{+}19a, KNHH18, MLDG12, MPV12, NHO^{+}13, PC11, PRN^{+}19, SB15, TLL^{+}18, TVT^{+}17, YJL16, YAA10, YMLP14, ZSW14, ZLMC14]. load-balanced [YMLP14]. Load-balancing [CKLW19, SBÇ12a, KAA^{+}19a, NHO^{+}13, YJL16]. loads [CG12, HV13, KVA18, LML^{+}10]. local [ADD17, GTGLSA12, GNZ18, LMJC11, ROB^{+}18, Sch13, WWW17a]. Locality [GXYZ13, JL11, SBRM19, EHL^{+}15, LK13, Ozt11, SKK14, SRT^{+}18, ZWQ^{+}16]. locality-aware [EHL^{+}15, SKK14, ZWQ^{+}16]. locality-cognizant [LK13]. Locality-sensitive [JL11, SRT^{+}18]. localization [AKBD10, CCW14, CRWX12, DLLL11, LDS16, MKM16, PD19, WDS^{+}18]. locally [LFZ^{+}17, XHZZ16]. locate [DWX10]. located [SBÇ12a]. Location [Li17, ABF^{+}14, BJL18, DBW^{+}18, HCM11, KHK18, LLDL15, OJP^{+}18, TZI11, TR16, TKR^{+}19, dOBG^{+}15]. location-based [ABF^{+}14]. location-free [dOBG^{+}15]. Lock [SSdlB^{+}10, HSY10]. Lock-free [SSdlB^{+}10, HSY10]. locking [LZLX11]. locks [AFA13, CG10, UBES10]. log [CJA^{+}19, ZFT^{+}18]. log-analysis [CJA^{+}19]. log-based [ZFT^{+}18]. logarithmic [AF17]. logging [MMCL^{+}17, MMCL^{+}17]. Logic [NKV14, FPM^{+}14, MLZY17, XYZW14]. logical [TPLY18]. Long [LKM12, KVNV17]. Long-term [LKM12]. look [SHL^{+}13, HZL18]. look-ahead [SHL^{+}13]. Look-Up [HZL18]. Looking [LKD14]. loops [ZLKK19]. losses [HZA^{+}15]. lossless [CW15]. lossy [GYP13]. lost [LdSB^{+}18]. Low [AEY12, CM12, Dav17, MC17, ABO^{+}17, FABG^{+}19, HZL18, KK11, KHK18, LW19, MGRRK14, NKV14, RM11, Sol13, YGZ^{+}10]. low-area [ABO^{+}17]. low-complexity [Sol13]. Low-contention [AEY12]. low-cost [YGZ^{+}10]. Low-Density [MC17]. low-latency [LW19]. low-power [KK11, MGRRK14]. low-rate [KHK18]. lower [Li19b, NDP13]. Lowest [MAKWZ13]. LQR [ZMZJ17]. LSB [DKJG19]. LSM [SLHS19]. LSM-tree-based [SLHS19]. LTI [AD12]. LTL [BBBC12]. LUT [HZL18, WD18]. LUT-based [WD18]. LXCloud [LACJ18]. LXCloud-CR [LACJ18]. LXCloudFT [LAC18].

M [GA18]. M2M [TKG^{+}17]. MAC [GZY14b, TLY12]. Machine [RDCQ17, TKG^{+}17, ZLZ^{+}19, AES11, CL14, FMIF18, GMRRG19, HTB19, KHT^{+}14, KS18, KAA^{+}19a, KCFP18, LCJ^{+}18, Upa13, WSX^{+}19, WLK^{+}19, ZG13, ZLCZ18, LST^{+}13, TKG^{+}17]. machine-learning [KAA^{+}19a]. Machines [YFS^{+}15, CLW^{+}19, CG11, Fu10, GG19, KR10a, KR10b, RT18]. magnetic [LdPLC^{+}19]. main [BBH^{+}17]. Maintaining [LMP10]. maintenance [GHIJ19, MAPF14]. make [AS19b]. makes [JXZ^{+}19]. makespan [GG19, TFMS15]. Making [LLT12, ZKZF18, AKK^{+}19]. Making-a-stop [LLT12]. Malicious [CDW^{+}19, HMY^{+}18]. Malleable [FZWL12]. malware [TY17]. manage [PST^{+}19]. manageable [GRZ^{+}18, dAMFdS13]. Management [AS13, AS15, AM11, AK18, BVGV14, CF19, CKMP17, Fu10, FX10, HCM11, HHS12, HHK15, JWH^{+}17, KK11, KLJ^{+}11, LC11, LAS^{+}19, LJQ^{+}19, MBS^{+}12, MLMSMG12, MCP^{+}18, NPS^{+}19, NAB^{+}11, OJP^{+}18, RWB^{+}13, RAN^{+}17, SNMB16, SB12, SA19, SLG^{+}18, TZI11, WYW15, WZZ^{+}17, XRB12, ZV12, dKG^{+}10, SHSH17]. Managing [AKBD10, SS17, SLZ^{+}19]. MANET [YAA10]. MANETs [Hu11, YA11]. Many [CHLL18, DDO^{+}18, AFA13, APRA18, AA16, ARI17, BBBC12, CKK^{+}13, FTM^{+}19, JHF^{+}17, Lai14, LWC^{+}18, LTG14, MZC18, PCMM^{+}17, PTK^{+}13, PR13, RLA^{+}16, RLA^{+}17, TCHC12, ZLS17, dCPD19]. Many-Core [DDO^{+}18, CHLL18, AFA13, APRA18, AA16, ARI17, BBBC12, CKK^{+}13, FTM^{+}19, JHF^{+}17, KSG13, LWC^{+}18, MBBD13, MZC18, PCMM^{+}17, PTK^{+}13, PR13, RLA^{+}16, RLA^{+}17, TCHC12]. many-cores [ZLS17]. manycore [ETS14, FCP^{+}15]. map [IZ12, CKML12]. Mapping [TBG^{+}17, BLMB13, CGM14, CDAN14, DFST13, FLL14, KG19, LW16a, MTL^{+}18a, PMAL11, YWJ^{+}18, YWG15, ZLKK19]. MapReduce [ALTV13, AM17, BK13, BD11, CCA18, CLOL17, GYY^{+}14, LYW^{+}16, LWWQ18, NMS^{+}18, NF16, Pla13, SMT15, VETT18, WTWZ16, uRIL^{+}18, WD13]. MapReduce-based [VETT18, WD13]. maps [DP12]. Marathon [LHNBB19]. MaRCO [ALTV13]. Marine [YWJ^{+}18]. maritime [WWA^{+}18]. Market [CKMP17]. Market-based [CKMP17]. Markov [ASKO16, GA18, LWW19]. Markovian [BC11]. MASC [TJCB10]. Mashup [CLW^{+}19]. massive [FCG^{+}14, JWH^{+}17]. Massively [DAG^{+}17, UGG^{+}11, FLM^{+}19, JJ12, LGRV19, RBB17, SMH^{+}14]. master [BMT12]. master-worker [BMT12]. Matching [HBS17, DKU15, GK10, HN19, KSJC17, KSSG14, MPN17]. material [LE19]. materials [CXX^{+}18, DAG^{+}17]. mathematical [DJH11]. matrices [JM15]. Matrix [BG16, ASA18, ASES15, BGO19, CP10b, GA18, LV15, MLW^{+}19, MBW16, MPG17b, OT19, PB15, PR13, SAJ13, SE15]. matrix-vector [ASES15, CP10b, MBW16, PR13]. maturity [CMC^{+}19]. maximal [BCH15, MPR19, SMT15, TSFZ14, WCH^{+}17]. Maximization [YZG18, LHX^{+}16, LL12b, VLL^{+}14]. maximize [SSFP11, UFF19]. maximizing [CDR12, DW12, Li14, LLCZ19, MA11]. Maximum [BLMB13, AFD^{+}11]. Maximum-throughput [BLMB13]. maxmin [ZLCJ12]. may [STKW12]. me [MPS16]. mean [BDL^{+}19]. Means [DBCF13, CDD^{+}19]. measurements [ASKTZ13, JKIE13]. measures [DGBN14]. Mechanism [TAM^{+}19, AS19b, CG11, CG12, CMR^{+}18, CCW14, GYY^{+}14, HCM11, KO11, MBO11, PMdO11, RA11, ZBW^{+}17]. mechanisms [ASKO16, KV10, ALLM11]. media [KLP10]. medical [KDO^{+}13]. medium [KGN11]. Meeting [AFH^{+}19]. meets [CRB19]. Membership [BMS19, LC14b]. membrane [YLZW18]. memories [GKK^{+}13, KR17]. Memory [CRB19, HMR15, SLL18, Tam18, VS16, ZLH^{+}18, AM13, ACHY18, BBH^{+}17, BBD18, CC16, CGM14, CDAN14, DT11, ETS14, Eij18, EKNS17, FWM^{+}10, FLC14, Gra10b, HDCM11, HGFF10, HZHS18, HHA14, Hus17, IIH16, IRRS16, KKR14, KRM14, KKLJ14, KMS10, LLB^{+}18, LHZ^{+}18, Lop18, MTM10, MSV19, MSK^{+}16, NHX^{+}19, No12, Pet19, QGZP17, QGZP19, RS19, RHH12, RDCQ17, SSGG18, SB15, SDS10, TGPUC16, VETT18, YGZ^{+}10, ZKZF18, ZPK^{+}14, ZLWL12, HZL18, MP10]. Memory-aware [HMR15]. memory-based [No12]. memory-distributed [Pet19]. memoryless [BKMT14]. mental [Eij18]. Mesh [DR19, GPJA10, DJDK19, DAB^{+}14, FLL14, GDL^{+}11, GA16, GNZ18, KKK11a, KHK18, LC11, LWLD12, MRJ^{+}19, NPGV10, SSZ10, SBRM19, WCXL11, WBRT13, FC14]. mesh-NoC-based [FLL14]. Mesh-of-Tree-based [DR19]. meshes [BG16, CZZ^{+}17, DV13, LWCC15, LXLS12, dMS18]. Mesos [LHNBB19]. Message [MMCL^{+}17, Pra16, CPA^{+}11, CJA^{+}19, DLM19, DDNT10, HZA^{+}15, IRRS16, JZZ^{+}17, Kak15, KMS10, KS13, LWK^{+}19, LWW19, PS14, TGPUC16]. message-optimal [DLM19]. message-passing [DDNT10, IRRS16, Kak15, KMS10, KS13]. messages [ASKTZ13, HBSASA19, XLL15]. meta [GVBB13, KKS^{+}12, LGZ^{+}10]. meta-heuristic [KKS^{+}12]. meta-learning [LGZ^{+}10]. meta-scheduling [GVBB13]. metadata [ZV14]. metaheuristic [MMK^{+}11, ROB^{+}18, TLW18, WMG13]. Metaheuristics [TH11, TH13]. metering [URK^{+}19]. Method [AC16, AST12, Ale19a, ATDH13, CLZ19, CLC^{+}17, CW15, DM17, GNZ18, HGX^{+}19, LR14, MVP17, MA19, MRT18, SHL^{+}13, WPC19, WHW^{+}19, YLL17, dlAMCFN12, PPTV^{+}10]. Method-Level [AC16]. methodologies [DMS^{+}16, PSGS17]. methodology [CSJ^{+}13, DSEP17, LdSB^{+}18, MSAZ10a, MSAZ10b, OMT^{+}17]. methods [DKUÇ15, LWCC15, PAS15, SSZ10, SHRM19]. metric [KC17, Luc18, Sta17, XCC^{+}19]. metrics [DKUÇ15, PARB14]. MGR [DAPR18]. MIC [WTWZ16]. micMR [WTWZ16]. micro [KKH17, KC17]. micro-benchmarks [KC17]. micro-clusters [KKH17]. microarchitecture [Zha11]. Microarray [BF13]. MicroClAn [BF13]. microrobot [LBMG15]. microscopic [WHW^{+}19]. Microwave [XTN12]. MIDAS [ECWV19]. middleware [AZW13, RPS19, SMPMLVLS11, dKG^{+}10]. midpoint [TW15]. midpoint-based [TW15]. Migration [ZXYO11, CLC^{+}17, FMIF18, GRJ^{+}15, HTB19, JTZZ11, LY12, MB19, WMES12, WSX^{+}19, ZLZ^{+}19]. Migration-aware [ZXYO11]. migrations [TKX^{+}13]. million [PGP^{+}12]. million-core [PGP^{+}12]. MIMD [YBM13]. MIMO [AD12, GZY14b, ZY12]. mini [BCD^{+}15]. mini-applications [BCD^{+}15]. minimal [MS15, SWLP19]. Minimization [THGY15, CPLY18, JZF^{+}15, KR10a, Li17, Li19b, LZLX11, RTZ11, TFMS15, YWG15]. minimize [GG19]. Minimizing [XLPL19, ZWW17, TKX^{+}13, WHS^{+}18]. minimum [BPBR11, BBD18, HS12, KO12, KSK15, LY10, YWW12, YYLC11]. Mining [SMT15, Cuz11, Cuz13, WD13, WZQ^{+}13, BMLLC^{+}19]. MIPS [NPVG^{+}19]. MIPS-based [NPVG^{+}19]. miss [CK13]. mitigate [ASSS19]. mitigating [EMC19]. Mitigation [BK18, WCF14]. mixed [ZZJ^{+}18, ZLWZ18]. mixed-criticality [ZZJ^{+}18]. MixHeter [ZLWZ18]. Mixing [FHL^{+}15, Li10]. mixture [SQQL19]. Mobile [LC14b, THGY15, TAM^{+}19, TPS^{+}18, AKBD10, AH12, Ana14, BWP^{+}11, CWD11, DWX10, EM11, FCML13, FP17, GQZ18, Kim11, LL19, LZ11, LL10, LC11, LHW14, Li17, LWW18, MXSL12, MKM16, NSA11, NMN^{+}14, PVP18, PMHM19, RB12, REZN17, SNCP12, SGAC14, SMO^{+}18, SJS11, TZI11, TC13, TY17, TWQS12, VLW18, VRM10, WW18a, WZX^{+}19, WZH^{+}19, YSS11, ZGW^{+}19, RBP^{+}11]. Mobility [KO12, BEN12, CKT11, LL19, LZN19, RBP^{+}11]. Mobility-assisted [KO12]. mobility-aware [LL19]. modal [AM11, BWP^{+}11, Kar19]. mode [YZX11]. Model [AM17, DKUÇ15, SS18, AAH17, ASKO16, AHZ11, ASES15, BBBC12, Ben19, CLW^{+}19, CAK13, CXX^{+}18, DZC17, DJH11, DKC14, DXS^{+}19, GJ12, GPSH19, HMY^{+}18, IEWK17, JLWX11, KyLPC17, KC17, LR14, LMGLGLG17, LZY11, LMXJ18, LWXX19, LCJ^{+}18, LGK^{+}12, LWWQ18, LXZ13, MS19, MMN^{+}18, MMVL11, NV19, NSKN17, PV19, RHH12, TLL^{+}18, TJCB10, WWW17b, gWW18, XYZW14, KR11]. model-based [DXS^{+}19]. Model-driven [SS18, ASES15, LGK^{+}12].

Modeling [DDO^{+}18, GHC^{+}17, LpJS^{+}18, PLD14, PMMMA15, SFT^{+}13, WPC19, HES11, JWH^{+}17, KKK^{+}11b, LWCC15, LC13, MCM^{+}11, MSAZ11, NSA11, ORWT^{+}18, RA11, UMM^{+}18, YL12, YZW^{+}15]. Modelling [STS19, Wu11, KKTZ13, RK18, Sie16]. Models [Cuz11, Cuz13, ASA18, AFH^{+}19, DXS^{+}19, Eij18, FTM^{+}19, GLGLBG12, JKIE13, KVNV17, Nes10, Rao16, SAR^{+}18, SRI14, SQQL19, TJCB10, YQTV12]. modern [EFG^{+}14, GS18, YFS^{+}15]. modified [CCC^{+}19, GLW14]. modular [BM17a, ZBW^{+}17]. modularity [LK15]. moldable [SBÇ12b]. molecular [LHWJ19, PARB14, PTK^{+}13, WYTX13, XLHT13]. molecules [BOT13]. moment [RMU14]. moments [TRS^{+}12, XLH18]. Monitoring [CSMML10, ST14, ASKO16, ACPT15, BOKS19, CL14, CF19, FEH^{+}14, KDSS18, LFS16, RSVW19, SB12, WSX^{+}19, WZQ^{+}13, YDZ^{+}18, ZCW19]. monotonic [MAHKZ12]. Monte [PAS15, ZS13]. MOOC [MBG^{+}17]. motifs [RSL12]. movement [AKBD10, KSB11]. movements [CKT11]. MPI [DMK19, GM13, MBBD13, Nes10, NCB^{+}17, PARB14, TPLY18, Zah12, dlAMCFN12]. MPI-CUDA [dlAMCFN12]. MPSoC [FLL14, LZLX11, OMT^{+}17, ZXYO11]. MPSoCBench [DMS^{+}16]. MPSoCs [LW16a, MTL^{+}18a, TBG^{+}17]. MR [uRIL^{+}18]. MR-Advisor [uRIL^{+}18]. MRI [GOH^{+}13]. MSA [BFKW13]. Multi [BBH^{+}17, FPF14, LK15, LWH^{+}19, MCZ14, OMT^{+}17, PKN10, PVRS17, VLL^{+}14, AS19a, AVAH18, AHZ11, ASSS19, ADDB18, BWP^{+}11, BLMB13, COV13, CDD^{+}19, CKC19, CMMT13, CMC^{+}19, COF^{+}17, CDW^{+}19, DBA^{+}18, DWYB10, FCW11, FCZ^{+}12, FTM^{+}19, GDL^{+}11, GS18, GKS15, GZY14b, GB11, GSASA19, HRM17, Hu11, Hus17, ICQO^{+}12, IIH^{+}17, JJ12, JLWX11, KVA18, KSG13, KKN13, KN18b, KN18a, KHK18, Kum17, LKS14, LSS^{+}11a, LSS^{+}11b, LZY11, LNAL17, LZWZ19, LSC^{+}15, LY13, LPLFMC^{+}12, LLS^{+}16, MS19, Man13, MB13, MPV12, MZC18, MPN17, MAHKZ12, MRJ^{+}19, MGRRK14, MZZC12, NDP13, NL19, NFHL13, NVK^{+}11, PYP^{+}10, PTN^{+}19, PKW^{+}10, RLA^{+}16, RLA^{+}17, RB12, RA11, ROB^{+}18, SNMB16, SGVRP19, SFT^{+}13, SHL^{+}13, SSZ10, SWLP19, SAJ13, SHRM19, SMB10, Sta17, Str12, Tal19, TGPUC16, TRS^{+}12, TCHC12, VBDRC13, VFAD17, WCL^{+}13, WQL14, WQZ^{+}13, WH17, gWW18, XL11]. multi [YZS15, YHWY18a, ZMG^{+}16, ZXB14, ZLS17, dCPD19, DAPR18]. multi- [KSG13, ZLS17]. multi-/many-core [KSG13]. multi-accelerator [ICQO^{+}12]. multi-agent [YZS15]. multi-attribute [LSS^{+}11a, LSS^{+}11b]. multi-block [MRJ^{+}19]. multi-budgeted [Sta17]. multi-channel [GDL^{+}11, GZY14b, SSZ10, ZMG^{+}16]. multi-chip [TCHC12]. multi-cloud [KVA18]. multi-cluster [NVK^{+}11, SHL^{+}13]. multi-core [AVAH18, BLMB13, CDD^{+}19, CMMT13, CMC^{+}19, DBA^{+}18, DWYB10, FTM^{+}19, GS18, GKS15, Hus17, LKS14, LNAL17, LSC^{+}15, LLS^{+}16, MAHKZ12, MGRRK14, RLA^{+}16, RLA^{+}17, SNMB16, SFT^{+}13, SAJ13, SHRM19, WQZ^{+}13, WH17, ZXB14]. multi-cores [TGPUC16]. multi-CPU [TRS^{+}12]. multi-device [VFAD17]. multi-dimensional [GB11, KKN13, KN18b, KN18a, LZY11, LZWZ19]. multi-document [SGVRP19]. multi-epidemic [AHZ11]. multi-GPU [LPLFMC^{+}12, MB13, NFHL13, ROB^{+}18, TRS^{+}12, VBDRC13]. multi-granularity [WCL^{+}13]. Multi-heuristic [PKN10]. multi-hop [FCW11, FCZ^{+}12, JLWX11, KHK18, MPV12, RB12, ZMG^{+}16]. Multi-level [LWH^{+}19, OMT^{+}17]. multi-link [FCZ^{+}12]. multi-modal [BWP^{+}11]. multi-model [gWW18]. Multi-objective [FPF14, ADDB18, COV13, COF^{+}17, CDW^{+}19, NL19, SGVRP19, SWLP19, Tal19]. multi-packet [CKC19]. Multi-parameter [DAPR18]. multi-pass [MPN17]. Multi-path [VLL^{+}14]. multi-phase [Man13]. multi-policy [SMB10]. multi-processor [LY13, SHRM19]. multi-processors [JJ12]. multi-radio [FCZ^{+}12, GDL^{+}11, SSZ10]. multi-railing [PKW^{+}10]. multi-rate [Hu11]. multi-robot [IIH^{+}17]. multi-secret [LWH^{+}19]. multi-sensory [HRM17]. multi-service [RA11]. multi-spectral [GSASA19]. multi-staged [AS19a]. multi-swarm [dCPD19]. multi-target [NDP13]. Multi-tenant [PVRS17, YHWY18a]. multi-thread [DWYB10]. Multi-threaded [BBH^{+}17, LK15, ASSS19, PYP^{+}10]. Multi-tier [MCZ14, MS19, MZZC12, WQL14]. multi-unit [XL11]. multi-valued [Str12]. multi-vehicle-type [PTN^{+}19]. multi-year [Kum17]. multiagent [JL11]. multicast [CWD11, DDG^{+}17, MAGL13, RA11, WW12, YCH^{+}10]. Multicasting [WE13, LSXX14]. multichannel [WIR^{+}18]. multicomputer [GJ12]. Multicore [PSGS17, BM17a, BSS^{+}13, CN14, CP17, DR19, DKU15, FWM^{+}10, FCP^{+}15, GZG^{+}17, KHT^{+}14, KyLPC17, KNHH18, LK13, LLLC15, LM16, Li19a, MBBD13, ND12, NZ17, PP13, SSFP11, SPC^{+}17, SSGG18, SP13, SC10, WLST16, PPP14]. multicore/many [MBBD13]. multicore/many-core [MBBD13]. multicores [CRSB13, LCB16, SS17]. multidimensional [DMK19, dADC18, SJG19]. multifluid [LW16b]. multigrid [MRS^{+}14, WH17]. multihop [ZLCJ12]. multilevel [LK15, PAS15]. Multilinear [ECWV19]. multimedia [AFG^{+}19, AM12a, ZVL11]. multipath [OM10]. Multiple [NSAS10, AFK14, BOT13, BFKW13, CDS10, DKUÇ15, IEWK17, JTZZ11, JM15, Kum17, KIH15, LY10, LDP^{+}14, Li19a, LSWC14, LWWQ18, PLK^{+}18, SPRG^{+}12, SI13, SRT^{+}18, XCC^{+}19, ZWWX16, TJCB10]. multiplex [ZXGD18]. multiplexed [HRG^{+}11]. multiplication [ASES15, BGO19, LV15, MBW16, MPG17b, PR13, SKH15]. multiplicity [PMHM19]. multiprocessing [ZLWL12]. multiprocessor [AA10, BYG^{+}18, CRJ10b, DMS^{+}16, JWSG14, Li16, ZQMM11]. multiprocessors [CRJ10a, HCM11, HRG^{+}11, KK11, LPK^{+}10, LWCG14, dOCS14]. multiresolution [SHRM19]. multiscale [BFL^{+}13]. multistage [BJ15, PW16, PW17]. Multistart [Cza13]. multiswapped [Ste17]. multitask [LST^{+}13]. multithreaded [ACD^{+}18, CN14, LLLC15, NZ17, TP18]. multithreading [LK13]. municipal [LHX^{+}16]. musical [WIR^{+}18]. Mutual [Kak15, Ara13, BDM18, LASS15, NTN12, RDA18]. mutually [WW18a]. myoelectric [BAT^{+}19].

N [BM17a, GPSH19]. N-modular [BM17a]. NAND [No12]. nanoarchitectures [FCG^{+}14]. nanophotonic [HRG^{+}11]. nanoscale [PLD14, ZRN^{+}14]. nanotechnology [MKN14, MNK12]. Near [SLHS19]. Near-Data [SLHS19]. nearest [AK19, BDL^{+}19, HN19, JHL^{+}18, NMN^{+}14, SDG17, YTZ19]. NEAT [LST17]. need [LTG14]. needed [IR12]. needs [CHLL18]. neighbor [AK19, CKC19, CRHC19, HN19, JHL^{+}18, YTZ19, ZMG^{+}16]. Neighborhood [JdSJC^{+}15]. neighbors [BDL^{+}19]. neighbours [NMN^{+}14, SDG17]. Nek5000 [OGM^{+}19]. NERSC [RÖE^{+}18]. nested [ZLKK19]. NET [KK17]. nets [BYT19, ESCV15]. Network [BJS18, DR18, ES12, LST17, LK10, AHA^{+}16, ARI17, AM11, AS19b, BFH^{+}17, BM14, BCO^{+}12, BWP^{+}11, BJ15, CCAAS19, CMMN10, CMR^{+}18, CLG^{+}16, CWP12, CLXX19, CDY^{+}19, DR19, DAPR18, DYL^{+}12, GJ12, HD10, HMY^{+}18, IZ12, JF12, KMC16, KO11, KO12, KRS15, KH12, LT10, LSS^{+}11a, LSS^{+}11b, LB12, LLT12, LÜ14, LY13, LRS18, LWCG14, NPS^{+}19, NM17, NGQM12, RCG18, RGAN18, RSL12, SMW18, SHK19, SCW^{+}18, STKW12, Sta17, TDP15, TCHC12, TYD^{+}19, VRM10, WL11, WW18b, WMC^{+}18, WLYS19, WG11, WLZ^{+}18, WWA^{+}18, WHS^{+}18, YGWJ19, YLZW18, ZY12, AA14, SLW10, SLG^{+}18, ZCF^{+}17]. network-aware [RCG18]. Network-on-Chip [BJS18, DR19, GJ12, LY13, AA14, ZCF^{+}17]. Network-on-Chips [LK10]. network-When [STKW12]. networked [FX10, JL11, XLL15]. networking [DWYB10]. Networks [GPJA10, LC14b, MSEM^{+}19, THGY15, TPJ^{+}19, ZHW19, AGMS16, Amm16, AH11, AH12, AHG12, Ana14, ACA^{+}19, AMT13, AYB^{+}15, ABLP17, AS18, BM11, BWP^{+}11, BOY10, BFVB19, BC11, BJL18, CG12, CB15, CFI^{+}18, CKC19, CC14, CCW14, CS10, CTC^{+}10, CRWX12, CGC16, CHCG18, CPA^{+}11, CRSB13, CKML12, CTT16, DF17, DLLL11, DK11, DGBN14, DBW^{+}18, DBCF13, DKM10, EMC19, ESGQ^{+}18, EM11, EDH^{+}17, FCW11, FCML13, FCZ^{+}12, FGL^{+}11, FZ14, GHY10, GJ12, GDCC18, GDL^{+}11, GHIJ19, GYP13, GZY14b, GM14a, GB11, GL12, HZA^{+}15, HS12, HRG^{+}11, HZDP12, HJLR12, HMY^{+}18, HBAD15, HS17, HAC17, ISAZ10, JF12, JLY12, JBA15, JBS14, JHPL13, JLWX11, JKV15, KTP17, KSSL16, Kar19, KKK11a, Kim11, KKKP12, KSK15, KHK18, KZ11, KKK^{+}11b, KKTZ13, KGN11, Lai15, LL19, LBMG15, LDZ^{+}17, LY10, LNA12, Li10, LC11, LMJC11]. networks [LWLD12, LL12b, LHW14, LSXX14, Li14, LpJS^{+}18, LGM18, LWXX19, LZC11, LHLM14, LDS16, LWW18, LHW^{+}19, MAGL13, MYM10, MAPF14, MPV12, MA11, MBMC19, MCS14, MYD^{+}11, MVP17, MBO11, MSAZ11, NPGV10, NPE^{+}19, NSA11, NFHL13, NMN^{+}14, NZA13, OWK14, OM10, PLY15, PCX^{+}11, PCX^{+}14, PSC^{+}16, PKW^{+}10, PW16, PW17, PMCC18, RM10, RM11, REK10a, REK10b, RLP14, RFS^{+}12, RBP^{+}11, RA11, SCN12, SMP15, SB12, SZMK13, SGAC14, SSZ10, Ste17, SCLL10, SK11, SJS11, TLY12, TODQ18, TC13, TMK^{+}17, TM10, TCS^{+}10, TWQS12, VRM10, WW12, WCL^{+}13, WYW15, WFLJ16, WW18a, WCXL11, WL10, WBRT13, WHW^{+}19, XHZ^{+}10, YpGyLlC13, YTZ19, YDZ^{+}18, YWW12, ZMG^{+}16, ZW11, ZBR11, ZLCJ12, ZCMY12, ZXGD18, ZSCX18, ZTGL17, ZLS17, dOBG^{+}15, ALLM11, LDZ^{+}14, LDP^{+}14, LK11, MLCFH^{+}18, MEMEMH17, RBP^{+}11]. Networks-on-Chip [MSEM^{+}19, HRG^{+}11, KKK^{+}11b, LHLM14, ALLM11, LK11, MEMEMH17]. Neural [MLCFH^{+}18, CLXX19, CDY^{+}19, OGRV^{+}12, PGP^{+}12, TDP15, TYD^{+}19]. neutrino [AAA^{+}15]. neutrosophic [MHLZ16]. Newest [AK17]. Newton [Pet19]. Next [NAB^{+}11, HPB^{+}10]. Next-generation [NAB^{+}11, HPB^{+}10]. nexus [LC14a]. niched [AS19a]. nine [DM17]. nm [HRF^{+}11]. NN [ZHT16]. no [IR12]. NoC [AA16, CZPP16, CAF^{+}11, DJDK19, FLL14, HRF^{+}11, LZI^{+}11, LW16a, LK11]. NoC-based [HRF^{+}11, CAF^{+}11, LZI^{+}11, LW16a]. NoCs [BK18, CG17, LK10, MP10]. Node [BGO19, HAC17, AKBD10, CFL^{+}19, DLLL11, DM17, GM13, KHN17, KVA18, Lai14, Lai15, Lai17, LDS16, PCX^{+}11, PCX^{+}14, RS19, RMHR17, SJG19, TAM^{+}19, Zah12]. node-disjoint [Lai14, Lai15, Lai17]. Node-independent [HAC17, CFL^{+}19]. node-powered [TAM^{+}19]. nodes [LKS14, LWW18, NM17, SI13]. noise [SFT^{+}13]. Non [Li19a, TVT^{+}17, BBFN14, BKMT14, GOH^{+}13, GTGLSA12, HZHS18, HLL^{+}19, KK10, KR17, LHWJ19, NKV14, WMY^{+}17, ZPK^{+}14]. non-blocking [KR17]. non-Cartesian [GOH^{+}13]. Non-clairvoyant [Li19a]. Non-cooperative [TVT^{+}17, KK10]. non-dedicated [HLL^{+}19]. non-deterministic [GTGLSA12]. non-functional [WMY^{+}17]. non-memoryless [BKMT14]. non-uniform [BBFN14, LHWJ19]. non-volatile [HZHS18, NKV14, ZPK^{+}14]. nonlinear [BNBR16, GMMP12, Kub17, Pet19]. nonscaling [Zha11]. Nontrivial [ACH18]. nonzero [ASA18]. nonzero-based [ASA18]. Normally [TOR^{+}14]. NoSQL [Luc18]. Note [NCRK19, RSVW19]. notice [PCX^{+}14]. notifications [APRA18]. Novel [GMSS^{+}11, BJS18, CLC^{+}17, COF^{+}17, CSW^{+}17, GB11, Hus17, JdSJC^{+}15, LMJC11, MSGS^{+}13, PLSM18, SDG17, WLL16, WXZ^{+}18, WXMZ19, ZVL11, ZBR11, ZWWX16, ZLCZ18]. NSGA [SMO14]. NUMA [FCP^{+}15, PB19]. number [GA18, KCP19, Li14, Pet19, Pet18, PH16]. numbers [Can18, JD12]. Numerical [Ben15, MRJ^{+}19, CDPS18, EFG^{+}14]. NVHT [HZHS18]. NVIDIA [JM15]. NVM [ZLH^{+}18].

O [DLW^{+}12, GFPC14, No12, WHW^{+}17]. O-intensive [HZZ^{+}19]. obfuscation [MMN^{+}18]. Object [Lop13, HD10, LLLC15, LC11, SA19, SRB^{+}19, SFHS19, TCS^{+}10]. objective [ADDB18, COV13, COF^{+}17, CDW^{+}19, FPF14, LÜ14, MMK^{+}11, NL19, SGVRP19, SJVRVVS19, SWLP19, Tal19, dCPD19]. objectives [FEH^{+}14]. objects [AEF11, MB19, SB15]. Oblivious [CRSB13, ABBD14, BFVB19].

OBQA [ESGQ^{+}11]. observability [MH18]. observations [RTZ11, WHW^{+}19]. observatory [AAA^{+}15]. obstacles [SJS11]. obstructed [DWX10]. Obtaining [VAS^{+}13]. ocean [Nes10]. ODEs [FKB17, KKR14]. OFDMA [UM17]. Off [JXZ^{+}19, TOR^{+}14, ACA^{+}19, ECLV12]. offloading [Ale19a, LYJ^{+}19, NHX^{+}19, ZGW^{+}19]. offs [LCB16]. OLAP [DKRC^{+}15]. OLSR [KKK11a]. OLSR-aware [KKK11a]. omnipotent [BBD18]. OmpSs [PSB^{+}19]. on-chip [BYG^{+}18, DJDK19, KH12, LNA12, LLKY13, LSXX14, LLT12, LWCG14, MYD^{+}11, PMCC18, UM17]. On-demand [YYLC11, HZDP12, LSZZ15, NKK16]. On-GPU [LW19]. on-machine [AES11]. on/off [ACA^{+}19]. once [ACHY18]. One [Bog17, BPBR11, Lai14]. one-to-many [Lai14]. Online [CRH11, DTK11b, JTC^{+}18, KKR14, LQM^{+}12, LHLM14, ZGW^{+}19, ZLMC14, AZC13, AFG^{+}19, BJL18, CXX^{+}18, DFLO17, SHC14, TZI11, WWY^{+}18]. only [SLKK12]. ONoC [TKKH17]. OnRamp [FKR^{+}17]. ontologies [ASHO19]. OP2 [GMS^{+}13]. opacity [KKW17]. Open [DDO^{+}18, Kar19, ZSW14]. open-source [ZSW14]. OpenACC [OGM^{+}19]. OpenCL [AB13, MC17, PHW^{+}13, PSB^{+}19, RBB17, Str12, dAT17]. OpenMP [LNW^{+}12, PARB14]. OpenMP-based [LNW^{+}12]. Operating [PTN^{+}19, ASSS19, FABG^{+}19]. operations [DT11, SLZ^{+}19]. operator [CKLW19]. operators [SMO14, WH17]. Opportunistic [LYJ^{+}19, DBW^{+}18, LWW18, WW18a, WWA^{+}18, dKG^{+}10]. Opportunities [PJ18, ATKT19]. opposition [WRW13]. opposition-based [WRW13]. Optical [DR18, CS10, KK17, KH12, LY13, PLD14, SQQL19]. Optimal [AH12, CCC^{+}19, HLL^{+}19, JLY12, Lai17, LL12b, Li14, Li19b, MPR19, YA11, ZV14, BPBR11, DKKV15, DLM19, EB13, Li10, LCB16, MPG17b, NZA13, PW16, RTZ11, Tam18, VS16, VAS^{+}13, WIB12, ZQMM11]. optimisation [AD12]. optimising [PVRS17]. Optimization [CGN^{+}13, CLA^{+}18, DDGK13, DDE19, MBW16, MC17, PMAL11, SLHS19, ALM^{+}16, AS19a, APK18, ADDB18, BNBR16, BDGR13, BHLT14, BMS19, BYH^{+}17, CMMT13, CCK11, DJH11, GZG^{+}17, GL12, HVW16, JZZ^{+}17, KS18, LL10, LQM^{+}12, LBT19, LGK^{+}12, MZC18, NS12, NST19, NL19, Ozt11, PTN^{+}19, RCG18, Ren11, SGVRP19, SS11, SWLP19, SPPA19, Str12, TCMB^{+}19, TPS^{+}18, WCL^{+}13, WRW13, WQL14, WMG13, XLHT13, XLH18, ZZJ^{+}18, ZV12, ZWWX16, dCPD19]. Optimization-based [PMAL11]. optimizations [CZPP16, LJZ^{+}19]. optimize [TdAR18]. Optimized [LMXJ18, WJ14, Ana14, DKC14, LHCC19, Pet18, SBRM19, TW15]. Optimizing [ASA18, CC16, JST12, LM16, MLW^{+}19, PGRP17, SBÇ12b, WCWO17, WG11, WSLC11, AFNT17, AHA^{+}16, DV13, FMIF18, GYY^{+}14, PB19, ZGG^{+}14]. Optimum [BHK17]. orchestration [BYT19, PVP18, RCG^{+}11]. order [BMLLC^{+}19, CB15, KKW17]. Ordered [HCR12, CG10, KKS^{+}12, SW18, YLB^{+}15]. ordering [Zah12]. organization [AAH17, HKK^{+}18]. organizing [IZ12, KO11, MYM10]. oriented [CF19, CSW^{+}17, DZC17, DWYB10, GYAB11, HdR13, HRM17, KHW13, Kum17, LWWQ18, MXSL12, PSGS17, SCG10, WWY^{+}18]. ORN [SK11]. Orthogonal [JD12]. other [Kum17]. OUSNs [LWW19]. out-of-core [KR11]. outcomes [NKSA17]. outer [CTKA17, CKLW19]. outerplanar [TSFZ14]. outsourced [XLC^{+}18]. outsourcing [CXY14, YZC^{+}19]. overall [XL11]. overhead [LMGLGLG17]. overlap [ALTV13]. overlapped [SWLZ17]. overlay [BMS19, BHK17, CMMN10, EDH^{+}17, LSS^{+}11a, LSS^{+}11b, RA11, SB12]. Overlays [HASB16]. oversubscription [KKLJ14]. overview [SSZ10].

P2P [AS19b, CFI^{+}18, DW12, EDH^{+}17, FZ14, GB11, LL19, LZY11, Luc18, MAPF14, RS19, SK11, WCXL11, YCH^{+}10]. PA [SRT^{+}18]. PA-Star [SRT^{+}18]. packages [DAB^{+}14]. Packet [ZY12, CKC19, CK13, EKNS17, HBS17, HDCM11, KK10, LW19, UM17]. Packing [UFF19, CRD12, TSFZ14]. Page [Ano18y, Ano18z, Ano18-27, Ano18-28, MTM10]. paging [Li17]. PAHON [DR18]. Pairs [DCA^{+}15]. PAME [YLZW18]. PaMeLA [GDL^{+}11]. pancyclicity [XHZZ16]. paper [OY13]. Papers [Ben15, BÇ19, Mue13, Phi13]. paradigm [KDSS18, LK15, Sie16]. Parallel [AS13, AK19, ASC^{+}18, BV13, BKCM17, BW18, COV13, CP10b, CRD17, Cuz11, DDO^{+}18, DFHH13, DDGK13, DDE19, FLL14, FZWL12, FGcF17, FTM^{+}14, FPS11, GMSS^{+}11, GNZ18, HES11, KKN13, KRS13, LWCC15, Las12, LSS^{+}11a, LST^{+}13, LK11, LKB^{+}15, MSGS^{+}13, MZC18, MSAZ10a, MNK12, NST19, NZA13, OY13, OT19, PDP17, PCX^{+}14, PBB^{+}17, PRS14, REK10a, RSL12, RPN19, SHRM19, SLKK13, STS19, TH11, Tát11, Upa13, WRW13, XP10, YBX^{+}13, YWAT13, dCPD19, AKDMN15, Ada17, AS19a, ABGV11, AFG^{+}19, AMM^{+}18, ADDB18, AB13, AJG18, BKC^{+}15, BNBR16, BKMT14, BCK^{+}13, BSH15, CP10a, CTS17, CDS10, CSMML10, CCE^{+}17, CVK^{+}18b, CZZ^{+}17, CLOL17, CFJW13, CKWT17, CMC^{+}19, CM12, CB11, DMK19, dADC18, DMG18, DLW^{+}12, DAG^{+}17, DRR13, EB13, FKB17, FKR^{+}17, FCG^{+}14, GMMP12, GVBB13, GG19, GMVRGS16, GAC^{+}17]. parallel [HSS10, HZZ^{+}19, HSH10, HdR13, HN19, HSS17, mH14, JJ12, JST12, JZF^{+}15, KKR14, KR10a, KR10b, KHT^{+}14, KKS^{+}12, KCR14, KN18b, KN18a, KSSG14, KBC^{+}10, KIH15, LBMG15, Las13, LPK^{+}10, LGRV19, LY12, LMB^{+}17, LJZ^{+}19, Li19a, LÜ14, LZZ^{+}11, LTG14, LGL13, LFEP19, MTL^{+}18b, Men18, MMK^{+}11, NVK^{+}11, NDW17, NSDZ18, NZY^{+}11, OTKT12, PMAL11, PPTV^{+}10, PA15, PQ19, PPSV15, PGKV18, Rao16, RAN^{+}17, ROB^{+}18, RSK19, RBB17, Sch14, SPH13, SCMH13, SWLP19, Ski16, SMH^{+}14, SSdlB^{+}10, SR16, Suk18, SHC14, SRT^{+}18, SSGZ13, Tal19, Tam18, Ter16, TLW18, UGG^{+}11, UFF19, VS16, WLL16, WLCZ15, WIB12, XL11, XS11, XYZW14, YÖ11, YDTZ18, YBM13, Zha11, ZBW^{+}17, ZHG^{+}19, Cza13, FTK14, KR11, YÖ11]. parallelisation [AD12]. Parallelism [Bog17, CGN^{+}13, XMMD17, MM15, Ozt11, RDCQ17, TBG^{+}17, VBF13, WYTX13, ZLWL12]. Parallelization [Kub17, LMY^{+}11, MPN17, Nes10, QGZP19, WCEA10]. parallelized [WZY^{+}19, ZMZJ17]. Parallelizing [LLS^{+}16, SGVRP19, WCH^{+}17]. Parameter [ZRN^{+}14, APK18, LZY^{+}18, DAPR18]. parameters [WRW13]. parametrisation [MLCFH^{+}18]. Parity [MC17]. Pars [BJ15]. parsimony [SJVRVVS19]. part [SQQL19, RLA^{+}16, RLA^{+}17]. Partial [HBS17, DGDF10, IR12, LÜ14, OT19, TR16, WHW^{+}19]. partially [KKS^{+}12]. participants [GHK^{+}12]. participation [AK18]. particle [LHWJ19, VBDRC13]. Partition [SCG10]. Partitionable [LC14b, PW17]. partitioned [GÖÖ16]. partitioning [ASA18, AHA^{+}16, DKUÇ15, ES12, GHC^{+}17, LSXX14, LZLX11, RMU14, STA12, SLKK13, LWC^{+}18]. partitions [SBÇ12a]. pass [MPN17]. passing [DDNT10, IRRS16, Kak15, KMS10, KS13, PS14, TGPUC16]. Password [Lop18, YTH^{+}19]. password-authenticated [YTH^{+}19]. Password-based [Lop18]. patch [GA16]. patch-based [GA16]. Path [MKM16, CHCG18, DGNW13, GHIJ19, LLFJ18, VLL^{+}14, YC12, DCA^{+}15]. paths [Lai14, Lai15, Lai17, MT14, NCA^{+}12, WFLJ16]. pattern [PDP17]. Patternlets [Ada17]. Patterns [AM17, AM13, Ada17, CTS17, ETS14, HHA14, HKK^{+}18, KIH15]. Paving [APV18]. payments [CSS11]. PBS [GPJA10]. PC3 [AHG12]. pCT [KDO^{+}13]. PDC [AYB^{+}15, Kum17]. peak [YJKD10]. Pedagogy [GAC^{+}17]. Peer [HBF12, LCCL10, NMN^{+}14, SJG19, TMK^{+}17, AS18, BBB11, CTC11, CGKY12, LKS14, LLW12, SAL10, ZCMY12]. Peer-to-Peer [LCCL10, SJG19, TMK^{+}17, HBF12, NMN^{+}14, AS18, CTC11, LKS14, SAL10, ZCMY12]. penalty [CK13]. people [HRM17]. per-core [LSC^{+}15]. per-object [LC11]. per-user [LC11]. perfectly [ZLKK19]. Performance [AMSÅ19, CGKY12, DWYB10, DB18, FRM15, GLGLBG12, HTB19, KRS13, KRS14, LGRV19, LNW^{+}12, LYW^{+}16, MSAZ11, MBG^{+}17, NSKN17, NSA11, Nee17, PARB14, PBB^{+}17, RSK19, TLY12, TPJ^{+}19, TdAR18, XMMD17, YAK15, AM13, AA10, AFH^{+}19, AS19a, ARI17, AD12, BCD^{+}15, BDGR13, CG17, CCE^{+}17, CVK^{+}18b, CKWT17, CSW^{+}17, Cuz11, Cuz13, DR19, DJH11, DXS^{+}19, DKK18, DF12, DYL^{+}12, ETS14, ECLV12, FHL^{+}15, FCP^{+}15, GJ12, GMSS^{+}11, GYY^{+}14, HES10, HBSASA19, HHS12, HRG^{+}11, HdR13, ICQO^{+}12, JST12, KVNV17, KyLPC17, KWZ19, KCR14, KZ11, KC17, LWC^{+}18, LWCC15, LC13, LSXX14, LJZ^{+}19, LB12, LZZ^{+}11, LGL13, LB18, LCB16, LGK^{+}12, LWWQ18, MC17, MSGS^{+}13, MZC18, MRS^{+}14, MBO11, MLK12, MRJ^{+}19, MGRRK14, ND12, No12, PCMM^{+}17, PB19, PHW^{+}13, PVRS17, PGKV18, RPS19, SPRG^{+}12, SSFP11, STMZ18, SA11, SE15, SR16]. performance [TTH12, UMM^{+}18, WG11, WLZ^{+}18, YAA10, YZW^{+}15, YYWZ19, ZWY^{+}15, ZKZF18, ZW13, ZWQ^{+}16, ZLCZ18, dAT17]. Performance-constrained [YAK15]. performance-portable [MRS^{+}14]. performance/power [RPS19]. periphery [ABLP17]. perishable [GAOHG17]. PERP [ZWY^{+}15]. persistency [CRB19]. persistent [ST14]. Person [XCC^{+}19]. personal [HBF12]. personalized [PW16]. perspective [FSP18, HRM17, LNC13, NXTK17, RBP^{+}11]. Perspectives [PRS14]. perturbation [CHX^{+}17]. Pervasive [NDW17, KKKP12, Kol19, Ksh12, Sie16]. Pessimistic [MMCL^{+}17]. Petascale [SWHB17, RFP^{+}19, WYTX13]. Petri [BYT19]. phase [Man13, SNCP12]. phases [GA18, SSGZ13]. PHAST [DGNW13]. Phi [CHLL18, RPN19]. philosophers [AFNT17]. Phi^{T M} [KVNV17]. Phoebus [KSB11]. photon [FLL14]. photon-mapping [FLL14]. Photonic [APRA18]. Photonic-based [APRA18]. Physical [QGB^{+}17, SNMB16, BC11, CSW^{+}17, DZC17, HRM17, JWH^{+}17, LLWC17, SLG^{+}18]. Physical-aware [SNMB16]. physics [CP10b]. Pi [EHKSS19]. PIC [YBX^{+}13]. pictures [FGcF17]. Piece [CTC11]. Pilot [LSZJ15]. Pilot-Data [LSZJ15]. pipeline [PW17]. Pipelined [GÖÖ16, BDGR13, LHHH11]. pipelines [WG11]. pivoting [ADV14]. placement [FMIF18, GM14b, ISAZ10, LBT19, LRS18, MSRB19, MTM10, PA15, WCWO17, WZX^{+}19, WSLC11]. planar [DCA^{+}15]. plane [CRJ10a]. plane-based [CRJ10a]. planning [CHX^{+}17, LHW^{+}19, MKM16]. plasticity [RBOH^{+}18]. platform [Ale19a, AMSÅ19, AM11, BSH15, CJYC19, CVK^{+}18b, CS17, CB11, Cza13, FLL14, LTG14, PLSM18]. platforms [Ale19b, BLMB13, CGL^{+}14, CDR12, FCP^{+}15, GAOHG17, LSC^{+}15, LLS^{+}16, MBBD13, MSV19, PP13, SSVC10, VLW18, WJ12, WJ14, YFS^{+}15, ZXMR18, ZHG^{+}19, dSS11]. playing [RPN19, WW18b]. plex [WCH^{+}17]. plucked [CKK^{+}13]. plus [WXMZ19]. PMSS [HHA14]. Point [REZN17, Can18, CNLGRL18, Dav17, SHRM19, SSGZ13]. point-to-point [SSGZ13]. point-value [SHRM19]. pointer [KHK18]. pointers [LS19]. Poisson [JGMY17, WJ14]. policies [BVGV14, CG17, FCJG^{+}18, GHC^{+}17, MP10, SS17]. policy [ARD14, EHL^{+}15, LWLD12, LL12a, SMB10]. polling [GHK^{+}12]. Poly [AF17]. Poly-logarithmic [AF17]. polymorphic [ETS14]. Polynomial [GPJA10]. pools [AFD^{+}11]. population [MS15]. Portability [SGdSS13, AFH^{+}19, ETS14, LGRV19, PHW^{+}13]. Portable [VAF19, LFGM17, MRS^{+}14, MLK12]. portal [FKR^{+}17]. possession [LZY^{+}18]. possibly [MCS14]. potential [ARD14]. Power [CG17, DR18, KCR14, MAHKZ12, WQL14, ARI17, AG12, CJYC19, CZPP16, CHCG18, DZC17, HZL18, HLL^{+}19, JHF^{+}17, KK11, LM16, LB12, MGRRK14, OJP^{+}18, RPS19, Ren11, SZL10, TJCB10, TVT^{+}17, YBX^{+}13, YA11, YZW^{+}15, YJKD10, ZLKK19, ZV12, ZCF^{+}17]. Power-Aware [DR18, KCR14, WQL14, CHCG18, OMT^{+}17, SZL10]. power-constrained [JHF^{+}17]. power-gating [CZPP16]. Power-performance [CG17]. powered [TAM^{+}19]. PPM [LW16b]. practical [FSP18, LXW^{+}11, Suk18]. pre [GDCC18, HMR15, SJS11]. pre-assigned [HMR15]. pre-detection [GDCC18]. Pre-Processed [SJS11]. precedence [XLL15]. Precise [KSJC17]. precision [BGBC^{+}16]. preconditioned [CP10b]. preconditioner [GLW14]. preconditioning [CASD18]. predicates [Ksh12, SKK14]. Predictability [SB12]. Predictable [SB12, YYWZ19]. predicting [BCD^{+}15, KBC19]. Prediction [ASKO16, AYB^{+}15, DBW^{+}18, WDS^{+}18, WWA^{+}18, ARVZ14, CXX^{+}18, CXQ^{+}18, DZC17, DKC14, KVA18, LGZ^{+}10, LC14a, LKM12, LWWQ18, MVP17, PMdO11, Udd19, WWY^{+}18, WZH^{+}19, uRIL^{+}18, WHW^{+}19]. Prediction-based [AYB^{+}15, DBW^{+}18]. predictions [LSH^{+}13, NVK^{+}11]. predictive [BYH^{+}17, SNMB16]. preemptable [LQM^{+}12]. preemption [SJB12]. preemptive [JTZZ11]. preferences [WMY^{+}17, WTY^{+}18]. prefetch [Zha11]. prefetching [HD10, LAK10, SSGG18]. prefix [EB13, SPH13]. prefix-based [SPH13]. Pregel [XYZW14]. Preparing [GS18]. presence [HZA^{+}15, PMHM19, WE13, WSLC11]. preservation [GSASA19]. preserved [SWW^{+}17, YGWJ19]. Preserving [WBS19, AKK^{+}19, CXY14, SRB^{+}19, TKR^{+}19, TH19, WML^{+}18, YTZ19, ZLT^{+}19]. prevention [BYT19]. pricing [AKSZ19, ZV12]. primitives [AF17, BBH^{+}17]. principal [VLW18, AHG12]. principle [GXYZ13]. principles [DAG^{+}17]. Prior [KHN17, SHK19]. prioritized [LASS15]. priority [MAKWZ13, SR16]. Privacy [AKK^{+}19, CXY14, TH19, YTZ19, ZLJ^{+}19, BJL18, GSASA19, LHW^{+}19, LLDL15, SRB^{+}19, SWW^{+}17, TKR^{+}19, URK^{+}19, WML^{+}18, WLYS19, YGWJ19, ZLT^{+}19]. Privacy-aware [ZLJ^{+}19]. privacy-based [WLYS19]. privacy-preservation [GSASA19]. privacy-preserved [SWW^{+}17]. Privacy-preserving [AKK^{+}19, CXY14, TH19, YTZ19, WML^{+}18, ZLT^{+}19]. Private [REK10a, REK10b, CKMP17, LTWW12, SHK19]. privatizing [QGZP19]. privileges [LS19]. Pro [KV10]. Pro-active [KV10]. Proactive [TXLL14, WMES12, DW12, FX10, KAA^{+}19b, WWY^{+}18]. Probabilistic [SCMS12, ESCV15, JHPL13, WMG13]. probabilities [XCC^{+}19]. probability [DJH11, GXYZ13, LNAL17, LXLS12, NGQM12]. probability-based [GXYZ13]. Problem [BSH15, BCMV15, BW18, DBA^{+}18, dADC18, EMC19, FZWL12, HC11, LCCL10, LLCZ19, LFEP19, OA10, PDB13, Sch13, Sta17, WLL16, WCEA10, WMG13, Cza13]. problems [CMMT13, GRS19, Men18, NST19, Pet19, PPSV15, WRW13, WMG13, YS11, ZTFK16, dCPD19]. procedure [Kub17]. process [Ale19b, GA18, HRF^{+}11, MEMEMH17, MSEM^{+}19, SDG17, TKX^{+}13, WMES12]. process-level [WMES12]. Processed [SJS11]. processes [GK15]. Processing [AGWY11, DDGK13, Eme13, GLGLBG12, OY13, SLHS19, VAF19, YL12, YJL16, AAA^{+}15, ATDH13, AM11, BK13, BAT^{+}19, BHS13, CLA^{+}18, CHLL18, CM12, DFLO17, EKNS17, HBS17, JKD^{+}15, KKN13, KN18b, KN18a, LB12, LW19, LL18, LKB^{+}15, MTL^{+}18b, NHX^{+}19, NLB^{+}18, PV19, PYP^{+}10, PGP^{+}12, RCG18, Ren11, RAN^{+}17, SIY14, SS18, SCLL10, SI13, VETT18, WL10, ZMCP11, ZHH15, PRS14]. processing-in-memory [NHX^{+}19]. processor [DK11, LC13, LY13, NPVG^{+}19, PR13, SHL^{+}13, SAJ13, SE15, SHRM19, TdAR18, WIR^{+}18, XP10, YBM13]. processors [ARI17, BM17a, CN14, CCK11, CHLL18, CKK^{+}13, CRSB13, CMC^{+}19, DDG^{+}17, DWYB10, FSP18, JJ12, JHF^{+}17, JZF^{+}15, Li19a, LV15, NS12, NZ17, SPC^{+}17, SNMB16, SP13, XTN12, ZXB14]. producer [KK11]. producer-consumer [KK11]. Product [AFG^{+}19, ISAZ10, JD12, MSAZ11]. production [TDBL13, TYD^{+}19]. productivity [VFAD17]. products [CP10b]. profiles [YWAT13]. profiling [KC17, uRIL^{+}18]. Profit [LWZZ12, KSSK16, LLCZ19, ZV12]. Profit-driven [LWZZ12]. programmability [KWZ19]. programmable [EAB^{+}19, HHA14, PYP^{+}10]. Programming [COV13, DRR13, VBF13, VFAD17, AB13, AJG18, BYG^{+}18, Bog17, BHS13, BLZ^{+}18, CTS17, CCE^{+}17, CMR19, Eij18, FGcF17, HdR13, HSS17, IEWK17, KSG13, KZ11, Pet19, RK18, SSdlB^{+}10, TFMS15, YQTV12]. programs [CC16, CAK13, DMG18, GÖÖ16, LPK^{+}10, LZZ^{+}11, SCMH13, UFF19, YDTZ18, ZXB14]. Project [BSH15]. projection [FGL^{+}11, NCA^{+}12]. prone [DDG^{+}17, GK15, OWK14]. proof [ZLT^{+}19]. proof-of-delivery [ZLT^{+}19]. proofs [AP16]. propagated [SHK19]. propagation [BEN12]. proper [NGQM12]. properties [WMY^{+}17, YDTZ18]. properties-aware [WMY^{+}17]. proportionality [KR12, KCR14]. proposal [ESGQ^{+}14, NKK16]. Protected [LS19]. protection [Lop13, Lop18, YGZ^{+}10]. protein [WDS^{+}18, YL12]. Protocol [BHK17, AMT13, ARD14, BDM18, BOY10, CCC^{+}19, EDH^{+}17, GZY14b, HLS12, HZDP12, MAGL13, MPG17a, NPGV10, NSA11, SMPMLVLS11, TLY12, WCCH18, YGWJ19, ZLCJ12, SJS11]. protocols [BJL18, CXY14, CDAN14, KLP10, MMCL^{+}17, MS15, RFS^{+}12, YTH^{+}19]. proton [KDO^{+}13]. provable [LZY^{+}18]. provably [LWK^{+}19]. providing [Zah12]. proving [SHSH17]. provisioning [AMU^{+}19, JAB12, KM17, Kim17, MZZC12, MCZ14, NF16]. proximity [CJDC10]. proxy [WXMZ19, ZVL11]. PS [dAAD^{+}19]. PSCR [MRT18]. pseudo [CVK^{+}18a]. pseudo-spectral [CVK^{+}18a]. PSO [ADDB18, dCPD19]. PTNet [BFH^{+}17]. public [AVAH18, SSX14]. public-key [AVAH18]. Publish [MS19, ZW13, dAAD^{+}19]. Publish/subscribe [MS19, ZW13, dAAD^{+}19]. PUF [BDM18]. PUF-based [BDM18]. Pull [DLLL11]. purpose [CW15, GRS19, LCB16]. pursuit [YpGyLlC13]. pursuit-evasion [YpGyLlC13]. Push [DLLL11]. PUT [HLS12]. pyrosequencing [SPRG^{+}12].

QDaS [LJQ^{+}19]. QoE [KS18, MSRB19]. QoS [BOY10, CLW^{+}19, CKML12, Kim11, Kim17, KKK^{+}11b, LZI^{+}11, RSVW19, SJB12, SA19, WZH^{+}19, YJKD10]. QoS-aware [CLW^{+}19, CKML12, LZI^{+}11, YJKD10]. Quadratic [Cza13]. qualitative [WMY^{+}17, WTY^{+}18]. Quality [LJQ^{+}19, MSRB19, NCRK19, NZY^{+}11, AKK^{+}19, AH11, AH12, DV13, FC14, LNA12, LZN19, PTN^{+}19, TCMB^{+}19]. quality-aware [AH12]. quality-of-service [LNA12]. Quantifying [FX10]. Quantitative [YZW^{+}15, GXYZ13, KC17, WMY^{+}17, WTY^{+}18]. Quantization [AFG^{+}19]. Quantized [FKB17]. quasi [Pet19]. quasi-Newton [Pet19]. queries [BBCQ13, LSZZ15, LKB^{+}15, PAG^{+}18, SSKS11]. Query [HASB16, CHLL18, GB11, JHL^{+}18, KKN13, LZWZ19, LL18, NSAS10, SCLL10, SJG19, WL10, YTZ19, ZHT16]. Querying [TT10, DTK11b]. queue [ESGQ^{+}11]. queueing [MGRRK14]. queues [ACH18].

R [KKN13, LMY^{+}11, TR16]. R-tree [TR16]. R-trees [KKN13]. Racetrack [HZL18]. radiation [KVNV17]. RADIC [CLMRL15]. radio [FCZ^{+}12, GDCC18, GDL^{+}11, RFS^{+}12, SSZ10]. radix [MRT18, VAS^{+}13]. radix-4 [MRT18]. RAFT [MYD^{+}11]. RAID [TTH12]. railing [PKW^{+}10]. Ramos [DBLB^{+}12]. Ranch [LMP10]. Random [BBS13, AGMS16, BBFN12, BCK^{+}13, DJH11, KCP19, Li10, Pet18, SMP15, SCMS12]. randomized [CKC19]. range [CDD^{+}19, GB11, KKN13, LHWJ19, LZWZ19, MKM16, PARB14, YWAT13]. range-free [MKM16]. ranges [CHCG18]. rank [YZC^{+}19]. rapid [XSYG18]. rapidly [Li10]. rare [BV13]. Rate [OJP^{+}18, AGWY11, GA18, Hu11, KHK18, MAHKZ12, SCW^{+}18]. Rate-based [OJP^{+}18]. RAW [SV19]. ray [CS17]. RBF [TYD^{+}19]. rCUDA [RPS19, RS19]. re [LMJC11, RCG18, WXMZ19, XCC^{+}19]. re-authentication [LMJC11]. re-encryption [WXMZ19]. re-identification [XCC^{+}19]. re-optimization [RCG18]. reaction [XLHT13]. Reactive [OOSGVG^{+}16, NPGV10]. Read [IRRS16, AM12b, CG10, IR12]. read-dominated [AM12b]. read-write [CG10]. Read/write [IRRS16, IR12]. readers [FKKR16]. reads [SPRG^{+}12]. Real [LML^{+}10, BVGV14, BDGR13, CF19, CCK11, CRJ10a, CRJ10b, DKRC^{+}15, FSP18, FC14, GZG^{+}17, HV13, JLWX11, JZZ^{+}17, JHL^{+}18, KKW17, MLDG12, MAKWZ13, MVP17, SA19, SRB^{+}19, TYD^{+}19, ZZJ^{+}18, ZHH15, ZQMM11, ZHLQ12]. Real-time [LML^{+}10, BVGV14, BDGR13, CCK11, CRJ10a, CRJ10b, DKRC^{+}15, FC14, GZG^{+}17, HV13, JZZ^{+}17, JHL^{+}18, KKW17, MLDG12, MAKWZ13, SA19, SRB^{+}19, TYD^{+}19, ZZJ^{+}18, ZHH15, ZQMM11, ZHLQ12]. realistic [FTM^{+}19, SJS11]. rear [CXQ^{+}18]. rear-end [CXQ^{+}18]. recall [BGBC^{+}16]. reception [CKC19]. reciprocity [HBF12]. Reco [EHKSS19]. Reco-Pi [EHKSS19]. recognition [CNLGRL18, CWZ^{+}18, CLXX19, CDY^{+}19, RK18, SAR^{+}18]. recommendation [CLW^{+}19, COF^{+}17, LMXJ18, WTY^{+}18, WLYS19]. recommendations [WZH^{+}19]. recommender [HWL18, ZLJ^{+}19]. reconfigurability [ZXYO11]. Reconfigurable [CGFH19, ORWT^{+}18, AM13, AHA^{+}16, EHKSS19, HZL18, Lla17, NL19, PPP14, TJCB10]. reconfiguration [DMG18, DBLB^{+}12, HBS17, JWSG14, LBMG15, LHX^{+}16, NPVG^{+}19, ZBW^{+}17]. Reconstructing [BDG^{+}15]. reconstruction [BDRB14, HES10, OGRV^{+}12]. recoverable [ZSCX18]. Recovery [LY10, TTH12, ZWY^{+}15]. rectangular [SBÇ12a]. recursive [AKDMN15, MM15]. red [BE13, DMI^{+}19]. red-black [BE13]. redirect [ACCP12]. redistribution [KNHH18]. RedSync [FFYH19]. reduce [CRD12, LMGLGLG17]. reduced [LC13]. Reducing [FFYH19, IIH16, ASA18, CK13, RWB^{+}13]. reduction [BV13, BW18, Li17, MS19, SPH13, YAK15]. redundancy [BM17a, RMHR17]. Redundant [CKT11]. ReduxSTM [PGRP17]. Reevaluating [SC10]. reference [SPRG^{+}12]. refinement [ASC^{+}18, DAB^{+}14, GA16, GNZ18]. reflectance [YWAT13]. reflective [KKKP12]. register [YPCW16]. registration [CDD^{+}19]. Regression [MZZC12]. Regression-based [MZZC12]. regressive [KBC19]. Regular [ZHW19, Li10]. Regularizing [SSKÇ15]. regulation [RDCQ17]. reinforcement [HHK15, OPR18, TXLL14, XRB12]. ReKonf [PPP14]. related [BBFN12, CDPS18]. Relating [TJCB10]. Relation [HCR12]. relational [CLW^{+}19, TR16]. relations [KLP10]. relationship [XS11]. relationships [CRWX12]. relaxation [ATDH13]. Relaxing [KKW17]. Reliability [BDGR13, TLLV10, HHK15, JST12, KHW13, TLQS12, TTH12, VRM10, WWW17b, WWY^{+}18, XS11, XLPL19]. Reliability-aware [TLLV10]. reliability-driven [TLQS12]. reliability-oriented [KHW13]. Reliable [AAH17, MPS16, AA16, ACPT15, ATKT19, OWK14, ZW13, DAPR18]. reloading [BBS13]. ReLog [ZTGL17]. Remapping [LW19, YGZ^{+}10]. Remote [PLK^{+}18, BVGV14]. renaming [AP16]. rendering [FLL14]. rendezvous [DJH11, GGY19, MP15]. renewable [AK18, URK^{+}19]. reordering [LMGLGLG17]. repeated [JTC^{+}18]. replacement [BV13]. replica [GM14b, JXZ^{+}19]. Replicable [AMM^{+}18]. replicas [TR16]. Replicated [RJKL11, STA12, ASB18]. Replication [MD13, ARDQ18, BCC^{+}18, KR12, MB19, WW12, WWW17b]. replication-based [WWW17b]. replications [ZV14]. report [Kum17]. repositories [VLGV^{+}18]. Reproducing [CDPS18]. reprogrammable [LLY15]. reprogramming [MAGL13, ZTGL17]. Reputation [HBC15, LS10]. reputation-driven [LS10]. request [ZV14]. require [AF17]. Requirement [HV13, WW18a]. Requirement-aware [HV13]. Requirements [CZPP16]. rerouting [JWSG14]. rescue [WWA^{+}18]. Research [GLW14, Kum17, MLZY17, WZ13, Hua17, LZ11, NPE^{+}19, PSGS17]. Research-oriented [Kum17, PSGS17]. reservations [CRH11]. reserve [SWLP19]. residue [PH16]. resilience [RMGM19]. resilient [DFHH13, TKKH17]. resistive [ZPK^{+}14]. resizable [SR16]. resizing [CPLY18]. resolution [GOH^{+}13, SQQL19]. Resolving [Zha11]. Resource [BVGV14, BSH15, BKK^{+}11, ISAZ10, KM17, SZMK13, SSVC10, ZAB18, AAA^{+}10, ADD17, AS19b, BMS19, BSS^{+}13, CCA18, CDS10, CRH11, CJA^{+}19, CKMP17, DW12, ESCV15, Fu10, HHK15, JAB12, JHF^{+}17, KBC19, LL10, LL12a, LS10, MAPF14, MZZC12, MCZ14, NF16, OJP^{+}18, RCG^{+}11, SSM^{+}16, SNCP12, SCMS12, TFMS15, TKX^{+}13, VD18, VMMB10, XL11, ZLL14]. resource-constrained [VMMB10]. Resource-efficient [SZMK13]. resources [ASKO16, AM11, CFI^{+}18, LKM12, LZI^{+}11, LDP^{+}14, NVK^{+}11, NSDZ18, URK^{+}19, YZS15]. respectable [GHK^{+}12]. Response [TPS^{+}18, HPB^{+}10]. Restart [LACJ18, NC13]. restarts [GK15]. restricted [JZF^{+}15]. resultants [Eme13]. results [HSH10]. Rethink [WW18a]. Retraction [PCX^{+}14]. retrieval [KTP17, SWW^{+}17]. REU [Hua17]. Reuse [BC11, DSEP17, DMI^{+}19]. revealing [AF17]. reverse [NMN^{+}14]. review [ZGJ^{+}18]. Reviewer [Ano10b, Ano11k, Ano12n, Ano14g, Ano15k]. reviewers [Gra10a]. revised [KP17]. revisited [DJ16, GXYZ13]. Revisiting [SPH13]. rewarding [CFI^{+}18]. RF [UM17]. RGA [LHCC19]. ring [LLKY13, LLDL15, RM10]. rings [DLM19, PR12, SMO^{+}18]. RISE [AZW13]. risk [FGL^{+}11, PVRS17, WHW^{+}19]. river [SQQL19]. RNS [PH16]. road [LHW^{+}19, SWLZ17, YTZ19]. robot [IIH^{+}17]. robots [PMHM19, ZBW^{+}17]. Robust [BSS^{+}13, KRS15, PVP18, ZMG^{+}16, BBCQ13, BAT^{+}19, LDS16, LZY^{+}18, MSF^{+}13, SSM^{+}16, SNCP12]. robustness [CKWT17, TdAR18]. Roe [dlAMCFN12]. role [BCD^{+}15]. roofline [KC17, NV19, NSKN17]. root [LXW^{+}11]. rooted [GHIJ19]. route [LHW^{+}19]. router [MYD^{+}11]. routers [ZCF^{+}17]. Routing [CS10, AA14, AA16, AD10, ABF^{+}14, BOY10, BFVB19, CC14, CHCG18, CMN12, CAF^{+}11, DJDK19, DJH11, DBW^{+}18, GHY10, GDL^{+}11, GTGLSA12, Hu11, HJLR12, JLWX11, KLP10, KSK15, LLT12, LY13, LLDL15, MPS16, MSAZ10a, MSAZ10b, MSEM^{+}19, OM10, RFS^{+}12, RB12, SW12, Sch13, SWLZ17, SJS11, TCHC12, WW12, WCL^{+}13, WHC^{+}18, WWA^{+}18, YMLP14, Zah12, ZW11]. RRAM [TOR^{+}14]. RRAM-based [TOR^{+}14]. Run [LLY15, LW16b, LTG14, NVK^{+}11, XL11]. Run-time [LLY15, XL11]. running [GRR13, dSS11]. runtime [BGA12, KNHH18, LFS16, LMY^{+}11, PQ19, SP13, TDBL13].

S [AGWY11]. SABA [ZVL15]. sacrificing [FKKR16]. safe [CDD^{+}15]. salesman [WMG13]. sampling [BBB11, SMP15]. SAT [SHA17]. satellite [SQQL19]. SAUCE [HSS17]. saving [SSGZ13]. SBCI [AS19b]. Scala [GKK^{+}13]. Scalability [PTK^{+}13, dSS11, AFH^{+}19, CLG^{+}16, CP10b, GA16, LdPLC^{+}19, NSKN17, QGZP17, RM10]. Scalable [AS13, AS15, BM17b, CSMML10, DKRC^{+}15, DKJG19, FPS12, JJ12, KG19, KH12, PGP^{+}12, QGB^{+}17, RBA^{+}18, SFC17, TFMS15, TCS^{+}10, VLGV^{+}18, ZXMR18, ZLS17, AKDMN15, ACPT15, ADDB18, CGL^{+}14, CAK13, CJ17, DKKV15, FPS11, GM13, GRZ^{+}18, HSY10, KHT^{+}14, KCFP18, LGRV19, LLB^{+}18, MVP17, NKK16, ND12, RBOH^{+}18, RSK19, Ter16, TCHC12, WCEA10, YQTV12, SLG^{+}18]. scalar [SKH15, Sol13]. scalar/vector [Sol13]. Scale [MYM10, AFG^{+}19, ACCP12, BM16, BDL^{+}19, BCC^{+}18, CC16, CLOL17, DB11, DBCF13, DLW^{+}12, ECWV19, IEWK17, KSSL16, KBC^{+}10, LGZ^{+}10, LZY11, Luc18, LWCG14, MBMC19, MRJ^{+}19, NAB^{+}11, PB19, SFT^{+}13, WCWO17, WBRT13, YZW^{+}15, ZVL11]. Scale-free [MYM10]. scaled [KNHH18]. scaler [VD18]. scales [PLK^{+}18]. Scaling [CVK^{+}18a, YFS^{+}15, FZ14, MBR19, VD18, YÖ11]. scan [PD19]. scatter [dSAJ15]. scatter-based [dSAJ15]. scattering [LPLFMC^{+}12]. SCC [LTG14]. SCDN [SLW10]. scenario [DBW^{+}18]. scene [OGRV^{+}12]. scheduler [ASSS19, HHA14, KAA^{+}19a, LS10, SCG10, ZLWZ18, MSK^{+}16]. schedules [CDR12, ZXYO11]. Scheduling [ATKT19, BD11, BSH15, GG19, JZF^{+}15, MD13, PR12, RAN^{+}17, ALM^{+}16, AAD10, ALLM11, AH12, AM12b, Ben19, BHLT14, BKMT14, CG11, CG12, CHLL18, CRJ10a, CRJ10b, CMR10, CDR12, DBA^{+}18, DK11, DP16, DXS^{+}19, DRR13, EHL^{+}15, FPF14, FCJG^{+}18, GYAB11, GVBB13, GK15, GMVRGS16, GFPC14, HSH10, HV13, JLY12, JHF^{+}17, JBS14, JTC^{+}18, KHN17, KVA18, KYS13, KKK11a, KM17, KV10, Kim17, KNHH18, KK10, KSSK16, KBC^{+}10, LDZ^{+}14, LDZ^{+}17, LWZZ12, LQM^{+}12, LW16a, Li16, LNAL17, LBT19, Li19a, LML^{+}10, LSC^{+}15, LYW^{+}16, LFEP19, MGSG12, MLDG12, MSV19, MCAS12, MMK^{+}11, MAHKZ12, NSAS10, NHO^{+}13, ND12, OA10, OPR18, PW17, PDB13, RBA^{+}18, SSFP11, SPC^{+}17, SJB12, SMO14, SP13, SS18, STK11, SZL10, SR16, SHC14, TLLL10, TLLV10, TLQS12, TDBL13, TXLL14, TDP15, UM17, UFF19, VMMB10, VS16, WL10, WBRT13, gWW18, XLL15]. scheduling [XLHT13, YWG15, ZVL15, ZTFK16, ZY12, ZS13, ZQMM11, ZHLQ12, ZLMC14, dOCS14, FZWL12]. schema [TMK^{+}17]. scheme [AK18, BBS13, CXQ^{+}18, CFL^{+}19, DBW^{+}18, ESGQ^{+}11, GPJA10, KHK18, Kol19, LAK10, LHX^{+}16, LMJC11, LWH^{+}19, LWK^{+}19, LSZZ15, LLDL15, SNCP12, TC13, TCHC12, WW12, WXMZ19, WZY^{+}19, XLHT13, YGZ^{+}10, YJL16, YAA10, YC12, ZCMY12, ZSCX18, ZWWX16, ZBR11]. schemes [CKML12, HDCM11, JWSG14, SHSH17]. science [APV18, BKK^{+}11]. scientific [BCD^{+}15, CXY14, EFG^{+}14, NV19, PB19, RMGM19, WHW^{+}17, YYLC11, ZKZF18]. Scripting [WXZ^{+}18, LMY^{+}11]. sculpture [LMB^{+}17]. SDFGs [BLMB13]. SDN [AK18, HTB19]. SDN-based [AK18]. SDN-enabled [HTB19]. sea [ZWW17]. seamless [ORWT^{+}18]. Search [BSH15, Cza13, AFG^{+}19, AMM^{+}18, Can18, CTT16, CSW^{+}17, ES12, GHY10, HN19, LSS^{+}11a, LSS^{+}11b, MB13, PSC^{+}16, PPSV15, RM10, RM11, ROB^{+}18, Sch13, SJG19, Tam18, WWA^{+}18, YZC^{+}19, ZCS^{+}18, CB11]. searchable [WCCH18]. Searching [SGAC14, KIH15, LTWW12]. secondary [BLZ^{+}18]. secret [LWH^{+}19, YTH^{+}19]. Section [HAC^{+}19]. Secure [CPA^{+}11, EAB^{+}19, PRN^{+}19, ZHT16, ZBR11, BK18, GTGLSA12, JZZ^{+}17, KTP17, LAK10, LZWZ19, LWK^{+}19, LLW12, REK10a, REK10b, SSX14, Sie16, WXMZ19, WCCH18, ZSCX18]. Security [FCJG^{+}18, NL19, DZC17, DKJG19, GSASA19, NZY^{+}11, OM10, TODQ18, TKG^{+}17, WLK^{+}19, XCC^{+}19, ZVL15, ZAAB17, ZZJ^{+}18]. Security-aware [NL19, ZVL15]. security-sensitive [ZZJ^{+}18]. SeeMore [LMB^{+}17]. Segment [MYYY17]. segmentation [KYS13, RK18]. segments [Lop18, SWLZ17]. seismic [KSSL16]. Selected [Ben15, BÇ19]. Selecting [NGQM12]. selection [CKML12, DMI^{+}19, GM14b, JXZ^{+}19, KHN17, LZY^{+}18, LCJ^{+}18, LGK^{+}12, MHLZ16, RTZ11, WLST16, CTC11]. selective [LHCC19, SSGG18]. selectivity [CTT16, GÖÖ16]. selectivity-driven [CTT16]. Self [BPBR11, CDD^{+}15, DPBNT12, FZ14, JM14, LLLC15, Lla17, ASKTZ13, BBS13, BBD18, CDDL10, CAK13, DLV11, DJ16, GK19, GK10, IZ12, KO11, LBMG15, LHX^{+}16, LSH^{+}13, dAMFdS13, MYM10, NPVG^{+}19, PPTV^{+}10, TWQS12, Tur12, WRW13, ZBW^{+}17]. self-adapting [WRW13]. self-adaptive [LHX^{+}16, PPTV^{+}10]. self-correction [LSH^{+}13]. self-deployment [TWQS12]. self-join [GK19]. self-manageable [dAMFdS13]. self-organizing [KO11, MYM10]. Self-reconfigurable [Lla17]. self-reconfiguration [LBMG15, NPVG^{+}19, ZBW^{+}17]. Self-scaling [FZ14]. self-similarity [ASKTZ13]. Self-stabilizing [BPBR11, CDD^{+}15, DPBNT12, JM14, BBS13, BBD18, CDDL10, CAK13, DLV11, DJ16, GK10, Tur12]. Semantic [SLG^{+}18, EHL^{+}15, KLJ^{+}11, LKB^{+}15, MLZY17, MYYY17, MA11, NSAS10]. semantics [MTL^{+}18b]. semi [CTT16]. semi-structured [CTT16]. sense [BC11, ZKZF18]. sensing [CCC^{+}19, GDCC18, ZRN^{+}14]. sensitive [ASSS19, Hu11, JL11, NLB^{+}18, OWK14, RCG^{+}11, SRT^{+}18, WCXL11, ZZJ^{+}18]. Sensor [LDZ^{+}14, LDP^{+}14, THGY15, Amm16, AHG12, Ana14, AMT13, AYB^{+}15, BWP^{+}11, BOY10, BEN12, BJL18, CCW14, CRWX12, DLLL11, DGBN14, DJH11, DKM10, EMC19, EM11, ECP^{+}18, GHY10, GYP13, GZY14b, GM14a, HZA^{+}15, HS12, HZDP12, HJLR12, JF12, JLY12, JBS14, JHPL13, KSSL16, KO11, KO12, KKKP12, KKTZ13, KGN11, LDZ^{+}17, LY10, LL12a, LL12b, Li14, LLB^{+}18, LÜ14, LZC11, LDS16, LWW18, MAGL13, MYM10, MBMC19, NSA11, PLY15, PCX^{+}11, PCX^{+}14, RM10, RM11, REK10a, REK10b, RLP14, RB12, SCN12, SZMK13, SCLL10, SJS11, TLY12, TAM^{+}19, TCS^{+}10, TWQS12, Udd19, VRM10, WL11, WL10, WWA^{+}18, XHZ^{+}10, YpGyLlC13, YDZ^{+}18, ZW11, ZSCX18, ZTGL17, dOBG^{+}15]. sensor-actuator [KKKP12, SCN12]. sensor-based [Udd19]. sensor-cloud [LLB^{+}18]. sensors [AKBD10, AD10, CJDC10, DWX10, REZN17]. sensory [HRM17]. sentiment [XLW^{+}18]. separable [MRT18]. separating [HSS10]. sequence [BCF14, BFKW13, DKKV15, SPRG^{+}12, SMB10, SRT^{+}18]. sequences [BNBR16]. Sequential [LWC^{+}18, SBÇ12b, SLKK13, ZXB14]. serializing [HHS12]. series [LLB^{+}18]. server [JTZZ11, WZX^{+}19, WSLC11, WLZ^{+}18, ZVL11, ZCS^{+}18]. server-side [ZVL11]. servers [AAA^{+}10, DLW^{+}12, LY12, LYW^{+}16, MZZC12]. Service [BK18, NCRK19, ABF^{+}14, BYT19, CCA18, CLW^{+}19, FZ14, JM14, KKKP12, LNA12, LZN19, LB18, MHLZ16, MXSL12, MCZ14, PTN^{+}19, RA11, SB12, SMB10, SSVC10, TR16, TKR^{+}19, WMY^{+}17, WTY^{+}18, WWY^{+}18, WZH^{+}19, YHWY18a, YHWY18b]. service-based [YHWY18a, YHWY18b]. service-oriented [WWY^{+}18]. services [AFG^{+}19, HBSASA19, KSSK16, LWZZ12, LMXJ18, LZN19, MCP^{+}18, SCW^{+}18, Suk18, YAK15]. session [LAK10, MZZC12]. sessions [FSP18]. Set [HCR12, AFD^{+}11, AP16, BYG^{+}18, HES10, HDCM11, JPD17, MHLZ16, MPR19, YSS11]. Sets [XMMD17, FSV14, FSV17, KCR14, SW18, SHC14, YWW12, dOCS14]. setting [Li19b, WLK^{+}19]. several [MCAS12]. shader [PYP^{+}10]. SHadoop [GYY^{+}14]. shallow [dlAMCFN12]. shape [KSJC17, NCA^{+}12]. share [KNHH18]. Shared [JHF^{+}17, AAC10, CCA18, CDAN14, DMI^{+}19, EKNS17, IRRS16, KKR14, KLP10, KMS10, LZI^{+}11, MSV19, SB15, SAJ13, SS17, TGPUC16, ZLWL12]. shared-coin [AAC10]. shared-memory [KKR14, KMS10, MSV19]. sharing [AS19b, CTC11, KK11, KKS^{+}12, LWH^{+}19, LS10, SSX14, WXMZ19, YTH^{+}19, YAK15]. Shield [SSX14]. shielded [CWCW18]. shift [BDL^{+}19]. ships [SQQL19]. shop [DBA^{+}18, LFEP19]. short [LHWJ19, MBS^{+}12, PARB14]. short-range [LHWJ19, PARB14]. short-term [MBS^{+}12]. Shortest [DCA^{+}15, DGNW13, GHIJ19, Lai15, Lai17]. shortest-path [GHIJ19]. Shot [TRS^{+}12]. shrink [REZN17]. Shuffled [KM17]. shuffling [BBB11]. side [ZVL11, WHW^{+}17]. SideIO [WHW^{+}17]. SIFT [LJZ^{+}19]. sign [PH16, RK18]. signal [BAT^{+}19, CLXX19]. signature [WML^{+}18]. signature-based [WML^{+}18]. Silent [DJ16, BCC^{+}18]. Silicon [SDS^{+}18, HRG^{+}11]. SIMD [Ren11]. SIMD/SPMD [Ren11]. similarities [CL14]. similarity [ASKTZ13, AFG^{+}19, BHK17, GK19, KSSG14, UGG^{+}11, XCC^{+}19]. Simple [Ara13, SE15, Nes10, YAA10]. simplex [ASC^{+}18]. Simplified [AS19b]. simulated [dADC18]. Simulating [NFHL13, AAK^{+}13, GN15, LE19, RBOH^{+}18]. Simulation [CZPP16, MRJ^{+}19, STS19, AZW13, AZC13, BBH^{+}17, CGL^{+}14, DAG^{+}17, FTM^{+}19, FCG^{+}14, NSKN17, NPE^{+}19, PARB14, PLD14, PQ19, PTK^{+}13, Rao16]. simulations [AM12a, CCAAS19, DB11, FC14, LHWJ19, VBDRC13]. simulator [CZPP16, dOCS14]. simultaneous [LY10, WIR^{+}18]. Single [HLBM16, KNHH18, Li19a, LPLFMC^{+}12, RFS^{+}12, SSFP11, SPC^{+}17, ZCS^{+}18, PR13]. Single-Chip [PR13]. single-hop [RFS^{+}12]. single-ISA [KNHH18, SSFP11, SPC^{+}17]. Sink [THGY15, LLDL15]. sink-location [LLDL15]. sinks [RB12]. SIoT [SA19]. SIR [ZXGD18]. SIR-based [ZXGD18]. Site [WXZ^{+}18, Hua17]. situation [HRH18]. size [AST12, ASC^{+}18, EB13, JM14, LCJ^{+}18]. size-independent [EB13]. sizes [SMT15]. skeletons [MPS16]. Skinny [BDG^{+}15]. skyline [SCLL10]. SLA [RT18, SMW18]. Slack [KR10b, KR10a]. Slackmin [PDP17]. Sleep [YZX11]. Sleep-aware [YZX11]. sleeping [GDCC18]. slices [DSEP17]. SLID [TFV19]. sliding [MTL^{+}18b]. sliding-window [MTL^{+}18b]. slimmed [YMLP14]. slot [PLY15]. slots [ABBD14]. Slow [HZA^{+}15]. slowdown [MZZC12]. slower [STKW12]. small [CTKA17, GA18, HBS17, JM15, MAGL13, Pet19]. small-large [CTKA17]. smallest [ASC^{+}18]. Smart [ESGQ^{+}11, NCRK19, AKK^{+}19, AMSÅ19, AKSZ19, CF19, DFLO17, HRM17, HRH18, KDSS18, LLWC17, LWK^{+}19, LZN19, MCP^{+}18, NPS^{+}19, NML^{+}19, SLZ^{+}19, Udd19, URK^{+}19, XCC^{+}19, YZS15, ZLJ^{+}19, ZCW19]. smartphone [CWZ^{+}18]. smartphones [LM16]. SMCA [HBSASA19]. smooth [ZBR11]. smoothed [CL14, VBDRC13]. smoothers [WH17]. Snap [BDP16, DDNT10, ADD17, FGcF17]. Snap-stabilization [DDNT10]. Snap-stabilizing [BDP16, ADD17]. snapshot [AEF11, IR12]. snapshots [AST12, KS13, WBS19]. SOAP [ASKTZ13, HBSASA19]. SoC [BLMB13, RBG17, ZAAB17]. social [CMMN10, CLXX19, MPS16, RGAN18, SHK19, WXMZ19, WLYS19, WBRT13, WHW^{+}19, YZC^{+}19]. Socially [LGM18]. Socially-conforming [LGM18]. societies [SA19]. SoCs [LZI^{+}11]. soft [BGBC^{+}16, NPVG^{+}19]. soft-core [NPVG^{+}19]. Software [GKK^{+}13, CV16, DP16, GS18, Kol19, LKD14, MCS^{+}19, NHO^{+}13, RMGM19, RDCQ17, ZMZJ17, ZLT^{+}19]. software-based [NHO^{+}13]. solid [GFPC14]. solid-state [GFPC14]. solution [DP16, GA18, HC11, KKR14, LWW19, LFGM17, YS11, ZAAB17]. Solutions [BCMV15, OT19, TKG^{+}17]. Solver [FKB17, ADV14, CVK^{+}18a, CP10b, Dav17, LPLFMC^{+}12, PP13, PPTV^{+}10, Pet19]. solvers [FHL^{+}15, SHA17]. Solving [BSH15, JGMY17, Men18, PDB13, BW18, CMMT13, CASD18, GRS19, Kub17, LFEP19, MRT18, Ter16, WLL16, WRW13, dCPD19]. Some [RTZ11, MS15, ZHW19]. Sophia [GTGLSA12]. sophomore [GAC^{+}17]. sort [BM14]. sorting [KR11, TW15]. sound [CKK^{+}13]. source [BJL18, LCCL10, MH18, NCB^{+}17, ZSW14]. source-to-source [MH18]. sources [AK18]. Space [PPSV15, SDS^{+}18, AD12, Ara13, BW18, CKK^{+}13, LLKY13, NV19, ST12, SZB16, YQTV12]. Space-efficient [PPSV15, Ara13]. spaces [LdPLC^{+}19]. spanning [BPBR11, BBD18, CFJW13, CFL^{+}19, GHY10, HAC17, KG10, MKW18, RDA18, Ten16, WFZJ12, WIB12]. spark [ZKZF18]. Sparse [PR13, ASA18, ASES15, BGO19, CP10b, CASD18, GMMP12, LHW14, LV15, MLW^{+}19, MBW16, PB15, SLV19]. spatial [CRWX12, JF12, TR16, TYD^{+}19, WCF14]. spatial-temporal [CRWX12]. spatially [Rao16, SBÇ12a]. spatially-explicit [Rao16]. Special [AS13, BÇ19, BS11, CDJL11, CGFH19, DF12, DDE19, DB18, FTM^{+}14, FPS11, FPS12, GMSS^{+}11, HAC^{+}19, KRS13, KRS14, LZ11, Las12, LK10, MSGS^{+}13, MNK12, QGB^{+}17, RLA^{+}16, RLA^{+}17, SLL18, STS19, SFC17, TH11, TFV^{+}15, TFV19, XMMD17, Cuz11, Gra10a, LK11, MKN14, PRS14]. Specific [KRS13, KRS14, MRS^{+}14, RMGM19]. specification [BFL^{+}13]. specify [LS19]. spectral [CVK^{+}18a, GSASA19]. spectrum [FCZ^{+}12, GDCC18]. Speculation [AC16, FKKR16]. speculative [PQ19]. speed [CDD^{+}19, Li16, Li19b]. Speeding [CCAAS19]. speeds [LFS16]. spin [FPM^{+}14]. spin-transistor [FPM^{+}14]. Spintronic [NKV14]. Spline [BNBR16]. Spline-based [BNBR16]. splitting [SJG19]. SPMD [LZZ^{+}11, Ren11]. SPMD-style [LZZ^{+}11]. SpMV [YLL17, ZGG^{+}14]. spoofing [GSASA19]. Sporadic [DKK18, MAPF14, dOCS14]. spot [EMC19, MB19]. Spread [REZN17, SIY14]. Spreading [MBMC19, LpJS^{+}18, ZXGD18]. SR [DYL^{+}12, GRJ^{+}15]. SR-IOV [DYL^{+}12]. SRAM [WCF14]. SS [CLOL17]. SSD [WBS19]. st [BCMV15]. st-connectivity [BCMV15]. stability [LXW^{+}11, WCF14]. stabilization [ADDP19, DDNT10]. Stabilizing [DGDF10, AFNT17, ADD17, BBS13, BPBR11, BBD18, BDP16, CDDL10, CDD^{+}15, CAK13, DLV11, DJ16, DPBNT12, GK10, JM14, Kar19, Tur12]. stable [SKK14, SLW10]. stack [HSY10]. stackable [SSX14]. stacked [TLL^{+}18]. Stackelberg [JTC^{+}18]. stacks [ACH18]. stage [CC14]. staged [AS19a]. Star [SRT^{+}18]. starvation [LASS15]. starvation-free [LASS15]. stash [YPCW16]. State [FKB17, LSH^{+}13, PSC^{+}16, ASKO16, ASB18, AD12, GÖÖ16, GFPC14, PQ19, WHW^{+}19]. State-based [LSH^{+}13]. State-of-the-art [PSC^{+}16]. Static [KKK^{+}11b, YMLP14, BSS^{+}13, PC11, ZXYO11]. statistical [CDPS18]. Statistically [SLZ^{+}19]. statuses [MB19]. stealing [DKKV15]. Steiner [LY10, Sta17]. step [Bog17, KKR14]. STLA [NKV14]. STM [HHS12, PGRP17]. Stochastic [SSM^{+}16, ZS13, BM11, CRHC19, MBO11, WW18b, WMG13]. Stochastic-based [SSM^{+}16]. stop [BCC^{+}18, LLT12]. stopping [AMT13]. storage [FLCB10, HZHS18, HGX^{+}19, JWH^{+}17, KBK^{+}19, KR12, LJQ^{+}19, Luc18, MB19, MAPF14, MPG17a, SSX14, SWW^{+}17, SFHS19, WCWO17, WWW17b, XSYG18, YYLC11, ZWY^{+}15, ZFT^{+}18, ZLZ^{+}19, ZGG^{+}14, ZWWX16]. Stores [SLHS19, JXZ^{+}19, ZWQ^{+}16]. Storm [KKH17]. Strategic [RA11]. strategies [BHS13, CGM14, GM14b, HV13, ROB^{+}18, SSGZ13, Wu11, dCPD19]. Strategy [MD13, ASES15, CTT16, DLW^{+}12, EM11, GOH^{+}13, GMVRGS16, JF12, KVA18, KS18, PLSM18, SRT^{+}18, TLLV10, TW15, WYW15, WSX^{+}19, ZVL11, ZV14, ZVL15, ZLCZ18]. Stream [WQZ^{+}13, AAK^{+}13, AM11, DFLO17, GÖÖ16, KKH17, MTL^{+}18b, RCG18, RAN^{+}17, SS18, ZHH15]. Streaming [PS14, BOKS19, CGKY12, GRR13, GHC^{+}17, JHL^{+}18, LCCL10, WCXL11]. streams [AGWY11, BMLLC^{+}19, ST14, VLGV^{+}18]. StreamTMC [WQZ^{+}13]. stretch [SBÇ12b]. stride [AM13]. string [CKK^{+}13]. structural [RBOH^{+}18]. structure [AFK14, GB11, JdSJC^{+}15, LLFJ18, NZA13, Tam18, XLHT13, YL12, ZLZ^{+}19]. structure-based [XLHT13]. structured [CGKY12, CTT16, DAB^{+}14, MRJ^{+}19, SRI14]. structured-mesh [MRJ^{+}19]. structures [AEY12, Zsa16]. students [Ada17, APV18, AJG18]. studies [CCE^{+}17]. study [BJ18, CHLL18, FRM15, HdR13, LWC^{+}18, LGZ^{+}10, LGRV19, MCAS12, MFT^{+}19, NXTK17, PCMM^{+}17, PP13, PTK^{+}13, RÖE^{+}18, SJVRVVS19, TdAR18, WLCZ15, WMG13, ZKZF18, ZLJ^{+}19]. style [LZZ^{+}11]. subcellular [WDS^{+}18]. subgraph [Pla13]. subgraphs [BCH15]. subproblem [SMT15]. subscribe [MS19, ZW13, dAAD^{+}19]. subscriptions [ST12]. subset [AVAH18, WLL16]. subset-sum [WLL16]. subunit [RK18]. Sufficient [Ste17]. suite [GN15]. sum [AVAH18, WLL16]. summarisation [LJQ^{+}19]. summarization [NML^{+}19, SGVRP19]. super [PW17, SE15]. super-matrix [SE15]. super-pipeline [PW17]. supercomputer [SWHB17]. supercomputers [GRS19, TDBL13]. supercomputing [KRS15]. superposition [dSAJ15]. superscalar [LC13]. supplier [SK11]. Support [YFS^{+}15, DRR13, GB11, HPB^{+}10, Hus17, Kim11, NSDZ18, PB19, PQ19, RS19, SDS10, SRI14, TGPUC16, ZBR11, LST^{+}13]. supported [NPS^{+}19, YPCW16]. supporting [LSZZ15, LHCC19, ZTGL17]. supportive [FCJG^{+}18]. suppression [DZC17]. surface [VBDRC13]. surveillance [NML^{+}19, PLSM18, SMP17]. survey [CPJ^{+}19, CGC16, DAB^{+}14, FEH^{+}14, FMIF18, GM14b, GK10, HLBM16, HBC15, JHL^{+}18, KWZ19, SCN12, SRI14, SHA17, TH19, TKG^{+}17, Upa13, WLK^{+}19, ZAB18]. susceptibility [DFST13]. sustainability [AK18]. sustainable [LS10]. sustained [RMHR17]. SW [RBG17]. Swarm [LdPLC^{+}19, ZGJ^{+}18, dCPD19]. sweep [GM14a, CMR10]. Switch [CRD12]. Switch-based [CRD12]. switched [LWCG14]. switches [VAS^{+}13]. switching [BKCM17, CC14, KG10, LCCL10, LWLD12, STKW12, ZPK^{+}14]. Sybil [YXX13]. Symbol [OWK14]. Symbol-level [OWK14]. symbolic [WD18]. Symmetric [GGY19, ABGV11, ADV14]. Symposium [OY13]. Synchronization [CTC^{+}10, FFYH19, HLS12, HZDP12, Tau16]. Synchronized [LNA12, XLL15]. synchronous [ARP18, ABBD14, DGDF10, KVNV17, MCS14, MEMEMH17, TBG^{+}17]. synchrony [CB15]. synthesis [BYG^{+}18, CKK^{+}13, TdAR18, WD18]. synthesize [DSEP17]. synthesized [MC17]. synthetic [AAK^{+}13]. System [FKB17, HWLR14, SLHS19, ZLH^{+}18, ASSS19, BGA12, BSS^{+}13, BYH^{+}17, BJ18, CLMRL15, DXS^{+}19, DTK11a, DLW^{+}12, DMS^{+}16, FFYH19, GTGLSA12, GSASA19, HLBM16, HWL18, HMY^{+}18, HGX^{+}19, HHA14, Hus17, KHN17, KAA^{+}19a, KSB11, KS13, LMSK18, LLWC17, LAS^{+}19, LY13, LHZ^{+}18, LAC18, No12, PKN10, PLD14, RV13, RBA^{+}18, RAN^{+}17, SPRG^{+}12, SSM^{+}16, SFT^{+}13, SSX14, SLG^{+}18, SV18, TKR^{+}19, Udd19, WHW^{+}17, WZQ^{+}13, WYTX13, gWW18, YCH^{+}10, YXW^{+}18, ZHH15, ZFT^{+}18, ZKZF18, ZGW^{+}19, ZW13, dAAD^{+}19, AGWY11, Sie16, Ski16]. system-on-chip [DMS^{+}16, LY13]. systematic [ZTGL17]. systemic [LZN19]. Systems [AS13, AS15, DDO^{+}18, QGB^{+}17, SFC17, ALM^{+}16, AA16, AAK^{+}13, AD12, ACCP12, AAI^{+}15, ABBD14, BBCQ13, BDGR13, BOKS19, BKMT14, CNLGRL18, CKWT17, CTC11, CRJ10b, CASD18, CJA^{+}19, CP17, CAF^{+}11, COF^{+}17, CSW^{+}17, CCC^{+}19, DZC17, DB11, DR19, DDNT10, DGDF10, DWYB10, DBLB^{+}12, FWM^{+}10, FPS11, FLCB10, FX10, GMMP12, GZG^{+}17, GMVRGS16, GRR13, HBC15, HLL^{+}19, HHK15, IRRS16, JKIE13, JST12, JL11, JZZ^{+}17, JWH^{+}17, Kak15, KKR14, KHW13, KVA18, KVNV17, KyLPC17, KSG13, KMS10, Kub17, LLLC15, LWC^{+}18, LFS16, LGZ^{+}10, LZ11, LAK10, LASS15, LQM^{+}12, LNAL17, LLCZ19, LAS^{+}19, LPLFMC^{+}12, Lop13, Lop18, LS19, Luc18, LXZ13, LLW12, LHCC19, MGSG12, MLMSMG12, MB13, MCS^{+}19, MP10, MMK^{+}11, MAHKZ12, MAKWZ13, MLK12, MSK^{+}16, MGRRK14, MRT18, NL19, NLB^{+}18, NFHL13, ND12, NZY^{+}11, OPR18, PLSM18, PB19, PTN^{+}19, PC11]. systems [PSB^{+}19, PH16, PMdO11, QGZP17, RLA^{+}16, RLA^{+}17, RPS19, RÖE^{+}18, SSFP11, SW12, SLV19, SPH13, SFT^{+}13, SCMS12, SHL^{+}13, SS18, Sie16, SLKK13, SI13, SFHS19, TLLL10, TLLV10, TLQS12, TFMS15, Ter16, TCHC12, VMMB10, VS16, WCWO17, WLST16, WZZ^{+}17, WWW17b, WWY^{+}18, Wu11, WSLC11, XLL15, XLHT13, XLPL19, YLL17, YHWY18a, YHWY18b, YQTV12, YZW^{+}15, YYLC11, YZX11, ZAB18, ZGJ^{+}18, ZLKK19, ZAAB17, ZZJ^{+}18, ZWY^{+}15, ZTFK16, ZLJ^{+}19, ZCW19, ZQMM11, ZBW^{+}17, dlAMCFN12, FPS12, ORWT^{+}18, TFV19]. Systems-on-Chip [ORWT^{+}18]. systolic [PYP^{+}10].

T [CRJ10a, PTK^{+}13]. T-L [CRJ10a]. Table [HZL18, LACJ18]. Tables [TT10]. Tabu [BSH15, Cza13, CB11]. TACD [HGX^{+}19]. Tackling [SMT15]. Tag [VRGS17]. tail [ASSS19, SLZ^{+}19]. tailing [YDZ^{+}18]. taint [WXZ^{+}18]. Talent [JL11]. Tall [BDG^{+}15]. Tall-Skinny [BDG^{+}15]. target [CJDC10, KO11, NDP13]. targeted [BKK^{+}11]. targets [CRWX12]. Task [FZWL12, JTZZ11, SMO14, UFF19, AAK^{+}13, AJG18, CDS10, DK11, DDG^{+}17, GQZ18, JL11, JTC^{+}18, KHW13, Kim17, Li16, Li19b, LSC^{+}15, LZLX11, MTL^{+}18a, OA10, PKN10, PA15, SP13, STK11, SZB16, TDP15, VS16, YWG15, ZTFK16, ZHG^{+}19, dOCS14]. task-based [AJG18]. task-parallel [ZHG^{+}19]. task-scheduling [Kim17]. tasks [AS19a, BFMT^{+}18, BHLT14, CCK11, DRR13, GK15, HMR15, HWLR14, LMGLGLG17, LQM^{+}12, Li19a, PDB13, SSM^{+}16, SBÇ12b, SNCP12, XLL15, ZHLQ12, dSS11]. Taxonomy [FEH^{+}14, HBC15]. TCP [BM11, VLL^{+}14]. Teaching [CTS17, Eij18, LB18, PBB^{+}17, PGKV18, Ada17, FKR^{+}17, GAC^{+}17, HSS17, Kum17]. teamwork [NKSA17]. TEASE [ZBR11]. technique [ASKTZ13, CRD12, HBSASA19, KK11, Nes10, RBB17, WCF14]. techniques [ARP18, BBR13, DJDK19, LPK^{+}10, MBW16, RM11, RSVW19]. technologies [SJVRVVS19]. Technology [CGFH19, PBB^{+}17, PGKV18]. TEES [ZWWX16]. Temperature [SWHB17, AS19a, ZWWX16]. temperature-constrained [ZWWX16]. template [EFG^{+}14]. Temporal [SHL^{+}13, CRWX12, WCF14, XYZW14, YDTZ18, DFLO17]. tenant [PVRS17, YHWY18a]. tensor [IEWK17, LGK^{+}12, MLW^{+}19, RSK19, SMH^{+}14]. term [BV13, LKM12, MBS^{+}12]. terminal [Li17]. terminals [HB11]. Terminating [MS15]. ternary [KRM14]. test [ALLM11, LTG14, NCA^{+}12, dMS18, ALLM11, KCP19]. testbed [HGFF10]. testing [KCP19]. tetrahedral [CZZ^{+}17, LWCC15]. text [BV13, LHCC19, PAG^{+}18, SGVRP19, SWW^{+}17, WD13]. their [CRWX12]. Theorem [SHSH17]. theoretic [KK10, MGRRK14, PC11]. Theoretical [LZC11, CKT11]. Theory [ZLCJ12]. Thermal [NHX^{+}19, SHSH17, LFS16, OJP^{+}18, SNMB16]. Thermal-aware [NHX^{+}19, LFS16]. thermally [TKKH17]. theta [LL18, STMZ18]. theta-join [LL18]. things [AMU^{+}19, TKR^{+}19, CCC^{+}19, CDPS18, DAPR18, ECP^{+}18, HAC^{+}19, HMY^{+}18, LAS^{+}19, LJQ^{+}19, MS19, NLB^{+}18, RSVW19, WSX^{+}19, WHC^{+}18, WCCH18, YWJ^{+}18, ZLT^{+}19]. thinking [CCE^{+}17]. Thinning [KLP10]. Thread [KCSS18, LW19, OTKT12, CGM14, CDAN14, DWYB10, LK13, RDCQ17]. Thread-Data [LW19]. thread-parallelism [RDCQ17]. threaded [ASSS19, BBH^{+}17, LK15, PYP^{+}10]. threads [PMdO11]. threat [HMY^{+}18]. threats [CWCW18, MMN^{+}18, TKG^{+}17]. three [ANEA13, LDS16, YJL16]. three-body [YJL16]. three-dimensional [ANEA13, LDS16]. Threshold [BFMT^{+}18, NKV14, LWXX19, YTH^{+}19]. Throughput [HB11, MMVL11, BLMB13, CLA^{+}18, DW12, GRR13, HVW16, HWLR14, HGX^{+}19, KSB11, LMSK18, LHX^{+}16, LNC13, SA11]. tickets [LMJC11]. tier [MS19, MZZC12, MCZ14, WQL14]. tile [LCJ^{+}18, ZLKK19]. tiled [JHF^{+}17, WQZ^{+}13]. Tilera [PCMM^{+}17]. Time [Ana14, CB15, HRG^{+}11, PLY15, ACCP12, BVGV14, BDGR13, Bog17, BW18, BKK^{+}11, CF19, CCK11, CRJ10a, CRJ10b, DLV11, DKRC^{+}15, FC14, GZG^{+}17, HV13, HZDP12, JZZ^{+}17, JHL^{+}18, KKR14, KSSL16, KKW17, LFS16, LR14, LST17, LLY15, Li16, LLB^{+}18, LML^{+}10, LWW19, MHLZ16, MLDG12, MAKWZ13, NVK^{+}11, SS11, SA19, SRB^{+}19, SZB16, TPS^{+}18, TYD^{+}19, XL11, ZZJ^{+}18, ZHH15, ZQMM11, ZHLQ12]. time-aware [MHLZ16]. Time-division-multiplexed [HRG^{+}11]. time-domain [SS11]. time-inhomogeneous [LWW19]. Time-optimized [Ana14]. time-scale [ACCP12]. time-space [BW18]. time-step-based [KKR14]. time-targeted [BKK^{+}11]. times [MLW^{+}19]. timetabling [PTN^{+}19]. timing [CSJ^{+}13, KKK^{+}11b]. Title [Ano18y, Ano18z, Ano18-27, Ano18-28]. TLA [SHL^{+}13]. TM [FKKR16, FWM^{+}10]. Toeplitz [GOH^{+}13, ABGV11, ADV14, Ter16]. Toeplitz-based [GOH^{+}13]. token [CRD12]. tokens [SGAC14]. tolerance [BJ15, CLMRL15, LFGM17, PAS15, SLZ^{+}19]. tolerant [AA14, AA16, ANEA13, AH11, ABBD14, BKMT14, DBCF13, DTK11a, FABG^{+}19, JBA15, JBS14, KG10, LDZ^{+}17, LFZ^{+}17, LAC18, MPG17b, MSEM^{+}19, NCB^{+}17, TCHC12, WW12, WYW15, XHZZ16]. tomography [BDRB14, GRS19, KSSL16, KDO^{+}13, XTN12]. tool [Ada17, ACD^{+}18, DXS^{+}19, uRIL^{+}18]. toolbox [EFG^{+}14]. tools [DMS^{+}16, FEH^{+}14, GAC^{+}17, RSVW19]. Top [SSKS11, IRRS16]. Top- [SSKS11]. Topic [YZC^{+}19, CLW^{+}19, dAAD^{+}19]. Topic-based [YZC^{+}19, dAAD^{+}19]. Topics [Ano16l, Kum17]. topologies [PD19]. Topology [TKKH17, WHS^{+}18, ZLKK19, AHA^{+}16, GL12, KBC^{+}10, LMP10, MBBD13, PMCC18, RCG18]. topology-aware [KBC^{+}10, MBBD13]. tori [LXLS12]. torus [ACA^{+}19, DM17, Lai15]. total [CB15]. Trace [JKIE13, CCAAS19, LC13]. traces [MTM10]. tracing [BM16, BM17b, CS17]. tracking [CJDC10, IIH^{+}17, KO11, NDP13, PLSM18, SRB^{+}19, TCS^{+}10, WXZ^{+}18]. trade [ECLV12, LCB16]. trade-off [ECLV12]. trade-offs [LCB16]. tradeoff [NLB^{+}18, RPS19]. Tradeoffs [MP15, CGKY12, PCMM^{+}17, SDS10, YZW^{+}15]. Trading [MPG17a, AKSZ19, ZLL14]. Traffic [ZCW19, BJ18, CRD12, LHLM14, MPG17a, OOSGVG^{+}16, YBM13, Zah12]. traffic-aware [LHLM14]. trails [PR12]. training [CDY^{+}19, FFYH19, ZHG^{+}19, ZLS17]. trajectory [WLYS19]. transaction [SI13]. Transactional [AM12b, CRB19, Gra10b, MP10, BGA12, CGM14, DT11, FWM^{+}10, GKK^{+}13, HGFF10, KR17, PB19, QGZP17, QGZP19, RDCQ17, SDS10]. transactions [CC16, FGG17, MLMSMG12, QGZP19, UBES10]. transfer [APK18, JKV15, LE19, WH17]. transferability [CSS11]. Transferring [SAR^{+}18]. transfers [GLGLBG12, LMGLGLG17, SCMH13]. Transform [Lla17, IIH16, CVK^{+}18a]. transformation [Tur12]. transformations [AM17]. Transformer [LLY15]. Transforming [LW16b]. transistor [FPM^{+}14]. transistors [LC14a]. transition [SP13]. transition-aware [SP13]. translation [NCB^{+}17]. transmissible [CWCW18, YHWY18b]. transmission [BDRB14, CHCG18, CPA^{+}11, DJDK19, WHC^{+}18, YA11]. Transparent [LMY^{+}11, GRZ^{+}18, LLY15]. Transparently [KLJ^{+}11]. transport [NPGV10, PKW^{+}10, WCL^{+}13]. transpose [BG16]. travel [KSSL16]. travel-time [KSSL16]. traveling [WMG13]. traversal [BBS13, CMN12, YFBY17]. traversals [HMR15]. treasure [MP15]. Tree [DR19, ASC^{+}18, AB13, BM14, BE13, BPBR11, BBD18, CG12, CRD17, DJ16, GHIJ19, HSS10, HMR15, HN19, KG10, KSK15, LY10, Li10, SV18, SLHS19, SJG19, TR16, WW12, Zah12, BBCQ13, GB11, SJG19]. trees [CFJW13, CFL^{+}19, DGNW13, ESGQ^{+}11, ESGQ^{+}14, GHY10, HAC17, JLY12, KKN13, LFZ^{+}17, MKW18, RDA18, WFZJ12, WIB12, YMLP14]. Trellis [SGdSS13]. Trends [ACB^{+}15, HAC^{+}19, KKKG14, MCS^{+}19, BHS13]. triangular [CASD18, dMS18]. tridiagonal [MRT18, PP13, SPH13, Ter16]. trimming [ZCS^{+}18]. triumph [Sch14]. Trojan [BK18]. Troodon [KAA^{+}19a]. Trust [TAM^{+}19, GTGLSA12, LZY11, LMXJ18, MLMSMG12]. trustworthy [MHLZ16]. TSGL [ACD^{+}18]. tsunami [NSKN17]. TT-XSS [WXZ^{+}18]. tumors [HES11]. tuned [PSB^{+}19]. Tuning [CSMML10, TdAR18, ABGV11, APK18, KKR14, MYD^{+}11, uRIL^{+}18]. Tunnel [ZBR11]. Tunnel-based [ZBR11]. Tuple [STKW12, LdPLC^{+}19]. turbulence [PLK^{+}18]. twig [LSZZ15]. twisted [LFZ^{+}17, WFZJ12, XHZZ16]. Two [BBH^{+}17, CMR19, DKU15, CB11, FSV17, JD12, MP10, SNCP12, WLL16, YYWZ19, dlAMCFN12]. Two-dimensional [CMR19]. two-layer [YYWZ19, dlAMCFN12]. Two-level [BBH^{+}17]. two-list [WLL16]. two-phase [SNCP12]. type [GA18, GNZ18, PTN^{+}19]. types [ASB18, MFT^{+}19, RJKL11].

ultra [BM16, FABG^{+}19]. ultra-low [FABG^{+}19]. ultra-scale [BM16]. ultrasound [BDRB14, GRS19]. unauthentic [MLMSMG12]. unbalanced [UFF19]. unbiased [BW18]. unchoking [ARD14]. uncoordinated [LDZ^{+}14]. undergraduate [AJG18, GAC^{+}17, Kum17]. Understanding [ECLV12, SFHS19, XS11, RÖE^{+}18]. underwater [LWW18, ZWW17]. undirected [STA12]. undo [LHCC19]. uneven [SMT15]. Unified [CP10a, Amm16, ABO^{+}17, IIH16, LAS^{+}19, LZN19, Tal19, XRB12]. Uniform [SMO^{+}18, BBFN14, KSG13, LHWJ19]. uniformity [BBB11]. Unifying [NSDZ18]. union [ST14]. unison [DPBNT12]. unit [ASC^{+}18, BHS13, JPD17, SIY14, SAJ13, XL11, ZMCP11]. Units [DDGK13, YJL16, ATDH13, BK13, CLA^{+}18, DP16, Eme13, GLGLBG12, YL12]. Universal [BBS13, LWXX19, ACH18]. universe [RFP^{+}19]. Unleashing [ARD14]. unmanned [SRB^{+}19]. unrelated [CG11]. unreliable [DDG^{+}17, KRS15]. unstructured [FZ14, LL19, LWCC15, SBRM19]. updatable [MLZY17]. update [BM11, KHK18, LL19]. updates [Kol19, YZG18, ZLT^{+}19]. updating [AEF11, JBA15]. Upper [LXLS12, NDP13]. uranium [YDZ^{+}18]. urban [PTN^{+}19]. URL [XRB12]. usage [IIH16, KBC19]. usages [CJYC19]. Use [NVK^{+}11, ACHY18, CJA^{+}19]. User [WCXL11, CFI^{+}18, LC11, LBT19, NGQM12]. users [RÖE^{+}18]. Using [BMLLC^{+}19, CASD18, DWX10, VRM10, XLW^{+}18, dOCS14, ASKO16, Ale19b, AZC13, AD12, BCMV15, BHLT14, BSH15, CL14, COV13, CJDC10, CRWX12, CLXX19, DMK19, DDG^{+}17, DJDK19, ES12, FTK14, GK19, HDCM11, HSH10, HWL18, HN19, HMY^{+}18, JTZZ11, JGMY17, KRKS11, KDO^{+}13, KKH17, KM17, KS18, KSJC17, KR12, KC17, LK15, LT10, LY10, LST^{+}13, LW19, LSWC14, LWW18, MHLZ16, MZC18, MRS^{+}14, Men18, MRT18, NV19, NMS^{+}18, OPR18, Ozt11, PKN10, PP13, QGZP19, RBN11, RB12, RPN19, SMO14, SBÇ12a, SHK19, SA19, ST12, SGAC14, SIY14, SDG17, SQQL19, SLKK13, SLHS19, SJG19, SMT15, Tam18, TP18, TRS^{+}12, TPLY18, TDP15, TKX^{+}13, VLGV^{+}18, WCF14, WZZ^{+}17, WDS^{+}18, WD18, WBRT13, XCC^{+}19, XLHT13, ZS13, ZBW^{+}17]. using/for [MZC18]. Utility [CRJ10b, VMMB10, VLL^{+}14]. utility-based [VMMB10]. Utilization [ZV12, UFF19]. Utilization-based [ZV12]. utilizing [PDP17].

V2G [TODQ18]. VA [PW17]. VA-DE [PW17]. VAF [PLSM18]. valid [BBCQ13]. validation [LST^{+}13]. Valuable [PW17]. Value [SLHS19, HZHS18, JXZ^{+}19, LÜ14, SHRM19, YS11, ZWQ^{+}16]. valued [Str12]. VANET [WZ13]. VANETs [ABF^{+}14, SWLZ17, YXX13]. variable [CHCG18]. variant [HGX^{+}19]. variation [HRF^{+}11, MEMEMH17, MSEM^{+}19, ZRN^{+}14]. variations [WCF14]. various [KIH15]. VAYU [RCG18]. VCube [dAAD^{+}19]. VCube-PS [dAAD^{+}19]. Vector [LST^{+}13, YFS^{+}15, ASES15, BV13, BGO19, CP10b, MBW16, PR13, Sol13, ZLMC14]. vector-core [Sol13]. vectorial [SSKS11]. vectorizable [Can18]. Vectorization [KCSS18]. vectorized [TP18]. vectors [BDG^{+}15]. VEF3 [CCAAS19]. vehicle [PTN^{+}19, Sch13]. Vehicles [CXQ^{+}18, AK18, AKSZ19, BJ18, SRB^{+}19, ZWW17, ZCW19]. verifiable [CXY14, XLC^{+}18, YZC^{+}19]. Verification [SHSH17, AM17, RPN19]. Verifying [YDTZ18]. Versatile [CGL^{+}14]. versatility [KGN11]. version [LHZ^{+}18]. Versioning [ZLH^{+}18]. versus [JL11]. Vertex [AK17, WFLJ16, XYZW14, XHZZ16]. Vertex-disjoint [WFLJ16]. vertex-pancyclicity [XHZZ16]. Vertical [LHNBB19]. very [MYM10, PDB13, YÖ11]. VForce [MLK12]. via [AM13, ASA18, AKBD10, AD10, BM17b, BYT19, CMR10, ECLV12, HVW16, HBF12, KNHH18, LÜ14, MTM10, MS15, PS14, YZS15, YYWZ19]. video [LY12, YAK15]. video-sharing [YAK15]. videos [NML^{+}19]. view [BBB11, Tal19]. viewshed [CSL15]. Viola [NHO^{+}13]. Virtual [NC13, ZLZ^{+}19, CL14, FMIF18, Fu10, HTB19, KS18, KNHH18, PVRS17, RT18, SA19, WSX^{+}19, YLZW18, ZG13, ZLCZ18, BBCQ13]. virtualization [DYL^{+}12, FLCB10]. virtualized [AAA^{+}10, CP17, KLJ^{+}11, KKLJ14, SJB12, SSVC10]. visible [RFP^{+}19]. vision [SRB^{+}19]. Visual [PLSM18, BLZ^{+}18, PLK^{+}18]. visualization [RV13]. visualizing [ACD^{+}18]. VLIW [NS12]. VLSI [CHX^{+}17, PD19]. VM [MA19]. VOD [SK11, YCH^{+}10]. VoIP [TCMB^{+}19]. volatile [CDR12, HZHS18, NKV14, ZPK^{+}14]. voltage [FABG^{+}19]. Volume [Ano10a, Ano10b, Ano11j, Ano11k, Ano12m, Ano12n, Ano14f, Ano14g, Ano15k, LWCC15, PST^{+}19]. volunteer [LKM12]. voting [AFD^{+}11, LWH^{+}19]. vs [MNR^{+}19]. VSI [PGKV18]. VSS [Pen11]. vulnerability [OTKT12].

Wait [FKKR16, ACH18]. Wait-free [FKKR16, ACH18]. wake [JLY12]. wake-up [JLY12]. walk [BBS13, RM11, SMP15]. walks [Li10]. warehouse [CJYC19]. warehousing [DTK11a]. warp [NHO^{+}13]. waste [JXZ^{+}19]. water [dlAMCFN12]. wave [BBS13, Kar19, KVNV17]. WaveCluster [YÖ11]. wavelength [CS10, TKKH17]. wavelength-based [TKKH17]. wavelet [IIH16]. way [APV18, Sch14]. WCET [LZLX11]. WDM [CS10, PR12]. Weak [RHH12, PMHM19]. weakly [YWW12]. weakly-connected [YWW12]. wearable [Udd19]. Weather [AFH^{+}19]. web [CCA18, CVK^{+}18b, FKR^{+}17, HBSASA19, HSS17, Suk18, ASKTZ13, UGG^{+}11]. web-based [CVK^{+}18b, Suk18]. web-portal [FKR^{+}17]. weight [JM14, WZZ^{+}17, WW18b, ZFT^{+}18]. weight-based [JM14]. Weighted [MD13, BFMT^{+}18, CDDL10, DM17, LWXX19, Sta17, SZB16]. WFR [FKKR16]. WFR-TM [FKKR16]. wide [Can18, JKV15]. wide-area [JKV15]. Wihidum [JKD^{+}15]. wildfire [DFST13]. Wimpy [LNC13]. window [BM11, MTL^{+}18b]. Wireless [BJS18, GPJA10, LDZ^{+}14, THGY15, Amm16, AHG12, AYB^{+}15, BM11, BWP^{+}11, BOY10, BC11, BJL18, CKC19, CCW14, CRWX12, CHCG18, DLLL11, DGBN14, DR19, DJH11, DKM10, EMC19, EM11, FCW11, FCML13, GHY10, GDL^{+}11, GYP13, GZY14b, GM14a, GL12, HZA^{+}15, HS12, HZDP12, JF12, JLY12, JBS14, JHPL13, JLWX11, KKK11a, KO11, KO12, KSK15, KHK18, KZ11, KK10, KKTZ13, KGN11, LZ11, LDZ^{+}17, LY10, LC11, LMJC11, LWLD12, LL12b, LÜ14, LZC11, LSWC14, LDS16, MAGL13, MPV12, MA11, MBMC19, NPGV10, NSA11, NM17, NGQM12, OWK14, PLY15, RM10, RM11, RLP14, REZN17, SCN12, SZMK13, SSZ10, SCLL10, TLY12, TM10, VRM10, WL11, WCXL11, WBRT13, WWA^{+}18, XHZ^{+}10, YpGyLlC13, YZX11, ZMG^{+}16, ZW11, ZBR11, ZLCJ12, ZSCX18, ZTGL17, dOBG^{+}15, LDP^{+}14]. within [BPBR11, ZGJ^{+}18]. without [FKKR16, RCG^{+}11, WW12, WPC19]. WLAN [HB11]. WLANs [GZY14a]. WMNs [LHX^{+}16]. Work [BKC^{+}15, DKKV15, KM17]. worker [BMT12]. workers [KRS15]. workflow [ALM^{+}16, FPF14, RCG^{+}11, WHW^{+}17, YYLC11, YWG15, ZVL15]. workflows [BKK^{+}11, KHN17, PVP18]. Workload [MA19, KyLPC17, LTG14, SSFP11, YJL16]. Workloads [FTK14, AM12b, LLY15, MLK^{+}16, WD13, ZLWZ18]. world [GS18, MAGL13, MVP17]. write [ACHY18, CG10, IR12, IRRS16, SLKK12]. write-once [ACHY18]. write-only [SLKK12]. writers [FKKR16]. writing [DBLB^{+}12]. WSAN [Wu11]. WSN [BCO^{+}12]. WSNs [LLDL15, NDP13, SMP17]. Wukong [MXSL12].

x86 [AG12, RHH12]. XDraw [CSL15]. Xeon [CHLL18, KVNV17, RPN19]. Xilinx [RBG17]. XML [DWYB10, LSZZ15]. XSS [WXZ^{+}18]. XT [YQTV12].

Yao [STMZ18]. year [Kum17]. yield [CAF^{+}11].

Zernike [TRS^{+}12, XLH18]. ZMesh [PMCC18]. Zynq [RBG17, BAT^{+}19, ZAAB17]. Zynq-based [BAT^{+}19].

References

Aci:2010:NCC

[AA10] Cigdem Inan Aci and Mehmet Fatih Akay. A new congestion control algorithm for improving the performance of a broadcast-based multiprocessor architecture. Journal of Parallel and Distributed Computing, 70(9):930–940, September 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ahmed:2014:GDF

[AA14] Akram Ben Ahmed and Abderazek Ben Abdallah. Graceful deadlock-free fault-tolerant routing algorithm for 3D Network-on-Chip architectures. Journal of Parallel and Distributed Computing, 74(4):2229–2240, April 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000045.

Ahmed:2016:AFT

[AA16] Akram Ben Ahmed and Abderazek Ben Abdallah. Adaptive fault-tolerant architecture and routing algorithm for reliable many-core 3D-NoC systems. Journal of Parallel and Distributed Computing, 93–94(??):30–43, July 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300053.

Almeida:2010:JAC

[AAA^{+}10] Jussara Almeida, Virgílio Almeida, Danilo Ardagna, Ítalo Cunha, Chiara Francalanci, and Marco Trubian. Joint admission control and resource allocation in virtualized servers. Journal of Parallel and Distributed Computing, 70(4):344–362, April 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Aartsen:2015:IFD

[AAA^{+}15] M. G. Aartsen, R. Abbasi, M. Ackermann, J. Adams, J. A. Aguilar, M. Ahlers, D. Altmann, C. Arguelles, J. Auffenberg, X. Bai, M. Baker, S. W. Barwick, V. Baum, R. Bay, J. J. Beatty, J. Becker Tjus, K.-H. Becker, S. BenZvi, P. Berghaus, D. Berley, E. Bernardini, A. Bernhard, D. Z. Besson, G. Binder, D. Bindig, M. Bissok, E. Blaufuss, J. Blumenthal, D. J. Boersma, C. Bohm, D. Bose, S. Böser, O. Botner, L. Brayeur, H.-P. Bretz, A. M. Brown, R. Bruijn, J. Casey, and M. Casier. The IceProd framework: Distributed data processing for the IceCube neutrino observatory. Journal of Parallel and Distributed Computing, 75(??):198–211, January 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001439.

Aspnes:2010:CSC

[AAC10] James Aspnes, Hagit Attiya, and Keren Censor. Combining shared-coin algorithms. Journal of Parallel and Distributed Computing, 70(3):317–322, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Al-Azzoni:2010:DSH

[AAD10] Issam Al-Azzoni and Douglas G. Down. Dynamic scheduling for heterogeneous Desktop Grids. Journal of Parallel and Distributed Computing, 70(12):1231–1240, December 2010. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Abdullah:2017:REH

[AAH17] Aref M. Abdullah, Hesham A. Ali, and Amira Y. Haikal. Reliable and efficient hierarchical organization model for computational grid. Journal of Parallel and Distributed Computing, 104(??):191–205, June 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300461.

Arevalo:2015:FDH

[AAI^{+}15] Sergio Arévalo, Antonio Fernández Anta, Damien Imbs, Ernesto Jiménez, and Michel Raynal. Failure detectors in homonymous distributed systems (with an application to consensus). Journal of Parallel and Distributed Computing, 83(??):83–95, September 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515000970.

Ajwani:2013:GST

[AAK^{+}13] Deepak Ajwani, Shoukat Ali, Kostas Katrinis, Cheng-Hong Li, Alfred J. Park, John P. Morrison, and Eugen Schenfeld. Generating synthetic task graphs for simulating stream computing systems. Journal of Parallel and Distributed Computing, 73(10):1362–1374, October 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001172.

Augusto:2013:APG

[AB13] Douglas A. Augusto and Helio J. C. Barbosa. Accelerated parallel genetic programming tree evaluation with OpenCL. Journal of Parallel and Distributed Computing, 73(1):86–100, January 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151200024X.

Ateniese:2014:FTO

[ABBD14] Giuseppe Ateniese, Roberto Baldoni, Silvia Bonomi, and Giuseppe Antonio Di Luna. Fault-tolerant oblivious assignment with m slots in synchronous systems. Journal of Parallel and Distributed Computing, 74(7):2648–2661, July 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514000367.

Ayaida:2014:JRL

[ABF^{+}14] Marwane Ayaida, Mohtadi Barhoumi, Hacène Fouchal, Yacine Ghamri-Doudane, and Lissan Afilal. Joint routing and location-based service in VANETs. Journal of Parallel and Distributed Computing, 74(2):2077–2087, February 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002153.

Alonso:2011:ITP

[ABGV11] Pedro Alonso, Miguel O. Bernabéu, Victor M. García, and Antonio M. Vidal. Implementation and tuning of a parallel symmetric Toeplitz eigensolver. Journal of Parallel and Distributed Computing, 71(3):485–494, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Avin:2017:DCC

[ABLP17] Chen Avin, Michael Borokhovich, Zvi Lotker, and David Peleg. Distributed computing on core-periphery networks: Axiom-based design. Journal of Parallel and Distributed Computing, 99(??):51–67, January 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516301046.

At:2017:LAU

[ABO^{+}17] Nuray At, Jean-Luc Beuchat, Eiji Okamoto, Ismail San, and Teppei Yamazaki. A low-area unified hardware architecture for the AES and the cryptographic hash function Grøstl. Journal of Parallel and Distributed Computing, 106(??):106–120, August 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517300485.

Anjo:2016:DML

[AC16] Ivo Anjo and João Cachopo. Design of a method-level speculation framework for boosting irregular JVM applications. Journal of Parallel and Distributed Computing, 87(??):13–25, January 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515001720.

Andujar:2019:EET

[ACA^{+}19] Francisco J. Andújar, Salvador Coll, Marina Alonso, Juan-Miguel Martínez, Pedro López, José L. Sánchez, Francisco J. Alfaro, and Raúl Martínez. Energy efficient torus networks with on/off links. Journal of Parallel and Distributed Computing, 130(??):37–49, August 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519302266.

Assuncao:2015:BDC

[ACB^{+}15] Marcos D. Assunção, Rodrigo N. Calheiros, Silvia Bianchi, Marco A. S. Netto, and Rajkumar Buyya. Big Data computing and clouds: Trends and future directions. Journal of Parallel and Distributed Computing, 79–80(??):3–15, May 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001452.

Ardagna:2012:DTS

[ACCP12] Danilo Ardagna, Sara Casolari, Michele Colajanni, and Barbara Panicucci. Dual time-scale distributed capacity allocation and load redirect algorithms for cloud systems. Journal of Parallel and Distributed Computing, 72(6):796–808, June 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000585.

Adams:2018:TTV

[ACD^{+}18] Joel C. Adams, Patrick A. Crain, Christopher P. Dilley, Christiaan D. Hazlett, Elizabeth R. Koning, Serita M. Nelesen, Javin B. Unger, and Mark B. Vande Stel. TSGL: A tool for visualizing multithreaded behavior. Journal of Parallel and Distributed Computing, 118 (part 1)(??):233–246, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0743731518301035.

Attiya:2018:NUH

[ACH18] Hagit Attiya, Armando Castañeda, and Danny Hendler. Nontrivial and universal helping for wait-free queues and stacks. Journal of Parallel and Distributed Computing, 121(??):ii, November 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518304258.

Aspnes:2018:CUW

[ACHY18] James Aspnes, Keren Censor-Hillel, and Eitan Yaakobi. Concurrent use of write-once memory. Journal of Parallel and Distributed Computing, 113(??):250–260, March 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303325.

Andreolini:2015:ASR

[ACPT15] Mauro Andreolini, Michele Colajanni, Marcello Pietri, and Stefania Tosi. Adaptive, scalable and reliable monitoring of big data on clouds. Journal of Parallel and Distributed Computing, 79–80(??):67–79, May 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151400149X.

Ammari:2010:FCG

[AD10] Habib M. Ammari and Sajal K. Das. Forwarding via checkpoints: Geographic routing on always-on sensors. Journal of Parallel and Distributed Computing, 70(7):719–731, July 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Apopei:2012:APL

[AD12] B. Apopei and T. J. Dodd. Automatic parallelisation for LTI MIMO state space systems using FPGAs. An optimisation for cost & performance. Journal of Parallel and Distributed Computing, 72(8):990–1007, August 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001050.

Adams:2017:PTT

[Ada17] Joel C. Adams. Patternlets — a teaching tool for introducing students to parallel design patterns. Journal of Parallel and Distributed Computing, 105(??):31–41, July 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151730014X.

Altisen:2017:CSS

[ADD17] Karine Altisen, Stéphane Devismes, and Anaïs Durand. Concurrency in snap-stabilizing local resource allocation. Journal of Parallel and Distributed Computing, 102(??):42–56, April 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301484.

Atashpendar:2018:SPC

[ADDB18] Arash Atashpendar, Bernabé Dorronsoro, Grégoire Danoy, and Pascal Bouvry. A scalable parallel cooperative coevolutionary PSO algorithm for multi-objective optimization. Journal of Parallel and Distributed Computing, 112 (part 2)(??):111–125, February 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301752.

Altisen:2019:GS

[ADDP19] Karine Altisen, Stéphane Devismes, Anaïs Durand, and Franck Petit. Gradual stabilization. Journal of Parallel and Distributed Computing, 123(??):26–45, January 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306464.

Alonso:2014:BPI

[ADV14] Pedro Alonso, Manuel F. Dolz, and Antonio M. Vidal. Block pivoting implementation of a symmetric Toeplitz solver. Journal of Parallel and Distributed Computing, 74(5):2392–2399, May 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514000318.

Attiya:2011:CUS

[AEF11] Hagit Attiya, Faith Ellen, and Panagiota Fatourou. The complexity of updating snapshot objects. Journal of Parallel and Distributed Computing, 71(12):1570–1577, December 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511001596.

Ali:2011:CML

[AES11] Shoukat Ali, Behdis Eslamnour, and Zehra Shah. A case for on-machine load balancing. Journal of Parallel and Distributed Computing, 71(4):556–564, April 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Aspnes:2012:LCD

[AEY12] James Aspnes, David Eisenstat, and Yitong Yin. Lowcontention data structures. Journal of Parallel and Distributed Computing, 72(5):705–715, May 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512000202.

Attiya:2017:PLA

[AF17] Hagit Attiya and Arie Fouren. Poly-logarithmic adaptive algorithms require revealing primitives. Journal of Parallel and Distributed Computing, 109(??):102–116, November 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517301612.

Abellan:2013:DEC

[AFA13] José L. Abellán, Juan Fernández, and Manuel E. Acacio. Design of an efficient communication infrastructure for highly contended locks in many-core CMPs. Journal of Parallel and Distributed Computing, 73(7):972–985, July 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001499.

Araujo:2011:MIS

[AFD^{+}11] Filipe Araujo, Jorge Farinha, Patricio Domingues, Gheorghe Cosmin Silaghi, and Derrick Kondo. A maximum independent set approach for collusion detection in voting pools. Journal of Parallel and Distributed Computing, 71(10):1356–1366, October 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001316.

Andrade:2019:LSP

[AFG^{+}19] Guilherme Andrade, André Fernandes, Jeremias M. Gomes, Renato Ferreira, and George Teodoro. Large-scale parallel similarity search with product quantization for online multimedia services. Journal of Parallel and Distributed Computing, 125(??):81–92, March 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518308566.

Adams:2019:LMC

[AFH^{+}19] S. V. Adams, R. W. Ford, M. Hambley, J. M. Hobson, I. Kavčič, C. M. Maynard, T. Melvin, E. H. Müller, S. Mullerworth, A. R. Porter, M. Rezny, B. J. Shipway, and R. Wong. LFRic: Meeting the challenges of scalability and performance portability in weather and climate models. Journal of Parallel and Distributed Computing, 132(??):383–396, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518305306.

Arbenz:2014:BSA

[AFK14] Peter Arbenz, Cyril Flaig, and Daniel Kellenberger. Bone structure analysis on multiple GPGPUs. Journal of Parallel and Distributed Computing, 74(10):2941–2950, October 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001166.

Adamek:2017:EOS

[AFNT17] Jordan Adamek, Giovanni Farina, Mikhail Nesterenko, and Sébastien Tixeuil. Evaluating and optimizing stabilizing dining philosophers. Journal of Parallel and Distributed Computing, 109(??):63–74, November 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301545.

Aroca:2012:TGD

[AG12] Rafael Vidal Aroca and Luiz Marcos Garcia Gonçalves. Towards green data centers: a comparison of x86 and ARM architectures power efficiency. Journal of Parallel and Distributed Computing, 72(12):1770–1780, December 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002122.

Akrida:2016:ENR

[AGMS16] Eleni C. Akrida, Leszek Gasieniec, George B. Mertzios, and Paul G. Spirakis. Ephemeral networks with random availability of links: the case of fast networks. Journal of Parallel and Distributed Computing, 87(??):109–120, January 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515001872.

Andrade:2011:PHD

[AGWY11] H. Andrade, B. Gedik, K.-L. Wu, and P. S. Yu. Processing high data rate streams in System S. Journal of Parallel and Distributed Computing, 71(2):145–156, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anagnostopoulos:2011:DTD

[AH11] Christos Anagnostopoulos and Stathes Hadjiefthymiades. Delay-tolerant delivery of quality information in ad hoc networks. Journal of Parallel and Distributed Computing, 71 (7):974–987, July 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373151000239X.

Anagnostopoulos:2012:OQA

[AH12] Christos Anagnostopoulos and Stathes Hadjiefthymiades. Optimal, quality-aware scheduling of data consumption in mobile ad hoc networks. Journal of Parallel and Distributed Computing, 72(10):1269–1279, October 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001293.

Ajwani:2016:COA

[AHA^{+}16] Deepak Ajwani, Adam Hackett, Shoukat Ali, John P. Morrison, and Stephen Kirkland. Co-optimizing application partitioning and network topology for a reconfigurable interconnect. Journal of Parallel and Distributed Computing, 96(??):12–26, October 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300168.

Anagnostopoulos:2012:PPC

[AHG12] Christos Anagnostopoulos, Stathes Hadjiefthymiades, and Panagiotis Georgas. PC3: Principal Component-based Context Compression: Improving energy efficiency in wireless sensor networks. Journal of Parallel and Distributed Computing, 72(2):155–170, February 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001900.

Anagnostopoulos:2011:AMM

[AHZ11] Christos Anagnostopoulos, Stathes Hadjiefthymiades, and Evangelos Zervas. An analytical model for multi-epidemic information dissemination. Journal of Parallel and Distributed Computing, 71(1):87–104, January 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ayguade:2018:ATB

[AJG18] Eduard Ayguadé and Daniel Jiménez-González. An approach to task-based parallel programming for undergraduate students. Journal of Parallel and Distributed Computing, 118 (part 1)(??):140–156, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S074373151830100X.

Alkamper:2017:DNV

[AK17] Martin Alkämper and Robert Klöfkorn. Distributed newest vertex bisection. Journal of Parallel and Distributed Computing, 104(??):1–11, June 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516301782.

Aujla:2018:SBE

[AK18] Gagangeet Singh Aujla and Neeraj Kumar. SDN-based energy management scheme for sustainability of data centers: An analysis on renewable energy sources and electric vehicles participation. Journal of Parallel and Distributed Computing, 117(??):228–245, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302149.

Anastasiu:2019:PCN

[AK19] David C. Anastasiu and George Karypis. Parallel cosine nearest neighbor graph construction. Journal of Parallel and Distributed Computing, 129(??):61–82, July 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303283.

Akcan:2010:MCM

[AKBD10] Hüseyin Akcan, Vassil Kriakov, Hervé Brönnimann, and Alex Delis. Managing cohort movement of mobile sensors via GPSfree and compass-free node localization. Journal of Parallel and Distributed Computing, 70(7):743–757, July 2010. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Abu-Khzam:2015:SPR

[AKDMN15] Faisal N. Abu-Khzam, Khuzaima Daudjee, Amer E. Mouawad, and Naomi Nishimura. On scalable parallel recursive backtracking. Journal of Parallel and Distributed Computing, 84 (??):65–75, October 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731515001240.

Alabdulatif:2019:PPA

[AKK^{+}19] Abdulatif Alabdulatif, Ibrahim Khalil, Heshan Kumarage, Albert Y. Zomaya, and Xun Yi. Privacy-preserving anomaly detection in the cloud for quality assured decisionmaking in smart cities. Journal of Parallel and Distributed Computing, 127(??):209–223, May 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517303428.

Aujla:2019:ETD

[AKSZ19] Gagangeet Singh Aujla, Neeraj Kumar, Mukesh Singh, and Albert Y. Zomaya. Energy trading with dynamic pricing for electric vehicles in a smart city environment. Journal of Parallel and Distributed Computing, 127(??):169–183, May 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518304416.

Alelaiwi:2019:EMC

[Ale19a] Abdulhameed Alelaiwi. An efficient method of computation offloading in an edge cloud platform. Journal of Parallel and Distributed Computing, 127(??):58–64, May 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300140.

Alelaiwi:2019:EDI

[Ale19b] Abdulhameed Alelaiwi. Evaluating distributed IoT databases for edge/cloud platforms using the analytic hierarchy process. Journal of Parallel and Distributed Computing, 124(??):41–46, February 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518307603.

Amory:2011:NTS

[ALLM11] Alexandre M. Amory, Cristiano Lazzari, Marcelo S. Lubaszewski, and Fernando G. Moraes. A new test scheduling algorithm based on Networks-on-Chip as Test Access Mechanisms. Journal of Parallel and Distributed Computing, 71(5):675–686, May 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731510001954.

Ahmad:2016:HGA

[ALM^{+}16] Saima Gulzar Ahmad, Chee Sun Liew, Ehsan Ullah Munir, Tan Fong Ang, and Samee U. Khan. A hybrid genetic algorithm for optimization of scheduling workflow applications in heterogeneous computing systems. Journal of Parallel and Distributed Computing, 87(??):80–90, January 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515001860.

Ahmad:2013:MCO

[ALTV13] Faraz Ahmad, Seyong Lee, Mithuna Thottethodi, and T. N. Vijaykumar. MapReduce with communication overlap (MaRCO). Journal of Parallel and Distributed Computing, 73(5):608–620, May 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002936.

Asaduzzaman:2011:DMB

[AM11] Shah Asaduzzaman and Muthucumaru Maheswaran. Decentralized management of bi-modal network resources in a distributed stream processing platform. Journal of Parallel and Distributed Computing, 71(6):774–787, June 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000177.

Angeli:2012:CEC

[AM12a] Daniele Angeli and Enrico Masala. A cost-effective cloud computing framework for accelerating multimedia communication simulations. Journal of Parallel and Distributed Computing, 72(10):1373–1385, October 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151200144X.

Attiya:2012:TSR

[AM12b] Hagit Attiya and Alessia Milani. Transactional scheduling for read-dominated workloads. Journal of Parallel and Distributed Computing, 72(10):1386–1396, October 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151200130X.

Abed:2013:IPC

[AM13] Khalid H. Abed and Gerald R. Morris. Improving performance of codes with large/irregular stride memory access patterns via high performance reconfigurable computers. Journal of Parallel and Distributed Computing, 73(11):1430–1438, November 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001724.

Amato:2017:MTM

[AM17] Flora Amato and Francesco Moscato. Model transformations of MapReduce design patterns for automatic development and verification. Journal of Parallel and Distributed Computing, 110(??):52–59, December 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301939.

Ammari:2016:UFC

[Amm16] Habib M. Ammari. A unified framework for k-coverage and data collection in heterogeneous wireless sensor networks. Journal of Parallel and Distributed Computing, 89(??):37–49, March 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515001768.

Archibald:2018:RPB

[AMM^{+}18] Blair Archibald, Patrick Maier, Ciaran McCreesh, Robert Stewart, and Phil Trinder. Replicable parallel branch and bound search. Journal of Parallel and Distributed Computing, 113(??):92–114, March 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302861.

Araujo:2019:PEF

[AMSÅ19] Victor Araujo, Karan Mitra, Saguna Saguna, and Christer Åhlund. Performance evaluation of FIWARE: a cloud-based IoT platform for smart cities. Journal of Parallel and Distributed Computing, 132(??):250–261, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300164.

Anta:2013:ESP

[AMT13] Antonio Fernández Anta, Miguel A. Mosteiro, and Christopher Thraves. An early-stopping protocol for computing aggregate functions in sensor networks. Journal of Parallel and Distributed Computing, 73(2):111–121, February 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002286.

Alam:2019:ETC

[AMU^{+}19] Md. Golam Rabiul Alam, Md. Shirajum Munir, Md. Zia Uddin, Mohammed Shamsul Alam, Tri Nguyen Dang, and Choong Seon Hong. Edge-of-things computing framework for cost-effective provisioning of healthcare data. Journal of Parallel and Distributed Computing, 123(??):54–60, January 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518306282.

Anagnostopoulos:2014:TOC

[Ana14] Christos Anagnostopoulos. Time-optimized contextual information forwarding in mobile sensor networks. Journal of Parallel and Distributed Computing, 74(5):2317–2332, May 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514000185.

Al-Naqi:2013:DFT

[ANEA13] Asmaa Al-Naqi, Ahmet T. Erdogan, and Tughrul Arslan. Dynamic fault-tolerant three-dimensional cellular genetic algorithms. Journal of Parallel and Distributed Computing, 73(2):122–136, February 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002262.

Anonymous:2010:EVA

[Ano10a] Anonymous. End of volume author index. Journal of Parallel and Distributed Computing, 70(12):1266, December 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2010:EVR

[Ano10b] Anonymous. End of volume reviewer index. Journal of Parallel and Distributed Computing, 70(12):1267–1271, December 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2011:EBa

[Ano11a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 71(4):??, April 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2011:EBb

[Ano11b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 71(5):??, May 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511000402.

Anonymous:2011:EBc

[Ano11c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 71(6):??, June 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511000608.

Anonymous:2011:EBd

[Ano11d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 71(7):??, July 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511000906.

Anonymous:2011:EBe

[Ano11e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 71(8):??, August 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001079.

Anonymous:2011:EBf

[Ano11f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 71(9):??, September 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001201.

Anonymous:2011:EBg

[Ano11g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 71(10):??, October 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001511.

Anonymous:2011:EBh

[Ano11h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 71(11):??, November 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001651.

Anonymous:2011:EBi

[Ano11i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 71(12):??, December 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151100178X.

Anonymous:2011:EVA

[Ano11j] Anonymous. End of volume author index. Journal of Parallel and Distributed Computing, 71(12):1622–1627, December 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511001870.

Anonymous:2011:EVR

[Ano11k] Anonymous. End of volume reviewer index. Journal of Parallel and Distributed Computing, 71(12):1628–1634, December 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511001845.

Anonymous:2012:EBa

[Ano12a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(1):??, January 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001973.

Anonymous:2012:EBb

[Ano12b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(2):??, February 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511002371.

Anonymous:2012:EBc

[Ano12c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(3):??, March 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151200007X.

Anonymous:2012:EBd

[Ano12d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(4):??, April 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512000457.

Anonymous:2012:EBe

[Ano12e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(5):??, May 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512000743.

Anonymous:2012:EBf

[Ano12f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(6):??, June 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512000883.

Anonymous:2012:EBg

[Ano12g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(7):??, July 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001116.

Anonymous:2012:EBh

[Ano12h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(8):??, August 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001335.

Anonymous:2012:EBi

[Ano12i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(9):??, September 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001542.

Anonymous:2012:EBj

[Ano12j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(10):??, October 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001785.

Anonymous:2012:EBk

[Ano12k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(11):??, November 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001931.

Anonymous:2012:EBl

[Ano12l] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 72(12):??, December 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002389.

Anonymous:2012:EVA

[Ano12m] Anonymous. End of volume author index. Journal of Parallel and Distributed Computing, 72(12):1782–1788, December 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512002444.

Anonymous:2012:EVR

[Ano12n] Anonymous. End of volume reviewer index. Journal of Parallel and Distributed Computing, 72(12):1789–1796, December 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512002456.

Anonymous:2013:EBa

[Ano13a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 73(1):??, January 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002584.

Anonymous:2013:EBb

[Ano13b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 73(2):??, February 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002730.

Anonymous:2013:EBc

[Ano13c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 73(3):??, March 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002973.

Anonymous:2013:EBd

[Ano13d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 73(4):??, April 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151300021X.

Anonymous:2013:EBe

[Ano13e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 73(5):??, May 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000373.

Anonymous:2013:EBf

[Ano13f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 73(6):??, June 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000725.

Anonymous:2013:EBg

[Ano13g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 73(7):??, July 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151300083X.

Anonymous:2013:EBh

[Ano13h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 73(8):??, August 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513001020.

Anonymous:2013:EBi

[Ano13i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 73(9):??, September 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001202.

Anonymous:2013:EBj

[Ano13j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 73(10):??, October 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513001445.

Anonymous:2013:EBk

[Ano13k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 73(11):??, November 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001810.

Anonymous:2014:EBa

[Ano14a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 74(8):ifc, August 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514000963.

Anonymous:2014:EBb

[Ano14b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 74(9):ifc, September 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001191.

Anonymous:2014:EBc

[Ano14c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 74(10):ifc, October 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001361.

Anonymous:2014:EBd

[Ano14d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 74(11):ifc, November 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001555.

Anonymous:2014:EBe

[Ano14e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 74(12):ifc, December 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001774.

Anonymous:2014:EVA

[Ano14f] Anonymous. End of volume author index. Journal of Parallel and Distributed Computing, 74(12):3254–3259, December 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001828.

Anonymous:2014:EVR

[Ano14g] Anonymous. End of volume reviewer index. Journal of Parallel and Distributed Computing, 74(12):3260–3271, December 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151400183X.

Anonymous:2015:EBa

[Ano15a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 75(??):ifc, January 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514002123.

Anonymous:2015:EBb

[Ano15b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 76(??):ifc, February 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515000076.

Anonymous:2015:EBc

[Ano15c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 77(??):ifc, March 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515000283.

Anonymous:2015:EBd

[Ano15d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 78(??):ifc, April 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515000532.

Anonymous:2015:EBe

[Ano15e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 79–80(??):ifc, May 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515000702.

Anonymous:2015:EBf

[Ano15f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 81–82(??):ifc, July 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515000854.

Anonymous:2015:EBg

[Ano15g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 83(??):ifc, September 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001094.

Anonymous:2015:EBh

[Ano15h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 84(??):ifc, October 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001318.

Anonymous:2015:EBi

[Ano15i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 85(??):ifc, November 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001549.

Anonymous:2015:EBj

[Ano15j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 86(??):ifc, December 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001781.

Anonymous:2015:EVR

[Ano15k] Anonymous. End of volume reviewer index. Journal of Parallel and Distributed Computing, 86(??):CXIII–CXXI, December 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515001835.

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 87(??):ifc, January 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001963.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 88(??):ifc, February 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515002142.

Anonymous:2016:EBc

[Ano16c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 89(??):ifc, March 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516000071.

Anonymous:2016:EBd

[Ano16d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 90–91(??):ifc, April 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516000289.

Anonymous:2016:EBe

[Ano16e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 92(??):ifc, May 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151630020X.

Anonymous:2016:EBf

[Ano16f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 93–94(??):ifc, July 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300430.

Anonymous:2016:EBg

[Ano16g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 95(??):ifc, September 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300557.

Anonymous:2016:EBh

[Ano16h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 96(??):ifc, October 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300806.

Anonymous:2016:EBi

[Ano16i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 97(??):ifc, November 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151630096X.

Anonymous:2016:EBj

[Ano16j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 98(??):ifc, December 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301083.

Anonymous:2016:NAE

[Ano16k] Anonymous. New Associate Editors. Journal of Parallel and Distributed Computing, 87(??):iv–xi, January 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515002051.

Anonymous:2016:TI

[Ano16l] Anonymous. Topics of interest. Journal of Parallel and Distributed Computing, 87(??):xii, January 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515002063.

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 99(??):ifc, January 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516301319.

Anonymous:2017:EBb

[Ano17b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 100(??):ifc, February 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301642.

Anonymous:2017:EBc

[Ano17c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 101(??):ifc, March 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151630199X.

Anonymous:2017:EBd

[Ano17d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 102(??):ifc, April 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300278.

Anonymous:2017:EBe

[Ano17e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 103(??):ifc, May 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300539.

Anonymous:2017:EBf

[Ano17f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 104(??):ifc, June 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300862.

Anonymous:2017:EBg

[Ano17g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 104(??):ifc, June 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300862.

Anonymous:2017:EBh

[Ano17h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 105(??):ifc, July 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301028.

Anonymous:2017:EBi

[Ano17i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 106(??):ifc, August 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301399.

Anonymous:2017:EBj

[Ano17j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 107(??):ifc, September 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301673.

Anonymous:2017:EBk

[Ano17k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 108(??):ifc, October 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151730179X.

Anonymous:2017:EBl

[Ano17l] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 109(??):ifc, November 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302228.

Anonymous:2017:EBm

[Ano17m] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 110(??):ifc, December 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302484.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 111(??):i, January 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302733.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 111(??):ifc, January 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302721.

Anonymous:2018:EBd

[Ano18c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 112 (part 1)(??):ifc, February 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303106.

Anonymous:2018:EBe

[Ano18d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 112 (part 2)(??):ifc, February 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303180.

Anonymous:2018:EBf

[Ano18e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 113(??):i, March 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300030.

Anonymous:2018:EBg

[Ano18f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 113(??):ii, March 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300042.

Anonymous:2018:EBh

[Ano18g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 114(??):i, April 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300121.

Anonymous:2018:EBi

[Ano18h] Anonymous. Editorial: Board. Journal of Parallel and Distributed Computing, 114(??):ii, April 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300133.

Anonymous:2018:EBj

[Ano18i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 115(??):i, May 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300789.

Anonymous:2018:EBk

[Ano18j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 115(??):ii, May 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300790.

Anonymous:2018:EBl

[Ano18k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 116(??):i, ???? 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518301655.

Anonymous:2018:EBm

[Ano18l] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 117(??):i, July 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518302521.

Anonymous:2018:EBn

[Ano18m] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 117(??):ii, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518302533.

Anonymous:2018:EBo

[Ano18n] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 118 (part 1)(??):i, August 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0743731518302892.

Anonymous:2018:EBp

[Ano18o] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 118 (part 1)(??):ii, August 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0743731518302909.

Anonymous:2018:EBq

[Ano18p] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):ii, August 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303113.

Anonymous:2018:EBr

[Ano18q] Anonymous. Editorial board. Journal of Parallel and Distributed Computing, 119(??):ii, September 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303654.

Anonymous:2018:EBs

[Ano18r] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 120(??):ii, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151830546X.

Anonymous:2018:EBt

[Ano18s] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 121(??):i, November 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306348.

Anonymous:2018:EBu

[Ano18t] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 121(??):1–90, November 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306336.

Anonymous:2018:EBv

[Ano18u] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 122(??):ii, December 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518307172.

Anonymous:2018:EBCa

[Ano18v] Anonymous. Editorial board (continued). Journal of Parallel and Distributed Computing, 112 (part 1)(??):i, February 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517303118.

Anonymous:2018:EBCb

[Ano18w] Anonymous. Editorial board (continued). Journal of Parallel and Distributed Computing, 112 (part 2)(??):i, February 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517303192.

Anonymous:2018:EBCc

[Ano18x] Anonymous. Editorial board (continued). Journal of Parallel and Distributed Computing, 116(??):ii, ???? 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518301667.

Anonymous:2018:FTPa

[Ano18y] Anonymous. Full title page. Journal of Parallel and Distributed Computing, 119(??):i, September 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303642.

Anonymous:2018:FTPb

[Ano18z] Anonymous. Full title page. Journal of Parallel and Distributed Computing, 122(??):i, December 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518307160.

Anonymous:2018:TPa

[Ano18-27] Anonymous. Title page. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):i, August 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303101.

Anonymous:2018:TPb

[Ano18-28] Anonymous. Title page. Journal of Parallel and Distributed Computing, 120(??):i, October 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518305458.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 123(??):i, January 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518308049.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 123(??):ii, January 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518308050.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 124(??):i, February 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308475.

Anonymous:2019:EBd

[Ano19d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 124(??):ii, February 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308487.

Anonymous:2019:EBe

[Ano19e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 125(??):i, March 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518309481.

Anonymous:2019:EBf

[Ano19f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 125(??):ii, March 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518309493.

Anonymous:2019:EBg

[Ano19g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 126(??):i, April 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519301145.

Anonymous:2019:EBh

[Ano19h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 126(??):ii, April 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519301157.

Anonymous:2019:EBi

[Ano19i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 127(??):i, May 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519301431.

Anonymous:2019:EBj

[Ano19j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 127(??):ii, May 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519301443.

Anonymous:2019:EBk

[Ano19k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 128(??):i, June 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519302552.

Anonymous:2019:EBl

[Ano19l] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 128(??):ii, June 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519302564.

Anonymous:2019:EBm

[Ano19m] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 129(??):i, July 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151930303X.

Anonymous:2019:EBn

[Ano19n] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 129(??):ii, July 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519303041.

Anonymous:2019:EBo

[Ano19o] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 130(??):i, August 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519303818.

Anonymous:2019:EBp

[Ano19p] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 130(??):ii, August 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151930382X.

Anonymous:2019:EBq

[Ano19q] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 131(??):i, September 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304204.

Anonymous:2019:EBr

[Ano19r] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 131(??):ii, September 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304216.

Anonymous:2019:EBs

[Ano19s] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 132(??):i, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151930468X.

Anonymous:2019:EBt

[Ano19t] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 132(??):ii, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304691.

Anonymous:2019:EBu

[Ano19u] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 133(??):i, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519305933.

Anonymous:2019:EBv

[Ano19v] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 133(??):ii, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519305945.

Anonymous:2019:EBw

[Ano19w] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 134(??):i, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519306483.

Anonymous:2019:EBx

[Ano19x] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 134(??):ii, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519306495.

Attiya:2016:CBI

[AP16] Hagit Attiya and Ami Paz. Counting-based impossibility proofs for set agreement and renaming. Journal of Parallel and Distributed Computing, 87(??):1–12, January 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001677.

Arslan:2018:BDT

[APK18] Engin Arslan, Bahadir A. Pehlivan, and Tevfik Kosar. Big data transfer optimization through adaptive parameter tuning. Journal of Parallel and Distributed Computing, 120(??):89–100, October 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518303289.

Abellan:2018:PBE

[APRA18] José L. Abellán, Eduardo Padierna, Alberto Ros, and Manuel E. Acacio. Photonic-based express coherence notifications for many-core CMPs. Journal of Parallel and Distributed Computing, 113(??):179–194, March 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303271.

Antonov:2018:CSH

[APV18] A. Antonov, N. Popova, and Vl. Voevodin. Computational science and HPC education for graduate students: Paving the way to exascale. Journal of Parallel and Distributed Computing, 118 (part 1)(??):157–165, August 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0743731518301011.

Aravind:2013:SSE

[Ara13] Alex A. Aravind. Simple, space-efficient, and fairness improved FCFS mutual exclusion algorithms. Journal of Parallel and Distributed Computing, 73(8):1029–1038, August 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000531.

Atlidakis:2014:EUP

[ARD14] V. Atlidakis, M. Roussopoulos, and A. Delis. Enhanced-Bit: Unleashing the potential of the unchoking policy in the BitTorrent protocol. Journal of Parallel and Distributed Computing, 74(1):1959–1970, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001639.

Azari:2018:DRA

[ARDQ18] Leila Azari, Amir Masoud Rahmani, Helder A. Daniel, and Nooruldeen Nasih Qader. A data replication algorithm for groups of files in data grids. Journal of Parallel and Distributed Computing, 113(??):115–126, March 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302848.

AlFaisal:2017:NPE

[ARI17] Faiz Al Faisal, M. M. Hafizur Rahman, and Yasushi Inoguchi. A new power efficient high performance interconnection network for many-core processors. Journal of Parallel and Distributed Computing, 101(??):92–102, March 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301629.

Agostini:2018:GAE

[ARP18] E. Agostini, D. Rossetti, and S. Potluri. GPUDirect Async: Exploring GPU synchronous communication techniques for InfiniBand clusters. Journal of Parallel and Distributed Computing, 114(??):28–45, April 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303386.

Aupy:2014:CAF

[ARVZ14] Guillaume Aupy, Yves Robert, Frédéric Vivien, and Dounia Zaidouni. Checkpointing algorithms and fault prediction. Journal of Parallel and Distributed Computing, 74(2):2048–2064, February 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513002219.

Aluru:2013:SIJ

[AS13] Srinivas Aluru and Yogesh Simmhan. A special issue of Journal of Parallel and Distributed Computing: Scalable systems for big data management and analytics. Journal of Parallel and Distributed Computing, 73(6):896, June 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000658.

Aluru:2015:ESS

[AS15] Srinivas Aluru and Yogesh Simmhan. Editorial: Scalable systems for big data management and analytics. Journal of Parallel and Distributed Computing, 79–80(??):1–2, May 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515000738.

Azimi:2018:DGB

[AS18] Rasool Azimi and Hedieh Sajedi. A decentralized gossip based approach for data clustering in peer-to-peer networks. Journal of Parallel and Distributed Computing, 119(??):64–80, September 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518302168.

Ahmad:2019:MSN

[AS19a] Ishfaq Ahmad and Hafiz Fahad Sheikh. A multi-staged niched evolutionary approach for allocating parallel tasks with joint optimization of performance, energy, and temperature. Journal of Parallel and Distributed Computing, 134(??):65–74, December 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518303332.

Awasthi:2019:SBC

[AS19b] Sateesh Kumar Awasthi and Yatindra Nath Singh. Simplified Biased Contribution Index (SBCI): a mechanism to make P2P network fair and efficient for resource sharing. Journal of Parallel and Distributed Computing, 124(??):106–118, February 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518307366.

Acer:2018:ONB

[ASA18] Seher Acer, Oguz Selvitopi, and Cevdet Aykanat. Optimizing nonzero-based sparse matrix partitioning models via reducing latency. Journal of Parallel and Distributed Computing, 122(??):145–158, December 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518305860.

Almeida:2018:DSR

[ASB18] Paulo Sérgio Almeida, Ali Shoker, and Carlos Baquero. Delta state replicated data types. Journal of Parallel and Distributed Computing, 111(??):162–173, January 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302332.

Aparicio:2018:PAC

[ASC^{+}18] G. Aparicio, J. M. G. Salmerón, L. G. Casado, R. Asenjo, and E. M. T. Hendrix. Parallel algorithms for computing the smallest binary tree size in unit simplex refinement. Journal of Parallel and Distributed Computing, 112 (part 2)(??):166–178, February 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517301739.

Ashari:2015:MDB

[ASES15] Arash Ashari, Naser Sedaghati, John Eisenlohr, and P. Sadayappan. A model-driven blocking strategy for load balanced sparse matrix-vector multiplication on GPUs. Journal of Parallel and Distributed Computing, 76(??):3–15, February 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514002081.

Al-Sayed:2019:TEC

[ASHO19] Mustafa M. Al-Sayed, Hesham A. Hassan, and Fatma A. Omara. Towards evaluation of cloud ontologies. Journal of Parallel and Distributed Computing, 126(??):82–106, April 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518306105.

Al-Sayed:2016:PMM

[ASKO16] Mustafa M. Al-Sayed, Sherif Khattab, and Fatma A. Omara. Prediction mechanisms for monitoring state of cloud resources using Markov chain model. Journal of Parallel and Distributed Computing, 96(??):163–171, October 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300181.

Al-Shammary:2013:FSS

[ASKTZ13] Dhiah Al-Shammary, Ibrahim Khalil, Zahir Tari, and Albert Y. Zomaya. Fractal self-similarity measurements based clustering technique for SOAP Web messages. Journal of Parallel and Distributed Computing, 73(5):664–676, May 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000063.

Asyabi:2019:COS

[ASSS19] Esmail Asyabi, Erfan Sharafzadeh, SeyedAlireza SanaeeKohroudi, and Mohsen Sharifi. CTS: an operating system CPU scheduler to mitigate tail latency for latency-sensitive multithreaded applications. Journal of Parallel and Distributed Computing, 133(??):232–243, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518302387.

Afek:2012:ISJ

[AST12] Yehuda Afek, Nir Shavit, and Moran Tzafrir. Interrupting snapshots and the Java size method. Journal of Parallel and Distributed Computing, 72(7):880–888, July 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151200072X.

Anzt:2013:BAR

[ATDH13] Hartwig Anzt, Stanimire Tomov, Jack Dongarra, and Vincent Heuveline. A block-asynchronous relaxation method for graphics processing units. Journal of Parallel and Distributed Computing, 73(12):1613–1626, December 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001147.

Arslan:2019:SOA

[ATKT19] Sanem Arslan, Haluk Rahmi Topcuoglu, Mahmut Taylan Kandemir, and Oguz Tosun. Scheduling opportunities for asymmetrically reliable caches. Journal of Parallel and Distributed Computing, 126(??):134–151, April 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300176.

AlBadawi:2018:ASS

[AVAH18] Ahmad Al Badawi, Bharadwaj Veeravalli, Khin Mi Mi Aung, and Brahim Hamadicharef. Accelerating subset sum and lattice based public-key cryptosystems with multicore CPUs and GPUs. Journal of Parallel and Distributed Computing, 119(??):179–190, September 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518302831.

Ashouri:2015:PPB

[AYB^{+}15] Majid Ashouri, Hamed Yousefi, Javad Basiri, Ali Mohammad Afshin Hemmatyar, and Ali Movaghar. PDC: Predictionbased data-aware clustering in wireless sensor networks. Journal of Parallel and Distributed Computing, 81–82(??):24–35, July 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515000428.

Amoretti:2013:EAC

[AZC13] Michele Amoretti, Francesco Zanichelli, and Gianni Conte. Efficient autonomic cloud computing using online discrete event simulation. Journal of Parallel and Distributed Computing, 73(6):767–776, June 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000336.

Al-Zoubi:2013:RGS

[AZW13] Khaldoon Al-Zoubi and Gabriel Wainer. RISE: a general simulation interoperability middleware container. Journal of Parallel and Distributed Computing, 73(5):580–594, May 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000154.

Boschmann:2019:ZBA

[BAT^{+}19] Alexander Boschmann, Andreas Agne, Georg Thombansen, Linus Witschen, Florian Kraus, and Marco Platzner. Zynqbased acceleration of robust high density myoelectric signal processing. Journal of Parallel and Distributed Computing, 123(??):77–89, January 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518304787.

Busnel:2011:UPS

[BBB11] Yann Busnel, Roberto Beraldi, and Roberto Baldoni. On the uniformity of peer sampling based on view shuffling. Journal of Parallel and Distributed Computing, 71(8):1165–1176, August 2011. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511000219.

Barnat:2012:DFL

[BBBC12] Jirí Barnat, Petr Bauch, Lubos Brim, and Milan Ceska. Designing fast LTL model checking algorithms for manycore GPUs. Journal of Parallel and Distributed Computing, 72(9):1083–1097, September 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511002140.

Baldoni:2013:VTR

[BBCQ13] Roberto Baldoni, Silvia Bonomi, Adriano Cerocchi, and Leonardo Querzoni. Virtual Tree: a robust architecture for interval valid queries in dynamic distributed systems. Journal of Parallel and Distributed Computing, 73(8):1135–1145, August 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513000610.

Blin:2018:SSM

[BBD18] Lélia Blin, Fadwa Boubekeur, and Swan Dubois. A selfstabilizing memory efficient algorithm for the minimum diameter spanning tree under an omnipotent daemon. Journal of Parallel and Distributed Computing, 117(??):50–62, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518300844.

Berenbrink:2012:BBR

[BBFN12] Petra Berenbrink, André Brinkmann, Tom Friedetzky, and Lars Nagel. Balls into bins with related random choices. Journal of Parallel and Distributed Computing, 72(2):246–253, February 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151100195X.

Berenbrink:2014:BNU

[BBFN14] Petra Berenbrink, André Brinkmann, Tom Friedetzky, and Lars Nagel. Balls into non-uniform bins. Journal of Parallel and Distributed Computing, 74(2):2065–2076, February 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513002190.

Bender:2017:TLM

[BBH^{+}17] Michael A. Bender, Jonathan W. Berry, Simon D. Hammond, K. Scott Hemmert, Samuel McCauley, Branden Moore, Benjamin Moseley, Cynthia A. Phillips, David Resnick, and Arun Rodrigues. Two-level main memory co-design: Multi-threaded algorithmic primitives, analysis, and simulation. Journal of Parallel and Distributed Computing, 102(??):213–228, April 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151630185X.

Bernaschi:2013:BCT

[BBR13] M. Bernaschi, M. Bisson, and D. Rossetti. Benchmarking of communication techniques for GPUs. Journal of Parallel and Distributed Computing, 73(2):250–255, February 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002213.

Bernard:2013:UAS

[BBS13] Thibault Bernard, Alain Bui, and Devan Sohier. Universal adaptive self-stabilizing traversal scheme: Random walk and reloading wave. Journal of Parallel and Distributed Computing, 73(2):137–149, February 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002560.

Borgonovo:2011:REW

[BC11] Flaminio Borgonovo and Matteo Cesana. Reuse efficiency of wireless access networks under physical carrier sense: a Markovian analysis. Journal of Parallel and Distributed Computing, 71(9):1179–1188, September 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731510002637.

Benoit:2019:SIS

[BÇ19] Anne Benoit and Ümit V. Çatalyürek. Special issue: Selected papers from IPDPS’18. Journal of Parallel and Distributed Computing, 133(??):174–175, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151930526X.

Benoit:2018:CSF

[BCC^{+}18] Anne Benoit, Aurélien Cavelan, Franck Cappello, Padma Raghavan, Yves Robert, and Hongyang Sun. Coping with silent and fail-stop errors at scale by combining replication and checkpointing. Journal of Parallel and Distributed Computing, 122(??):209–225, December 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305835.

Barrett:2015:ARM

[BCD^{+}15] R. F. Barrett, P. S. Crozier, D. W. Doerfler, M. A. Heroux, P. T. Lin, H. K. Thornquist, T. G. Trucano, and C. T. Vaughan. Assessing the role of mini-applications in predicting key performance characteristics of scientific and engineering applications. Journal of Parallel and Distributed Computing, 75(??):107–122, January 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001695.

Baumann:2014:FDG

[BCF14] Hervé Baumann, Pierluigi Crescenzi, and Pierre Fraigniaud. Flooding in dynamic graphs with arbitrary degree sequence.

Journal of Parallel and Distributed Computing, 74(5):2433– 2437, May 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514000173.

Bhowmick:2015:NAB

[BCH15] Sanjukta Bhowmick, Tzu-Yi Chen, and Mahantesh Halappanavar. A new augmentation based algorithm for extracting maximal chordal subgraphs. Journal of Parallel and Distributed Computing, 76(??):132–144, February 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151400197X.

Bressan:2013:EEP

[BCK^{+}13] Stéphane Bressan, Alfredo Cuzzocrea, Panagiotis Karras, Xuesong Lu, and Sadegh Heyrani Nobari. An effective and efficient parallel approach for random graph generation over GPUs. Journal of Parallel and Distributed Computing, 73(3):303–316, March 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002250.

Bernaschi:2015:SSC

[BCMV15] Massimo Bernaschi, Giancarlo Carbone, Enrico Mastrostefano, and Flavio Vella. Solutions to the st-connectivity problem using a GPU-based distributed BFS. Journal of Parallel and Distributed Computing, 76(??):145–153, February 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001865.

Bagaa:2012:EDA

[BCO^{+}12] Miloud Bagaa, Yacine Challal, Abdelraouf Ouadjaout, Noureddine Lasla, and Nadjib Badache. Efficient data aggregation with in-network integrity control for WSN. Journal of Parallel and Distributed Computing, 72(10):1157–1170, October 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001451.

Berlinska:2011:SDM

[BD11] J. Berlińska and M. Drozdowski. Scheduling divisible MapReduce computations. Journal of Parallel and Distributed Computing, 71(3):450–459, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ballard:2015:RHV

[BDG^{+}15] G. Ballard, J. Demmel, L. Grigori, M. Jacquelin, N. Knight, and H. D. Nguyen. Reconstructing Householder vectors from tall-skinny QR. Journal of Parallel and Distributed Computing, 85(??):3–31, November 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151500101X.

Benoit:2013:RPO

[BDGR13] Anne Benoit, Fanny Dufossé, Alain Girault, and Yves Robert. Reliability and performance optimization of pipelined real-time systems. Journal of Parallel and Distributed Computing, 73(6):851–865, June 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000348.

Beck:2019:DAN

[BDL^{+}19] Gaël Beck, Tarn Duong, Mustapha Lebbah, Hanane Azzag, and Christophe Cérin. A distributed approximate nearest neighbors algorithm for efficient large scale mean shift clustering. Journal of Parallel and Distributed Computing, 134(??):128–139, December 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518308724.

Barbareschi:2018:PBH

[BDM18] Mario Barbareschi, Alessandra De Benedictis, and Nicola Mazzocca. A PUF-based hardware mutual authentication protocol. Journal of Parallel and Distributed Computing, 119(??):107–120, September 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518302582.

Bonakdarpour:2016:SSC

[BDP16] Borzoo Bonakdarpour, Stéphane Devismes, and Franck Petit. Snap-stabilizing committee coordination. Journal of Parallel and Distributed Computing, 87(??):26–42, January 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515001719.

Birk:2014:GBI

[BDRB14] Matthias Birk, Robin Dapp, N. V. Ruiter, and J. Becker. GPU-based iterative transmission reconstruction in 3D ultrasound computer tomography. Journal of Parallel and Distributed Computing, 74(1):1730–1743, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002037.

Besa:2013:CRB

[BE13] Juan Besa and Yadran Eterovic. A concurrent red-black tree. Journal of Parallel and Distributed Computing, 73(4):434–449, April 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512002912.

Boukerche:2012:DBA

[BEN12] Azzedine Boukerche, Dionysios Efstathiou, and Sotiris Nikoletseas. Direction-based adaptive data propagation for heterogeneous sensor mobility. Journal of Parallel and Distributed Computing, 72(6):778–790, June 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000548.

Benoit:2015:ISP

[Ben15] Anne Benoit. IPDPS 2014 selected papers on numerical and combinatorial algorithms. Journal of Parallel and Distributed Computing, 85(??):1–2, November 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001525.

BenHafaiedh:2019:GFM

[Ben19] Imene Ben Hafaiedh. A generic formal model for the comparison and analysis of distributed job-scheduling algorithms in grid environment. Journal of Parallel and Distributed Computing, 132(??):331–343, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303703.

Bruno:2013:MMC

[BF13] Giulia Bruno and Alessandro Fiori. MicroClAn: Microarray clustering analysis. Journal of Parallel and Distributed Computing, 73(3):360–370, March 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002237.

Baccour:2017:PEG

[BFH^{+}17] E. Baccour, S. Foufou, R. Hamila, Z. Tari, and A. Y. Zomaya. PTNet: an efficient and green data center network. Journal of Parallel and Distributed Computing, 107(??):3–18, September 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300953.

Blazewicz:2013:MGB

[BFKW13] Jacek Blazewicz, Wojciech Frohmberg, Michal Kierzynka, and Pawel Wojciechowski. G-MSA — a GPU-based, fast and accurate algorithm for multiple sequence alignment. Journal of Parallel and Distributed Computing, 73(1):32–41, January 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512001001.

Borgdorff:2013:FDM

[BFL^{+}13] Joris Borgdorff, Jean-Luc Falcone, Eric Lorenz, Carles Bona-Casas, Bastien Chopard, and Alfons G. Hoekstra. Foundations of distributed multiscale computing: Formalization, specification, and analysis. Journal of Parallel and Distributed Computing, 73(4):465–483, April 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002924.

Berenbrink:2018:TLB

[BFMT^{+}18] Petra Berenbrink, Tom Friedetzky, Frederik Mallmann-Trenn, Sepehr Meshkinfamfard, and Chris Wastell. Threshold load balancing with weighted tasks. Journal of Parallel and Distributed Computing, 113(??):218–226, March 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302885.

Benito:2019:AAC

[BFVB19] M. Benito, P. Fuentes, E. Vallejo, and R. Beivide. ACOR: Adaptive congestion-oblivious routing in dragonfly networks. Journal of Parallel and Distributed Computing, 131(??):173– 188, September 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731519303508.

Bekesi:2016:MTM

[BG16] József Békési and Gábor Galambos. Matrix transpose on meshes with buses. Journal of Parallel and Distributed Computing, 96(??):194–201, October 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151630051X.

Baldassin:2012:TRS

[BGA12] Alexandro Baldassin, Felipe Goldstein, and Rodolfo Azevedo. A transactional runtime system for the Cell/BE architecture. Journal of Parallel and Distributed Computing, 72(12):1535–1546, December 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001888.

Bautista-Gomez:2016:CRP

[BGBC^{+}16] Leonardo Bautista-Gomez, Anne Benoit, Aurélien Cavelan, Saurabh K. Raina, Yves Robert, and Hongyang Sun. Coping with recall and precision of soft error detectors. Journal of Parallel and Distributed Computing, 98(??):8–24, December 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516300922.

Bienz:2019:NAS

[BGO19] Amanda Bienz, William D. Gropp, and Luke N. Olson. Node aware sparse matrix–vector multiplication. Journal of Parallel and Distributed Computing, 130(??):166–178, August 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519302321.

Bukhari:2017:OBP

[BHK17] I. F. Bukhari, A. Harwood, and S. Karunasekera. Optimum benefit protocol: a fast converging, bandwidth-efficient decentralized similarity overlay. Journal of Parallel and Distributed Computing, 109(??):129–141, November 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301648.

Bertin:2014:FSB

[BHLT14] Rémi Bertin, Sascha Hunold, Arnaud Legrand, and Corinne Touati. Fair scheduling of bag-of-tasks applications using distributed Lagrangian optimization. Journal of Parallel and Distributed Computing, 74(1):1914–1929, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001779.

Brodtkorb:2013:GPU

[BHS13] André R. Brodtkorb, Trond R. Hagen, and Martin L. Sætra. Graphics processing unit (GPU) programming strategies and trends in GPU computing. Journal of Parallel and Distributed Computing, 73(1):4–13, January 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000998.

Bistouni:2015:PNM

[BJ15] Fathollah Bistouni and Mohsen Jahanshahi. Pars network: a multistage interconnection network with faulttolerance capability. Journal of Parallel and Distributed Computing, 75(??):168–183, January 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001476.

Bui:2018:IAF

[BJ18] Khac-Hoai Nam Bui and Jason J. Jung. Internet of Agents framework for connected vehicles: A case study on distributed traffic control system. Journal of Parallel and Distributed Computing, 116(??):89–95, ???? 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302952.

Bradbury:2018:HOP

[BJL18] Matthew Bradbury, Arshad Jhumka, and Matthew Leeke. Hybrid online protocols for source location privacy in wireless sensor networks. Journal of Parallel and Distributed Computing, 115(??):67–81, May 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300236.

Bahrami:2018:NHA

[BJS18] Bahareh Bahrami, Mohammad Ali Jabraeil Jamali, and Shahram Saeidi. A novel hierarchical architecture for wireless network-on-chip. Journal of Parallel and Distributed Computing, 120(??):307–321, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518301102.

Basaran:2013:GEM

[BK13] Can Basaran and Kyoung-Don Kang. Grex: an efficient MapReduce framework for graphics processing units. Journal of Parallel and Distributed Computing, 73(4):522–533, April 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000051.

Boraten:2018:MHT

[BK18] Travis Boraten and Avinash Kodi. Mitigation of hardware Trojan based denial-of-service attack for secure NoCs. Journal of Parallel and Distributed Computing, 111(??):24–38, January 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302034.

Banerjee:2015:WEP

[BKC^{+}15] Dip Sankar Banerjee, Ashutosh Kumar, Meher Chaitanya, Shashank Sharma, and Kishore Kothapalli. Work efficient parallel algorithms for large graph exploration on emerging heterogeneous architectures. Journal of Parallel and Distributed Computing, 76(??):81–93, February 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151400224X.

Bhuiyan:2017:PAS

[BKCM17] Hasanuzzaman Bhuiyan, Maleq Khan, Jiangzhuo Chen, and Madhav Marathe. Parallel algorithms for switching edges in heterogeneous graphs. Journal of Parallel and Distributed Computing, 104(??):19–35, June 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516301800.

Byun:2011:BRC

[BKK^{+}11] Eun-Kyu Byun, Yang-Suk Kee, Jin-Soo Kim, Ewa Deelman, and Seungryoul Maeng. BTS: Resource capacity estimate for time-targeted science workflows. Journal of Parallel and Distributed Computing, 71(6):848–862, June 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000207.

Bouguerra:2014:FTS

[BKMT14] Mohamed Slim Bouguerra, Derrick Kondo, Fernando Mendonca, and Denis Trystram. Fault-tolerant scheduling on parallel systems with non-memoryless failure distributions. Journal of Parallel and Distributed Computing, 74(5):2411–2422, May 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151400015X.

Bonfietti:2013:MTM

[BLMB13] Alessio Bonfietti, Michele Lombardi, Michela Milano, and Luca Benini. Maximum-throughput mapping of SDFGs on multi-core SoC platforms. Journal of Parallel and Distributed Computing, 73(10):1337–1350, October 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001093.

Broll:2018:VPE

[BLZ^{+}18] Brian Broll, Ákos Lédeczi, Hamid Zare, Dung Nguyen Do, János Sallai, Péter Völgyesi, Miklós Maróti, Lesa Brown, and Chris Vanags. A visual programming environment for introducing distributed computing to secondary education. Journal of Parallel and Distributed Computing, 118 (part 1)(??):189– 200, August 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0743731518300996.

Badarla:2011:LTS

[BM11] Venkataramana Badarla and C. Siva Ram Murthy. Learning-TCP: a stochastic approach for efficient update in TCP congestion window in ad hoc wireless networks. Journal of Parallel and Distributed Computing, 71(6):863–878, June 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731510002819.

Baddar:2014:BSC

[BM14] Sherenaz W. Al-Haj Baddar and Basel A. Mahafzah. Bitonic sort on a chained-cubic tree interconnection network. Journal of Parallel and Distributed Computing, 74(1):1744–1761, January 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513002049.

Bahmani:2016:ECU

[BM16] Amir Bahmani and Frank Mueller. Efficient clustering for ultra-scale application tracing. Journal of Parallel and Distributed Computing, 98(??):25–39, December 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300934.

Baharvand:2017:AAA

[BM17a] Farshad Baharvand and S. Ghassem Miremadi. ANMR: Aging-aware adaptive n-modular redundancy for homogeneous multicore embedded processors. Journal of Parallel and Distributed Computing, 109(??):29–41, November 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301491.

Bahmani:2017:SCE

[BM17b] Amir Bahmani and Frank Mueller. Scalable communication event tracing via clustering. Journal of Parallel and Distributed Computing, 109(??):230–244, November 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301971.

Bustio-Martinez:2019:UHL

[BMLLC^{+}19] Lázaro Bustio-Martínez, Martín Letras-Luna, René Cumplido, Raudel Hernández-León, Claudia Feregrino-Uribe, and José M. Bande-Serrano. Using hashing and lexicographic order for Frequent Itemsets Mining on data streams. Journal of Parallel and Distributed Computing, 125(??):58–71, March 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308244.

Boschetti:2019:MOD

[BMS19] Marco Antonio Boschetti, Vittorio Maniezzo, and Francesco Strappaveccia. Membership overlay design optimization with resource constraints (accelerated on GPU). Journal of Parallel and Distributed Computing, 133(??):286–296, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518304908.

Bendjoudi:2012:AHM

[BMT12] A. Bendjoudi, N. Melab, and E-G. Talbi. An adaptive hierarchical master-worker (AHMW) framework for grids — application to B&B algorithms. Journal of Parallel and Distributed Computing, 72(2):120–131, February 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001912.

Ben-Nun:2016:SBP

[BNBR16] Tal Ben-Nun, Amnon Barak, and Uri Raviv. Spline-based parallel nonlinear optimization of function sequences. Journal of Parallel and Distributed Computing, 93–94(??):132–145, July 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151630017X.

Bogaerts:2017:OST

[Bog17] Steven A. Bogaerts. One step at a time: Parallelism in an introductory programming course. Journal of Parallel and Distributed Computing, 105(??):4–17, July 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300023.

Bernstein:2019:LCM

[BOKS19] Ran Bernstein, Margarita Osadchy, Daniel Keren, and Assaf Schuster. LDA classifier monitoring in distributed streaming systems. Journal of Parallel and Distributed Computing, 123(??):156–167, January 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518306993.

Bertier:2013:AAC

[BOT13] Marin Bertier, Marko Obrovac, and Cédric Tedeschi. Adaptive atomic capture of multiple molecules. Journal of Parallel and Distributed Computing, 73(9):1251–1266, September 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000543.

Ben-Othman:2010:EEQ

[BOY10] Jalel Ben-Othman and Bashir Yahya. Energy efficient and QoS based routing protocol for wireless sensor networks. Journal of Parallel and Distributed Computing, 70(8):849–857, August 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Blin:2011:SSM

[BPBR11] Lélia Blin, Maria Gradinariu Potop-Butucaru, and Stephane Rovedakis. Self-stabilizing minimum degree spanning tree within one from the optimal degree. Journal of Parallel and Distributed Computing, 71(3):438–449, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Byna:2011:SID

[BS11] Surendra Byna and Xian-He Sun. Special issue on data intensive computing. Journal of Parallel and Distributed Computing, 71(2):143–144, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bukata:2015:SRC

[BSH15] Libor Bukata, Premysl Sucha, and Zdenek Hanzálek. Solving the resource constrained project scheduling problem using the parallel tabu search designed for the CUDA platform. Journal of Parallel and Distributed Computing, 77(??):58–68, March 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514002226.

Briceno:2013:RSR

[BSS^{+}13] Luis Diego Briceño, Jay Smith, Howard Jay Siegel, Anthony A. Maciejewski, Paul Maxwell, Russ Wakefield, Abdulla Al-Qawasmeh, Ron C. Chiang, and Jiayin Li. Robust static resource allocation of DAGs in a heterogeneous multicore system. Journal of Parallel and Distributed Computing, 73(12):1705–1717, December 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001627.

Berka:2013:PRT

[BV13] T. Berka and M. Vajtersic. Parallel rare term vector replacement: Fast and effective dimensionality reduction for text. Journal of Parallel and Distributed Computing, 73(3):341–351, March 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512002158.

Basanta-Val:2014:RMP

[BVGV14] Pablo Basanta-Val and Marisol García-Valls. Resource management policies for real-time Java remote invocations. Journal of Parallel and Distributed Computing, 74(1):1930–1944, January 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513001378.

Buayen:2018:PTS

[BW18] Patcharin Buayen and Jeeraporn Werapun. Parallel timespace reduction by unbiased filtering for solving the 0/1knapsack problem. Journal of Parallel and Distributed Computing, 122(??):195–208, December 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518305847.

Bein:2011:DNC

[BWP^{+}11] Doina Bein, Yicheng Wen, Shashi Phoha, Bharat B. Madan, and Asok Ray. Distributed network control for mobile multimodal wireless sensor networks. Journal of Parallel and Distributed Computing, 71(3):460–470, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bobda:2018:HLS

[BYG^{+}18] Christophe Bobda, Franck Yonga, Martin Gebser, Harold Ishebabi, and Torsten Schaub. High-level synthesis of on-chip multiprocessor architectures based on answer set programming. Journal of Parallel and Distributed Computing, 117(??):161– 179, July 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151830087X.

Bui:2017:EEC

[BYH^{+}17] Dinh-Mao Bui, YongIk Yoon, Eui-Nam Huh, SungIk Jun, and Sungyoung Lee. Energy efficiency for cloud computing system based on predictive optimization. Journal of Parallel and Distributed Computing, 102(??):103–114, April 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301708.

Bi:2019:DPS

[BYT19] Jing Bi, Haitao Yuan, and Wei Tan. Deadlock prevention for service orchestration via controlled Petri nets. Journal of Parallel and Distributed Computing, 124(??):92–105, February 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518306749.

Concatto:2011:IYN

[CAF^{+}11] Caroline Concatto, João Almeida, Guilherme Fachini, Marcos Hervé, Fernanda Kastensmidt, Érika Cota, and Marcelo Lubaszewski. Improving the yield of NoC-based systems through fault diagnosis and adaptive routing. Journal of Parallel and Distributed Computing, 71(5):664–674, May 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731510002248.

Chen:2013:TSM

[CAK13] Jingshu Chen, Fuad Abujarad, and Sandeep Kulkarni. Towards scalable model checking of self-stabilizing programs. Journal of Parallel and Distributed Computing, 73(4):400– 410, April 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512002900.

Cannizzo:2018:FVA

[Can18] Fabio Cannizzo. A fast and vectorizable alternative to binary search in O(1) with wide applicability to arrays of floating point numbers. Journal of Parallel and Distributed Computing, 113(??):37–54, March 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302836.

Chow:2018:UJI

[CASD18] Edmond Chow, Hartwig Anzt, Jennifer Scott, and Jack Dongarra. Using Jacobi iterations and blocking for solving sparse triangular systems in incomplete factorization preconditioning. Journal of Parallel and Distributed Computing, 119(??):219–230, September 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518303034.

Czapinski:2011:TST

[CB11] Michal Czapiński and Stuart Barnes. Tabu Search with two approaches to parallel flowshop evaluation on CUDA platform.

Journal of Parallel and Distributed Computing, 71(6):802–811, June 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511000384.

Cason:2015:THT

[CB15] Daniel Cason and Luiz E. Buzato. Time hybrid total order broadcast: Exploiting the inherent synchrony of broadcast networks. Journal of Parallel and Distributed Computing, 77(??):26–40, March 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514002056.

Chakrabarty:2014:RAS

[CC14] Amitabha Chakrabarty and Martin Collier. O(log m̄, log N) routing algorithm for 2 log N − 1-stage switching networks and beyond. Journal of Parallel and Distributed Computing, 74(10):3045–3055, October 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001063.

Carvalho:2016:OMT

[CC16] Fernando Miguel Carvalho and João Cachopo. Optimizing memory transactions for large-scale programs. Journal of Parallel and Distributed Computing, 89(??):13–24, March 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515002099.

Cano:2018:FJR

[CCA18] Lorela Cano, Giuliana Carello, and Danilo Ardagna. A framework for joint resource allocation of MapReduce and web service applications in a shared cloud cluster. Journal of Parallel and Distributed Computing, 120(??):127–147, October 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518303903.

Cano-Cano:2019:SEI

[CCAAS19] Javier Cano-Cano, Francisco J. Andújar, Francisco J. Alfaro, and José L. Sánchez. Speeding up exascale interconnection network simulations with the VEF3 trace framework. Journal of Parallel and Distributed Computing, 133(??):124–135, November 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519304782.

Cui:2019:OLP

[CCC^{+}19] Zhihua Cui, Yang Cao, Xingjuan Cai, Jianghui Cai, and Jinjun Chen. Optimal LEACH protocol with modified bat algorithm for big data sensing systems in Internet of Things. Journal of Parallel and Distributed Computing, 132(??):217–229, October 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517303453.

Cesar:2017:ICT

[CCE^{+}17] Eduardo Cesar, Ana Cortés, Antonio Espinosa, Tomàs Margalef, Juan Carlos Moure, Anna Sikora, and Remo Suppi. Introducing computational thinking, parallel programming and performance engineering in interdisciplinary studies. Journal of Parallel and Distributed Computing, 105(??):116–126, July 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300059.

Chen:2011:ART

[CCK11] Hua Chen, Albert Mo Kim Cheng, and Ying-Wei Kuo. Assigning real-time tasks to heterogeneous processors by applying ant colony optimization. Journal of Parallel and Distributed Computing, 71(1):132–142, January 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chang:2014:AAL

[CCW14] Chao-Tsun Chang, Chih-Yung Chang, and Tzu-Lin Wang. Accident aware localization mechanism for wireless sensor networks. Journal of Parallel and Distributed Computing, 74(9):2831–2844, September 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151400094X.

Czumaj:2019:CB

[CD19] Artur Czumaj and Peter Davies. Communicating with beeps. Journal of Parallel and Distributed Computing, 130(??):98– 109, August 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519302369.

Cruz:2014:DTM

[CDAN14] Eduardo H. M. Cruz, Matthias Diener, Marco A. Z. Alves, and Philippe O. A. Navaux. Dynamic thread mapping of shared memory applications by exploiting cache coherence protocols. Journal of Parallel and Distributed Computing, 74(3):2215– 2228, March 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513002335.

Carrier:2015:SSA

[CDD^{+}15] Fabienne Carrier, Ajoy K. Datta, Stéphane Devismes, Lawrence L. Larmore, and Yvan Rivierre. Self-stabilizing (f, g)-alliances with safe convergence. Journal of Parallel and Distributed Computing, 81–82(??):11–23, July 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515000398.

Casella:2019:EMC

[CDD^{+}19] A. Casella, I. De Falco, A. Della Cioppa, U. Scafuri, and E. Tarantino. Exploiting multi-core and GPU hardware to speed up the registration of range images by means of differential evolution. Journal of Parallel and Distributed Computing, 133(??):307–318, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304738.

Caron:2010:SSC

[CDDL10] Eddy Caron, Ajoy K. Datta, Benjamin Depardon, and Lawrence L. Larmore. A self-stabilizing k-clustering algorithm for weighted graphs. Journal of Parallel and Distributed Computing, 70(11):1159–1173, November 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chockler:2011:SIC

[CDJL11] Gregory Chockler, Eliezer Dekel, Joseph JaJa, and Jimmy Lin. Special issue on cloud computing. Journal of Parallel and Distributed Computing, 71(6):731, June 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000578.

Cuomo:2018:RDR

[CDPS18] Salvatore Cuomo, Pasquale De Michele, Francesco Piccialli, and Arun Kumar Sangaiah. Reproducing dynamics related to an Internet of Things framework: A numerical and statistical approach. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):359–368, August 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302095.

Cordasco:2012:SDV

[CDR12] Gennaro Cordasco, Rosario De Chiara, and Arnold L. Rosenberg. On scheduling DAGs for volatile computing platforms: Area-maximizing schedules. Journal of Parallel and Distributed Computing, 72(10):1347–1360, October 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001463.

Casanova:2010:CRA

[CDS10] Henri Casanova, Frédéric Desprez, and Frédéric Suter. On cluster resource allocation for multiple parallel task graphs. Journal of Parallel and Distributed Computing, 70(12):1193– 1203, December 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cui:2019:MCD

[CDW^{+}19] Zhihua Cui, Lei Du, Penghong Wang, Xingjuan Cai, and Wensheng Zhang. Malicious code detection based on CNNs and multi-objective algorithm. Journal of Parallel and Distributed Computing, 129(??):50–58, July 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308529.

Cong:2019:FNN

[CDY^{+}19] Guojing Cong, Giacomo Domeniconi, Chih-Chieh Yang, Joshua Shapiro, Fan Zhou, and Barry Chen. Fast neural network training on a cluster of GPUs for action recognition with high accuracy. Journal of Parallel and Distributed Computing, 134(??):153–165, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300061.

Celesti:2019:FRT

[CF19] Antonio Celesti and Maria Fazio. A framework for real time end to end monitoring and big data oriented management of smart environments. Journal of Parallel and Distributed Computing, 132(??):262–273, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308256.

Cassavia:2018:DCL

[CFI^{+}18] Nunziato Cassavia, Sergio Flesca, Michele Ianni, Elio Masciari, and Chiara Pulice. Distributed computing by leveraging and rewarding idling user resources from P2P networks. Journal of Parallel and Distributed Computing, 122(??):81–94, December 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151830515X.

Cheng:2013:DAT

[CFJW13] Baolei Cheng, Jianxi Fan, Xiaohua Jia, and Jin Wang. Dimension-adjacent trees and parallel construction of independent spanning trees on crossed cubes. Journal of Parallel and Distributed Computing, 73(5):641–652, May 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000105.

Cheng:2019:CNI

[CFL^{+}19] Baolei Cheng, Jianxi Fan, Cheng-Kuan Lin, Xiaohua Jia, and Xiaoyan Li. Constructing node-independent spanning trees on the line graph of the hypercube by an independent forest scheme. Journal of Parallel and Distributed Computing, 134(??):104–115, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306129.

Clauss:2010:ICO

[CG10] Pierre-Nicolas Clauss and Jens Gustedt. Iterative computations with ordered read-write locks. Journal of Parallel and Distributed Computing, 70(5):496–504, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Carroll:2011:DAM

[CG11] Thomas E. Carroll and Daniel Grosu. Distributed algorithmic mechanism design for scheduling on unrelated machines. Journal of Parallel and Distributed Computing, 71(3):397–406, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Carroll:2012:IBD

[CG12] T. E. Carroll and D. Grosu. An incentive-based distributed mechanism for scheduling divisible loads in tree networks. Journal of Parallel and Distributed Computing, 72(3):389–401, March 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511002243.

Casu:2017:PPA

[CG17] Mario R. Casu and Paolo Giaccone. Power-performance assessment of different DVFS control policies in NoCs. Journal of Parallel and Distributed Computing, 109(??):193–207, November 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517301934.

Chen:2016:FHA

[CGC16] Tao Chen, Xiaofeng Gao, and Guihai Chen. The features, hardware, and architectures of data center networks: a survey. Journal of Parallel and Distributed Computing, 96(??):45–74, October 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516300399.

Cumplido:2019:GES

[CGFH19] Rene Cumplido, Maya Gokhale, Claudia Feregrino, and Michael Huebner. Guest editorial: Special issue on reconfigurable computing and FPGA technology. Journal of Parallel and Distributed Computing, 133(??):359–361, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519305465.

Chow:2012:PTS

[CGKY12] Alix L. H. Chow, Leana Golubchik, Samir Khuller, and Yuan Yao. Performance tradeoffs in structured peer to peer streaming. Journal of Parallel and Distributed Computing, 72(3):323– 337, March 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511002462.

Casanova:2014:VSA

[CGL^{+}14] Henri Casanova, Arnaud Giersch, Arnaud Legrand, Martin Quinson, and Frédéric Suter. Versatile, scalable, and accurate simulation of distributed applications and platforms. Journal of Parallel and Distributed Computing, 74(10):2899–2917, October 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001105.

Castro:2014:ATM

[CGM14] Márcio Castro, Luís Fabrício W. Góes, and Jean-François Méhaut. Adaptive thread mapping strategies for transactional memory applications. Journal of Parallel and Distributed Computing, 74(9):2845–2859, September 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001026.

Cecilia:2013:EDP

[CGN^{+}13] José M. Cecilia, José M. García, Andy Nisbet, Martyn Amos, and Manuel Ujaldón. Enhancing data parallelism for ant colony optimization on GPUs. Journal of Parallel and Distributed Computing, 73(1):42–51, January 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000032.

Chen:2018:PAC

[CHCG18] Da-Ren Chen, Chiun-Chieh Hsu, Mu-Yen Chen, and Chun-Fu Guo. A power-aware 2-covered path routing for wireless body area networks with variable transmission ranges. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):379–397, August 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517302368.

Cheng:2018:MCN

[CHLL18] Xuntao Cheng, Bingsheng He, Mian Lu, and Chiew Tong Lau. Many-core needs fine-grained scheduling: a case study of query processing on Intel Xeon Phi processors. Journal of Parallel and Distributed Computing, 120(??):395–404, October 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517302599.

Chen:2017:CFB

[CHX^{+}17] Xiaodao Chen, Xiaohui Huang, Yang Xiang, Dongmei Zhang, Rajiv Ranjan, and Chen Liao. A CPS framework based perturbation constrained buffer planning approach in VLSI design. Journal of Parallel and Distributed Computing, 103(??): 3–10, May 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301733.

Coetzee:2017:GBC

[CJ17] P. Coetzee and S. A. Jarvis. Goal-based composition of scalable hybrid analytics for heterogeneous architectures. Journal of Parallel and Distributed Computing, 108(??):59–73, October 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301666.

Chuah:2019:TCD

[CJA^{+}19] Edward Chuah, Arshad Jhumka, Samantha Alt, Daniel Balouek-Thomert, James C. Browne, and Manish Parashar. Towards comprehensive dependability-driven resource use and message log-analysis for HPC systems diagnosis. Journal of Parallel and Distributed Computing, 132(??):95–112, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519304137.

Cao:2010:CTT

[CJDC10] Donglei Cao, Beihong Jin, Sajal K. Das, and Jiannong Cao. On collaborative tracking of a target group using binary proximity sensors. Journal of Parallel and Distributed Computing, 70(8): 825–838, August 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chang:2019:CBD

[CJYC19] Chih-Hung Chang, Fuu-Cheng Jiang, Chao-Tung Yang, and Sheng-Cang Chou. On construction of a big data warehouse accessing platform for campus power usages. Journal of Parallel and Distributed Computing, 133(??):40–50, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518307081.

Chang:2013:HBC

[CK13] Yeim-Kuan Chang and Fang-Chen Kuo. Hint-based cache design for reducing miss penalty in HBS packet classification algorithm. Journal of Parallel and Distributed Computing, 73(8):1170–1182, August 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151300049X.

Chae:2019:RAN

[CKC19] Sung-Yoon Chae, Kyungran Kang, and Young-Jong Cho. A randomized adaptive neighbor discovery for wireless networks with multi-packet reception capability. Journal of Parallel and Distributed Computing, 131(??):235–244, September 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518308578.

Choi:2013:DSE

[CKK^{+}13] Jiwon Choi, Myeongsu Kang, Yongmin Kim, Cheol-Hong Kim, and Jong-Myon Kim. Design space exploration in many-core processors for sound synthesis of plucked string instruments. Journal of Parallel and Distributed Computing, 73(11):1506–1522, November 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001748.

Cheng:2019:LBD

[CKLW19] Long Cheng, Spyros Kotoulas, Qingzhi Liu, and Ying Wang. Load-balancing distributed outer joins through operator decomposition. Journal of Parallel and Distributed Computing, 132(??):21–35, October 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518301989.

Chung:2012:QAD

[CKML12] WonSik Chung, Mun-Suk Kim, JeongHoon Mo, and SuKyoung Lee. QoS-aware dynamic MAP selection schemes in HMIPv6 networks. Journal of Parallel and Distributed Computing, 72(7):838–855, July 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512000834.

Costache:2017:MBA

[CKMP17] Stefania Costache, Samuel Kortas, Christine Morin, and Nikos Parlavantzas. Market-based autonomous resource and application management in private clouds. Journal of Parallel and Distributed Computing, 100(??):85–102, February 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301204.

Chechina:2011:RMA

[CKT11] Natalia Chechina, Peter King, and Phil Trinder. Redundant movements in autonomous mobility: Experimental and theoretical analysis. Journal of Parallel and Distributed Computing, 71(10):1278–1292, October 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001389.

Cheng:2017:IRP

[CKWT17] Long Cheng, Spyros Kotoulas, Tomas E. Ward, and Georgios Theodoropoulos. Improving the robustness and performance of parallel joins over distributed systems. Journal of Parallel and Distributed Computing, 109(??):310–323, November 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302058.

Canali:2014:DSV

[CL14] Claudia Canali and Riccardo Lancellotti. Detecting similarities in virtual machine behavior for cloud monitoring using smoothed histograms. Journal of Parallel and Distributed Computing, 74(8):2757–2769, August 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000343.

Cecilia:2018:HTA

[CLA^{+}18] José M. Cecilia, Antonio Llanes, José L. Abellán, Juan Gómez-Luna, Li-Wen Chang, and Wen-Mei W. Hwu. Highthroughput ant colony optimization on graphics processing units. Journal of Parallel and Distributed Computing, 113(??):261–274, March 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517303337.

Chen:2017:ENF

[CLC^{+}17] Yunliang Chen, Fangyuan Li, Jia Chen, Bo Du, Kim-Kwang Raymond Choo, and Houcine Hassan. EPLS: a novel feature extraction method for migration data clustering. Journal of Parallel and Distributed Computing, 103(??):96–103, May 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301654.

Chechina:2016:INS

[CLG^{+}16] Natalia Chechina, Huiqing Li, Amir Ghaffari, Simon Thompson, and Phil Trinder. Improving the network scalability of Erlang. Journal of Parallel and Distributed Computing, 90–91(??):22–34, April 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516000034.

Castro-Leon:2015:FTS

[CLMRL15] Marcela Castro-León, Hugo Meyer, Dolores Rexachs, and Emilio Luque. Fault tolerance at system level based on RADIC architecture. Journal of Parallel and Distributed Computing, 86(??):98–111, December 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001434.

Chen:2017:PAS

[CLOL17] Cen Chen, Kenli Li, Aijia Ouyang, and Keqin Li. A parallel approximate SS–ELM algorithm based on MapReduce for large-scale datasets. Journal of Parallel and Distributed Computing, 108(??):85–94, October 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300138.

Cao:2019:QAS

[CLW^{+}19] Buqing Cao, Jianxun Liu, Yiping Wen, Hongtao Li, Qiaoxiang Xiao, and Jinjun Chen. QoS-aware service recommendation based on relational topic model and factorization machines for IoT mashup applications. Journal of Parallel and Distributed Computing, 132(??):177–189, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518302375.

Chen:2019:ASS

[CLXX19] Jingying Chen, Yongqiang Lv, Ruyi Xu, and Can Xu. Automatic social signal analysis: Facial expression recognition using difference convolution neural network. Journal of Parallel and Distributed Computing, 131(??):97–102, September 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518308827.

Chang:2019:TIA

[CLZ19] Victor Chang, Taiyu Li, and Zhiyang Zeng. Towards an improved Adaboost algorithmic method for computational financial analysis. Journal of Parallel and Distributed Computing, 134(??):219–232, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518307068.

Cugola:2012:LLC

[CM12] Gianpaolo Cugola and Alessandro Margara. Low latency complex event processing on parallel hardware. Journal of Parallel and Distributed Computing, 72(2):205–218, February 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511002188.

Chronaki:2019:MPA

[CMC^{+}19] Kallia Chronaki, Miquel Moretó, Marc Casas, Alejandro Rico, Rosa M. Badia, Eduard Ayguadé, and Mateo Valero. On the maturity of parallel applications for asymmetric multi-core processors. Journal of Parallel and Distributed Computing, 127(??):105–115, May 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519300267.

Carchiolo:2010:AON

[CMMN10] Vincenza Carchiolo, Michele Malgeri, Giuseppe Mangioni, and Vincenzo Nicosia. An adaptive overlay network inspired by social behaviour. Journal of Parallel and Distributed Computing, 70(3):282–295, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chakroun:2013:CMC

[CMMT13] I. Chakroun, N. Melab, M. Mezmaz, and D. Tuyttens. Combining multi-core and GPU computing for solving combinatorial optimization problems. Journal of Parallel and Distributed Computing, 73(12):1563–1577, December 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001615.

Clouser:2012:CFT

[CMN12] Thomas Clouser, Mark Miyashita, and Mikhail Nesterenko. Concurrent face traversal for efficient geometric routing. Journal of Parallel and Distributed Computing, 72(5):627–636, May 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000214.

Cordasco:2010:EIS

[CMR10] Gennaro Cordasco, Grzegorz Malewicz, and Arnold L. Rosenberg. Extending IC-scheduling via the Sweep Algorithm. Journal of Parallel and Distributed Computing, 70(3):201–211, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cerrato:2018:EDE

[CMR^{+}18] Ivano Cerrato, Guido Marchetto, Fulvio Risso, Riccardo Sisto, Matteo Virgilio, and Roberto Bonafiglia. An efficient data exchange mechanism for chained network functions. Journal of Parallel and Distributed Computing, 114(??):1–15, April 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517303349.

Charlton:2019:TDB

[CMR19] John Charlton, Steve Maddock, and Paul Richmond. Twodimensional batch linear programming on the GPU. Journal of Parallel and Distributed Computing, 126(??):152–160, April 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151830399X.

Che:2014:ALM

[CN14] Hao Che and Minh Nguyen. Amdahl’s Law for multithreaded multicore processors. Journal of Parallel and Distributed Computing, 74(10):3056–3069, October 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001142.

Canto-Navarro:2018:FPA

[CNLGRL18] E. Cantó-Navarro, M. López-García, and R. Ramos-Lara. Floating-point accelerator for biometric recognition on FPGA embedded systems. Journal of Parallel and Distributed Computing, 112 (part 1)(??):20–34, February 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302642.

Cui:2017:NMO

[COF^{+}17] Laizhong Cui, Peng Ou, Xianghua Fu, Zhenkun Wen, and Nan Lu. A novel multi-objective evolutionary algorithm for recommendation systems. Journal of Parallel and Distributed Computing, 103(??):53–63, May 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301423.

Cano:2013:PMO

[COV13] Alberto Cano, Juan Luis Olmo, and Sebastián Ventura. Parallel multi-objective ant programming for classification using GPUs. Journal of Parallel and Distributed Computing, 73(6):713–728, June 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151300018X.

Caminiti:2010:UPE

[CP10a] Saverio Caminiti and Rossella Petreschi. Unified parallel encoding and decoding algorithms for Dandelion-like codes. Journal of Parallel and Distributed Computing, 70(11):1119– 1127, November 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:2010:PIE

[CP10b] Wenwu Chen and Bill Poirier. Parallel implementation of an efficient preconditioned linear solver for grid-based applications in chemical physics. III: Improved parallel scalability for sparse matrix-vector products. Journal of Parallel and Distributed Computing, 70(7):779–782, July 2010. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cinque:2017:IHF

[CP17] Marcello Cinque and Antonio Pecchia. On the injection of hardware faults in virtualized multicore systems. Journal of Parallel and Distributed Computing, 106(??):50–61, August 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300849.

Choudhury:2011:SMT

[CPA^{+}11] Ashish Choudhury, Arpita Patra, B. V. Ashwinkumar, Kannan Srinathan, and C. Pandu Rangan. Secure message transmission in asynchronous networks. Journal of Parallel and Distributed Computing, 71(8):1067–1074, August 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151100058X.

Chauhan:2019:BIC

[CPJ^{+}19] Sameer Singh Chauhan, Emmanuel S. Pilli, R. C. Joshi, Girdhari Singh, and M. C. Govil. Brokering in interconnected cloud computing environments: a survey. Journal of Parallel and Distributed Computing, 133(??):193–209, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518305719.

Chen:2018:CMD

[CPLY18] Wuhui Chen, Incheon Paik, Zhenni Li, and Neil Y. Yen. A cost minimization data allocation algorithm for dynamic datacenter resizing. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):280–295, August 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300989.

Castro:2019:HTM

[CRB19] Daniel Castro, Paolo Romano, and João Barreto. Hardware transactional memory meets memory persistency. Journal of Parallel and Distributed Computing, 130(??):63–79, August 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518303952.

Cuesta:2012:SBP

[CRD12] Blas Cuesta, Antonio Robles, and José Duato. Switchbased packing technique to reduce traffic and latency in token coherence. Journal of Parallel and Distributed Computing, 72(3):409–423, March 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511002279.

Cohen:2017:PCT

[CRD17] Jaime Cohen, Luiz A. Rodrigues, and Elias P. Duarte, Jr. Parallel cut tree algorithms. Journal of Parallel and Distributed Computing, 109(??):1–14, November 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301296.

Castillo:2011:OAA

[CRH11] C. Castillo, G. N. Rouskas, and K. Harfoush. Online algorithms for advance resource reservations. Journal of Parallel and Distributed Computing, 71(7):963–973, July 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000153.

Chan:2019:GAD

[CRHC19] David M. Chan, Roshan Rao, Forrest Huang, and John F. Canny. GPU accelerated t-distributed stochastic neighbor embedding. Journal of Parallel and Distributed Computing, 131(??):1–13, September 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151830875X.

Cho:2010:PBR

[CRJ10a] Hyeonjoong Cho, Binoy Ravindran, and E. Douglas Jensen. T-L plane-based real-time scheduling for homogeneous multiprocessors. Journal of Parallel and Distributed Computing, 70 (3):225–236, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cho:2010:UAR

[CRJ10b] Hyeonjoong Cho, Binoy Ravindran, and E. Douglas Jensen. Utility accrual real-time scheduling for multiprocessor embedded systems. Journal of Parallel and Distributed Computing, 70(2):101–110, February 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chowdhury:2013:OAM

[CRSB13] Rezaul Alam Chowdhury, Vijaya Ramachandran, Francesco Silvestri, and Brandon Blakeley. Oblivious algorithms for multicores and networks of processors. Journal of Parallel and Distributed Computing, 73(7):911–925, July 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000695.

Chen:2012:ILA

[CRWX12] Xiao Chen, Neil C. Rowe, Jie Wu, and Kaiqi Xiong. Improving the localization accuracy of targets by using their spatialtemporal relationships in wireless sensor networks. Journal of Parallel and Distributed Computing, 72(8):1008–1018, August 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512001074.

Chen:2010:RWA

[CS10] Yawen Chen and Hong Shen. Routing and wavelength assignment for hypercube in array-based WDM optical networks. Journal of Parallel and Distributed Computing, 70(1):59–68, January 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Collinson:2017:CAF

[CS17] S. Collinson and O. Sinnen. Caching architecture for flexible FPGA ray tracing platform. Journal of Parallel and Distributed Computing, 104(??):61–72, June 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300072.

Chandrasekaran:2013:CDT

[CSJ^{+}13] Sunita Chandrasekaran, Shilpa Shanbagh, Ramkumar Jayaraman, Douglas L. Maskell, and Hui Yan Cheah. C2FPGA — a dependency-timing graph design methodology. Journal of Parallel and Distributed Computing, 73(11):1417–1429, November 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151200216X.

Cauchi-Saunders:2015:GEX

[CSL15] Aran J. Cauchi-Saunders and Ian J. Lewis. GPU enabled XDraw viewshed analysis. Journal of Parallel and Distributed Computing, 84(??):87–93, October 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001197.

Caymes-Scutari:2010:SDM

[CSMML10] P. Caymes-Scutari, A. Morajko, T. Margalef, and E. Luque. Scalable dynamic monitoring, analysis and tuning environment for parallel applications. Journal of Parallel and Distributed Computing, 70(4):330–337, April 2010. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Carbunar:2011:CPT

[CSS11] Bogdan Carbunar, Weidong (Larry) Shi, and Radu Sion. Conditional e-payments with transferability. Journal of Parallel and Distributed Computing, 71(1):16–26, January 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cui:2017:NOC

[CSW^{+}17] Zhihua Cui, Bin Sun, Gaige Wang, Yu Xue, and Jinjun Chen. A novel oriented cuckoo search algorithm to improve DV-Hop performance for cyber-physical systems. Journal of Parallel and Distributed Computing, 103(??):42–52, May 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301393.

Chen:2010:SFC

[CTC^{+}10] Dan Chen, Stephen J. Turner, Wentong Cai, Georgios K. Theodoropoulos, Muzhou Xiong, and Michael Lees. Synchronization in federation community networks. Journal of Parallel and Distributed Computing, 70(2):144–159, February 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chiang:2011:IIP

[CTC11] Jeng-Long Chiang, Yin-Yeh Tseng, and Wen-Tsuen Chen. Interest-Intended Piece Selection in BitTorrent-like peer-to-peer file sharing systems. Journal of Parallel and Distributed Computing, 71(6):879–888, June 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731510002807.

Cheng:2017:DES

[CTKA17] Long Cheng, Ilias Tachmazidis, Spyros Kotoulas, and Grigoris Antoniou. Design and evaluation of small-large outer joins in cloud computing environments. Journal of Parallel and Distributed Computing, 110(??):2–15, December 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300758.

Capel:2017:TCP

[CTS17] Manuel I. Capel, Antonio J. Tomeu, and Alberto G. Salguero. Teaching concurrent and parallel programming by patterns: an interactive ICT approach. Journal of Parallel and Distributed Computing, 105(??):42–52, July 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300163.

Comito:2016:DSD

[CTT16] Carmela Comito, Domenico Talia, and Paolo Trunfio. A distributed selectivity-driven search strategy for semi-structured data over DHT-based networks. Journal of Parallel and Distributed Computing, 93–94(??):10–29, July 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300065.

Cuzzocrea:2011:SIJ

[Cuz11] Alfredo Cuzzocrea. A special issue of Journal of Parallel and Distributed Computing: Models and algorithms for highperformance distributed data mining. Journal of Parallel and Distributed Computing, 71(5):729–730, May 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000372.

Cuzzocrea:2013:MAH

[Cuz13] Alfredo Cuzzocrea. Models and algorithms for highperformance distributed data mining. Journal of Parallel and Distributed Computing, 73(3):281–283, March 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002717.

Chen:2016:EHS

[CV16] Jie Chen and Guru Venkataramani. enDebug: a hardwaresoftware framework for automated energy debugging. Journal of Parallel and Distributed Computing, 96(??):121–133, October 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516300351.

Chatterjee:2018:SFF

[CVK^{+}18a] Anando G. Chatterjee, Mahendra K. Verma, Abhishek Kumar, Ravi Samtaney, Bilel Hadri, and Rooh Khurram. Scaling of a Fast Fourier Transform and a pseudo-spectral fluid solver up to 196608 cores. Journal of Parallel and Distributed Computing, 113(??):77–91, March 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302903.

Chaudhury:2018:LHW

[CVK^{+}18b] Bhaskar Chaudhury, Akshar Varma, Yashwant Keswani, Yashodhan Bhatnagar, and Samarth Parikh. Let’s HPC: A web-based platform to aid parallel, distributed and high performance computing education. Journal of Parallel and Distributed Computing, 118 (part 1)(??):213–232, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0743731518301205.

Chlopkowski:2015:GPL

[CW15] Marek Chlopkowski and Rafal Walkowiak. A general purpose lossless data compression method for GPU. Journal of Parallel and Distributed Computing, 75(??):40–52, January 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001956.

Cheng:2018:DSE

[CWCW18] Yuxia Cheng, Qing Wu, Wenzhi Chen, and Bei Wang. Distributed shielded execution for transmissible cyber threats analysis. Journal of Parallel and Distributed Computing, 122(??):70–80, December 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731518305069.

Cui:2011:DDM

[CWD11] Yong Cui, Shengling Wang, and Sajal K. Das. Distributed dynamic mobile multicast. Journal of Parallel and Distributed Computing, 71(9):1215–1224, September 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731510002649.

Chen:2012:EIG

[CWP12] Lizhong Chen, Ruisheng Wang, and Timothy M. Pinkston. Efficient implementation of globally-aware network flow control. Journal of Parallel and Distributed Computing, 72(11):1412–1422, November 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151200038X.

Cao:2018:GEG

[CWZ^{+}18] Liang Cao, Yufeng Wang, Bo Zhang, Qun Jin, and Athanasios V. Vasilakos. GCHAR: An efficient group-based contextaware human activity recognition on smartphone. Journal of Parallel and Distributed Computing, 118 (part 1)(??):67–80, August 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0743731517301582.

Chen:2018:REC

[CXQ^{+}18] Chen Chen, Hongyu Xiang, Tie Qiu, Cong Wang, Yang Zhou, and Victor Chang. A rear-end collision prediction scheme based on deep learning in the Internet of vehicles. Journal of Parallel and Distributed Computing, 117(??):192–204, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302447.

Chen:2018:FAC

[CXX^{+}18] Hao Chen, Jianglong Xu, Guangyi Xiao, Qi Wu, and Shiqin Zhang. Fast auto-clean CNN model for online prediction of food materials. Journal of Parallel and Distributed Computing, 117(??):218–227, July 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302162.

Chen:2014:PPV

[CXY14] Fei Chen, Tao Xiang, and Yuanyuan Yang. Privacypreserving and verifiable protocols for scientific computation outsourcing to the cloud. Journal of Parallel and Distributed Computing, 74(3):2141–2151, March 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002347.

Czapinski:2013:EPM

[Cza13] Michal Czapiński. An effective Parallel Multistart Tabu Search for Quadratic Assignment Problem on CUDA platform. Journal of Parallel and Distributed Computing, 73(11):1461–1468, November 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151200175X.

Chen:2016:SNP

[CZPP16] Lizhong Chen, Di Zhu, Massoud Pedram, and Timothy M. Pinkston. Simulation of NoC power-gating: Requirements, optimizations, and the agate simulator. Journal of Parallel and Distributed Computing, 95(??):69–78, September 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516000356.

Chen:2017:DDA

[CZZ^{+}17] Jianjun Chen, Dawei Zhao, Yao Zheng, Yan Xu, Chenfeng Li, and Jianjing Zheng. Domain decomposition approach for parallel improvement of tetrahedral meshes. Journal of Parallel and Distributed Computing, 107(??):101–113, September 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517301302.

deAraujo:2019:VPC

[dAAD^{+}19] João Paulo de Araujo, Luciana Arantes, Elias P. Duarte, Luiz A. Rodrigues, and Pierre Sens. VCube-PS: a causal broadcast topic-based publish/subscribe system. Journal of Parallel and Distributed Computing, 125(??):18–30, March 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518307822.

Dubey:2014:SHL

[DAB^{+}14] Anshu Dubey, Ann Almgren, John Bell, Martin Berzins, Steve Brandt, Greg Bryan, Phillip Colella, Daniel Graves, Michael Lijewski, Frank Löffler, Brian O’Shea, Erik Schnetter, Brian Van Straalen, and Klaus Weide. A survey of high level frameworks in block-structured adaptive mesh refinement packages. Journal of Parallel and Distributed Computing, 74(12):3217–3227, December 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001178.

Dantas:2018:EEP

[dADC18] Bianca de Almeida Dantas and Edson Norberto Cáceres. An experimental evaluation of a parallel simulated annealing approach for the 0–1 multidimensional knapsack problem. Journal of Parallel and Distributed Computing, 120(??):211–221, October 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518301096.

Draeger:2017:MPF

[DAG^{+}17] Erik W. Draeger, Xavier Andrade, John A. Gunnels, Abhinav Bhatele, André Schleife, and Alfredo A. Correa. Massively parallel first-principles simulation of electron dynamics in materials. Journal of Parallel and Distributed Computing, 106(??):205–214, August 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300734.

Macedo:2013:EGC

[dAMFdS13] Raimundo José de Araújo Macêdo, Allan Edgard Silva Freitas, and Alírio Santos de Sá. Enhancing group communication with self-manageable behavior. Journal of Parallel and Distributed Computing, 73(4):420–433, April 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002869.

Din:2018:MMP

[DAPR18] Sadia Din, Awais Ahmad, Anand Paul, and Seungmin Rho. MGR: Multi-parameter Green Reliable communication for Internet of Things in 5G network. Journal of Parallel and Distributed Computing, 118 (part 1)(??):34–45, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S074373151730343X.

deAndrade:2017:OFH

[dAT17] Douglas Coimbra de Andrade and Luís Gonzaga Trabasso. An OpenCL framework for high performance extraction of image features. Journal of Parallel and Distributed Computing, 109(??):75–88, November 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301624.

David:2017:LLD

[Dav17] Jean Pierre David. Low latency and division free Gauss– Jordan solver in floating point arithmetic. Journal of Parallel and Distributed Computing, 106(??):185–193, August 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301897.

DeGrande:2011:DBC

[DB11] Robson E. De Grande and Azzedine Boukerche. Dynamic balancing of communication and computation load for HLA-based simulations on large-scale distributed systems. Journal of Parallel and Distributed Computing, 71(1):40–52, January 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Drummond:2018:SIC

[DB18] Lúcia Drummond and Edson Borin. Special issue on computer architecture and high performance computing. Journal of Parallel and Distributed Computing, 120(??):195, October 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518305148.

Dabah:2018:HMC

[DBA^{+}18] Adel Dabah, Ahcène Bendjoudi, Abdelhakim AitZai, Didier El-Baz, and Nadia Nouali Taboudjemat. Hybrid multicore CPU and GPU-based B&B approaches for the blocking job shop scheduling problem. Journal of Parallel and Distributed Computing, 117(??):73–86, July 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300820.

DiFatta:2013:FTD

[DBCF13] Giuseppe Di Fatta, Francesco Blasa, Simone Cafiero, and Giancarlo Fortino. Fault tolerant decentralised K-means clustering for asynchronous large-scale networks. Journal of Parallel and Distributed Computing, 73(3):317–329, March 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512002249.

Droz-Bartholet:2012:RCW

[DBLB^{+}12] L. Droz-Bartholet, J.-C. Lapayre, F. Bouquet, E. Garcia, and A. Heinisch. Ramos: Concurrent writing and reconfiguration for collaborative systems. Journal of Parallel and Distributed Computing, 72(5):637–649, May 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000561.

Dhurandher:2018:LPB

[DBW^{+}18] Sanjay K. Dhurandher, Satya J. Borah, I. Woungang, Aman Bansal, and Apoorv Gupta. A location predictionbased routing scheme for opportunistic networks in an IoT scenario. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):369–378, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302381.

Djidjev:2015:APS

[DCA^{+}15] Hristo Djidjev, Guillaume Chapuis, Rumen Andonov, Sunil Thulasidasan, and Dominique Lavenier. All-Pairs Shortest Path algorithms for planar graph for GPU-accelerated clusters. Journal of Parallel and Distributed Computing, 85(??):91–103, November 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515001069.

deCampos:2019:PMS

[dCPD19] Arion de Campos, Aurora T. R. Pozo, and Elias P. Duarte. Parallel multi-swarm PSO strategies for solving many objective optimization problems. Journal of Parallel and Distributed Computing, 126(??):13–33, April 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308554.

Dorronsoro:2019:SIA

[DDE19] Bernabé Dorronsoro, Grégoire Danoy, and Didier El Baz. Special issue on advances in parallel and distributed combinatorial optimization. Journal of Parallel and Distributed Computing, 133(??):176–178, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519305271.

Davtyan:2017:CCT

[DDG^{+}17] Seda Davtyan, Roberto De Prisco, Chryssis Georgiou, Theophanis Hadjistasi, and Alexander A. Schwarzmann. Coordinated cooperative task computing using crash-prone processors with unreliable multicast. Journal of Parallel and Distributed Computing, 109(??):272–285, November 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302022.

Delevacq:2013:PAC

[DDGK13] Audrey Delévacq, Pierre Delisle, Marc Gravel, and Michaël Krajecki. Parallel ant colony optimization on graphics processing units. Journal of Parallel and Distributed Computing, 73(1):52–61, January 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512000044.

Delaet:2010:SSM

[DDNT10] Sylvie Delaët, Stéphane Devismes, Mikhail Nesterenko, and Sébastien Tixeuil. Snap-stabilization in message-passing systems. Journal of Parallel and Distributed Computing, 70(12): 1220–1230, December 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

DAmbrosio:2018:OCA

[DDO^{+}18] Donato D’Ambrosio, Alessio De Rango, Marco Oliverio, Davide Spataro, William Spataro, Rocco Rongo, Giuseppe Mendicino, and Alfonso Senatore. The open computing abstraction layer for parallel complex systems modeling on many-core systems. Journal of Parallel and Distributed Computing, 121(??):42–52, November 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304799.

Doallo:2012:SIE

[DF12] Ramón Doallo and Basilio B. Fraguela. Special issue editorial: Accelerators for high-performance computing. Journal of Parallel and Distributed Computing, 72(9):1055–1056, September 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001219.

DAngelo:2017:HID

[DF17] Gabriele D’Angelo and Stefano Ferretti. Highly intensive data dissemination in complex networks. Journal of Parallel and Distributed Computing, 99(??):28–50, January 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301058.

Dai:2013:PDE

[DFHH13] Yuan Dai, Yong Fang, Dongjian He, and Bormin Huang. Parallel design for error-resilient entropy coding algorithm on GPU. Journal of Parallel and Distributed Computing, 73(4):411–419, April 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002894.

DeMaio:2017:DOT

[DFLO17] Carmen De Maio, Giuseppe Fenza, Vincenzo Loia, and Francesco Orciuoli. Distributed online Temporal Fuzzy Concept Analysis for stream processing in smart cities. Journal of Parallel and Distributed Computing, 110(??):31–41, December 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517300503.

DiGregorio:2013:AWS

[DFST13] Salvatore Di Gregorio, Giuseppe Filippone, William Spataro, and Giuseppe A. Trunfio. Accelerating wildfire susceptibility mapping through GPGPU. Journal of Parallel and Distributed Computing, 73(8):1183–1194, August 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000580.

Dash:2014:LCM

[DGBN14] Dinesh Dash, Arobinda Gupta, Arijit Bishnu, and Subhas C. Nandy. Line coverage measures in wireless sensor networks. Journal of Parallel and Distributed Computing, 74(7):2596– 2614, July 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514000604.

Delporte-Gallet:2010:SLE

[DGDF10] C. Delporte-Gallet, S. Devismes, and H. Fauconnier. Stabilizing leader election in partial synchronous systems with crash failures. Journal of Parallel and Distributed Computing, 70 (1):45–58, January 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Delling:2013:PHA

[DGNW13] Daniel Delling, Andrew V. Goldberg, Andreas Nowatzyk, and Renato F. Werneck. PHAST: Hardware-accelerated shortest path trees. Journal of Parallel and Distributed Computing, 73(7):940–952, July 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151200041X.

Devismes:2016:SSS

[DJ16] Stéphane Devismes and Colette Johnen. Silent self-stabilizing BFS tree algorithms revisited. Journal of Parallel and Distributed Computing, 97(??):11–23, November 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300685.

Deb:2019:CER

[DJDK19] Dipika Deb, John Jose, Shirshendu Das, and Hemangee K. Kapoor. Cost effective routing techniques in 2D mesh NoC using on-chip transmission lines. Journal of Parallel and Distributed Computing, 123(??):118–129, January 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306737.

Dhanapala:2011:RRW

[DJH11] Dulanjalie C. Dhanapala, Anura P. Jayasumana, and Qi Han. On random routing in wireless sensor grids: a mathematical model for rendezvous probability and performance optimization. Journal of Parallel and Distributed Computing, 71(3): 369–380, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Daoud:2011:HHG

[DK11] Mohammad I. Daoud and Nawwaf Kharma. A hybrid heuristic-genetic algorithm for task scheduling in heterogeneous processor networks. Journal of Parallel and Distributed Computing, 71(11):1518–1531, November 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001018.

Di:2014:GHP

[DKC14] Sheng Di, Derrick Kondo, and Walfredo Cirne. Google hostload prediction based on Bayesian model with optimized feature combination. Journal of Parallel and Distributed Computing, 74(1):1820–1832, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002128.

daSilvaeSilva:2010:AEM

[dKG^{+}10] Francisco José da Silva e Silva, Fabio Kon, Alfredo Goldman, Marcelo Finger, Raphael Y. de Camargo, Fernando Castor Filho, and Fábio M. Costa. Application execution management on the InteGrade opportunistic grid middleware. Journal of Parallel and Distributed Computing, 70(5):573–583, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dorri:2019:LLS

[DKJG19] Ali Dorri, Salil S. Kanhere, Raja Jurdak, and Praveen Gauravaram. LSB: a Lightweight Scalable Blockchain for IoT security and anonymity. Journal of Parallel and Distributed Computing, 134(??):180–197, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518307688.

Do:2018:ACA

[DKK18] Cong Thuan Do, Jong Myon Kim, and Cheol Hong Kim. Application characteristics-aware sporadic cache bypassing for high performance GPGPUs. Journal of Parallel and Distributed Computing, 122(??):238–250, December 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306373.

Daily:2015:WSB

[DKKV15] Jeff Daily, Ananth Kalyanaraman, Sriram Krishnamoorthy, and Abhinav Vishnu. A work stealing based approach for enabling scalable optimal sequence homology detection. Journal of Parallel and Distributed Computing, 79–80(??):132–142, May 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001518.

Dimokas:2010:EED

[DKM10] N. Dimokas, D. Katsaros, and Y. Manolopoulos. Energyefficient distributed clustering in wireless sensor networks. Journal of Parallel and Distributed Computing, 70(4):371–383, April 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dehne:2015:SRT

[DKRC^{+}15] F. Dehne, Q. Kong, A. Rau-Chaplin, H. Zaboli, and R. Zhou. Scalable real-time OLAP on cloud architectures. Journal of Parallel and Distributed Computing, 79–80(??):31–41, May 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001488.

Dufosse:2015:TAA

[DKU15] Fanny Dufossé, Kamer Kaya, and Bora Uçar. Two approximation algorithms for bipartite matching on multicore architectures. Journal of Parallel and Distributed Computing, 85(??):62–78, November 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001070.

Deveci:2015:HPM

[DKUÇ15] Mehmet Deveci, Kamer Kaya, Bora Uçar, and Ümit V. Çatalyürek. Hypergraph partitioning for multiple communication cost metrics: Model and methods. Journal of Parallel and Distributed Computing, 77(??):69–83, March 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514002275.

delaAsuncion:2012:MCI

[dlAMCFN12] Marc de la Asunción, José M. Mantas, Manuel J. Castro, and E. D. Fernández-Nieto. An MPI-CUDA implementation of an improved Roe method for two-layer shallow water systems. Journal of Parallel and Distributed Computing, 72(9):1065–1072, September 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151100147X.

Dang:2011:DPP

[DLLL11] Viet-Hung Dang, Viet-Duc Le, Young-Koo Lee, and Sungyoung Lee. Distributed Push–Pull Estimation for node localization in wireless sensor networks. Journal of Parallel and Distributed Computing, 71(3):471–484, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

DeMarco:2019:DMO

[DLM19] Gianluca De Marco, Mauro Leoncini, and Manuela Montangero. A distributed message-optimal assignment on rings. Journal of Parallel and Distributed Computing, 132(??): 284–298, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731519303946.

Datta:2011:TSS

[DLV11] Ajoy K. Datta, Lawrence L. Larmore, and Priyanka Vemula. An O(n)-time self-stabilizing leader election algorithm. Journal of Parallel and Distributed Computing, 71(11):1532–1544, November 2011. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511001043.

Dong:2012:DAL

[DLW^{+}12] Bin Dong, Xiuqiao Li, Qimeng Wu, Limin Xiao, and Li Ruan. A dynamic and adaptive load balancing strategy for parallel file system with large-scale I/O servers. Journal of Parallel and Distributed Computing, 72(10):1254–1268, October 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512001244.

Dimitrakopoulou:2017:NNE

[DM17] K. A. Dimitrakopoulou and N. M. Missirlis. The nine node extrapolated diffusion method for weighted torus graphs. Journal of Parallel and Distributed Computing, 106(??):62–78, August 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300801.

Dias:2018:JDR

[DMG18] Vinicius Dias, Wagner Meira, and Dorgival Guedes. Janus: Diagnostics and reconfiguration of data parallel programs. Journal of Parallel and Distributed Computing, 120(??):196–210, October 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518301084.

Diaz:2019:RRD

[DMI^{+}19] Javier Díaz, Teresa Monreal, Pablo Ibáñez, José M. Llabería, and Víctor Viñals. ReD: a reuse detector for content selection in exclusive shared last-level caches. Journal of Parallel and Distributed Computing, 125(??):106–120, March 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308414.

Dalcin:2019:FPM

[DMK19] Lisandro Dalcin, Mikael Mortensen, and David E. Keyes. Fast parallel multidimensional FFT using advanced MPI. Journal of Parallel and Distributed Computing, 128(??):137–150, June 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151830306X.

Duenha:2016:MBH

[DMS^{+}16] Liana Duenha, Guilherme Madalozzo, Thiago Santiago, Fernando Moraes, and Rodolfo Azevedo. MPSoCBench: a benchmark for high-level evaluation of multiprocessor system-on-chip tools and methodologies. Journal of Parallel and Distributed Computing, 95(??):138–157, September 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516000381.

deMiras:2018:GIT

[dMS18] Juan Ruiz de Miras and Mario Salazar. GPU inclusion test for triangular meshes. Journal of Parallel and Distributed Computing, 120(??):170–181, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304246.

deOliveira:2015:ELF

[dOBG^{+}15] Horacio A. B. F. de Oliveira, Azzedine Boukerche, Daniel L. Guidoni, Eduardo F. Nakamura, Raquel A. F. Mini, and Antonio A. F. Loureiro. An enhanced location-free Greedy Forward algorithm with hole bypass capability in wireless sensor networks. Journal of Parallel and Distributed Computing, 77(??): 1–10, March 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514002007.

deOliveira:2014:UAS

[dOCS14] Romulo Silva de Oliveira, Andreu Carminati, and Renan Augusto Starke. Using an adversary simulator to evaluate global EDF scheduling of sporadic task sets on multiprocessors. Journal of Parallel and Distributed Computing, 74(10):3037–3044, October 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001130.

Dereniowski:2012:DMA

[DP12] Dariusz Dereniowski and Andrzej Pelc. Drawing maps with advice. Journal of Parallel and Distributed Computing, 72(2):132–143, February 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001936.

Defour:2016:SSS

[DP16] David Defour and Eric Petit. A software scheduling solution to avoid corrupted units on GPUs. Journal of Parallel and Distributed Computing, 90–91(??):1–8, April 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516000022.

Dubois:2012:SSB

[DPBNT12] Swan Dubois, Maria Potop-Butucaru, Mikhail Nesterenko, and Sébastien Tixeuil. Self-stabilizing Byzantine asynchronous unison. Journal of Parallel and Distributed Computing, 72(7):917–923, July 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151200086X.

Dizaji:2018:PPA

[DR18] Lida Ghaemi Dizaji and Akbar Ghaffarpour Rahbar. PA-HON: Power-aware hybrid optical network. Journal of Parallel and Distributed Computing, 117(??):1–16, July 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300248.

Dehghani:2019:DPE

[DR19] Abbas Dehghani and Keyvan RahimiZadeh. Design and performance evaluation of mesh-of-tree-based hierarchical wireless network-on-chip for multicore systems. Journal of Parallel and Distributed Computing, 123(??):100–117, January 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518306592.

Dummler:2013:PSS

[DRR13] Jörg Dümmler, Thomas Rauber, and Gudula Rünger. Programming support and scheduling for communicating parallel tasks. Journal of Parallel and Distributed Computing, 73(2):220–234, February 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002328.

daSilva:2015:ESB

[dSAJ15] Joakim da Silva, Richard Ansorge, and Rajesh Jena. Efficient scatter-based kernel superposition on GPU. Journal of Parallel and Distributed Computing, 84(??):15–23, October 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515001215.

DeLucas:2017:DMA

[DSEP17] Enrique De Lucas, Marcos Sanchez-Elez, and Inmaculada Pardines. DSPONE48: a methodology for automatically synthesize HDL focus on the reuse of DSP slices. Journal of Parallel and Distributed Computing, 106(??):132–142, August 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300400.

daSilva:2011:SLB

[dSS11] Fabrício A. B. da Silva and Hermes Senger. Scalability limits of Bag-of-Tasks applications running on hierarchical platforms. Journal of Parallel and Distributed Computing, 71(6):788– 801, June 2011. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511000141.

Demsky:2011:IFO

[DT11] Brian Demsky and Navid Farri Tehrany. Integrating file operations into transactional memory. Journal of Parallel and Distributed Computing, 71(10):1293–1304, October 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511000803.

Doka:2011:BDF

[DTK11a] Katerina Doka, Dimitrios Tsoumakos, and Nectarios Koziris. Brown Dwarf: a fully-distributed, fault-tolerant data warehousing system. Journal of Parallel and Distributed Computing, 71(11):1434–1446, November 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001432.

Doka:2011:OQD

[DTK11b] Katerina Doka, Dimitrios Tsoumakos, and Nectarios Koziris. Online querying of d-dimensional hierarchies. Journal of Parallel and Distributed Computing, 71(3):424–437, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

DAmato:2013:CGF

[DV13] J. P. D’Amato and M. Vénere. A CPU-GPU framework for optimizing the quality of large meshes. Journal of Parallel and Distributed Computing, 73(8):1127–1134, August 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000518.

Di:2012:DPR

[DW12] Sheng Di and Cho-Li Wang. Decentralized proactive resource allocation for maximizing throughput of P2P Grid. Journal of Parallel and Distributed Computing, 72(2):308–321, February 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511002097.

Ding:2010:UMB

[DWX10] Yong Ding, Chen Wang, and Li Xiao. Using mobile beacons to locate sensors in obstructed environments. Journal of Parallel and Distributed Computing, 70(6):644–656, June 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ding:2010:PCM

[DWYB10] Jason Jianxun Ding, Abdul Waheed, Jingnan Yao, and Laxmi N. Bhuyan. Performance characterization of multithread and multi-core processors based XML application oriented networking systems. Journal of Parallel and Distributed Computing, 70(5):584–597, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ding:2019:APM

[DXS^{+}19] Nan Ding, Wei Xue, Zhenya Song, Haohuan Fu, Shiming Xu, and Weimin Zheng. An automatic performance model-based scheduling tool for coupled climate system models. Journal of Parallel and Distributed Computing, 132(??):204–216, October 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151830008X.

Dong:2012:HPN

[DYL^{+}12] Yaozu Dong, Xiaowei Yang, Jianhui Li, Guangdeng Liao, Kun Tian, and Haibing Guan. High performance network virtualization with SR-IOV. Journal of Parallel and Distributed Computing, 72(11):1471–1480, November 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000329.

Dai:2017:COP

[DZC17] Hongjun Dai, Shulin Zhao, and Kang Chen. A chaosoriented prediction and suppression model to enhance the security for cyber physical power systems. Journal of Parallel and Distributed Computing, 103(??):87–95, May 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301757.

Erbagci:2019:SHE

[EAB^{+}19] Burak Erbagci, Nail Etkin Can Akkaya, Mudit Bhargava, Rachel Dondero, and Ken Mai. Secure hardware-entangled field programmable gate arrays. Journal of Parallel and Distributed Computing, 131(??):81–96, September 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519302618.

El-Boghdadi:2013:CAO

[EB13] Hatem M. El-Boghdadi. A class of almost-optimal sizeindependent parallel prefix circuits. Journal of Parallel and Distributed Computing, 73(6):888–894, June 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000567.

Eckroth:2018:CBD

[Eck18] Joshua Eckroth. A course on big data analytics. Journal of Parallel and Distributed Computing, 118 (part 1)(??):166–176, August 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0743731518300972.

Etinski:2012:UFE

[ECLV12] M. Etinski, J. Corbalan, J. Labarta, and M. Valero. Understanding the future of energy-performance trade-off via DVFS in HPC environments. Journal of Parallel and Distributed Computing, 72(4):579–590, April 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000172.

Esposito:2018:EBS

[ECP^{+}18] Christian Esposito, Aniello Castiglione, Francesco Palmieri, Massimo Ficco, Ciprian Dobre, George V. Iordache, and Florin Pop. Event-based sensor data exchange and fusion in the Internet of Things environments. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):328–343, August 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517303416.

Ekanayake:2019:MMD

[ECWV19] Saliya Ekanayake, Jose Cadena, Udayanga Wickramasinghe, and Anil Vullikanti. MIDAS: Multilinear detection at scale. Journal of Parallel and Distributed Computing, 132(??):363–382, October 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518305239.

Evropeytsev:2017:ECG

[EDH^{+}17] Grigory Evropeytsev, Eduardo López Domínguez, Saul E. Pomares Hernandez, Marco Antonio López Trinidad, and José Roberto Perez Cruz. An efficient causal group communication protocol for P2P hierarchical overlay networks. Journal of Parallel and Distributed Computing, 102(??):149–162, April 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301836.

Esterie:2014:NTT

[EFG^{+}14] Pierre Estérie, Joel Falcou, Mathias Gaunard, Jean-Thierry Lapresté, and Lionel Lacassagne. The numerical template toolbox: a modern C++ design for scientific computing. Journal of Parallel and Distributed Computing, 74(12):3240–3253, December 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001245.

El-Hadedy:2019:RPR

[EHKSS19] Mohamed El-Hadedy, Amit Kulkarni, Dirk Stroobandt, and Kevin Skadron. Reco-Pi: a reconfigurable cryptoprocessor for π-cipher. Journal of Parallel and Distributed Computing, 133(??):420–431, November 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301636.

Eom:2015:EKL

[EHL^{+}15] Youngmoon Eom, Deukyeon Hwang, Junyong Lee, Jonghwan Moon, Minho Shin, and Beomseok Nam. EM-KDE: a locality-aware job scheduling policy with distributed semantic caches. Journal of Parallel and Distributed Computing, 83(??):119–132, September 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001008.

Eijkhout:2018:TDM

[Eij18] Victor Eijkhout. Teaching distributed memory programming from mental models. Journal of Parallel and Distributed Computing, 118 (part 1)(??):120–127, August 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0743731518301072.

Eugster:2017:HPP

[EKNS17] Patrick Eugster, Kirill Kogan, Sergey I. Nikolenko, and Alexander V. Sirotkin. Heterogeneous packet processing in shared memory buffers. Journal of Parallel and Distributed Computing, 99(??):1–13, January 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300879.

Esnaashari:2011:CLA

[EM11] M. Esnaashari and M. R. Meybodi. A cellular learning automata-based deployment strategy for mobile wireless sensor networks. Journal of Parallel and Distributed Computing, 71(7):988–1001, July 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151000225X.

Elkamel:2019:ELW

[EMC19] Rabiaa Elkamel, Asma Messouadi, and Adnane Cherif. Extending the lifetime of wireless sensor networks through mitigating the hot spot problem. Journal of Parallel and Distributed Computing, 133(??):159–169, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304460.

Emeliyanenko:2013:CRG

[Eme13] Pavel Emeliyanenko. Computing resultants on Graphics Processing Units: Towards GPU-accelerated computer algebra. Journal of Parallel and Distributed Computing, 73(11):1494–1505, November 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001761.

Ezhilarasi:2012:NPU

[ES12] G. Angeline Ezhilarasi and K. S. Swarup. Network partitioning using harmony search and equivalencing for distributed computing. Journal of Parallel and Distributed Computing, 72(8):936–943, August 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001025.

Esteves:2015:IDE

[ESCV15] Sérgio Esteves, João Nuno Silva, João Paulo Carvalho, and Luís Veiga. Incremental dataflow execution, resource efficiency and probabilistic guarantees with Fuzzy Boolean nets. Journal of Parallel and Distributed Computing, 79–80(??):52– 66, May 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515000507.

Escudero-Sahuquillo:2011:OSC

[ESGQ^{+}11] Jesus Escudero-Sahuquillo, Pedro J. Garcia, Francisco J. Quiles, Jose Flich, and Jose Duato. OBQA: Smart and cost-efficient queue scheme for Head-of-Line blocking elimination in fat-trees. Journal of Parallel and Distributed Computing, 71(11):1460–1472, November 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001420.

Escudero-Sahuquillo:2014:NPD

[ESGQ^{+}14] Jesus Escudero-Sahuquillo, Pedro J. Garcia, Francisco J. Quiles, Sven-Arne Reinemo, Tor Skeie, Olav Lysne, and Jose Duato. A new proposal to deal with congestion in InfiniBand-based fat-trees. Journal of Parallel and Distributed Computing, 74(1):1802–1819, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001986.

Escudero-Sahuquillo:2018:FEC

[ESGQ^{+}18] Jesus Escudero-Sahuquillo, Pedro J. Garcia, Francisco J. Quiles, German Maglione-Mathey, and Jose Duato. Feasible enhancements to congestion control in InfiniBandbased networks. Journal of Parallel and Distributed Computing, 112 (part 1)(??):35–52, February 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302629.

Edwards:2014:KEM

[ETS14] H. Carter Edwards, Christian R. Trott, and Daniel Sunderland. Kokkos: Enabling manycore performance portability through polymorphic memory access patterns. Journal of Parallel and Distributed Computing, 74(12):3202–3216, December 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001257.

Ferreron:2019:FTL

[FABG^{+}19] Alexandra Ferrerón, Jesús Alastruey-Benedé, Darío Suárez Gracia, Teresa Monreal Arnal, Pablo Ibáñez Marín, and Víctor Viñals Yúfera. A fault-tolerant last level cache for CMPs operating at ultra-low voltage. Journal of Parallel and Distributed Computing, 125(??):31–44, March 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518307810.

Foteinos:2014:HQR

[FC14] Panagiotis A. Foteinos and Nikos P. Chrisochoides. High quality real-time Image-to-Mesh conversion for finite element simulations. Journal of Parallel and Distributed Computing, 74(2):2123–2140, February 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513002232.

Frache:2014:EDS

[FCG^{+}14] S. Frache, D. Chiabrando, M. Graziano, M. Vacca, L. Boarino, and M. Zamboni. Enabling design and simulation of massive parallel nanoarchitectures. Journal of Parallel and Distributed Computing, 74(6):2530–2541, June 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001366.

Fernandez-Cerero:2018:SSE

[FCJG^{+}18] Damián Fernández-Cerero, Agnieszka Jakóbik, Daniel Grzonka, Joanna Kolodziej, and Alejandro Fernández-Montes. Security supportive energy-aware scheduling and energy policies for cloud environments. Journal of Parallel and Distributed Computing, 119(??):191–202, September 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518302843.

Fan:2013:GBC

[FCML13] Xiaopeng Fan, Jiannong Cao, Haixia Mao, and Yunhuai Liu. Gossip-based cooperative caching for mobile applications in mobile wireless networks. Journal of Parallel and Distributed Computing, 73(5):653–663, May 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000075.

Francesquini:2015:EEP

[FCP^{+}15] Emilio Francesquini, Márcio Castro, Pedro H. Penna, Fabrice Dupros, Henrique C. Freitas, Philippe O. A. Navaux, and Jean-François Méhaut. On the energy efficiency and performance of irregular application executions on multicore, NUMA and manycore platforms. Journal of Parallel and Distributed Computing, 76(??):32–48, February 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514002093.

Fan:2011:CAD

[FCW11] Xiaopeng Fan, Jiannong Cao, and Weigang Wu. Contentionaware data caching in wireless multi-hop ad hoc networks. Journal of Parallel and Distributed Computing, 71(4):603–614, April 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Feng:2012:CML

[FCZ^{+}12] Wei Feng, Jiannong Cao, Chisheng Zhang, Jun Zhang, and Qin Xin. Coordination of multi-link spectrum handoff in multi-radio multi-hop cognitive networks. Journal of Parallel and Distributed Computing, 72(4):613–625, April 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511002206.

Fatema:2014:SCM

[FEH^{+}14] Kaniz Fatema, Vincent C. Emeakaroha, Philip D. Healy, John P. Morrison, and Theo Lynn. A survey of cloud monitoring tools: Taxonomy, capabilities and objectives. Journal of Parallel and Distributed Computing, 74(10):2918–2933, October 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001099.

Fang:2019:RRS

[FFYH19] Jiarui Fang, Haohuan Fu, Guangwen Yang, and Cho-Jui Hsieh. RedSync: Reducing synchronization bandwidth for distributed deep learning training system. Journal of Parallel and Distributed Computing, 133(??):30–39, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518308657.

Feng:2017:PPP

[FGcF17] Annette Feng, Mark Gardner, and Wu chun Feng. Parallel programming with pictures is a Snap! Journal of Parallel and Distributed Computing, 105(??):150–162, July 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300242.

Felber:2017:ET

[FGG17] Pascal Felber, Vincent Gramoli, and Rachid Guerraoui. Elastic transactions. Journal of Parallel and Distributed Computing, 100(??):103–127, February 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516301290.

Fu:2011:GGR

[FGL^{+}11] Cai Fu, Xiang Gao, Ming Liu, Xiaoyang Liu, Lansheng Han, and Jing Chen. GRAP: Grey risk assessment based on projection in ad hoc networks. Journal of Parallel and Distributed Computing, 71(9):1249–1260, September 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731510002613.

Faverge:2015:MFA

[FHL^{+}15] Mathieu Faverge, Julien Herrmann, Julien Langou, Bradley Lowery, Yves Robert, and Jack Dongarra. Mixing LU and QR factorization algorithms to design high-performance dense linear algebra solvers. Journal of Parallel and Distributed Computing, 85(??):32–46, November 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001057.

Fernandez:2017:PQS

[FKB17] Joaquín Fernandez, Ernesto Kofman, and Federico Bergero. A parallel quantized state system solver for ODEs. Journal of Parallel and Distributed Computing, 106(??):14–30, August 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300795.

Fatourou:2016:WTW

[FKKR16] Panagiota Fatourou, Eleni Kanellou, Eleftherios Kosmas, and Md Forhad Rabbi. WFR-TM: Wait-free readers without sacrificing speculation of writers. Journal of Parallel and Distributed Computing, 96(??):134–151, October 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300326.

Foley:2017:OWP

[FKR^{+}17] Samantha S. Foley, Daniel Koepke, Justin Ragatz, Christa Brehm, Jason Regina, and Joshua Hursey. OnRamp: a webportal for teaching parallel and distributed computing. Journal of Parallel and Distributed Computing, 105(??):138–149, July 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300205.

Fu:2014:DDM

[FLC14] Weiwei Fu, Li Liu, and Tianzhou Chen. Direct distributed memory access for CMPs. Journal of Parallel and Distributed Computing, 74(2):2109–2122, February 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002311.

Flouris:2010:EBL

[FLCB10] Michail D. Flouris, Renaud Lachaize, Konstantinos Chasapis, and Angelos Bilas. Extensible block-level storage virtualization in cluster-based systems. Journal of Parallel and Distributed Computing, 70(8):800–824, August 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fallahpour:2014:PPM

[FLL14] Mehrdad Fallahpour, Ming-Bo Lin, and Chang-Hong Lin. Parallel photon-mapping rendering on a mesh-NoC-based MPSoC platform. Journal of Parallel and Distributed Computing, 74(7):2626–2638, July 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514000616.

Funke:2019:CFM

[FLM^{+}19] Daniel Funke, Sebastian Lamm, Ulrich Meyer, Manuel Penschuck, Peter Sanders, Christian Schulz, Darren Strash, and Moritz von Looz. Communication-free massively distributed graph generation. Journal of Parallel and Distributed Computing, 131(??):200–217, September 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304684.

Filho:2018:AOV

[FMIF18] Manoel C. Silva Filho, Claudio C. Monteiro, Pedro R. M. Inácio, and Mário M. Freire. Approaches for optimizing virtual machine placement and migration in cloud environments: a survey. Journal of Parallel and Distributed Computing, 111(??):222–250, January 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151730240X.

Fraigniaud:2017:DCM

[FP17] Pierre Fraigniaud and Andrzej Pelc. Decidability classes for mobile agents computing. Journal of Parallel and Distributed Computing, 109(??):117–128, November 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301259.

Fard:2014:MOL

[FPF14] Hamid Mohammadi Fard, Radu Prodan, and Thomas Fahringer. Multi-objective list scheduling of workflow applications in distributed computing infrastructures. Journal of Parallel and Distributed Computing, 74(3):2152–2165, March 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513002384.

Friedman:2014:ECS

[FPM^{+}14] Joseph S. Friedman, John A. Peters, Gokhan Memik, Bruce W. Wessels, and Alan V. Sahakian. Emitter-coupled spin-transistor logic. Journal of Parallel and Distributed Computing, 74(6):2461–2469, June 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513001780.

Flich:2011:SIJ

[FPS11] José Flich, Scott Pakin, and Craig Stunkel. Special issue of Journal of Parallel and Distributed Computing: Communication architectures for scalable systems (CASS). Journal of Parallel and Distributed Computing, 71(3):523, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Flich:2012:SIC

[FPS12] José Flich, Scott Pakin, and Craig Stunkel. Special issue on Communication Architectures for Scalable Systems. Journal of Parallel and Distributed Computing, 72(11):1399–1400, November 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001256.

Feller:2015:PEE

[FRM15] Eugen Feller, Lavanya Ramakrishnan, and Christine Morin. Performance and energy efficiency of big data applications in cloud environments: a Hadoop case study. Journal of Parallel and Distributed Computing, 79–80(??):80–89, May 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515000027.

Feliu:2018:DLS

[FSP18] Josué Feliu, Julio Sahuquillo, and Salvador Petit. Designing lab sessions focusing on real processors for computer architecture courses: A practical perspective. Journal of Parallel and Distributed Computing, 118 (part 1)(??):128–139, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0743731518301047.

Fort:2014:FES

[FSV14] Marta Fort, J. Antoni Sellarès, and Nacho Valladares. Finding extremal sets on the GPU. Journal of Parallel and Distributed Computing, 74(1):1891–1899, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001305.

Fort:2017:ITF

[FSV17] Marta Fort, J. Antoni Sellarès, and Nacho Valladares. Intersecting two families of sets on the GPU. Journal of Parallel and Distributed Computing, 104(??):167–178, June 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151730045X.

Feitelson:2014:EUP

[FTK14] Dror G. Feitelson, Dan Tsafrir, and David Krakov. Experience with using the Parallel Workloads Archive. Journal of Parallel and Distributed Computing, 74(10):2967–2982, October 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001154.

Feo:2014:SIJ

[FTM^{+}14] John T. Feo, Antonino Tumeo, Timothy G. Mattson, Oreste Villa, and Simone Secchi. Special issue of Journal of Parallel and Distributed Computing: Architectures and Algorithms for Irregular Applications. Journal of Parallel and Distributed Computing, 74(1):2027–2028, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002244.

Florimbi:2019:EMC

[FTM^{+}19] Giordana Florimbi, Emanuele Torti, Stefano Masoli, Egidio D’Angelo, Giovanni Danese, and Francesco Leporati. Exploiting multi-core and many-core architectures for efficient simulation of biologically realistic models of Golgi cells. Journal of Parallel and Distributed Computing, 126(??):48–66, April 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518309109.

Fu:2010:FAR

[Fu10] Song Fu. Failure-aware resource management for highavailability computing clusters with distributed virtual machines. Journal of Parallel and Distributed Computing, 70 (4):384–393, April 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ferri:2010:ETE

[FWM^{+}10] Cesare Ferri, Samantha Wood, Tali Moreshet, R. Iris Bahar, and Maurice Herlihy. Embedded-TM: Energy and complexityeffective hardware transactional memory for embedded multicore systems. Journal of Parallel and Distributed Computing, 70(10):1042–1052, October 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fu:2010:QEC

[FX10] Song Fu and Cheng-Zhong Xu. Quantifying event correlations for proactive failure management in networked computing systems. Journal of Parallel and Distributed Computing, 70(11): 1100–1109, November 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Furno:2014:SSC

[FZ14] Angelo Furno and Eugenio Zimeo. Self-scaling cooperative discovery of service compositions in unstructured P2P networks. Journal of Parallel and Distributed Computing, 74(10):2994–3025, October 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001087.

Fan:2012:EAA

[FZWL12] Liya Fan, Fa Zhang, Gongming Wang, and Zhiyong Liu. An effective approximation algorithm for the Malleable Parallel Task Scheduling problem. Journal of Parallel and Distributed Computing, 72(5):693–704, May 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000238.

Gunney:2016:APB

[GA16] Brian T. N. Gunney and Robert W. Anderson. Advances in patch-based adaptive mesh refinement scalability. Journal of Parallel and Distributed Computing, 89(??):65–84, March 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515002129.

Garimella:2018:ESR

[GA18] Rama Murthy Garimella and Rumyantsev Alexander. On an exact solution of the rate matrix of G/M/1-type Markov process with small number of phases. Journal of Parallel and Distributed Computing, 119(??):172–178, September 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151830282X.

Grossman:2017:PTT

[GAC^{+}17] Max Grossman, Maha Aziz, Heng Chi, Anant Tibrewal, Shams Imam, and Vivek Sarkar. Pedagogy and tools for teaching parallel computing at the sophomore undergraduate level. Journal of Parallel and Distributed Computing, 105(??):18– 30, July 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517300047.

Gutierrez-Alcoba:2017:AAP

[GAOHG17] Alejandro Gutierrez-Alcoba, Gloria Ortega, Eligius M. T. Hendrix, and Inmaculada García. Accelerating an algorithm for perishable inventory control on heterogeneous platforms. Journal of Parallel and Distributed Computing, 104(??):12– 18, June 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516302076.

Gu:2011:HTN

[GB11] Yunfeng Gu and Azzedine Boukerche. HD Tree: a novel data structure to support multi-dimensional range query for P2P networks. Journal of Parallel and Distributed Computing, 71(8):1111–1124, August 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511000785.

Gao:2018:CPD

[GDCC18] Yuan Gao, Zhixiang Deng, Dongmin Choi, and Chang Choi. Combined pre-detection and sleeping for energy-efficient spectrum sensing in cognitive radio networks. Journal of Parallel and Distributed Computing, 114(??):85–94, April 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303441.

Gardellin:2011:GPD

[GDL^{+}11] Vanessa Gardellin, Sajal K. Das, Luciano Lenzini, Claudio Cicconetti, and Enzo Mingozzi. G-PaMeLA: a divide-and-conquer approach for joint channel assignment and routing in multiradio multi-channel wireless mesh networks. Journal of Parallel and Distributed Computing, 71(3):381–396, March 2011.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gonzalez-Ferez:2014:GFD

[GFPC14] Pilar González-Férez, Juan Piernas, and Toni Cortes. A general framework for dynamic and automatic I/O scheduling in hard and solid-state drives. Journal of Parallel and Distributed Computing, 74(5):2380–2391, May 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000227.

Ghalami:2019:SPI

[GG19] Laleh Ghalami and Daniel Grosu. Scheduling parallel identical machines to minimize makespan: a parallel approximation algorithm. Journal of Parallel and Distributed Computing, 133(??):221–231, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518303691.

Georgiou:2019:SRA

[GGY19] Konstantinos Georgiou, Jay Griffiths, and Yuval Yakubov. Symmetric rendezvous with advice: How to rendezvous in a disk. Journal of Parallel and Distributed Computing, 134(??):13–24, December 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518307251.

Guo:2017:MAE

[GHC^{+}17] Yong Guo, Sungpack Hong, Hassan Chafi, Alexandru Iosup, and Dick Epema. Modeling, analysis, and experimental comparison of streaming graph-partitioning policies. Journal of Parallel and Distributed Computing, 108(??):106–121, October 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516000186.

Glacet:2019:DCD

[GHIJ19] Christian Glacet, Nicolas Hanusse, David Ilcinkas, and Colette Johnen. Disconnected components detection and rooted shortest-path tree maintenance in networks. Journal of Parallel and Distributed Computing, 132(??):299–309, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519303934.

Guerraoui:2012:DPR

[GHK^{+}12] Rachid Guerraoui, Kévin Huguenin, Anne-Marie Kermarrec, Maxime Monod, and Ýmir Vigfússon. Decentralized polling with respectable participants. Journal of Parallel and Distributed Computing, 72(1):13–26, January 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001754.

Gagarin:2010:DHS

[GHY10] Andrei Gagarin, Sajid Hussain, and Laurence T. Yang. Distributed hierarchical search for balanced energy consumption routing spanning trees in wireless sensor networks. Journal of Parallel and Distributed Computing, 70(9):975–982, September 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gajin:2012:APM

[GJ12] Slavko Gajin and Zoran Jovanovic. An accurate performance model for network-on-chip and multicomputer interconnection networks. Journal of Parallel and Distributed Computing, 72(10):1280–1294, October 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001232.

Guellati:2010:SSS

[GK10] Nabil Guellati and Hamamache Kheddouci. A survey on selfstabilizing algorithms for independence, domination, coloring, and matching in graphs. Journal of Parallel and Distributed Computing, 70(4):406–415, April 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Georgiou:2015:CSD

[GK15] Chryssis Georgiou and Dariusz R. Kowalski. On the competitiveness of scheduling dynamically injected tasks on processes prone to crashes and restarts. Journal of Parallel and Distributed Computing, 84(??):94–107, October 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001252.

Gowanlock:2019:ASS

[GK19] Michael Gowanlock and Ben Karsin. Accelerating the similarity self-join using the GPU. Journal of Parallel and Distributed Computing, 133(??):107–123, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304209.

Goodman:2013:STM

[GKK^{+}13] Daniel Goodman, Behram Khan, Salman Khan, Mikel Luján, and Ian Watson. Software transactional memories for Scala. Journal of Parallel and Distributed Computing, 73(2):150–163, February 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512002304.

Gibson:2015:IEI

[GKS15] Michael J. Gibson, Edward C. Keedwell, and Dragan A. Savić. An investigation of the efficient implementation of cellular automata on multi-core CPU and GPU hardware. Journal of Parallel and Distributed Computing, 77(??):11–25, March 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514002044.

Gui:2012:NDT

[GL12] Jinsong Gui and Anfeng Liu. A new distributed topology control algorithm based on optimization of delay and energy in wireless networks. Journal of Parallel and Distributed Computing, 72(8):1032–1044, August 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001037.

Garcia:2014:HIA

[GLC14] Marta Garcia, Jesus Labarta, and Julita Corbalan. Hints to improve automatic load balancing with LeWI for hybrid applications. Journal of Parallel and Distributed Computing, 74(9):2781–2794, September 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514000926.

Gomez-Luna:2012:PMA

[GLGLBG12] Juan Gómez-Luna, José María González-Linares, José Ignacio Benavides, and Nicolás Guil. Performance models for asynchronous data transfers on consumer Graphics Processing Units. Journal of Parallel and Distributed Computing, 72(9):1117–1126, September 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001468.

Gao:2014:RCG

[GLW14] Jiaquan Gao, Ronghua Liang, and Jun Wang. Research on the conjugate gradient algorithm with a modified incomplete Cholesky preconditioner on GPU. Journal of Parallel and Distributed Computing, 74(2):2088–2098, February 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151300213X.

Goglin:2013:KGS

[GM13] Brice Goglin and Stéphanie Moreaud. KNEM: a generic and scalable kernel-assisted intra-node MPI communication framework. Journal of Parallel and Distributed Computing, 73(2):176–188, February 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002316.

Gorain:2014:AAS

[GM14a] Barun Gorain and Partha Sarathi Mandal. Approximation algorithms for sweep coverage in wireless sensor networks. Journal of Parallel and Distributed Computing, 74(8):2699–2707, August 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514000379.

Grace:2014:DRP

[GM14b] R. Kingsy Grace and R. Manimegalai. Dynamic replica placement and selection strategies in data grids — a comprehensive survey. Journal of Parallel and Distributed Computing, 74(2):2099–2108, February 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513002207.

Galiano:2012:GBP

[GMMP12] V. Galiano, H. Migallón, V. Migallón, and J. Penadés. GPUbased parallel algorithms for sparse nonlinear systems. Journal of Parallel and Distributed Computing, 72(9):1098–1105, September 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511002152.

Garcia-Martin:2019:EEC

[GMRRG19] Eva García-Martín, Crefeda Faviola Rodrigues, Graham Riley, and Håkan Grahn. Estimation of energy consumption in machine learning. Journal of Parallel and Distributed Computing, 134(??):75–88, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308773.

Giles:2013:DOG

[GMS^{+}13] M. B. Giles, G. R. Mudalige, B. Spencer, C. Bertolli, and I. Reguly. Designing OP2 for GPU architectures. Journal of Parallel and Distributed Computing, 73(11):1451–1460, November 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001694.

Gillan:2011:SIJ

[GMSS^{+}11] Charles Gillan, Simon McIntosh-Smith, Nico Sanna, Stan Scott, and Thomas Steinke. Special issue of the Journal of Parallel and Distributed Computing: Novel architectures for high-performance computing. Journal of Parallel and Distributed Computing, 71(2):333, February 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gomez-Martin:2016:FBI

[GMVRGS16] César Gómez-Martín, Miguel A. Vega-Rodríguez, and José-Luis González-Sánchez. Fattened backfilling: an improved strategy for job scheduling in parallel systems. Journal of Parallel and Distributed Computing, 97(??):69–77, November 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516300788.

Gupta:2015:IBS

[GN15] Suyash Gupta and V. Krishna Nandivada. IMSuite: a benchmark suite for simulating distributed algorithms. Journal of Parallel and Distributed Computing, 75(??):1–19, January 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514002032.

Guo:2018:PAM

[GNZ18] Wei Guo, Yufeng Nie, and Weiwei Zhang. Parallel adaptive mesh refinement method based on bubble-type local mesh generation. Journal of Parallel and Distributed Computing, 117(??):37–49, July 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300856.

Gai:2013:MIG

[GOH^{+}13] Jiading Gai, Nady Obeid, Joseph L. Holtrop, Xiao-Long Wu, Fan Lam, Maojing Fu, Justin P. Haldar, Wen mei W. Hwu, Zhi-Pei Liang, and Bradley P. Sutton. More IM-PATIENT: a gridding-accelerated toeplitz-based strategy for non-cartesian high-resolution 3D MRI on GPUs. Journal of Parallel and Distributed Computing, 73(5):686–697, May 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000026.

Gedik:2016:PFS

[GÖÖ16] B. Gedik, H. G. Özsema, and Ö. Öztürk. Pipelined fission for stream programs with dynamic selectivity and partitioned state. Journal of Parallel and Distributed Computing, 96(??):106–120, October 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300338.

Gaur:2010:PBS

[GPJA10] Amit Gaur, Abhinav Prakash, Saugat Joshi, and Dharma P. Agrawal. Polynomial based scheme (PBS) for establishing authentic associations in wireless mesh networks. Journal of Parallel and Distributed Computing, 70(4):338–343, April 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Guo:2019:HCM

[GPSH19] Chaopeng Guo, Jean-Marc Pierson, Jie Song, and Christina Herzog. Hot-N-Cold model for energy aware cloud databases. Journal of Parallel and Distributed Computing, 123(??): 130–144, January 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731518306919.

Gai:2018:EAT

[GQZ18] Keke Gai, Meikang Qiu, and Hui Zhao. Energy-aware task assignment for mobile cyber-enabled applications in heterogeneous cloud computing. Journal of Parallel and Distributed Computing, 111(??):126–135, January 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302319.

Grahn:2010:ASI

[Gra10a] Håkan Grahn. Acknowledgment to special issue reviewers. Journal of Parallel and Distributed Computing, 70(10):1085, October 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Grahn:2010:TM

[Gra10b] Håkan Grahn. Transactional memory. Journal of Parallel and Distributed Computing, 70(10):993–1008, October 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Guay:2015:EEL

[GRJ^{+}15] Wei Lin Guay, Sven-Arne Reinemo, Bjørn Dag Johnsen, Chien-Hua Yen, Tor Skeie, Olav Lysne, and Ola Tørudbakken. Early experiences with live migration of SR–IOV enabled InfiniBand. Journal of Parallel and Distributed Computing, 78(??):39–52, April 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515000052.

Guirado:2013:ETS

[GRR13] Fernando Guirado, Concepció Roig, and Ana Ripoll. Enhancing throughput for streaming applications running on cluster systems. Journal of Parallel and Distributed Computing, 73(8):1092–1105, August 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000671.

Goncharsky:2019:CCG

[GRS19] A. V. Goncharsky, S. Yu. Romanov, and S. Yu. Seryozhnikov. Comparison of the capabilities of GPU clusters and general-purpose supercomputers for solving 3D inverse problems of ultrasound tomography. Journal of Parallel and Distributed Computing, 133(??):77–92, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304514.

Guo:2018:SMI

[GRZ^{+}18] Hui Guo, Ju Ren, Deyu Zhang, Yaoxue Zhang, and Junying Hu. A scalable and manageable IoT architecture based on transparent computing. Journal of Parallel and Distributed Computing, 118 (part 1)(??):5–13, August 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0743731517302150.

Giacaman:2018:PSE

[GS18] Nasser Giacaman and Oliver Sinnen. Preparing the software engineer for a modern multi-core world. Journal of Parallel and Distributed Computing, 118 (part 1)(??):247–263, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0743731518301060.

Gumaei:2019:ASC

[GSASA19] Abdu Gumaei, Rachid Sammouda, Abdul Malik S. Al-Salman, and Ahmed Alsanad. Anti-spoofing cloud-based multi-spectral biometric identification system for enterprise security and privacy-preservation. Journal of Parallel and Distributed Computing, 124(??):27–40, February 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518307391.

Gracia-Tinedo:2012:SLT

[GTGLSA12] Raúl Gracia-Tinedo, Pedro García-López, and Marc Sánchez-Artigas. Sophia: a local trust system to secure key-based routing in non-deterministic DHTs. Journal of Parallel and Distributed Computing, 72(12):1696–1712, December 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001700.

Garg:2013:DAI

[GVBB13] Saurabh Kumar Garg, Srikumar Venugopal, James Broberg, and Rajkumar Buyya. Double auction-inspired metascheduling of parallel applications on global grids. Journal of Parallel and Distributed Computing, 73(4):450–464, April 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512002274.

Wu:2018:MME

[gWW18] Chu ge Wu and Ling Wang. A multi-model estimation of distribution algorithm for energy efficient scheduling under cloud computing system. Journal of Parallel and Distributed Computing, 117(??):63–72, July 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300868.

Gupta:2013:LPR

[GXYZ13] Saurabh Gupta, Ping Xiang, Yi Yang, and Huiyang Zhou. Locality principle revisited: a probability-based quantitative approach. Journal of Parallel and Distributed Computing, 73(7):1011–1027, July 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000117.

Garg:2011:ECS

[GYAB11] Saurabh Kumar Garg, Chee Shin Yeo, Arun Anandasivam, and Rajkumar Buyya. Environment-conscious scheduling of HPC applications on distributed cloud-oriented data centers. Journal of Parallel and Distributed Computing, 71(6):732–749, June 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731510000936.

Gong:2013:EEC

[GYP13] Dawei Gong, Yuanyuan Yang, and Zhexi Pan. Energy-efficient clustering in lossy wireless sensor networks. Journal of Parallel and Distributed Computing, 73(9):1323–1336, September 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513001123.

Gu:2014:SIM

[GYY^{+}14] Rong Gu, Xiaoliang Yang, Jinshuang Yan, Yuanhao Sun, Bing Wang, Chunfeng Yuan, and Yihua Huang. SHadoop: Improving MapReduce performance by optimizing job execution mechanism in Hadoop clusters. Journal of Parallel and Distributed Computing, 74(3):2166–2179, March 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002141.

Gan:2017:DAO

[GZG^{+}17] Zhihua Gan, Mingquan Zhang, Zhimin Gu, Hai Tan, and Jizan Zhang. Delay analysis and optimization for inter-core interference in real-time embedded multicore systems. Journal of Parallel and Distributed Computing, 103(??):77–86, May 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301794.

Gong:2014:DCA

[GZY14a] Dawei Gong, Miao Zhao, and Yuanyuan Yang. Distributed channel assignment algorithms for 802.11n WLANs with heterogeneous clients. Journal of Parallel and Distributed Computing, 74(5):2365–2379, May 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514000197.

Gong:2014:MCC

[GZY14b] Dawei Gong, Miao Zhao, and Yuanyuan Yang. A multichannel cooperative MIMO MAC protocol for clustered wireless sensor networks. Journal of Parallel and Distributed Computing, 74(11):3098–3114, November 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001348.

Hussain:2017:NIS

[HAC17] Zaid Hussain, Bader AlBdaiwi, and Anton Cerny. Nodeindependent spanning trees in Gaussian networks. Journal of Parallel and Distributed Computing, 109(??):324–332, November 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302071.

Hassan:2019:SSC

[HAC^{+}19] Mohammad Mehedi Hassan, Jemal Abawajy, Min Chen, Meikang Qiu, and Sheng Chen. Special section on Cloudof-Things and edge computing: Recent advances and future trends. Journal of Parallel and Distributed Computing, 133(??):170–173, November 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519305258.

Hidalgo:2016:EEC

[HASB16] Nicolas Hidalgo, Luciana Arantes, Pierre Sens, and Xavier Bonnaire. ECHO: Efficient complex query over DHT overlays.

Journal of Parallel and Distributed Computing, 88(??):31–45, February 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515001926.

Hung:2011:TAB

[HB11] Ka-Lok Hung and Brahim Bensaou. Throughput analysis and bandwidth allocation for IEEE 802.11 WLAN with hidden terminals. Journal of Parallel and Distributed Computing, 71(9):1201–1214, September 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731510002601.

Hussain:2015:EDH

[HBAD15] Zaid A. Hussain, Bella Bose, and Abdullah Al-Dhelaan. Edge disjoint Hamiltonian cycles in Eisenstein–Jacobi networks. Journal of Parallel and Distributed Computing, 86(??):62–70, December 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515001410.

Hendrikx:2015:RSS

[HBC15] Ferry Hendrikx, Kris Bubendorfer, and Ryan Chard. Reputation systems: a survey and taxonomy. Journal of Parallel and Distributed Computing, 75(??):184–197, January 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001464.

Hu:2012:PPI

[HBF12] Yi Hu, Laxmi N. Bhuyan, and Min Feng. Peer-to-peer indirect reciprocity via personal currency. Journal of Parallel and Distributed Computing, 72(8):1045–1054, August 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512001049.

Hager:2017:MCC

[HBS17] Sven Hager, Daniel Bendyk, and Björn Scheuermann. Matching circuits can be small: Partial evaluation and reconfiguration for FPGA-based packet processing. Journal of Parallel and Distributed Computing, 109(??):42–49, November 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517301557.

Haroune-Belkacem:2019:SES

[HBSASA19] Nassima Haroune-Belkacem, Fouzi Semchedine, Ahmed Al-Shammari, and Djamil Aissani. SMCA: an efficient SOAP messages compression and aggregation technique for improving web services performance. Journal of Parallel and Distributed Computing, 133(??):149–158, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304940.

Hsieh:2011:NSB

[HC11] Hui-Ching Hsieh and Mao-Lun Chiang. A new solution for the Byzantine agreement problem. Journal of Parallel and Distributed Computing, 71(10):1261–1277, October 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511001298.

Hammoud:2011:CAL

[HCM11] Mohammad Hammoud, Sangyeun Cho, and Rami Melhem. C-AMTE: a location mechanism for flexible cache management in chip multiprocessors. Journal of Parallel and Distributed Computing, 71(6):889–896, June 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731510002418.

Hernandez:2012:HBR

[HCR12] S. E. Pomares Hernandez, J. R. Perez Cruz, and M. Raynal. From the Happened-Before Relation to the Causal Ordered Set Abstraction. Journal of Parallel and Distributed Computing, 72(6):791–795, June 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512000597.

Hu:2010:IIC

[HD10] Zhi-Hua Hu and Yong-Sheng Ding. An immune inspired coevolutionary affinity network for prefetching of distributed object. Journal of Parallel and Distributed Computing, 70(2): 92–100, February 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hanna:2011:AHS

[HDCM11] Michel Hanna, Socrates Demetriades, Sangyeun Cho, and Rami Melhem. Advanced hashing schemes for packet forwarding using set associative memory architectures. Journal of Parallel and Distributed Computing, 71(1):1–15, January 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

HerondeCarvalhoJunior:2013:CSE

[HdR13] Francisco Heron de Carvalho Junior and Cenez Araújo de Rezende. A case study on expressiveness and performance of component-oriented parallel programming. Journal of Parallel and Distributed Computing, 73(5):557–569, May 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512002882.

Hajihashemi:2010:HPC

[HES10] Mohammad R. Hajihashemi and Magda El-Shenawee. High performance computing for the level-set reconstruction algorithm. Journal of Parallel and Distributed Computing, 70(6): 671–679, June 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hassan:2011:PID

[HES11] Ahmed M. Hassan and Magda El-Shenawee. Parallel implementation of the diffusion-drift algorithm for modeling the electrophysiological activity of breast tumors. Journal of Parallel and Distributed Computing, 71(7):1011–1023, July 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511000797.

Harmanci:2010:ETM

[HGFF10] Derin Harmanci, Vincent Gramoli, Pascal Felber, and Christof Fetzer. Extensible transactional memory testbed. Journal of Parallel and Distributed Computing, 70(10):1053–1067, October 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Huo:2019:TTA

[HGX^{+}19] Zhisheng Huo, Minyi Guo, Limin Xiao, Zhenxue He, Xiaoling Rong, and Bing Wei. TACD: a throughput allocation method based on variant of Cobb–Douglas for hybrid storage system. Journal of Parallel and Distributed Computing, 128(??):43– 56, June 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518306117.

Hussain:2014:PPM

[HHA14] Tassadaq Hussain, Amna Haider, and Eduard Ayguadé. PMSS: a programmable memory system and scheduler for complex memory patterns. Journal of Parallel and Distributed Computing, 74(10):2983–2993, October 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001075.

Hussin:2015:IRR

[HHK15] Masnida Hussin, Nor Asilah Wati Abdul Hamid, and Khairul Azhar Kasmiran. Improving reliability in resource management through adaptive reinforcement learning for distributed systems. Journal of Parallel and Distributed Computing, 75(??):93–100, January 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001919.

Heber:2012:ISC

[HHS12] Tomer Heber, Danny Hendler, and Adi Suissa. On the impact of serializing contention management on STM performance. Journal of Parallel and Distributed Computing, 72(6):739– 750, June 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512000433.

Huc:2012:ERS

[HJLR12] Florian Huc, Aubin Jarry, Pierre Leone, and José Rolim. On the efficiency of routing in sensor networks. Journal of Parallel and Distributed Computing, 72(7):889–901, July 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000664.

Hussain:2018:IDL

[HKK^{+}18] Shahid Hussain, Jacky Keung, Arif Ali Khan, Awais Ahmad, Salvatore Cuomo, Francesco Piccialli, Gwanggil Jeon, and Adnan Akhunzada. Implications of deep learning for the automation of design patterns organization. Journal of Parallel and Distributed Computing, 117(??):256–266, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302113.

Healy:2016:SSI

[HLBM16] Philip Healy, Theo Lynn, Enda Barrett, and John P. Morrison. Single system image: a survey. Journal of Parallel and Distributed Computing, 90–91(??):35–51, April 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516000058.

Huang:2019:OPA

[HLL^{+}19] Jing Huang, Yan Liu, Renfa Li, Keqin Li, JiYao An, Yang Bai, Fan Yang, and Guoqi Xie. Optimal power allocation and load balancing for non-dedicated heterogeneous distributed embedded computing systems. Journal of Parallel and Distributed Computing, 130(??):24–36, August 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519302357.

Hori:2012:ANS

[HLS12] Atsushi Hori, Jinpil Lee, and Mitsuhisa Sato. Audit: a new synchronization API for the GET/PUT protocol. Journal of Parallel and Distributed Computing, 72(11):1464–1470, November 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512000317.

Herrmann:2015:MAT

[HMR15] Julien Herrmann, Loris Marchal, and Yves Robert. Memoryaware tree traversals with pre-assigned tasks. Journal of Parallel and Distributed Computing, 75(??):53–66, January 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001944.

Huda:2018:MTD

[HMY^{+}18] Shamsul Huda, Suruz Miah, John Yearwood, Sultan Alyahya, Hmood Al-Dossari, and Robin Doss. A malicious threat detection model for cloud assisted Internet of Things (CoT) based industrial control system (ICS) networks using deep belief network. Journal of Parallel and Distributed Computing, 120(??):23–31, October 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518302442.

Hu:2019:HDI

[HN19] Linjia Hu and Saeid Nooshabadi. High-dimensional image descriptor matching using highly parallel KD-tree construction and approximate nearest neighbor search. Journal of Parallel and Distributed Computing, 132(??):127–140, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519304319.

Hollingsworth:2017:E

[Hol17] Jeff Hollingsworth. Editorial. Journal of Parallel and Distributed Computing, 106(??):194, August 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151730134X.

Han:2010:FIN

[HPB^{+}10] Liangxiu Han, Stephen Potter, George Beckett, Gavin Pringle, Stephen Welch, Sung-Han Koo, Gerhard Wickler, Asif Usmani, José L. Torero, and Austin Tate. FireGrid: An einfrastructure for next-generation emergency response support. Journal of Parallel and Distributed Computing, 70(11): 1128–1141, November 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hernandez:2011:CIP

[HRF^{+}11] C. Hernández, A. Roca, J. Flich, F. Silla, and J. Duato. Characterizing the impact of process variation on 45 nm NoC-based CMPs. Journal of Parallel and Distributed Computing, 71(5):651–663, May 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731510001930.

Hendry:2011:TDM

[HRG^{+}11] Gilbert Hendry, Eric Robinson, Vitaliy Gleyzer, Johnnie Chan, Luca P. Carloni, Nadya Bliss, and Keren Bergman. Time-division-multiplexed arbitration in silicon nanophotonic networks-on-chip for high-performance chip multiprocessors. Journal of Parallel and Distributed Computing, 71(5):641– 650, May 2011. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731510001966.

Hossain:2018:ECF

[HRH18] SK Alamgir Hossain, Md. Anisur Rahman, and M. Anwar Hossain. Edge computing framework for enabling situation awareness in IoT based smart city. Journal of Parallel and Distributed Computing, 122(??):226–237, December 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518306130.

Hossain:2017:CPC

[HRM17] M. Shamim Hossain, Md. Abdur Rahman, and Ghulam Muhammad. Cyber-physical cloud-oriented multi-sensory smart home framework for elderly people: an energy efficiency perspective. Journal of Parallel and Distributed Computing, 103(??):11–21, May 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151630123X.

He:2012:CSB

[HS12] Jun He and Hongchi Shi. Constructing sensor barriers with minimum cost in wireless sensor networks. Journal of Parallel and Distributed Computing, 72(12):1654–1663, December 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512001657.

Hussain:2017:HDE

[HS17] Zaid Hussain and Arash Shamaei. Higher dimensional Eisenstein–Jacobi networks. Journal of Parallel and Distributed Computing, 102(??):91–102, April 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516301617.

He:2010:IRS

[HSH10] Yuxiong He, Hongyang Sun, and Wen-Jing Hsu. Improved results for scheduling batched parallel jobs by using a generalized analysis framework. Journal of Parallel and Distributed Computing, 70(2):173–182, February 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Han:2010:EPA

[HSS10] Yijie Han, Sanjeev Saxena, and Xiaojun Shen. An efficient parallel algorithm for building the separating tree. Journal of Parallel and Distributed Computing, 70(6):625–629, June 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hundt:2017:SWA

[HSS17] Christian Hundt, Moritz Schlarb, and Bertil Schmidt. SAUCE: a web application for interactive teaching and learning of parallel programming. Journal of Parallel and Distributed Computing, 105(??):163–173, July 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300060.

Hendler:2010:SLF

[HSY10] Danny Hendler, Nir Shavit, and Lena Yerushalmi. A scalable lock-free stack algorithm. Journal of Parallel and Distributed Computing, 70(1):1–12, January 2010. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

He:2019:PEL

[HTB19] TianZhang He, Adel N. Toosi, and Rajkumar Buyya. Performance evaluation of live virtual machine migration in SDNenabled cloud data centers. Journal of Parallel and Distributed Computing, 131(??):55–68, September 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151830474X.

Hu:2011:DSR

[Hu11] Chia-Cheng Hu. Delay-sensitive routing in multi-rate MANETs. Journal of Parallel and Distributed Computing, 71 (1):53–61, January 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Huang:2017:RSB

[Hua17] Chun-Hsi Huang. REU Site: Bio-grid initiatives for interdisciplinary research and education. Journal of Parallel and Distributed Computing, 105(??):174–182, July 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300187.

Hussain:2017:NHS

[Hus17] Tassadaq Hussain. A novel hardware support for heterogeneous multi-core memory system. Journal of Parallel and Distributed Computing, 106(??):31–49, August 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151730076X.

Hu:2013:RAS

[HV13] Menglan Hu and Bharadwaj Veeravalli. Requirementaware strategies for scheduling real-time divisible loads on clusters. Journal of Parallel and Distributed Computing, 73(8):1083–1091, August 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000579.

Hsieh:2016:HTD

[HVW16] Cheng-Liang Hsieh, Lucas Vespa, and Ning Weng. A highthroughput DPI engine on GPU via algorithm/implementation co-optimization. Journal of Parallel and Distributed Computing, 88(??):46–56, February 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001938.

Hsieh:2018:KAR

[HWL18] Meng-Yen Hsieh, Tien-Hsiung Weng, and Kuan-Ching Li. A keyword-aware recommender system using implicit feedback on Hadoop. Journal of Parallel and Distributed Computing, 116(??):63–73, ???? 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517303398.

Hua:2014:ETH

[HWLR14] Xiayu Hua, Hao Wu, Zheng Li, and Shangping Ren. Enhancing throughput of the Hadoop distributed file system for interaction-intensive tasks. Journal of Parallel and Distributed Computing, 74(8):2770–2779, August 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000665.

Huang:2010:DBA

[HWY^{+}10] Xiaomeng Huang, Yongwei Wu, Guangwen Yang, Weiming Zheng, and Jinlei Jiang. Distributed bandwidth allocation based on alternating evolution algorithm. Journal of Parallel and Distributed Computing, 70(5):547–557, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hajisheykhi:2015:SFW

[HZA^{+}15] Reza Hajisheykhi, Ling Zhu, Mahesh Arumugam, Murat Demirbas, and Sandeep Kulkarni. “Slow is Fast” for wireless sensor networks in the presence of message losses. Journal of Parallel and Distributed Computing, 77(??):41–57, March 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514002214.

Huang:2012:ADT

[HZDP12] Ge Huang, Albert Y. Zomaya, Flávia C. Delicato, and Paulo F. Pires. An accurate on-demand time synchronization protocol for wireless sensor networks. Journal of Parallel and Distributed Computing, 72(10):1332–1346, October 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001426.

Huang:2018:NEK

[HZHS18] Kaixin Huang, Jie Zhou, Linpeng Huang, and Yanyan Shen. NVHT: an efficient key–value storage library for non-volatile memory. Journal of Parallel and Distributed Computing, 120(??):339–354, October 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151830090X.

Huang:2018:RMB

[HZL18] Kejie Huang, Rong Zhao, and Yong Lian. Racetrack Memory based hybrid Look-Up Table (LUT) for low power reconfigurable computing. Journal of Parallel and Distributed Computing, 117(??):127–137, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300960.

Hao:2019:AGB

[HZZ^{+}19] Meng Hao, Weizhe Zhang, You Zhang, Marc Snir, and Laurence T. Yang. Automatic generation of benchmarks for I/O-intensive parallel applications. Journal of Parallel and Distributed Computing, 124(??):1–13, February 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151830738X.

Igual:2012:FAD

[ICQO^{+}12] Francisco D. Igual, Ernie Chan, Enrique S. Quintana-Ortí, Gregorio Quintana-Ortí, Robert A. van de Geijn, and Field G. Van Zee. The FLAME approach: From dense linear algebra algorithms to high-performance multiaccelerator implementations. Journal of Parallel and Distributed Computing, 72(9):1134–1143, September 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511002139.

Ibrahim:2017:CSE

[IEWK17] Khaled Z. Ibrahim, Evgeny Epifanovsky, Samuel Williams, and Anna I. Krylov. Cross-scale efficient tensor contractions for coupled cluster computations through multiple programming model backends. Journal of Parallel and Distributed Computing, 106(??):92–105, August 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300783.

Ikuzawa:2016:RMU

[IIH16] Takuya Ikuzawa, Fumihiko Ino, and Kenichi Hagihara. Reducing memory usage by the lifting-based discrete wavelet transform with a unified buffer on a GPU. Journal of Parallel and Distributed Computing, 93–94(??):44–55, July 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516300016.

Irwansyah:2017:FBM

[IIH^{+}17] Arif Irwansyah, Omar W. Ibraheem, Jens Hagemeyer, Mario Porrmann, and Ulrich Rueckert. FPGA-based multi-robot tracking. Journal of Parallel and Distributed Computing, 107(??):146–161, September 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300965.

Imbs:2012:HWN

[IR12] Damien Imbs and Michel Raynal. Help when needed, but no more: Efficient read/write partial snapshot. Journal of Parallel and Distributed Computing, 72(1):1–12, January 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511001626.

Imbs:2016:RWS

[IRRS16] Damien Imbs, Sergio Rajsbaum, Michel Raynal, and Julien Stainer. Read/write shared memory abstraction on top of asynchronous Byzantine message-passing systems. Journal of Parallel and Distributed Computing, 93–94(??):1–9, July 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151630003X.

Imani:2010:RPC

[ISAZ10] N. Imani, H. Sarbazi-Azad, and A. Y. Zomaya. Resource placement in Cartesian product of networks. Journal of Parallel and Distributed Computing, 70(5):481–495, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ippoliti:2012:GAG

[IZ12] Dennis Ippoliti and Xiaobo Zhou. A-GHSOM: an adaptive growing hierarchical self organizing map for network anomaly detection. Journal of Parallel and Distributed Computing, 72(12):1576–1590, December 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002195.

Javadi:2012:FAR

[JAB12] Bahman Javadi, Jemal Abawajy, and Rajkumar Buyya. Failure-aware resource provisioning for hybrid cloud infrastructure. Journal of Parallel and Distributed Computing, 72(10):1318–1331, October 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001517.

Jesus:2015:FUF

[JBA15] Paulo Jesus, Carlos Baquero, and Paulo Sérgio Almeida. Flow updating: Fault-tolerant aggregation for dynamic networks. Journal of Parallel and Distributed Computing, 78(??):53–64, April 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515000416.

Jhumka:2014:EFT

[JBS14] Arshad Jhumka, Matthew Bradbury, and Sain Saginbekov. Efficient fault-tolerant collision-free data aggregation scheduling for wireless sensor networks. Journal of Parallel and Distributed Computing, 74(1):1789–1801, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002074.

Jha:2012:ODC

[JD12] Pranava K. Jha and Savitri Devisetty. Orthogonal drawings and crossing numbers of the Kronecker product of two cycles. Journal of Parallel and Distributed Computing, 72(2):195–204, February 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511002267.

Joselli:2015:NGN

[JdSJC^{+}15] Mark Joselli, José Ricardo da S. Junior, Esteban W. Clua, Anselmo Montenegro, Marcos Lage, and Paulo Pagliosa. Neighborhood grid: a novel data structure for fluids animation with GPU computing. Journal of Parallel and Distributed Computing, 75(??):20–28, January 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514002020.

Jabeen:2012:ASN

[JF12] Farhana Jabeen and Alvaro A. A. Fernandes. An algorithmic strategy for in-network distributed spatial analysis in wireless sensor networks. Journal of Parallel and Distributed Computing, 72(12):1628–1653, December 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002201.

Jodra:2017:SPE

[JGMY17] Jose L. Jodra, Ibai Gurrutxaga, Javier Muguerza, and Ainhoa Yera. Solving Poisson’s equation using FFT in a GPU cluster. Journal of Parallel and Distributed Computing, 102(??): 28–36, April 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301678.

Jha:2017:SRA

[JHF^{+}17] Sudhanshu Shekhar Jha, Wim Heirman, Ayose Falcón, Jordi Tubella, Antonio González, and Lieven Eeckhout. Shared resource aware scheduling on power-constrained tiled many-core processors. Journal of Parallel and Distributed Computing, 100(??):30–41, February 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516301186.

Jiang:2018:SRT

[JHL^{+}18] Xiaohui Jiang, Peng Hu, Yanchao Li, Chi Yuan, Isma Masood, Hamed Jelodar, Mahdi Rabbani, and Yongli Wang. A survey of real-time approximate nearest neighbor query over streaming data for fog computing. Journal of Parallel and Distributed Computing, 116(??):50–62, ???? 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300182.

Ji:2013:CDA

[JHPL13] Shouling Ji, Jing (Selena) He, Yi Pan, and Yingshu Li. Continuous data aggregation and capacity in probabilistic wireless sensor networks. Journal of Parallel and Distributed Computing, 73(6):729–745, June 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000300.

Jan:2012:SCA

[JJ12]  Yahya Jan and Lech Jóźwiak.  Scalable communica-

tion architectures for massively parallel hardware multiprocessors. Journal of Parallel and Distributed Computing, 72(11):1450–1463, November 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512000299.

Jayasekara:2015:WDC

[JKD^{+}15] Sachini Jayasekara, Sameera Kannangara, Tishan Dahanayakage, Isuru Ranawaka, Srinath Perera, and Vishaka Nanayakkara. Wihidum: Distributed complex event processing. Journal of Parallel and Distributed Computing, 79–80(??): 42–51, May 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515000519.

Javadi:2013:FTA

[JKIE13] Bahman Javadi, Derrick Kondo, Alexandru Iosup, and Dick Epema. The Failure Trace Archive: Enabling the comparison of failure measurements and models of distributed systems. Journal of Parallel and Distributed Computing, 73(8):1208–1223, August 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000634.

Jung:2015:CCD

[JKV15] Eun-Sung Jung, Rajkumar Kettimuthu, and Venkatram Vishwanath. Cluster-to-cluster data transfer with data compression over wide-area networks. Journal of Parallel and Distributed Computing, 79–80(??):90–103, May 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001713.

Jiang:2011:LST

[JL11] Yichuan Jiang and Zhaofeng Li. Locality-sensitive task allocation and load balancing in networked multiagent systems: Talent versus centrality. Journal of Parallel and Distributed Computing, 71(6):822–836, June 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000189.

Jiang:2011:CIC

[JLWX11] Zhen Jiang, Zhigang Li, Jie Wu, and Nong Xiao. Capability information: a cost-effective information model for multi-hop routing of wireless ad hoc networks in the real environment. Journal of Parallel and Distributed Computing, 71(8):1085–1097, August 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511000773.

Jang:2012:OWS

[JLY12] Ungjin Jang, Sunggu Lee, and Sungjoo Yoo. Optimal wake-up scheduling of data gathering trees for wireless sensor networks. Journal of Parallel and Distributed Computing, 72(4):536–546, April 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512000196.

Johnen:2014:SSS

[JM14] Colette Johnen and Fouzi Mekhaldi. Self-stabilizing with service guarantee construction of 1-hop weight-based bounded size clusters. Journal of Parallel and Distributed Computing, 74(1):1900–1913, January 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513002001.

Jhurani:2015:GII

[JM15] Chetan Jhurani and Paul Mullowney. A GEMM interface and implementation on NVIDIA GPUs for multiple small matrices. Journal of Parallel and Distributed Computing, 75(??):133–140, January 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151400166X.

Jallu:2017:DCC

[JPD17] Ramesh K. Jallu, Prajwal R. Prasad, and Gautam K. Das. Distributed construction of connected dominating set in unit disk graphs. Journal of Parallel and Distributed Computing, 104(??):159–166, June 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300424.

Jeannot:2012:OPR

[JST12] Emmanuel Jeannot, Erik Saule, and Denis Trystram. Optimizing performance and reliability on heterogeneous parallel systems: Approximation algorithms and heuristics. Journal of Parallel and Distributed Computing, 72(2):268–280, February 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151100219X.

Jie:2018:OTS

[JTC^{+}18] Yingmo Jie, Xinyu Tang, Kim-Kwang Raymond Choo, Shenghao Su, Mingchu Li, and Cheng Guo. Online task scheduling for edge computing based on repeated Stackelberg game. Journal of Parallel and Distributed Computing, 122(??):159–172, December 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518305409.

Jayasinghe:2011:TAM

[JTZZ11] Malith Jayasinghe, Zahir Tari, Panlop Zeephongsekul, and Albert Y. Zomaya. Task assignment in multiple server farms using preemptive migration and flow control. Journal of Parallel and Distributed Computing, 71(12):1608–1621, December 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511001365.

Jing:2017:MLD

[JWH^{+}17] Yuxin Jing, Hanpin Wang, Yu Huang, Lei Zhang, Jiang Xu, and Yongzhi Cao. A modeling language to describe massive data storage management in cyber-physical systems. Journal of Parallel and Distributed Computing, 103(??):113–120, May 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301848.

Jiang:2014:FRS

[JWSG14] Guiyuan Jiang, Jigang Wu, Jizhou Sun, and Yiyi Gao. Flexible rerouting schemes for reconfiguration of multiprocessor arrays. Journal of Parallel and Distributed Computing, 74(10):3026–3036, October 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001117.

Jiang:2019:HMW

[JXZ^{+}19] Wanchun Jiang, Haiming Xie, Xiangqian Zhou, Liyuan Fang, and Jianxin Wang. Haste makes waste: the on–off algorithm for replica selection in key–value stores. Journal of Parallel and Distributed Computing, 130(??):80–90, August 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519302333.

Jin:2015:SEM

[JZF^{+}15] Xibo Jin, Fa Zhang, Liya Fan, Ying Song, and Zhiyong Liu. Scheduling for energy minimization on restricted parallel processors. Journal of Parallel and Distributed Computing, 81–82(??):36–46, July 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515000647.

Jiang:2017:DOS

[JZZ^{+}17] Wei Jiang, Xia Zhang, Jinyu Zhan, Yue Ma, and Ke Jiang. Design optimization of secure message communication for energy-constrained distributed real-time systems. Journal of Parallel and Distributed Computing, 100(??):1–15, February 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301228.

Khalid:2019:TML

[KAA^{+}19a] Yasir Noman Khalid, Muhammad Aleem, Usman Ahmed, Muhammad Arshad Islam, and Muhammad Azhar Iqbal. Troodon: a machine-learning based load-balancing application scheduler for CPU–GPU system. Journal of Parallel and Distributed Computing, 132(??):79–94, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306841.

Khan:2019:TAP

[KAA^{+}19b] Tanveer Khan, Masoom Alam, Adnan Akhunzada, Ali Hur, Muhammad Asif, and Muhammad Khurram Khan. Towards augmented proactive cyberthreat intelligence. Journal of Parallel and Distributed Computing, 124(??):47–59, February 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518307408.

Kakugawa:2015:MIA

[Kak15] Hirotsugu Kakugawa. Mutual inclusion in asynchronous message-passing distributed systems. Journal of Parallel and Distributed Computing, 77(??):95–104, March 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515000040.

Karaata:2019:OMS

[Kar19] Mehmet Hakan Karaata. An open modal stabilizing mwave algorithm for arbitrary networks. Journal of Parallel and Distributed Computing, 128(??):15–29, June 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300905.

Kravtsov:2010:SFL

[KBC^{+}10] Valentin Kravtsov, Pavel Bar, David Carmeli, Assaf Schuster, and Martin Swain. A scheduling framework for large-scale, parallel, and topology-aware applications. Journal of Parallel and Distributed Computing, 70(9):983–992, September 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kaur:2019:IRE

[KBC19] Gurleen Kaur, Anju Bala, and Inderveer Chana. An intelligent regressive ensemble approach for predicting resource usage in cloud computing. Journal of Parallel and Distributed Computing, 123(??):1–12, January 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306063.

Khelaifa:2019:CAA

[KBK^{+}19] Abdennacer Khelaifa, Saber Benharzallah, Laid Kahloul, Reinhardt Euler, Abdelkader Laouid, and Ahcène Bounceur. A comparative analysis of adaptive consistency approaches in cloud storage. Journal of Parallel and Distributed Computing, 129(??):36–49, July 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518301795.

Konstantinidis:2017:QRM

[KC17] Elias Konstantinidis and Yiannis Cotronis. A quantitative roofline model for GPU kernel performance estimation using micro-benchmarks and hardware metric profiling. Journal of Parallel and Distributed Computing, 107(??):37–56, September 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517301247.

Kozik:2018:SDM

[KCFP18] Rafal Kozik, Michal Choraś, Massimo Ficco, and Francesco Palmieri. A scalable distributed machine learning approach for attack detection in edge computing environments. Journal of Parallel and Distributed Computing, 119(??):18–26, September 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518302004.

Kim:2019:GBA

[KCP19] HyungGyoon Kim, Hyungmin Cho, and Changwoo Pyo. GPUbased acceleration of the Linear Complexity Test for random number generator testing. Journal of Parallel and Distributed Computing, 128(??):115–125, June 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300784.

Kim:2014:IPA

[KCR14] Jinoh Kim, Jerry Chou, and Doron Rotem. iPACS: Poweraware covering sets for energy proportionality and performance in data parallel computing clusters. Journal of Parallel and Distributed Computing, 74(1):1762–1774, January 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513002025.

Kalathingal:2018:DDI

[KCSS18] Sajith Kalathingal, Sylvain Collange, Bharath N. Swamy, and André Seznec. DITVA: Dynamic Inter-Thread Vectorization Architecture. Journal of Parallel and Distributed Computing, 120(??):267–281, October 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517303040.

Karonis:2013:DHA

[KDO^{+}13] Nicholas T. Karonis, Kirk L. Duffin, Caesar E. Ordoñez, Bela Erdelyi, Thomas D. Uram, Eric C. Olson, George Coutrakon, and Michael E. Papka. Distributed and hardware accelerated computing for clinical medical imaging using proton computed tomography (pCT). Journal of Parallel and Distributed Computing, 73(12):1605–1612, December 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513001548.

Kumar:2018:IDC

[KDSS18] Santosh Kumar, Deepanwita Datta, Sanjay Kumar Singh, and Arun Kumar Sangaiah. An intelligent decision computing paradigm for crowd monitoring in the smart city. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):344– 358, August 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517300825.

Karmakar:2010:ABF

[KG10] Sushanta Karmakar and Arobinda Gupta. Adaptive broadcast by fault-tolerant spanning tree switching. Journal of Parallel and Distributed Computing, 70(9):889–906, September 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kalms:2019:SCM

[KG19] Lester Kalms and Diana Göhringer. Scalable clustering and mapping algorithm for application distribution on heterogeneous and irregular FPGA clusters. Journal of Parallel and Distributed Computing, 133(??):367–376, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518301199.

Kuntz:2011:VAA

[KGN11] Romain Kuntz, Antoine Gallais, and Thomas Noël. From versatility to auto-adaptation of the medium access control in wireless sensor networks. Journal of Parallel and Distributed Computing, 71(9):1236–1248, September 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731510001991.

Koohi:2012:SAC

[KH12] Somayyeh Koohi and Shaahin Hessabi. Scalable architecture for a contention-free optical network on-chip. Journal of Parallel and Distributed Computing, 72(11):1493–1506, November 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000378.

Kim:2018:MHP

[KHK18] Seong Hoon Kim, Minkeun Ha, and Daeyoung Kim. A multi-hop pointer forwarding scheme for efficient location update in low-rate wireless mesh networks. Journal of Parallel and Distributed Computing, 122(??):109–121, December 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518304969.

Kanemitsu:2017:PNS

[KHN17] Hidehiro Kanemitsu, Masaki Hanada, and Hidenori Nakazato. Prior node selection for scheduling workflows in a heterogeneous system. Journal of Parallel and Distributed Computing, 109(??):155–177, November 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301946.

Kannan:2014:HSC

[KHT^{+}14] R. Kannan, V. Harrand, X. G. Tan, H. Q. Yang, and A. J. Przekwas. Highly scalable computational algorithms on emerging parallel machine multicore architectures II: Development and implementation in the CSD and FSI contexts. Journal of Parallel and Distributed Computing, 74(9):2808–2817, September 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514000896.

Kang:2013:EIG

[KHW13] Qinma Kang, Hong He, and Jun Wei. An effective iterated greedy algorithm for reliability-oriented task allocation in distributed computing systems. Journal of Parallel and Distributed Computing, 73(8):1106–1115, August 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151300052X.

Kusudo:2015:BPA

[KIH15] Ko Kusudo, Fumihiko Ino, and Kenichi Hagihara. A bitparallel algorithm for searching multiple patterns with various lengths. Journal of Parallel and Distributed Computing, 76(??):49–57, February 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151400210X.

Kim:2011:DCA

[Kim11] Kyungjun Kim. A distributed channel assignment control for QoS support in mobile ad hoc networks. Journal of Parallel and Distributed Computing, 71(3):335–342, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kim:2017:QPT

[Kim17] Seungku Kim. QoS provisioning of a task-scheduling algorithm for lightweight devices. Journal of Parallel and Distributed Computing, 107(??):67–75, September 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301326.

Kong:2010:EWP

[KK10] Zhen Kong and Yu-Kwong Kwok. Efficient wireless packet scheduling in a non-cooperative environment: Game theoretic analysis and algorithms. Journal of Parallel and Distributed Computing, 70(8):790–799, August 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kim:2011:LAL

[KK11] Hyunhee Kim and Jihong Kim. A leakage-aware L2 cache management technique for producer-consumer sharing in lowpower chip multiprocessors. Journal of Parallel and Distributed Computing, 71(12):1545–1557, December 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001638.

Kennedy:2017:CNB

[KK17] Matthew Kennedy and Avinash Karanth Kodi. CLAP-NET: Bandwidth adaptive optical crossbar architecture. Journal of Parallel and Distributed Computing, 100(??):130–139, February 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151630034X.

Karunaratne:2017:DSC

[KKH17] Pasan Karunaratne, Shanika Karunasekera, and Aaron Harwood. Distributed stream clustering using micro-clusters on Apache Storm. Journal of Parallel and Distributed Computing, 108(??):74–84, October 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300697.

Kas:2011:OAC

[KKK11a] Miray Kas, Ibrahim Korpeoglu, and Ezhan Karasan. OLSRaware channel access scheduling in wireless mesh networks. Journal of Parallel and Distributed Computing, 71(9):1225– 1235, September 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731510002625.

Krimer:2011:STA

[KKK^{+}11b] Evgeni Krimer, Isaac Keslassy, Avinoam Kolodny, Isask’har Walter, and Mattan Erez. Static timing analysis for modeling QoS in networks-on-chip. Journal of Parallel and Distributed Computing, 71(5):687–699, May 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731510002017.

Kambatla:2014:TBD

[KKKG14] Karthik Kambatla, Giorgos Kollias, Vipin Kumar, and Ananth Grama. Trends in big data analytics. Journal of Parallel and Distributed Computing, 74(7):2561–2573, July 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000057.

Kim:2012:RSG

[KKKP12] Seong Hoon Kim, Daeyoung Kim, Jeong-Seok Kang, and Hong Seong Park. A reflective service gateway for integrating evolvable sensor-actuator networks with pervasive infrastructure. Journal of Parallel and Distributed Computing, 72(10):1237–1253, October 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001505.

Kim:2014:GBM

[KKLJ14] Sangwook Kim, Hwanju Kim, Joonwon Lee, and Jinkyu Jeong. Group-based memory oversubscription for virtualized clouds. Journal of Parallel and Distributed Computing, 74(4):2241– 2256, April 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514000033.

Kim:2013:PMD

[KKN13] Jinwoong Kim, Sul-Gi Kim, and Beomseok Nam. Parallel multi-dimensional range query processing with Rtrees on GPU. Journal of Parallel and Distributed Computing, 73(8):1195–1207, August 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000592.

Kalinnik:2014:OAT

[KKR14] Natalia Kalinnik, Matthias Korch, and Thomas Rauber. Online auto-tuning for the time-step-based parallel solution of ODEs on shared-memory systems. Journal of Parallel and Distributed Computing, 74(8):2722–2744, August 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000628.

Kim:2012:PDM

[KKS^{+}12] Jinwoo Kim, Minyoung Kim, Mark-Oliver Stehr, Hyunok Oh, and Soonhoi Ha. A parallel and distributed meta-heuristic framework based on partially ordered knowledge sharing. Journal of Parallel and Distributed Computing, 72(4):564– 578, April 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512000184.

Kumarage:2013:DAD

[KKTZ13] Heshan Kumarage, Ibrahim Khalil, Zahir Tari, and Albert Zomaya. Distributed anomaly detection for industrial wireless sensor networks based on fuzzy data modelling. Journal of Parallel and Distributed Computing, 73(6):790–806, June 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000294.

Kobus:2017:RRT

[KKW17] Tadeusz Kobus, Maciej Kokociński, and Paweł T. Wojciechowski. Relaxing real-time order in opacity and linearizability. Journal of Parallel and Distributed Computing, 100(??):57–70, February 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516301253.

Kim:2011:TBS

[KLJ^{+}11] Hwanju Kim, Hyeontaek Lim, Jinkyu Jeong, Heeseung Jo, Joonwon Lee, and Seungryoul Maeng. Transparently bridging semantic gap in CPU management for virtualized environments. Journal of Parallel and Distributed Computing, 71(6):758–773, June 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731510002376.

Kautonen:2010:TPR

[KLP10] Anssi Kautonen, Ville Leppänen, and Martti Penttonen. Thinning protocols for routing h-relations over shared media. Journal of Parallel and Distributed Computing, 70(8):783–789, August 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kaur:2017:RPW

[KM17] Parmeet Kaur and Shikha Mehta. Resource provisioning and work flow scheduling in clouds using augmented Shuffled Frog Leaping Algorithm. Journal of Parallel and Distributed Computing, 101(??):41–50, March 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301460.

Khanna:2016:NDL

[KMC16] Gaurav Khanna, Rajesh Mishra, and S. K. Chaturvedi. 4DGIN-3: a new design layout of 4-disjoint gamma interconnection network. Journal of Parallel and Distributed Computing, 98(??):40–47, December 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300946.

Kowalski:2010:ESM

[KMS10] Dariusz R. Kowalski, Mariam Momenzadeh, and Alexander A. Shvartsman. Emulating shared-memory Do-All algorithms in asynchronous message-passing systems. Journal of Parallel and Distributed Computing, 70(6):699–705, June 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kim:2018:CPH

[KN18a] Jinwoong Kim and Beomseok Nam. Co-processing heterogeneous parallel index for multi-dimensional datasets. Journal of Parallel and Distributed Computing, 113(??):195–203, March 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517302915.

Kim:2018:CCP

[KN18b] Jinwoong Kim and Beomseok Nam. Corrigendum to “Coprocessing heterogeneous parallel index for multi-dimensional datasets” [J. Parallel Distrib. Comput. 113 (2018) 195–203]. Journal of Parallel and Distributed Computing, 117(??):17, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518300522.

Kim:2018:FSS

[KNHH18] Myungsun Kim, Soonhyun Noh, Jinhwa Hyeon, and Seongsoo Hong. Fair-share scheduling in single-ISA asymmetric multicore architecture via scaled virtual runtime and load redistribution. Journal of Parallel and Distributed Computing, 111(??):174–186, January 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302423.

Khedr:2011:ETT

[KO11] Ahmed M. Khedr and Walid Osamy. Effective target tracking mechanism in a self-organizing wireless sensor network. Journal of Parallel and Distributed Computing, 71(10):1318–1326, October 2011. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511001183.

Khedr:2012:MAM

[KO12] Ahmed M. Khedr and Walid Osamy. Mobility-assisted minimum connected cover in a wireless sensor network. Journal of Parallel and Distributed Computing, 72(7):827–837, July 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000846.

Kolomvatsos:2019:ESA

[Kol19] Kostas Kolomvatsos. An efficient scheme for applying software updates in pervasive computing applications. Journal of Parallel and Distributed Computing, 128(??):1–14, June 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519300772.

Kaczmarski:2017:FLL

[KP17] Krzysztof Kaczmarski and Piotr Przymus. Fixed length lightweight compression for GPU revised. Journal of Parallel and Distributed Computing, 107(??):19–36, September 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300990.

Kang:2010:DSA

[KR10a] Jaeyeon Kang and Sanjay Ranka. Dynamic slack allocation algorithms for energy minimization on parallel machines. Journal of Parallel and Distributed Computing, 70(5):417–430, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kang:2010:SAA

[KR10b] Jaeyeon Kang and Sanjay Ranka. Slack allocation algorithm for parallel machines. Journal of Parallel and Distributed Computing, 70(1):23–34, January 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kundeti:2011:ECS

[KR11] Vamsi Kundeti and Sanguthevar Rajasekaran. Efficient outof-core sorting algorithms for the Parallel Disks Model. Journal of Parallel and Distributed Computing, 71(11):1427–1433, November 2011. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511001390.

Kim:2012:FEP

[KR12] Jinoh Kim and Doron Rotem. FREP: Energy proportionality for disk storage using replication. Journal of Parallel and Distributed Computing, 72(8):960–974, August 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000858.

Kuznetsov:2017:GGB

[KR17] Petr Kuznetsov and Srivatsan Ravi. Grasping the gap between blocking and non-blocking transactional memories. Journal of Parallel and Distributed Computing, 101(??):1–16, March 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301265.

Kalentev:2011:CCL

[KRKS11] Oleksandr Kalentev, Abha Rai, Stefan Kemnitz, and Ralf Schneider. Connected component labeling on a 2D grid using CUDA. Journal of Parallel and Distributed Computing, 71 (4):615–620, April 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Khasanvis:2014:HGC

[KRM14] Santosh Khasanvis, Mostafizur Rahman, and Csaba Andras Moritz. Heterogeneous graphene-CMOS ternary content addressable memory. Journal of Parallel and Distributed Computing, 74(6):2497–2503, June 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151300138X.

Krishnamoorthy:2013:SIJ

[KRS13] Sriram Krishnamoorthy, J. Ramanujam, and P. Sadayappan. A special issue of Journal of Parallel and Distributed Computing: Domain-specific languages and high-level frameworks for high-performance computing. Journal of Parallel and Distributed Computing, 73(6):895, June 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151300066X.

Krishnamoorthy:2014:IJS

[KRS14] Sriram Krishnamoorthy, J. Ramanujam, and P. Sadayappan. Introduction to the JPDC Special Issue on Domain-Specific Languages and High-Level Frameworks for High-Performance Computing. Journal of Parallel and Distributed Computing, 74(12):3175, December 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001749.

Konwar:2015:RNS

[KRS15] Kishori M. Konwar, Sanguthevar Rajasekaran, and Alexander A. Shvartsman. Robust network supercomputing with unreliable workers. Journal of Parallel and Distributed Computing, 75(??):81–92, January 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001920.

Kshemkalyani:2013:EDS

[KS13] Ajay D. Kshemkalyani and Mukesh Singhal. Efficient distributed snapshots in an anonymous asynchronous messagepassing system. Journal of Parallel and Distributed Computing, 73(5):621–629, May 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373151300004X.

Kesavaraja:2018:QEC

[KS18] D. Kesavaraja and A. Shenbagavalli. QoE enhancement in cloud virtual machine allocation using Eagle strategy of hybrid krill herd optimization. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):267–279, August 2018.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302459.

Kissel:2011:PSH

[KSB11] Ezra Kissel, Martin Swany, and Aaron Brown. Phoebus: a system for high throughput data movement. Journal of Parallel and Distributed Computing, 71(2):266–279, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kegel:2013:DTU

[KSG13] Philipp Kegel, Michel Steuwer, and Sergei Gorlatch. dOpenCL: Towards uniform programming of distributed heterogeneous multi-/many-core systems. Journal of Parallel and Distributed Computing, 73(12):1639–1648, December 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001597.

Kshemkalyani:2012:IDP

[Ksh12] Ajay D. Kshemkalyani. Immediate detection of predicates in pervasive environments. Journal of Parallel and Distributed Computing, 72(2):219–230, February 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001766.

Khalid:2017:PSM

[KSJC17] Shehzad Khalid, Bushra Sabir, Sohail Jabbar, and Naveen Chilamkurti. Precise shape matching of large shape datasets using hybrid approach. Journal of Parallel and Distributed Computing, 110(??):16–30, December 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301260.

Kim:2015:DFD

[KSK15] Taehong Kim, Seog Chung Seo, and Daeyoung Kim. Distributed formation of degree constrained minimum routing cost tree in wireless ad-hoc networks. Journal of Parallel and Distributed Computing, 83(??):143–158, September 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515000969.

Kollias:2014:FPA

[KSSG14] Giorgos Kollias, Madan Sathe, Olaf Schenk, and Ananth Grama. Fast parallel algorithms for graph similarity and matching. Journal of Parallel and Distributed Computing, 74 (5):2400–2410, May 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731513002529.

Koutsandria:2016:CEH

[KSSK16] Georgia Koutsandria, Emmanouil Skevakis, Amir A. Sayegh, and Polychronis Koutsakis. Can everybody be happy in the cloud? Delay, profit and energy-efficient scheduling for cloud services. Journal of Parallel and Distributed Computing, 96(??):202–217, October 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300491.

Kamath:2016:DTT

[KSSL16] Goutham Kamath, Lei Shi, Wen-Zhan Song, and Jonathan Lees. Distributed travel-time seismic tomography in largescale sensor networks. Journal of Parallel and Distributed Computing, 89(??):50–64, March 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515002117.

K:2017:ESI

[KTP17] Ashwin Kumar T. K., Johnson P. Thomas, and Saikiran Parepally. An efficient and secure information retrieval framework for content centric networks. Journal of Parallel and Distributed Computing, 104(??):223–233, June 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300436.

Kubica:2017:PBC

[Kub17] Bartlomiej Jacek Kubica. Parallelization of a boundconsistency enforcing procedure and its application in solving nonlinear systems. Journal of Parallel and Distributed Computing, 107(??):57–66, September 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300977.

Kumar:2017:ROT

[Kum17] Sandeep Kumar. Research-oriented teaching of PDC topics in integration with other undergraduate courses at multiple levels: a multi-year report. Journal of Parallel and Distributed Computing, 105(??):92–104, July 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300102.

Khoo:2010:PAF

[KV10] B. T. Benjamin Khoo and Bharadwaj Veeravalli. Pro-active failure handling mechanisms for scheduling in grid computing environments. Journal of Parallel and Distributed Computing, 70(3):189–200, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kang:2018:DSS

[KVA18] Seungmin Kang, Bharadwaj Veeravalli, and Khin Mi Mi Aung. Dynamic scheduling strategy with efficient node availability prediction for handling divisible loads in multi-cloud systems. Journal of Parallel and Distributed Computing, 113(??):1–16, March 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517302824.

Kashyap:2017:ASM

[KVNV17] Amlesh Kashyap, Sathish S. Vadhiyar, Ravi S. Nanjundiah, and P. N. Vinayachandran. Asynchronous and synchronous models of executions on Intel(R) Xeon Phi^{T M} coprocessor systems for high performance of long wave radiation calculations in atmosphere models. Journal of Parallel and Distributed Computing, 102(??):199–212, April 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516302040.

Khairy:2019:SAA

[KWZ19] Mahmoud Khairy, Amr G. Wassal, and Mohamed Zahran. A survey of architectural approaches for improving GPGPU performance, programmability and heterogeneity. Journal of Parallel and Distributed Computing, 127(??):65–88, May 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518308669.

Kee:2017:AMB

[KyLPC17] Minkwan Kee, Hong yeol Lim, Gi-Ho Park, and Sangyeun Cho. An analytical model based on performance demand of workload for energy-efficient heterogeneous multicore systems. Journal of Parallel and Distributed Computing, 100(??):172–180, February 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151630048X.

Karimi:2013:SAF

[KYS13] Hamid Karimi, Saleh Yousefi, and Maghsud Solimanpur. A segmentation approach for file broadcast scheduling. Journal of Parallel and Distributed Computing, 73(10):1375–1388, October 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513001184.

Kolli:2011:DPA

[KZ11] Sandeep Kolli and Maciej Zawodniok. A dynamic programming approach: Improving the performance of wireless networks. Journal of Parallel and Distributed Computing, 71(11):1447–1459, November 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001419.

Louati:2018:LTH

[LAC18] Thouraya Louati, Heithem Abbes, and Christophe Cérin. LX-CloudFT: Towards high availability, fault tolerant cloud system based Linux containers. Journal of Parallel and Distributed Computing, 122(??):51–69, December 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305136.

Louati:2018:LCT

[LACJ18] Thouraya Louati, Heithem Abbes, Christophe Cérin, and Mohamed Jemni. LXCloud-CR: Towards LinuX containers distributed hash table based checkpoint-restart. Journal of Parallel and Distributed Computing, 111(??):187–205, January 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302411.

Lai:2014:ECO

[Lai14] Cheng-Nan Lai. An efficient construction of one-to-many node-disjoint paths in folded hypercubes. Journal of Parallel and Distributed Computing, 74(4):2310–2316, April 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002396.

Lai:2015:CAS

[Lai15] Cheng-Nan Lai. Constructing all shortest node-disjoint paths in torus networks. Journal of Parallel and Distributed Computing, 75(??):123–132, January 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001671.

Lai:2017:OCN

[Lai17] Cheng-Nan Lai. Optimal construction of node-disjoint shortest paths in folded hypercubes. Journal of Parallel and Distributed Computing, 102(??):37–41, April 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301770.

Lee:2010:SKC

[LAK10] Manhee Lee, Baik Song An, and Eun Jung Kim. A session key caching and prefetching scheme for secure communication in cluster systems. Journal of Parallel and Distributed Computing, 70(7):732–742, July 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lastovetsky:2012:SIJ

[Las12] Alexey Lastovetsky. Special issue of Journal of Parallel and Distributed Computing: Heterogeneity in parallel and distributed computing. Journal of Parallel and Distributed Computing, 72(10):1397, October 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001414.

Lastovetsky:2013:HPD

[Las13] Alexey Lastovetsky. Heterogeneity in parallel and distributed computing. Journal of Parallel and Distributed Computing, 73(12):1523–1524, December 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513001652.

Li:2019:ICU

[LAS^{+}19] Yinhao Li, Awatif Alqahtani, Ellis Solaiman, Charith Perera, Prem Prakash Jayaraman, Rajkumar Buyya, Graham Morgan, and Rajiv Ranjan. IoT-CANE: a unified knowledge management system for data-centric Internet of Things application systems. Journal of Parallel and Distributed Computing, 131(??):161–172, September 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518309742.

Lejeune:2015:FSF

[LASS15] Jonathan Lejeune, Luciana Arantes, Julien Sopena, and Pierre Sens. A fair starvation-free prioritized mutual exclusion algorithm for distributed systems. Journal of Parallel and Distributed Computing, 83(??):13–29, September 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515000659.

Liao:2012:APP

[LB12] Guangdeng Liao and Laxmi Bhuyan. Analyzing performance and power efficiency of network processing over 10 GbE. Journal of Parallel and Distributed Computing, 72(11):1442–1449, November 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512000603.

Lopez:2017:CCC

[LB17] Pedro López and Elvira Baydal. On a course on computer cluster configuration and administration. Journal of Parallel and Distributed Computing, 105(??):127–137, July 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300151.

Lopez:2018:THP

[LB18] Pedro López and Elvira Baydal. Teaching high-performance service in a cluster computing course. Journal of Parallel and Distributed Computing, 117(??):138–147, July 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518301059.

Lakhlef:2015:EAP

[LBMG15] Hicham Lakhlef, Julien Bourgeois, Hakim Mabed, and Seth Copen Goldstein. Energy-aware parallel self-reconfiguration for chains [of] microrobot networks. Journal of Parallel and Distributed Computing, 75(??):67–80, January 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001932.

Li:2019:JOD

[LBT19] Chunlin Li, Jingpan Bai, and JianHang Tang. Joint optimization of data placement and scheduling for improving user experience in edge computing. Journal of Parallel and Distributed Computing, 125(??):93–105, March 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518302661.

Li:2011:APU

[LC11] Yinan Li and Ing-Ray Chen. Adaptive per-user perobject cache consistency management for mobile data access in wireless mesh networks. Journal of Parallel and Distributed Computing, 71(7):1034–1046, July 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000542.

Lee:2013:AMS

[LC13] Kiyeon Lee and Sangyeun Cho. Accurately modeling superscalar processor performance with reduced trace. Journal of Parallel and Distributed Computing, 73(4):509–521, April 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512002833.

Laurenciu:2014:CTN

[LC14a] N. Cucu Laurenciu and S. D. Cotofana. Critical transistors nexus based circuit-level aging assessment and prediction. Journal of Parallel and Distributed Computing, 74(6):2512– 2520, June 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513001421.

Lim:2014:PGM

[LC14b] L. Lim and D. Conan. Partitionable group membership for mobile ad hoc networks. Journal of Parallel and Distributed Computing, 74(8):2708–2721, August 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000598.

Lorenzon:2016:IDG

[LCB16] Arthur Francisco Lorenzon, Márcia Cristina Cera, and Antonio Carlos Schneider Beck. Investigating different generalpurpose and embedded multicores to achieve optimal tradeoffs between performance and energy. Journal of Parallel and Distributed Computing, 95(??):107–123, September 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516300090.

Li:2010:SSP

[LCCL10] Zhenhua Li, Jiannong Cao, Guihai Chen, and Yan Liu. On the source switching problem of peer-to-peer streaming. Journal of Parallel and Distributed Computing, 70(5):537–546, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Liu:2018:ETS

[LCJ^{+}18] Song Liu, Yuanzhen Cui, Qing Jiang, Qian Wang, and Weiguo Wu. An efficient tile size selection model based on machine learning. Journal of Parallel and Distributed Computing, 121(??):1–14, November 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151830426X.

Li:2014:EAR

[LDP^{+}14] Wei Li, Flávia C. Delicato, Paulo F. Pires, Young Choon Lee, Albert Y. Zomaya, Claudio Miceli, and Luci Pirmez. Efficient allocation of resources in multiple heterogeneous Wireless Sensor Networks. Journal of Parallel and Distributed Computing, 74(1):1775–1788, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002104.

Lima:2019:BSS

[LdPLC^{+}19] Henrique D. Lima, Luiz A. de P. Lima, Alcides Calsavara, Henri F. Eberspächer, Ricardo C. Nabhen, and Elias P. Duarte. Beyond scalability: Swarm intelligence affected by magnetic fields in distributed tuple spaces. Journal of Parallel and Distributed Computing, 123(??):90–99, January 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518306488.

Liu:2016:AFA

[LDS16] Wei Liu, Enqing Dong, and Yang Song. Analysis of flip ambiguity for robust three-dimensional node localization in wireless sensor networks. Journal of Parallel and Distributed Computing, 97(??):57–68, November 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300776.

Lemeire:2018:EAM

[LdSB^{+}18] Jan Lemeire, Bruno da Silva, An Braeken, Jan G. Cornelis, and Abdellah Touhafi. Efficiency analysis methodology of FPGAs based on lost frequencies, area and cycles. Journal of Parallel and Distributed Computing, 113(??):204–217, March 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517303246.

Le:2014:LLA

[LDZ^{+}14] Duc Tai Le, Thang Le Duc, Vyacheslav V. Zalyubovskiy, Dongsoo S. Kim, and Hyunseung Choo. LABS: Latency aware broadcast scheduling in uncoordinated Duty-Cycled Wireless Sensor Networks. Journal of Parallel and Distributed Computing, 74(11):3141–3152, November 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001336.

Le:2017:CTB

[LDZ^{+}17] Duc Tai Le, Thang Le Duc, Vyacheslav V. Zalyubovskiy, Dongsoo S. Kim, and Hyunseung Choo. Collision-tolerant broadcast scheduling in duty-cycled wireless sensor networks. Journal of Parallel and Distributed Computing, 100(??):42–56, February 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301241.

Loeb:2019:AHC

[LE19] Andrew Loeb and Christopher Earls. Analysis of heterogeneous computing approaches to simulating heat transfer in heterogeneous material. Journal of Parallel and Distributed Computing, 133(??):1–17, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304423.

Luo:2019:GBP

[LFEP19] Jia Luo, Shigeru Fujimura, Didier El Baz, and Bastien Plazolles. GPU based parallel genetic algorithm for solving an energy efficient dynamic flexible flow shop scheduling problem. Journal of Parallel and Distributed Computing, 133(??):244–257, November 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518305628.

Losada:2017:PAF

[LFGM17] Nuria Losada, Basilio B. Fraguela, Patricia González, and María J. Martín. A portable and adaptable fault tolerance solution for heterogeneous applications. Journal of Parallel and Distributed Computing, 104(??):146–158, June 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300394.

Lampka:2016:KIC

[LFS16] Kai Lampka, Björn Forsberg, and Vasileios Spiliopoulos. Keep it cool and in time: With runtime monitoring to thermal-aware execution speeds for deadline constrained systems. Journal of Parallel and Distributed Computing, 95(??):79–91, September 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151600023X.

Liu:2017:FTE

[LFZ^{+}17] Zhao Liu, Jianxi Fan, Jingya Zhou, Baolei Cheng, and Xiaohua Jia. Fault-tolerant embedding of complete binary trees in locally twisted cubes. Journal of Parallel and Distributed Computing, 101(??):69–78, March 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301496.

Lu:2012:EPM

[LGK^{+}12] Qingda Lu, Xiaoyang Gao, Sriram Krishnamoorthy, Gerald Baumgartner, J. Ramanujam, and P. Sadayappan. Empirical performance model-driven data layout optimization and library call selection for tensor contraction expressions. Journal of Parallel and Distributed Computing, 72(3):338–352, March 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511002255.

Lobachev:2013:EPP

[LGL13] Oleg Lobachev, Michael Guthe, and Rita Loogen. Estimating parallel performance. Journal of Parallel and Distributed Computing, 73(6):876–887, June 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000129.

Li:2018:SCC

[LGM18] Tao Li, Brij Bhooshan Gupta, and Roberto Metere. Sociallyconforming cooperative computation in cloud networks. Journal of Parallel and Distributed Computing, 117(??):274–280, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517301958.

Lee:2019:PPS

[LGRV19] Seyong Lee, John Gounley, Amanda Randles, and Jeffrey S. Vetter. Performance portability study for massively parallel computational fluid dynamics application on scalable heterogeneous architectures. Journal of Parallel and Distributed Computing, 129(??):1–13, July 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519301571.

Lan:2010:SDM

[LGZ^{+}10] Zhiling Lan, Jiexing Gu, Ziming Zheng, Rajeev Thakur, and Susan Coghlan. A study of dynamic meta-learning for failure prediction in large-scale systems. Journal of Parallel and Distributed Computing, 70(6):630–643, June 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lv:2019:ORS

[LHCC19] Xiao Lv, Fazhi He, Weiwei Cai, and Yuan Cheng. An optimized RGA supporting selective undo for collaborative text editing systems. Journal of Parallel and Distributed Computing, 132(??):310–330, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303922.

Li:2011:EPA

[LHHH11] Hui-Ya Li, Chia-Lung Hung, Wen-Jyi Hwang, and Yi-Tsan Hung. Efficient pipelined architecture for competitive learning. Journal of Parallel and Distributed Computing, 71(2):236–244, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Liu:2014:OTA

[LHLM14] Junxiu Liu, Jim Harkin, Yuhua Li, and Liam Maguire. Online traffic-aware fault detection for networks-on-chip. Journal of Parallel and Distributed Computing, 74(1):1984–1993, January 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513001974.

Lopez-Huguet:2019:VEM

[LHNBB19] Sergio López-Huguet, Igor Natanael, Andrey Brito, and Ignacio Blanquer. Vertical elasticity on Marathon and Chronos Mesos frameworks. Journal of Parallel and Distributed Computing, 133(??):179–192, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300085.

Li:2014:ABC

[LHW14] Chuanyou Li, Michel Hurfin, and Yun Wang. Approximate Byzantine consensus in sparse, mobile ad-hoc networks. Journal of Parallel and Distributed Computing, 74(9):2860–2871, September 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514000938.

Liu:2019:IRP

[LHW^{+}19] Qin Liu, Panlin Hou, Guojun Wang, Tao Peng, and Shaobo Zhang. Intelligent route planning on large road networks with efficiency and privacy. Journal of Parallel and Distributed Computing, 133(??):93–106, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300814.

Law:2019:ACS

[LHWJ19] T. R. Law, J. Hancox, S. A. Wright, and S. A. Jarvis. An algorithm for computing short-range forces in molecular dynamics simulations with non-uniform particle densities. Journal of Parallel and Distributed Computing, 130(??):1–11, August 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519302047.

Leng:2016:SAR

[LHX^{+}16] Bing Leng, Liusheng Huang, Hongli Xu, Chenkai Yang, and Xinglong Wang. A self-adaptive reconfiguration scheme for throughput maximization in municipal WMNs. Journal of Parallel and Distributed Computing, 92(??):50–59, May 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516000253.

Liu:2018:HHM

[LHZ^{+}18] Hao Liu, Linpeng Huang, Yanmin Zhu, Shengan Zheng, and Yanyan Shen. HMFS: A hybrid in-memory file system with version consistency. Journal of Parallel and Distributed Computing, 117(??):18–36, July 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300662.

Li:2010:AOD

[Li10] Keqin Li. Asymptotically optimal dynamic tree evolution by rapidly mixing random walks on regular networks. Journal of Parallel and Distributed Computing, 70(9):907–916, September 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2014:ONA

[Li14] Keqin Li. Optimal number of annuli for maximizing the lifetime of sensor networks. Journal of Parallel and Distributed Computing, 74(1):1719–1729, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002062.

Li:2016:ETC

[Li16] Keqin Li. Energy and time constrained task scheduling on multiprocessor computers with discrete speed levels. Journal of Parallel and Distributed Computing, 95(??):15–28, September 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516000216.

Li:2017:LDM

[Li17] Keqin Li. Location distribution of a mobile terminal and its application to paging cost reduction and minimization. Journal of Parallel and Distributed Computing, 99(??):68–89, January 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301174.

Li:2019:NCS

[Li19a] Keqin Li. Non-clairvoyant scheduling of independent parallel tasks on single and multiple multicore processors. Journal of Parallel and Distributed Computing, 133(??):210–220, November 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518303988.

Li:2019:OTE

[Li19b] Keqin Li. Optimal task execution speed setting and lower bound for delay and energy minimization. Journal of Parallel and Distributed Computing, 123(??):13–25, January 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306476.

Liono:2019:QQD

[LJQ^{+}19] Jonathan Liono, Prem Prakash Jayaraman, A. K. Qin, Thuong Nguyen, and Flora D. Salim. QDaS: Quality driven data summarisation for effective storage management in Internet of Things. Journal of Parallel and Distributed Computing, 127(??):196–208, May 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151830220X.

Li:2019:EPO

[LJZ^{+}19] Zhihao Li, Haipeng Jia, Yunquan Zhang, Shice Liu, Shigang Li, Xiao Wang, and Hao Zhang. Efficient parallel optimizations of a high-performance SIFT on GPUs. Journal of Parallel and Distributed Computing, 124(??):78–91, February 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518307858.

Louri:2010:SIN

[LK10] Ahmed Louri and Avinash Kodi. Special issue on network-on-chips (NoCs). Journal of Parallel and Distributed Computing, 70(1):69, January 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Louri:2011:ISI

[LK11] Ahmed Louri and Avinash Karanth Kodi. Introduction to the special issue on Networks-on-Chip (NoC) of the Journal of Parallel and Distributed Computing (JPDC). Journal of Parallel and Distributed Computing, 71(5):623–624, May 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511000220.

LaFratta:2013:EEM

[LK13] Patrick A. La Fratta and Peter M. Kogge. Energy-efficient multithreading for a hierarchical heterogeneous multicore through locality-cognizant thread generation. Journal of Parallel and Distributed Computing, 73(12):1551–1562, December 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513001494.

LaSalle:2015:MTM

[LK15] Dominique LaSalle and George Karypis. Multi-threaded modularity based graph clustering using the multilevel paradigm. Journal of Parallel and Distributed Computing, 76(??):66–80, February 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001750.

Lugowski:2015:PPF

[LKB^{+}15] Adam Lugowski, Shoaib Kamil, Aydin Buluç, Samuel Williams, Erika Duriakova, Leonid Oliker, Armando Fox, and John R. Gilbert. Parallel processing of filtered queries in attributed semantic graphs. Journal of Parallel and Distributed Computing, 79–80(??):115–131, May 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151400152X.

Luszczek:2014:LBD

[LKD14] Piotr Luszczek, Jakub Kurzak, and Jack Dongarra. Looking back at dense linear algebra software. Journal of Parallel and Distributed Computing, 74(7):2548–2560, July 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002165.

Lazaro:2012:LTA

[LKM12] Daniel Lázaro, Derrick Kondo, and Joan Manuel Marquès. Long-term availability prediction for groups of volunteer resources. Journal of Parallel and Distributed Computing, 72(2):281–296, February 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511002061.

Lee:2014:EMC

[LKS14] Jaehwan Lee, Pete Keleher, and Alan Sussman. Exploiting multi-core nodes in peer-to-peer grids. Journal of Parallel and Distributed Computing, 74(4):2286–2303, April 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002359.

Li:2010:ECR

[LL10] Chunlin Li and Layuan Li. Energy constrained resource allocation optimization for mobile grids. Journal of Parallel and Distributed Computing, 70(3):245–258, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2012:FLC

[LL12a] Chunlin Li and Layuan Li. A flexible layered control policy for resource allocation in a sensor grid. Journal of Parallel and Distributed Computing, 72(8):925–935, August 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000986.

Li:2012:OEA

[LL12b] Keqin Li and Jie Li. Optimal energy allocation in heterogeneous wireless sensor networks for lifetime maximization.

Journal of Parallel and Distributed Computing, 72(7):902– 916, July 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512000652.

Liu:2018:ETJ

[LL18] Wenjie Liu and Zhanhuai Li. An efficient theta-join query processing in distributed environment. Journal of Parallel and Distributed Computing, 121(??):27–41, November 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304878.

Lai:2019:MAA

[LL19] Chuan-Chi Lai and Chuan-Ming Liu. A mobility-aware approach for distributed data update on unstructured mobile P2P networks. Journal of Parallel and Distributed Computing, 123(??):168–179, January 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518306956.

Llamocca:2017:SRA

[Lla17] Daniel Llamocca. Self-reconfigurable architectures for HEVC forward and inverse transform. Journal of Parallel and Distributed Computing, 109(??):178–192, November 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301740.

Li:2018:FES

[LLB^{+}18] Chen Li, Bo Li, Md Zakirul Alam Bhuiyan, Lihong Wang, Jinghui Si, Guanyu Wei, and Jianxin Li. FluteDB: an efficient and scalable in-memory time series database for sensor-cloud. Journal of Parallel and Distributed Computing, 122(??):95–108, December 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518305422.

Li:2019:AAE

[LLCZ19] Weidong Li, Xi Liu, Xiaobo Cai, and Xuejie Zhang. Approximation algorithm for the energy-aware profit maximizing problem in heterogeneous computing systems. Journal of Parallel and Distributed Computing, 124(??):70–77, February 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151830786X.

Long:2015:EES

[LLDL15] Jun Long, Anfeng Liu, Mianxiong Dong, and Zhi Li. An energy-efficient and sink-location privacy enhanced scheme for WSNs through ring based routing. Journal of Parallel and Distributed Computing, 81–82(??):47–65, July 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515000660.

Lv:2018:HCP

[LLFJ18] Yali Lv, Cheng-Kuan Lin, Jianxi Fan, and Xiaohua Jia. Hamiltonian cycle and path embeddings in 3-ary n-cubes based on K_{1,3}-structure faults. Journal of Parallel and Distributed Computing, 120(??):148–158, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304283.

Lee:2013:DSE

[LLKY13] Jaekyu Lee, Si Li, Hyesoon Kim, and Sudhakar Yalamanchili. Design space exploration of on-chip ring interconnection for a CPU–GPU heterogeneous architecture. Journal of Parallel and Distributed Computing, 73(12):1525–1538, December 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513001524.

Lai:2015:SAM

[LLLC15] Bo-Cheng Charles Lai, Kun-Chun Li, Guan-Ru Li, and Chin-Hsuan Chiang. Self adaptable multithreaded object detection on embedded multicore systems. Journal of Parallel and Distributed Computing, 78(??):25–38, April 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515000192.

Lu:2016:PIF

[LLS^{+}16] Yunping Lu, Yi Li, Bo Song, Weihua Zhang, Haibo Chen, and Lu Peng. Parallelizing image feature extraction algorithms on multi-core platforms. Journal of Parallel and Distributed Computing, 92(??):1–14, May 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516000228.

Lin:2012:MSN

[LLT12] Jing Lin, Xiaola Lin, and Liang Tang. Making-a-stop: a new bufferless routing algorithm for on-chip network. Journal of Parallel and Distributed Computing, 72(4):515–524, April 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000020.

Lv:2012:GKA

[LLW12] Xixiang Lv, Hui Li, and Baocang Wang. Group key agreement for secure group communication in dynamic peer systems. Journal of Parallel and Distributed Computing, 72(10):1195–1200, October 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001438.

Li:2017:DHB

[LLWC17] Beibei Li, Rongxing Lu, Wei Wang, and Kim-Kwang Raymond Choo. Distributed host-based collaborative detection for false data injection attacks in smart grid cyber-physical system. Journal of Parallel and Distributed Computing, 103(??): 32–41, May 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301885.

Li:2015:TRT

[LLY15] Peilong Li, Yan Luo, and Jun Yang. Transformer: Runtime reprogrammable heterogeneous architecture for transparent acceleration of dynamic workloads. Journal of Parallel and Distributed Computing, 86(??):45–61, December 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001409.

Li:2016:OPC

[LM16] Shaosong Li and Shivakant Mishra. Optimizing power consumption in multicore smartphones. Journal of Parallel and Distributed Computing, 95(??):124–137, September 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516000198.

Li:2017:SKP

[LMB^{+}17] Bo Li, John Mooring, Sam Blanchard, Aditya Johri, Melinda Leko, and Kirk W. Cameron. SeeMore: a kinetic parallel computer sculpture for educating broad audiences on parallel computation. Journal of Parallel and Distributed Computing, 105(??):183–199, July 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300230.

Lazaro-Munoz:2017:TRM

[LMGLGLG17] A. J. Lázaro-Muñoz, J. M. González-Linares, J. Gómez-Luna, and N. Guil. A tasks reordering model to reduce transfers overhead on GPUs. Journal of Parallel and Distributed Computing, 109(??):258–271, November 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302046.

Li:2011:NRA

[LMJC11] Guangsong Li, Jianfeng Ma, Qi Jiang, and Xi Chen. A novel re-authentication scheme based on tickets in wireless local area networks. Journal of Parallel and Distributed Computing, 71(7):906–914, July 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511000554.

Lin:2010:RTS

[LML^{+}10] Xuan Lin, Anwar Mamat, Ying Lu, Jitender Deogun, and Steve Goddard. Real-time scheduling of divisible loads in cluster computing environments. Journal of Parallel and Distributed Computing, 70(3):296–308, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2010:MRT

[LMP10] Xiaozhou Li, Jayadev Misra, and C. Greg Plaxton. Maintaining the Ranch topology. Journal of Parallel and Distributed Computing, 70(11):1142–1158, November 2010. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:2018:HAS

[LMSK18] Dajung Lee, Nirja Mehta, Alexandria Shearer, and Ryan Kastner. A hardware accelerated system for high throughput cellular image analysis. Journal of Parallel and Distributed Computing, 113(??):167–178, March 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303258.

Li:2018:OTM

[LMXJ18] Weimin Li, Jun Mo, Minjun Xin, and Qun Jin. An optimized trust model integrated with linear features for cyberenabled recommendation services. Journal of Parallel and Distributed Computing, 118 (part 1)(??):81–88, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0743731517302794.

Li:2011:TRP

[LMY^{+}11] Jiangtian Li, Xiaosong Ma, Srikanth Yoginath, Guruprasad Kora, and Nagiza F. Samatova. Transparent runtime parallelization of the R scripting language. Journal of Parallel and Distributed Computing, 71(2):157–168, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:2012:GSF

[LNA12] Jae W. Lee, Man Cheuk Ng, and Krste Asanović. Globally Synchronized Frames for guaranteed quality-of-service in on-chip networks. Journal of Parallel and Distributed Computing, 72(11):1401–1411, November 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512000251.

Li:2017:EAS

[LNAL17] Ying Li, Jianwei Niu, Mohammed Atiquzzaman, and Xiang Long. Energy-aware scheduling on heterogeneous multicore systems with guaranteed probability. Journal of Parallel and Distributed Computing, 103(??):64–76, May 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301745.

Liang:2013:WBC

[LNC13] Xiangyang Liang, Minh Nguyen, and Hao Che. Wimpy or brawny cores: a throughput perspective. Journal of Parallel and Distributed Computing, 73(10):1351–1361, October 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513001160.

Lima:2012:PEO

[LNW^{+}12] Antonio M. Lima, Marco A. S. Netto, Thais Webber, Ricardo M. Czekster, Cesar A. F. De Rose, and Paulo Fernandes. Performance evaluation of OpenMP-based algorithms for handling Kronecker descriptors. Journal of Parallel and Distributed Computing, 72(5):678–692, May 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000354.

Lopriore:2013:OPD

[Lop13] Lanfranco Lopriore. Object protection in distributed systems. Journal of Parallel and Distributed Computing, 73(5):570–579, May 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000099.

Lopriore:2018:PBP

[Lop18] Lanfranco Lopriore. Password-based protection of clustered segments in distributed memory systems. Journal of Parallel and Distributed Computing, 115(??):29–40, May 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300169.

Li:2018:MAE

[LpJS^{+}18] Chanchan Li, Guo ping Jiang, Yurong Song, Lingling Xia, Yinwei Li, and Bo Song. Modeling and analysis of epidemic spreading on community networks with heterogeneity. Journal of Parallel and Distributed Computing, 119(??):136–145, September 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518302685.

Lee:2010:AET

[LPK^{+}10] Jaejin Lee, Jung-Ho Park, Honggyu Kim, Changhee Jung, Daeseob Lim, and SangYong Han. Adaptive execution techniques of parallel programs for multiprocessors. Journal of Parallel and Distributed Computing, 70(5):467–480, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lopez-Portugues:2012:ASS

[LPLFMC^{+}12] Miguel López-Portugués, Jesús A. López-Fernández, Jonatan Menéndez-Canal, Alberto Rodríguez-Campa, and José Ranilla. Acoustic scattering solver based on single level FMM for multi-GPU systems. Journal of Parallel and Distributed Computing, 72(9):1057–1064, September 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001481.

Li:2012:OOS

[LQM^{+}12] Jiayin Li, Meikang Qiu, Zhong Ming, Gang Quan, Xiao Qin, and Zonghua Gu. Online optimization for scheduling preemptable tasks on IaaS cloud systems. Journal of Parallel and Distributed Computing, 72(5):666–677, May 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000366.

Lang:2014:ETE

[LR14] Jens Lang and Gudula Rünger. An execution time and energy model for an energy-aware execution of a conjugate gradient method with CPU/GPU collaboration. Journal of Parallel and Distributed Computing, 74(9):2884–2897, September 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001038.

Lukovszki:2018:AIN

[LRS18] Tamás Lukovszki, Matthias Rost, and Stefan Schmid. Approximate and incremental network function placement. Journal of Parallel and Distributed Computing, 120(??):159–169, October 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518304271.

Liang:2010:RDS

[LS10] Zhengqiang Liang and Weisong Shi. A reputation-driven scheduler for autonomic and sustainable resource sharing in Grid computing. Journal of Parallel and Distributed Computing, 70(2):111–125, February 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lopriore:2019:PPS

[LS19] Lanfranco Lopriore and Antonella Santone. Protected pointers to specify access privileges in distributed systems. Journal of Parallel and Distributed Computing, 126(??):1–12, April 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518308876.

Lin:2015:EET

[LSC^{+}15] Ching-Chi Lin, You-Cheng Syu, Chao-Jui Chang, Jan-Jan Wu, Pangfeng Liu, Po-Wen Cheng, and Wei-Te Hsu. Energy-efficient task scheduling for multi-core platforms with per-core DVFS. Journal of Parallel and Distributed Computing, 86(??):71–81, December 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001422.

Lerida:2013:SBP

[LSH^{+}13] Josep L. Lerida, Francesc Solsona, Porfidio Hernandez, Francesc Gine, Mauricio Hanzich, and Josep Conde. Statebased predictions with self-correction on Enterprise Desktop Grid environments. Journal of Parallel and Distributed Computing, 73(6):777–789, June 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000324.

Li:2011:CFO

[LSS^{+}11a] Ruixuan Li, Wei Song, Haiying Shen, Weijun Xiao, and Zhengding Lu. Corrigendum to “A flabellate overlay network for multi-attribute search” [J. Parallel Distrib. Comput. 71 (2011) 407–423]. Journal of Parallel and Distributed Computing, 71(7):1065, July 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511000232.

Li:2011:FON

[LSS^{+}11b] Ruixuan Li, Wei Song, Haiying Shen, Weijun Xiao, and Zhengding Lu. A flabellate overlay network for multi-attribute search. Journal of Parallel and Distributed Computing, 71(3): 407–423, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2013:PMC

[LST^{+}13] Qi Li, Raied Salman, Erik Test, Robert Strack, and Vojislav Kecman. Parallel multitask cross validation for Support Vector Machine using GPU. Journal of Parallel and Distributed Computing, 73(3):293–302, March 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151200055X.

Lee:2017:NNL

[LST17] Hee Won Lee, Mihail L. Sichitiu, and David Thuente. NEAT: Network link emulation with adaptive time dilation. Journal of Parallel and Distributed Computing, 104(??):88–98, June 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300199.

Liu:2014:DBD

[LSWC14] Chuan-Ming Liu, Ta-Chih Su, Jenq-Haur Wang, and Yen-Lin Chen. Data broadcasting for dependent information using multiple channels in wireless broadcast environments. Journal of Parallel and Distributed Computing, 74(9):2795–2807, September 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514000902.

Li:2014:DPM

[LSXX14] Jianhua Li, Liang Shi, Chun Jason Xue, and Yinlong Xu. Dual partitioning multicasting for high-performance onchip networks. Journal of Parallel and Distributed Computing, 74(1):1858–1871, January 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513001287.

Luckow:2015:PDA

[LSZJ15] Andre Luckow, Mark Santcroos, Ashley Zebrowski, and Shantenu Jha. Pilot-Data: an abstraction for distributed data. Journal of Parallel and Distributed Computing, 79–80(??):16– 30, May 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001725.

Liu:2015:ABI

[LSZZ15] Peng Liu, Weiwei Sun, Jian Zhang, and Baihua Zheng. An automaton-based index scheme supporting twig queries for on-demand XML data broadcast. Journal of Parallel and Distributed Computing, 86(??):82–97, December 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001288.

Lahiri:2010:IFI

[LT10] Bibudh Lahiri and Srikanta Tirthapura. Identifying frequent items in a network using gossip. Journal of Parallel and Distributed Computing, 70(12):1241–1253, December 2010. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Liu:2014:HMC

[LTG14] Chen Liu, Pollawat Thanarungroj, and Jean-Luc Gaudiot. How many cores do we need to run a parallel workload: a test drive of the Intel SCC platform? Journal of Parallel and Distributed Computing, 74(7):2582–2595, July 2014.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000021.

Liu:2012:CPS

[LTWW12] Qin Liu, Chiu C. Tan, Jie Wu, and Guojun Wang. Cooperative private searching in clouds. Journal of Parallel and Distributed Computing, 72(8):1019–1031, August 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001086.

Lin:2014:PAE

[LÜ14] Hui Lin and Halit Üster. A parallel algorithm with enhancements via partial objective value cuts for cluster-based wireless sensor network design. Journal of Parallel and Distributed Computing, 74(7):2615–2625, July 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000641.

Lucchese:2018:PNC

[Luc18] Fabiano Lucchese. From P2P to NoSQL: a continuous metric for classifying large-scale storage systems. Journal of Parallel and Distributed Computing, 113(??):227–249, March 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517303295.

Liu:2015:FGS

[LV15] Weifeng Liu and Brian Vinter. A framework for general sparse matrix–matrix multiplication on GPUs and heterogeneous processors. Journal of Parallel and Distributed Computing, 85(??):47–61, November 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001185.

Li:2016:EEC

[LW16a] Dawei Li and Jie Wu. Energy-efficient contention-aware application mapping and scheduling on NoC-based MPSoCs. Journal of Parallel and Distributed Computing, 96(??):1–11, October 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516300120.

Lin:2016:TMP

[LW16b] Pei-Hung Lin and Paul R. Woodward. Transforming the multifluid PPM algorithm to run on GPUs. Journal of Parallel and Distributed Computing, 93–94(??):56–65, July 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516300119.

Lin:2019:ELL

[LW19] Huanxin Lin and Cho-Li Wang. Efficient low-latency packet processing using On-GPU thread-data remapping. Journal of Parallel and Distributed Computing, 133(??):51–62, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518305495.

Lai:2018:THP

[LWC^{+}18] Bo-Cheng Lai, Tung-Yu Wu, Tsou-Han Chiu, Kun-Chun Li, Chia-Ying Lee, Wei-Chen Chien, and Wing Hung Wong. Towards high performance data analytic on heterogeneous manycore systems: a study on Bayesian Sequential Partitioning. Journal of Parallel and Distributed Computing, 122(??):36–50, December 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518304921.

Langguth:2015:PPM

[LWCC15] J. Langguth, N. Wu, J. Chai, and X. Cai. Parallel performance modeling of irregular applications in cell-centered finite volume methods over unstructured tetrahedral meshes. Journal of Parallel and Distributed Computing, 76(??):120–131, February 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001968.

Luo:2014:HCS

[LWCG14] Hongyin Luo, Shaojun Wei, Deming Chen, and Donghui Guo. Hybrid circuit-switched network for on-chip communication in large-scale chip-multiprocessors. Journal of Parallel and Distributed Computing, 74(9):2818–2830, September 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514000914.

Li:2019:MLM

[LWH^{+}19] Jing Li, Xianmin Wang, Zhengan Huang, Licheng Wang, and Yang Xiang. Multi-level multi-secret sharing scheme for decentralized e-voting in cloud computing. Journal of Parallel and Distributed Computing, 130(??):91–97, August 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151930262X.

Li:2019:PSA

[LWK^{+}19] Xiong Li, Fan Wu, Saru Kumari, Lili Xu, Arun Kumar Sangaiah, and Kim-Kwang Raymond Choo. A provably secure and anonymous message authentication scheme for smart grids. Journal of Parallel and Distributed Computing, 132(??):242–249, October 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517303064.

Li:2012:CSC

[LWLD12] Xiaoguang Li, Jie Wu, Shan Lin, and Xiaojiang Du. Channel switching control policy for wireless mesh networks. Journal of Parallel and Distributed Computing, 72(10):1295–1305, October 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001475.

Liu:2018:ADF

[LWW18] Linfeng Liu, Ran Wang, and Jiagao Wu. On the adaptive data forwarding in opportunistic underwater sensor networks using GPS-free mobile nodes. Journal of Parallel and Distributed Computing, 122(??):131–144, December 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305859.

Liu:2019:TIM

[LWW19] Linfeng Liu, Ran Wang, and Jiagao Wu. A timeinhomogeneous Markov chain and its distributed solution for message dissemination in OUSNs. Journal of Parallel and Distributed Computing, 130(??):179–192, August 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518302077.

Lu:2018:IIB

[LWWQ18] Zhihui Lu, Nini Wang, Jie Wu, and Meikang Qiu. IoTDeM: An IoT big data-oriented MapReduce performance prediction extended model in multiple edge clouds. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):316–327, August 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517302988.

Li:2019:UBL

[LWXX19] Xiaolin Li, Peng Wang, Xin-Jian Xu, and Gaoxi Xiao. Universal behavior of the linear threshold model on weighted networks. Journal of Parallel and Distributed Computing, 123(??):223–229, January 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731518307378.

Lee:2012:PDS

[LWZZ12] Young Choon Lee, Chen Wang, Albert Y. Zomaya, and Bing Bing Zhou. Profit-driven scheduling for cloud services with data access awareness. Journal of Parallel and Distributed Computing, 72(4):591–602, April 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511002310.

Liang:2012:UBC

[LXLS12] Meilian Liang, Xiaodong Xu, Jiarong Liang, and Zehui Shao. Upper bounds on the connection probability for 2-D meshes and tori. Journal of Parallel and Distributed Computing, 72(2):185–194, February 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151100222X.

Li:2011:TPA

[LXW^{+}11] Qi Li, Mingwei Xu, Jianping Wu, Patrick P. C. Lee, and Dah Ming Chiu. Toward a practical approach for BGP stability with root cause check. Journal of Parallel and Distributed Computing, 71(8):1098–1110, August 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000840.

Luo:2013:BMH

[LXZ13] Jiaqing Luo, Bin Xiao, and Shijie Zhou. A bottom-up model for heterogeneous BitTorrent systems. Journal of Parallel and Distributed Computing, 73(8):1116–1126, August 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000646.

Lee:2010:RMS

[LY10] Sookyoung Lee and Mohamed Younis. Recovery from multiple simultaneous failures in wireless sensor networks using minimum Steiner tree. Journal of Parallel and Distributed Computing, 70(5):525–536, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2012:CCE

[LY12] Mingfu Li and Hsun-Hao Yang. CEA: a Cyclic Expansion Algorithm for data migration in parallel video servers. Journal of Parallel and Distributed Computing, 72(7):868–879, July 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000676.

Liu:2013:EAR

[LY13] Lin Liu and Yuanyuan Yang. Energy-aware routing in hybrid optical network-on-chip for future multi-processor system-on-chip. Journal of Parallel and Distributed Computing, 73(2):189–197, February 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151200233X.

Lopes:2019:FBD

[LYIP19] Paulo A. C. Lopes, Satyendra Singh Yadav, Aleksandar Ilic, and Sarat Kumar Patra. Fast block distributed CUDA implementation of the Hungarian algorithm. Journal of Parallel and Distributed Computing, 130(??):50–62, August 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519302254.

Li:2019:OCO

[LYJ^{+}19] Wei Li, Xinghui You, Yingying Jiang, Jun Yang, and Long Hu. Opportunistic computing offloading in edge clouds. Journal of Parallel and Distributed Computing, 123(??):69–76, January 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518306506.

Ling:2016:JSM

[LYW^{+}16] Xiao Ling, Yi Yuan, Dan Wang, Jiangchuan Liu, and Jiahai Yang. Joint scheduling of MapReduce jobs with servers: Performance bounds and experiments. Journal of Parallel and Distributed Computing, 90–91(??):52–66, April 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516000174.

Landfeldt:2011:SIA

[LZ11] Bjorn Landfeldt and Albert Y. Zomaya. Special issue on advancement of research in wireless access and mobile systems. Journal of Parallel and Distributed Computing, 71(9):1177–1178, September 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001055.

Liu:2011:TAL

[LZC11] An-Feng Liu, Peng-Hui Zhang, and Zhi-Gang Chen. Theoretical analysis of the lifetime and energy hole in cluster based wireless sensor networks. Journal of Parallel and Distributed Computing, 71(10):1327–1355, October 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000992.

Li:2011:CCQ

[LZI^{+}11] Bin Li, Li Zhao, Ravi Iyer, Li-Shiuan Peh, Michael Leddige, Michael Espig, Seung Eun Lee, and Donald Newell. CoQoS: Coordinating QoS-aware shared resources in NoC-based SoCs. Journal of Parallel and Distributed Computing, 71(5):700–713, May 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151000211X.

Liu:2011:JTA

[LZLX11] Tiantian Liu, Yingchao Zhao, Minming Li, and Chun Jason Xue. Joint task assignment and cache partitioning with cache locking for WCET minimization on MPSoC. Journal of Parallel and Distributed Computing, 71(11):1473–1483, November 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151100102X.

Longo:2019:UCM

[LZN19] Antonella Longo, Marco Zappatore, and Shamkant B. Navathe. The unified chart of mobility services: Towards a systemic approach to analyze service quality in smart mobility ecosystem. Journal of Parallel and Distributed Computing, 127(??):118–133, May 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519300097.

Li:2019:ESM

[LZWZ19] Xingxin Li, Youwen Zhu, Jian Wang, and Ji Zhang. Efficient and secure multi-dimensional geometric range query over encrypted data in cloud. Journal of Parallel and Distributed Computing, 131(??):44–54, September 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306294.

Li:2011:MDT

[LZY11] Xiaoyong Li, Feng Zhou, and Xudong Yang. A multidimensional trust evaluation model for large-scale P2P computing. Journal of Parallel and Distributed Computing, 71 (6):837–847, June 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731511000190.

Liu:2018:CFB

[LZY^{+}18] Meng Liu, Xuyun Zhang, Chi Yang, Qiang He, and Jianbing Zhang. Curve fitting based efficient parameter selection for robust provable data possession. Journal of Parallel and Distributed Computing, 120(??):62–76, October 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303514.

Liu:2011:APD

[LZZ^{+}11] Xu Liu, Jianfeng Zhan, Kunlin Zhan, Weisong Shi, Lin Yuan, Dan Meng, and Lei Wang. Automatic performance debugging of SPMD-style parallel programs. Journal of Parallel and Distributed Computing, 71(7):925–937, July 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000748.

Mershad:2011:CCD

[MA11] Khaleel Mershad and Hassan Artail. CODISC: Collaborative and distributed semantic caching for maximizing cache effectiveness in wireless networks. Journal of Parallel and Distributed Computing, 71(3):495–511, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mohiuddin:2019:WAV

[MA19] Irfan Mohiuddin and Ahmad Almogren. Workload aware VM consolidation method in edge/cloud computing for IoT applications. Journal of Parallel and Distributed Computing, 123(??):204–214, January 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518306762.

Maia:2013:MRP

[MAGL13] Guilherme Maia, Andre L. L. Aquino, Daniel L. Guidoni, and Antonio A. F. Loureiro. A multicast reprogramming protocol for wireless sensor networks based on small world concepts. Journal of Parallel and Distributed Computing, 73(9):1277–1291, September 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513001111.

Min-Allah:2012:PER

[MAHKZ12] Nasro Min-Allah, Hameed Hussain, Samee Ullah Khan, and Albert Y. Zomaya. Power efficient rate monotonic scheduling for multi-core systems. Journal of Parallel and Distributed Computing, 72(1):48–57, January 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001407.

Min-Allah:2013:LPF

[MAKWZ13] Nasro Min-Allah, Samee U. Khan, Xiuli Wang, and Albert Y. Zomaya. Lowest priority first based feasibility analysis of real-time systems. Journal of Parallel and Distributed Computing, 73(8):1066–1075, August 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000609.

Manjunathaiah:2013:FGM

[Man13] M. Manjunathaiah. Fine-grained multi-phase array designs. Journal of Parallel and Distributed Computing, 73 (8):1076–1082, August 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731513000476.

Martalo:2014:SDR

[MAPF14] M. Martalò, M. Amoretti, M. Picone, and G. Ferrari. Sporadic decentralized resource maintenance for P2P distributed storage networks. Journal of Parallel and Distributed Computing, 74(2):2029–2038, February 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002220.

Mastrostefano:2013:EBF

[MB13] Enrico Mastrostefano and Massimo Bernaschi. Efficient breadth first search on multi-GPU systems. Journal of Parallel and Distributed Computing, 73(9):1292–1305, September 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513001135.

Mansouri:2019:DRM

[MB19] Yaser Mansouri and Rajkumar Buyya. Dynamic replication and migration of data objects with hot-spot and cold-spot statuses across storage data centers. Journal of Parallel and Distributed Computing, 126(??):121–133, April 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518309092.

Ma:2013:KAT

[MBBD13] Teng Ma, George Bosilca, Aurelien Bouteiller, and Jack J. Dongarra. Kernel-assisted and topology-aware MPI collective communications on multicore/many-core platforms. Journal of Parallel and Distributed Computing, 73(7):1000–1010, July 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000166.

Mullen:2017:LDH

[MBG^{+}17] Julia Mullen, Chansup Byun, Vijay Gadepally, Siddharth Samsi, Albert Reuther, and Jeremy Kepner. Learning by doing, high performance computing education in the MOOC era. Journal of Parallel and Distributed Computing, 105(??):105– 115, July 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517300217.

Merzoug:2019:SAD

[MBMC19] M. A. Merzoug, A. Boukerche, A. Mostefaoui, and S. Chouali. Spreading aggregation: a distributed collision-free approach for data aggregation in large-scale wireless sensor networks. Journal of Parallel and Distributed Computing, 125(??):121– 134, March 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518308542.

Mokdad:2011:ACM

[MBO11] Lynda Mokdad and Jalel Ben-Othman. Admission control mechanism and performance analysis based on stochastic automata networks formalism. Journal of Parallel and Distributed Computing, 71(4):594–602, April 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Moghaddam:2019:AAB

[MBR19] Sara Kardani Moghaddam, Rajkumar Buyya, and Kotagiri Ramamohanarao. ACAS: an anomaly-based cause aware auto-scaling framework for clouds. Journal of Parallel and Distributed Computing, 126(??):107–120, April 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518309080.

Maciel:2012:BDS

[MBS^{+}12] Paulo Ditarso Maciel, Jr., Francisco Brasileiro, Ricardo Araújo Santos, David Candeia, Raquel Lopes, Marcus Carvalho, Renato Miceli, Nazareno Andrade, and Miranda Mowbray. Business-driven short-term management of a hybrid IT infrastructure. Journal of Parallel and Distributed Computing, 72(2):106–119, February 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511002176.

Maggioni:2016:OTS

[MBW16] Marco Maggioni and Tanya Berger-Wolf. Optimization techniques for sparse matrix-vector multiplication on GPUs. Journal of Parallel and Distributed Computing, 93–94(??):66–86, July 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516300028.

Maier:2017:OLD

[MC17] Andrew J. Maier and Bruce F. Cockburn. Optimization of low-density parity check decoder performance for OpenCL designs synthesized to FPGAs. Journal of Parallel and Distributed Computing, 107(??):134–145, September 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301004.

Martinez:2012:HIS

[MCAS12] Raúl Martínez, José M. Claver, Francisco J. Alfaro, and José L. Sánchez. Hardware implementation study of several new egress link scheduling algorithms. Journal of Parallel and Distributed Computing, 72(8):975–989, August 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001062.

Marcon:2011:CFI

[MCM^{+}11] César Marcon, Ney Calazans, Edson Moreno, Fernando Moraes, Fabiano Hessel, and Altamiro Susin. CAFES: a framework for intrachip application modeling and communication architecture design. Journal of Parallel and Distributed Computing, 71(5):714–728, May 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731510002005.

Musat:2018:ASE

[MCP^{+}18] George-Alexandru Musat, Mădălin Colezea, Florin Pop, Catalin Negru, Mariana Mocanu, Christian Esposito, and Aniello Castiglione. Advanced services for efficient management of smart farms. Journal of Parallel and Distributed Computing, 116(??):3–17, ???? 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302939.

Michail:2014:CIC

[MCS14] Othon Michail, Ioannis Chatzigiannakis, and Paul G. Spirakis. Causality, influence, and computation in possibly disconnected synchronous dynamic networks. Journal of Parallel and Distributed Computing, 74(1):2016–2026, January 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513001330.

Melo:2019:THI

[MCS^{+}19] Alba Melo, Jesus Carretero, Per Stenstrom, Sanjay Ranka, and Eduard Ayguade. Trends on heterogeneous and innovative hardware and software systems. Journal of Parallel and Distributed Computing, 133(??):362–364, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519305490.

Muppala:2014:MTS

[MCZ14] Sireesha Muppala, Guihai Chen, and Xiaobo Zhou. Multitier service differentiation by coordinated learning-based resource provisioning and admission control. Journal of Parallel and Distributed Computing, 74(5):2351–2364, May 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000069.

Mansouri:2013:EDH

[MD13] Najme Mansouri and Gholam Hosein Dastghaibyfard. Enhanced dynamic hierarchical replication and weighted scheduling strategy in data grid. Journal of Parallel and Distributed Computing, 73(4):534–543, April 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000038.

Muhammad:2017:ALA

[MEMEMH17] Sayed T. Muhammad, Magdy A. El-Moursy, Ali A. El-Moursy, and Hesham F. A. Hamed. Architecture level analysis for process variation in synchronous and asynchronous Networks-on-Chip. Journal of Parallel and Distributed Computing, 102(??):175–185, April 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516302052.

Menouer:2018:SCP

[Men18] Tarek Menouer. Solving combinatorial problems using a parallel framework. Journal of Parallel and Distributed Computing, 112 (part 2)(??):140–153, February 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301764.

Merelli:2019:EDC

[MFT^{+}19] Ivan Merelli, Federico Fornari, Fabio Tordini, Daniele D’Agostino, Marco Aldinucci, and Daniele Cesini. Exploiting Docker containers over grid computing for a comprehensive study of chromatin conformation in different cell types. Journal of Parallel and Distributed Computing, 134(??):116–127, December 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519305593.

Munir:2014:QTA

[MGRRK14] Arslan Munir, Ann Gordon-Ross, Sanjay Ranka, and Farinaz Koushanfar. A queueing theoretic approach for performance evaluation of low-power multi-core embedded systems. Journal of Parallel and Distributed Computing, 74(1):1872–1890, January 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513001299.

Ma:2012:EED

[MGSG12] Yan Ma, Bin Gong, Ryo Sugihara, and Rajesh Gupta. Energyefficient deadline scheduling for heterogeneous systems. Journal of Parallel and Distributed Computing, 72(12):1725–1740, December 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001670.

Hwu:2014:WAP

[mH14] Wen mei Hwu. What is ahead for parallel computing. Journal of Parallel and Distributed Computing, 74(7):2574–2581, July 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514000331.

Monson:2018:EDO

[MH18] Joshua S. Monson and Brad L. Hutchings. Enhancing debug observability for HLS-based FPGA circuits through source-to-source compilation. Journal of Parallel and Distributed Computing, 117(??):148–160, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300893.

Ma:2016:TTC

[MHLZ16] Hua Ma, Zhigang Hu, Keqin Li, and Hongyu Zhang. Toward trustworthy cloud service selection: a time-aware approach using interval neutrosophic set. Journal of Parallel and Distributed Computing, 96(??):75–94, October 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300387.

Mondal:2016:PPA

[MKM16] Kaushik Mondal, Arindam Karmakar, and Partha Sarathi Mandal. Path planning algorithms for mobile anchors towards range-free localization. Journal of Parallel and Distributed Computing, 97(??):35–46, November 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300661.

Moritz:2014:IJS

[MKN14] Csaba Andras Moritz, Santosh Khasanvis, and Pritish Narayanan. Introduction to JPDC special issue on computing with future nanotechnology. Journal of Parallel and Distributed Computing, 74(6):2439–2440, June 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000355.

Mane:2018:CST

[MKW18] S. A. Mane, S. A. Kandekar, and B. N. Waphare. Constructing spanning trees in augmented cubes. Journal of Parallel and Distributed Computing, 122(??):188–194, December 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518305872.

Martin:2018:ENE

[MLCFH^{+}18] Alejandro Martín, Raúl Lara-Cabrera, Félix Fuentes-Hurtado, Valery Naranjo, and David Camacho. EvoDeep: A new evolutionary approach for automatic Deep Neural Networks parametrisation. Journal of Parallel and Distributed Computing, 117(??):180–191, July 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302605.

Mamat:2012:ERT

[MLDG12] Anwar Mamat, Ying Lu, Jitender Deogun, and Steve Goddard. Efficient real-time divisible load scheduling. Journal of Parallel and Distributed Computing, 72(12):1603–1616, December 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512002183.

Moore:2012:VEP

[MLK12] Nicholas Moore, Miriam Leeser, and Laurie Smith King. VForce: an environment for portable applications on high performance systems with accelerators. Journal of Parallel and Distributed Computing, 72(9):1144–1156, September 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511001493.

Meng:2016:CCA

[MLK^{+}16] Jie Meng, Eduard Llamosí, Fulya Kaplan, Chulian Zhang, Jiayi Sheng, Martin Herbordt, Gunar Schirner, and Ayse K. Coskun. Communication and cooling aware job allocation in data centers for communication-intensive workloads. Journal of Parallel and Distributed Computing, 96(??):181–193, October 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516300521.

Manzanares-Lopez:2012:ICU

[MLMSMG12] Pilar Manzanares-Lopez, Josemaria Malgosa-Sanahuja, and Juan Pedro Muñoz-Gea. The importance of considering unauthentic transactions in trust management systems. Journal of Parallel and Distributed Computing, 72(6):809–818, June 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000718.

Ma:2019:OST

[MLW^{+}19] Yuchen Ma, Jiajia Li, Xiaolong Wu, Chenggang Yan, Jimeng Sun, and Richard Vuduc. Optimizing sparse tensor times matrix on GPUs. Journal of Parallel and Distributed Computing, 129(??):99–109, July 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518305161.

Ma:2017:RSU

[MLZY17] Li Ma, Peng Leng, Yong Zhong, and Wenyin Yang. Research on semantic of updatable distributed logic and its application in access control. Journal of Parallel and Distributed Computing, 103(??):104–112, May 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301812.

Mastoras:2015:ADB

[MM15] Aristeidis Mastoras and George Manis. Ariadne — directivebased parallelism extraction from recursive functions. Journal of Parallel and Distributed Computing, 86(??):16–28, December 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515001276.

Meyer:2017:HMP

[MMCL^{+}17] Hugo Meyer, Ronal Muresano, Marcela Castro-León, Dolores Rexachs, and Emilio Luque. Hybrid Message Pessimistic Logging. Improving current pessimistic message logging protocols. Journal of Parallel and Distributed Computing, 104(??):206– 222, June 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517300515.

Mezmaz:2011:PBO

[MMK^{+}11] M. Mezmaz, N. Melab, Y. Kessaci, Y. C. Lee, E.-G. Talbi, A. Y. Zomaya, and D. Tuyttens. A parallel biobjective hybrid metaheuristic for energy-aware scheduling for cloud computing systems. Journal of Parallel and Distributed Computing, 71(11):1497–1508, November 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000827.

Martinelli:2018:EMC

[MMN^{+}18] Fabio Martinelli, Francesco Mercaldo, Vittoria Nardone, Antonella Santone, Arun Kumar Sangaiah, and Aniello Cimitile. Evaluating model checking for cyber threats code obfuscation identification. Journal of Parallel and Distributed Computing, 119(??):203–218, September 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518302636.

Montero:2011:EMH

[MMVL11] Ruben S. Montero, Rafael Moreno-Vozmediano, and Ignacio M. Llorente. An elasticity model for High Throughput Computing clusters. Journal of Parallel and Distributed Computing, 71(6):750–757, June 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731510000985.

Moritz:2012:SIJ

[MNK12] Csaba Andras Moritz, Pritish Narayanan, and Santosh Khasanvis. Special issue of Journal of Parallel and Distributed Computing: Computing with Future Nanotechnology. Journal of Parallel and Distributed Computing, 72(12):1781, December 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512002365.

Malik:2019:BVL

[MNR^{+}19] Maria Malik, Katayoun Neshatpour, Setareh Rafatirad, Rajiv V. Joshi, Tinoosh Mohsenin, Hassan Ghasemzadeh, and Houman Homayoun. Big vs little core for energyefficient Hadoop computing. Journal of Parallel and Distributed Computing, 129(??):110–124, July 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300959.

Meunier:2010:LTM

[MP10] Quentin L. Meunier and Frédéric Pétrot. Lightweight Transactional Memory systems for NoCs based architectures: Design, implementation and comparison of two policies. Journal of Parallel and Distributed Computing, 70(10):1024–1041, October 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Miller:2015:TBC

[MP15] Avery Miller and Andrzej Pelc. Tradeoffs between cost and information for rendezvous and treasure hunt. Journal of Parallel and Distributed Computing, 83(??):159–167, September 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515001021.

Menezo:2017:ACC

[MPG17a] Lucia G. Menezo, Valentin Puente, and Jose-Angel Gregorio. An adaptive cache coherence protocol: Trading storage for traffic. Journal of Parallel and Distributed Computing, 102(??):163–174, April 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516302064.

Moldaschl:2017:FTC

[MPG17b] Michael Moldaschl, Karl E. Prikopa, and Wilfried N. Gansterer. Fault tolerant communication-optimal 2.5D matrix multiplication. Journal of Parallel and Distributed Computing, 104(??):179–190, June 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300412.

Mestre:2017:TEP

[MPN17] Demetrio Gomes Mestre, Carlos Eduardo Santos Pires, and Dimas Cassimiro Nascimento. Towards the efficient parallelization of multi-pass adaptive blocking for entity matching. Journal of Parallel and Distributed Computing, 101(??):27– 40, March 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301459.

Molla:2019:ODD

[MPR19] Anisur Rahaman Molla, Supantha Pandit, and Sasanka Roy. Optimal deterministic distributed algorithms for maximal independent set in geometric graphs. Journal of Parallel and Distributed Computing, 132(??):36–47, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304113.

Mei:2016:CMR

[MPS16] Alessandro Mei, Natascia Piroso, and Julinda Stefa. Count on me: Reliable broadcast and efficient routing in DTNs through social skeletons. Journal of Parallel and Distributed Computing, 96(??):95–105, October 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300417.

Mei:2012:FGL

[MPV12] Alessandro Mei, Natascia Piroso, and Bruno Vavala. Fine grained load balancing in multi-hop wireless networks. Journal of Parallel and Distributed Computing, 72(4):475–488, April 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000226.

Mudalige:2019:LSP

[MRJ^{+}19] G. R. Mudalige, I. Z. Reguly, S. P. Jammy, C. T. Jacobs, M. B. Giles, and N. D. Sandham. Large-scale performance of a DSL-based multi-block structured-mesh application for direct numerical simulation. Journal of Parallel and Distributed Computing, 131(??):130–146, September 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518305690.

Membarth:2014:TPP

[MRS^{+}14] Richard Membarth, Oliver Reiche, Christian Schmitt, Frank Hannig, Jürgen Teich, Markus Stürmer, and Harald Köstler. Towards a performance-portable description of geometric multigrid algorithms using a domain-specific language. Journal of Parallel and Distributed Computing, 74(12):3191–3201, December 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001506.

Myllykoski:2018:SSB

[MRT18] M. Myllykoski, T. Rossi, and J. Toivanen. On solving separable block tridiagonal linear systems using a GPU implementation of radix-4 PSCR method. Journal of Parallel and Distributed Computing, 115(??):56–66, May 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300170.

Michail:2015:TPP

[MS15] Othon Michail and Paul G. Spirakis. Terminating population protocols via some minimal global knowledge assumptions. Journal of Parallel and Distributed Computing, 81–82(??):1– 10, July 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151500043X.

M:2019:PSB

[MS19] Veeramanikandan M. and Suresh Sankaranarayanan. Publish/ subscribe based multi-tier edge computational model in Internet of Things for latency reduction. Journal of Parallel and Distributed Computing, 127(??):18–27, May 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300152.

Moraveji:2010:CGM

[MSAZ10a] R. Moraveji, H. Sarbazi-Azad, and A. Y. Zomaya. Corrigendum to “A general methodology for direction-based irregular routing algorithms” [J. Parallel Distrib. Comput. 70 (2010) 363–370]. Journal of Parallel and Distributed Computing, 70(5):623, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). See [MSAZ10b].

Moraveji:2010:GMD

[MSAZ10b] R. Moraveji, H. Sarbazi-Azad, and A. Y. Zomaya. A general methodology for direction-based irregular routing algorithms. Journal of Parallel and Distributed Computing, 70(4):363–370, April 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). See corrigendum [MSAZ10a].

Moraveji:2011:PMC

[MSAZ11] Reza Moraveji, Hamid Sarbazi-Azad, and Albert Y. Zomaya. Performance modeling of Cartesian product networks. Journal of Parallel and Distributed Computing, 71(1):105–113, January 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Muhammad:2019:CCA

[MSEM^{+}19] Sayed T. Muhammad, Mohamed Saad, Ali A. El-Moursy, Magdy A. El-Moursy, and Hesham F. A. Hamed. CFPA: Congestion aware, fault tolerant and process variation aware adaptive routing algorithm for asynchronous networks-on-chip. Journal of Parallel and Distributed Computing, 128(??):151– 166, June 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518303794.

Matos:2013:LER

[MSF^{+}13] Miguel Matos, Valerio Schiavoni, Pascal Felber, Rui Oliveira, and Etienne Rivière. Lightweight, efficient, robust epidemic dissemination. Journal of Parallel and Distributed Computing, 73(7):987–999, July 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000269.

McIntosh-Smith:2013:SIJ

[MSGS^{+}13] Simon McIntosh-Smith, Charles Gillan, Nico Sanna, Stan Scott, and Thomas Steinke. Special issue of the Journal of Parallel and Distributed Computing (JDPC) on novel architectures for high-performance computing. Journal of Parallel and Distributed Computing, 73(11):1415–1416, November 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513001792.

Morad:2016:EEF

[MSK^{+}16] Tomer Y. Morad, Noam Shalev, Idit Keidar, Avinoam Kolodny, and Uri C. Weiser. EFS: Energy-Friendly Scheduler for memory bandwidth constrained systems. Journal of Parallel and Distributed Computing, 95(??):3–14, September 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516000368.

Mahmud:2019:QEQ

[MSRB19] Redowan Mahmud, Satish Narayana Srirama, Kotagiri Ramamohanarao, and Rajkumar Buyya. Quality of experience (QoE)-aware placement of applications in Fog computing environments. Journal of Parallel and Distributed Computing, 132(??):190–203, October 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518301771.

Marchal:2019:LMF

[MSV19] Loris Marchal, Bertrand Simon, and Frédéric Vivien. Limiting the memory footprint when dynamically scheduling DAGs on shared-memory platforms. Journal of Parallel and Distributed Computing, 128(??):30–42, June 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305112.

Maurer:2014:BBF

[MT14] Alexandre Maurer and Sébastien Tixeuil. Byzantine broadcast with fixed disjoint paths. Journal of Parallel and Distributed Computing, 74(11):3153–3160, November 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001324.

Maqsood:2018:ECA

[MTL^{+}18a] Tahir Maqsood, Nikos Tziritas, Thanasis Loukopoulos, Sajjad A. Madani, Samee U. Khan, Cheng-Zhong Xu, and Albert Y. Zomaya. Energy and communication aware task mapping for MPSoCs. Journal of Parallel and Distributed Computing, 121(??):15–26, November 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151830217X.

Mencagli:2018:HSW

[MTL^{+}18b] Gabriele Mencagli, Massimo Torquati, Fabio Lucattini, Salvatore Cuomo, and Marco Aldinucci. Harnessing slidingwindow execution semantics for parallel stream processing. Journal of Parallel and Distributed Computing, 116(??):74– 88, ???? 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517302976.

Marathe:2010:FDP

[MTM10] Jaydeep Marathe, Vivek Thakkar, and Frank Mueller. Feedback-directed page placement for ccNUMA via hardwaregenerated memory traces. Journal of Parallel and Distributed Computing, 70(12):1204–1219, December 2010. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mueller:2013:BPI

[Mue13] Frank Mueller. Best papers, IPDPS 2011. Journal of Parallel and Distributed Computing, 73(7):939, July 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000816.

Mohan:2017:SML

[MVP17] Anuraj Mohan, R. Venkatesan, and K. V. Pramod. A scalable method for link prediction in large real world networks. Journal of Parallel and Distributed Computing, 109(??):89–101, November 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517301600.

Mao:2012:WCO

[MXSL12] Huajian Mao, Nong Xiao, Weisong Shi, and Yutong Lu. Wukong: a cloud-oriented file service for mobile Internet devices. Journal of Parallel and Distributed Computing, 72(2):171–184, February 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511002164.

Mishra:2011:RRA

[MYD^{+}11] Asit K. Mishra, Aditya Yanamandra, Reetuparna Das, Soumya Eachempati, Ravi Iyer, N. Vijaykrishnan, and Chita R. Das. RAFT: a router architecture with frequency tuning for on-chip networks. Journal of Parallel and Distributed Computing, 71(5):625–640, May 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731510001929.

Marinescu:2010:SFS

[MYM10] Dan C. Marinescu, Chen Yu, and Gabriela M. Marinescu. Scale-free, self-organizing very large sensor networks. Journal of Parallel and Distributed Computing, 70(5):612–622, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ma:2017:SAA

[MYYY17] Kun Ma, Bo Yang, Zhe Yang, and Ziqiang Yu. Segment access-aware dynamic semantic cache in cloud computing environment. Journal of Parallel and Distributed Computing, 110(??):42–51, December 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301338.

Melab:2018:POU

[MZC18] Nouredine Melab, Albert Y. Zomaya, and Imen Chakroun. Parallel optimization using/for multi and many-core high performance computing. Journal of Parallel and Distributed Computing, 112 (part 2)(??):109–110, February 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303234.

Muppala:2012:RBR

[MZZC12] Sireesha Muppala, Xiaobo Zhou, Liqiang Zhang, and Guihai Chen. Regression-based resource provisioning for session slowdown guarantee in multi-tier Internet servers. Journal of Parallel and Distributed Computing, 72(3):362–375, March 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511002280.

Nicolae:2011:BNG

[NAB^{+}11] Bogdan Nicolae, Gabriel Antoniu, Luc Bougé, Diana Moise, and Alexandra Carpen-Amarie. BlobSeer: Next-generation data management for large scale infrastructures. Journal of Parallel and Distributed Computing, 71(2):169–184, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nicolae:2013:BVD

[NC13] Bogdan Nicolae and Franck Cappello. BlobCR: Virtual disk based checkpoint-restart for HPC applications on IaaS clouds. Journal of Parallel and Distributed Computing, 73(5):698–711, May 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000142.

Navarro:2012:DDT

[NCA^{+}12] Angeles Navarro, Francisco Corbera, Rafael Asenjo, Rosa Castillo, and Emilio L. Zapata. A data dependence test based on the projection of paths over shape graphs. Journal of Parallel and Distributed Computing, 72(12):1547–1564, December 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001918.

Nguyen:2017:ATM

[NCB^{+}17] Tan Nguyen, Pietro Cicotti, Eric Bylaska, Dan Quinlan, and Scott Baden. Automatic translation of MPI source into a latency-tolerant, data-driven form. Journal of Parallel and Distributed Computing, 106(??):1–13, August 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300771.

Nepal:2019:NQS

[NCRK19] Surya Nepal, Mohan Baruwal Chhetri, Rajiv Ranjan, and Ryszard Kowalczyk. A note on quality of service issues in smart cities. Journal of Parallel and Distributed Computing, 127(??):116–117, May 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519301212.

Nie:2012:ESS

[ND12] Pengcheng Nie and Zhenhua Duan. Efficient and scalable scheduling for performance heterogeneous multicore systems. Journal of Parallel and Distributed Computing, 72(3):353–361, March 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511002358.

Naderan:2013:ULB

[NDP13] Marjan Naderan, Mehdi Dehghan, and Hossein Pedram. Upper and lower bounds for dynamic cluster assignment for multi-target tracking in heterogeneous WSNs. Journal of Parallel and Distributed Computing, 73(10):1389–1399, October 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000683.

Newhall:2017:PPD

[NDW17] Tia Newhall, Andrew Danner, and Kevin C. Webb. Pervasive parallel and distributed computing in a liberal arts college curriculum. Journal of Parallel and Distributed Computing, 105(??):53–62, July 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300114.

Neelima:2017:HPC

[Nee17] B. Neelima. High performance computing education in an Indian engineering institute. Journal of Parallel and Distributed Computing, 105(??):73–82, July 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300254.

Nesterov:2010:SPT

[Nes10] Oleksandr Nesterov. A simple parallelization technique with MPI for ocean circulation models. Journal of Parallel and Distributed Computing, 70(1):35–44, January 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nghiem:2016:TER

[NF16] Peter P. Nghiem and Silvia M. Figueira. Towards efficient resource provisioning in MapReduce. Journal of Parallel and Distributed Computing, 95(??):29–41, September 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300077.

Nere:2013:SCN

[NFHL13] Andrew Nere, Sean Franey, Atif Hashmi, and Mikko Lipasti. Simulating cortical networks on heterogeneous multi-GPU systems. Journal of Parallel and Distributed Computing, 73(7):953–971, July 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512000408.

Niu:2012:SPW

[NGQM12] Jianwei Niu, Yuhang Gao, Meikang Qiu, and Zhong Ming. Selecting proper wireless network interfaces for user experience enhancement with guaranteed probability. Journal of Parallel and Distributed Computing, 72(12):1565–1575, December 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512002134.

Nguyen:2013:SBD

[NHO^{+}13] Tan Nguyen, Daniel Hefenbrock, Jason Oberg, Ryan Kastner, and Scott Baden. A software-based dynamic-warp scheduling approach for load-balancing the Viola–Jones face detection algorithm on GPUs. Journal of Parallel and Distributed Computing, 73(5):677–685, May 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000130.

Nai:2019:TAP

[NHX^{+}19] Lifeng Nai, Ramyad Hadidi, He Xiao, Hyojong Kim, Jaewoong Sim, and Hyesoon Kim. Thermal-aware processingin-memory instruction offloading. Journal of Parallel and Distributed Computing, 130(??):193–207, August 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305938.

Nicolae:2016:TSD

[NKK16] Bogdan Nicolae, Andrzej Kochut, and Alexei Karve. Towards scalable on-demand collective data access in IaaS clouds: an adaptive collaborative content exchange proposal. Journal of Parallel and Distributed Computing, 87(??):67–79, January 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515001732.

Neumann:2017:ITH

[NKSA17] Philipp Neumann, Christoph Kowitz, Felix Schranner, and Dmitrii Azarnykh. Interdisciplinary teamwork in HPC education: Challenges, concepts, and outcomes. Journal of Parallel and Distributed Computing, 105(??):83–91, July 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300035.

Nukala:2014:STL

[NKV14] Nishant S. Nukala, Niranjan Kulkarni, and Sarma Vrudhula. Spintronic Threshold Logic Array (STLA) — a compact, low leakage, non-volatile gate array architecture. Journal of Parallel and Distributed Computing, 74(6):2452–2460, June 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513002116.

Nam:2019:SAM

[NL19] Hyunsuk Nam and Roman Lysecky. Security-aware multiobjective optimization of distributed reconfigurable embedded systems. Journal of Parallel and Distributed Computing, 133(??):377–390, November 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300923.

Nan:2018:DTD

[NLB^{+}18] Yucen Nan, Wei Li, Wei Bao, Flavia C. Delicato, Paulo F. Pires, and Albert Y. Zomaya. A dynamic tradeoff data processing framework for delay-sensitive applications in Cloud of Things systems. Journal of Parallel and Distributed Computing, 112 (part 1)(??):53–66, February 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302630.

Nichols:2017:DDB

[NM17] J. M. Nichols and J. V. Michalowicz. Distance distribution between nodes in a 3D wireless network. Journal of Parallel and Distributed Computing, 102(??):71–79, April 2017.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301721.

Nasir:2019:FCE

[NML^{+}19] Mansoor Nasir, Khan Muhammad, Jaime Lloret, Arun Kumar Sangaiah, and Muhammad Sajjad. Fog computing enabled cost-effective distributed summarization of surveillance videos for smart cities. Journal of Parallel and Distributed Computing, 126(??):161–170, April 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308402.

Nghiem:2014:PPB

[NMN^{+}14] Thao P. Nghiem, Kiki Maulana, Kinh Nguyen, David Green, Agustinus Borgy Waluyo, and David Taniar. Peer-to-peer bichromatic reverse nearest neighbours in mobile adhoc networks. Journal of Parallel and Distributed Computing, 74(11):3128–3140, November 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001294.

Neshatpour:2018:EEA

[NMS^{+}18] Katayoun Neshatpour, Maria Malik, Avesta Sasan, Setareh Rafatirad, Tinoush Mohsenin, Hassan Ghasemzadeh, and Houman Homayoun. Energy-efficient acceleration of MapReduce applications using FPGAs. Journal of Parallel and Distributed Computing, 119(??):1–17, September 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300753.

No:2012:NFM

[No12] Jaechun No. NAND flash memory-based hybrid file system for high I/O performance. Journal of Parallel and Distributed Computing, 72(12):1680–1695, December 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151200189X.

Navaridas:2019:IIN

[NPE^{+}19] Javier Navaridas, Jose A. Pascual, Alejandro Erickson, Iain A. Stewart, and Mikel Luján. INRFlow: an interconnection networks research flow-level simulation framework. Journal of Parallel and Distributed Computing, 130(??):140–152, August 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519302242.

Natalizio:2010:RDT

[NPGV10] Enrico Natalizio, Pasquale Pace, Francesca Guerriero, and Antonio Violi. A reactive and dependable transport protocol for wireless mesh networks. Journal of Parallel and Distributed Computing, 70(5):431–442, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Naranjo:2019:FFS

[NPS^{+}19] Paola G. Vinueza Naranjo, Zahra Pooranian, Mohammad Shojafar, Mauro Conti, and Rajkumar Buyya. FOCAN: a fogsupported smart city network architecture for management of applications in the Internet of Everything environments. Journal of Parallel and Distributed Computing, 132(??):274–283, October 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518304775.

Nolting:2019:DSR

[NPVG^{+}19] Stephan Nolting, Guillermo Payá-Vayá, Florian Giesemann, Holger Blume, Sebastian Niemann, and Christian Müller-Schloer. Dynamic self-reconfiguration of a MIPS-based softcore processor architecture. Journal of Parallel and Distributed Computing, 133(??):391–406, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302678.

Nagpal:2012:CAE

[NS12] Rahul Nagpal and Y. N. Srikant. Compiler-assisted energy optimization for clustered VLIW processors. Journal of Parallel and Distributed Computing, 72(8):944–959, August 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512001013.

Nayebi:2011:PML

[NSA11] Abbas Nayebi and Hamid Sarbazi-Azad. Performance modeling of the LEACH protocol for mobile wireless sensor networks. Journal of Parallel and Distributed Computing, 71(6):812–821, June 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511000360.

Nam:2010:MQS

[NSAS10] Beomseok Nam, Minho Shin, Henrique Andrade, and Alan Sussman. Multiple query scheduling for distributed semantic caches. Journal of Parallel and Distributed Computing, 70 (5):598–611, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ngo:2018:UCR

[NSDZ18] Linh B. Ngo, Ashwin Trikuta Srinath, Jeffrey Denton, and Marcin Ziolkowski. Unifying computing resources and access interface to support parallel and distributed computing education. Journal of Parallel and Distributed Computing, 118 (part 1)(??):201–212, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0743731518300984.

Nagasu:2017:FBT

[NSKN17] Kohei Nagasu, Kentaro Sano, Fumiya Kono, and Naohito Nakasato. FPGA-based tsunami simulation: Performance comparison with GPUs, and roofline model for scalability analysis. Journal of Parallel and Distributed Computing, 106(??):153–169, August 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516301915.

Nakib:2019:PFD

[NST19] A. Nakib, L. Souquet, and E.-G. Talbi. Parallel fractal decomposition based algorithm for big continuous optimization problems. Journal of Parallel and Distributed Computing, 133(??):297–306, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518304003.

Neamatollahi:2012:IBA

[NTN12] Peyman Neamatollahi, Hoda Taheri, and Mahmoud Naghibzadeh. Info-based approach in distributed mutual exclusion algorithms. Journal of Parallel and Distributed Computing, 72 (5):650–665, May 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731512000160.

Nabi:2019:FDS

[NV19] Syed Waqar Nabi and Wim Vanderbauwhede. FPGA design space exploration for scientific HPC applications using a fast and accurate cost model based on roofline analysis. Journal of Parallel and Distributed Computing, 133(??):407–419, November 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151730165X.

Netto:2011:URT

[NVK^{+}11] Marco A. S. Netto, Christian Vecchiola, Michael Kirley, Carlos A. Varela, and Rajkumar Buyya. Use of run time predictions for automatic co-allocation of multi-cluster resources for iterative parallel applications. Journal of Parallel and Distributed Computing, 71(10):1388–1399, October 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511001031.

Nai:2017:EBG

[NXTK17] Lifeng Nai, Yinglong Xia, Ilie G. Tanase, and Hyesoon Kim. Exploring big graph computing — an empirical study from architectural perspective. Journal of Parallel and Distributed Computing, 108(??):122–137, October 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300910.

Nutaro:2017:HAA

[NZ17] James Nutaro and Bernard Zeigler. How to apply Amdahl’s law to multithreaded multicore processors. Journal of Parallel and Distributed Computing, 107(??):1–2, September 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300941.

Nikolova:2013:PGO

[NZA13] Olga Nikolova, Jaroslaw Zola, and Srinivas Aluru. Parallel globally optimal structure learning of Bayesian networks. Journal of Parallel and Distributed Computing, 73 (8):1039–1048, August 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731513000622.

Nijim:2011:QSA

[NZY^{+}11] Mais Nijim, Ziliang Zong, Shu Yin, Kiranmai Bellam, and Xiao Qin. Quality of security adaptation in parallel disk systems. Journal of Parallel and Distributed Computing, 71(2):288–301, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Omara:2010:GAT

[OA10] Fatma A. Omara and Mona M. Arafa. Genetic algorithms for task scheduling problem. Journal of Parallel and Distributed Computing, 70(1):13–22, January 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Otero:2019:OAA

[OGM^{+}19] Evelyn Otero, Jing Gong, Misun Min, Paul Fischer, Philipp Schlatter, and Erwin Laure. OpenACC acceleration for the P_{N}–P_{N−2} algorithm in Nek5000. Journal of Parallel and Distributed Computing, 132(??):69–78, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305549.

Orts:2012:GIG

[OGRV^{+}12] Sergio Orts, Jose Garcia-Rodriguez, Diego Viejo, Miguel Cazorla, and Vicente Morell. GPGPU implementation of growing neural gas: Application to 3D scene reconstruction. Journal of Parallel and Distributed Computing, 72(10):1361–1372, October 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001268.

Oxley:2018:RBT

[OJP^{+}18] Mark A. Oxley, Eric Jonardi, Sudeep Pasricha, Anthony A. Maciejewski, Howard Jay Siegel, Patrick J. Burns, and Gregory A. Koenig. Rate-based thermal, power, and colocation aware resource management for heterogeneous data centers. Journal of Parallel and Distributed Computing, 112 (part 2)(??):126–139, February 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151730151X.

Othman:2010:EDS

[OM10] Jalel Ben Othman and Lynda Mokdad. Enhancing data security in ad hoc networks based on multipath routing. Journal of Parallel and Distributed Computing, 70(3):309–316, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ouni:2017:MLE

[OMT^{+}17] Bassem Ouni, Imen Mhedbi, Chiraz Trabelsi, Rabie Ben Atitallah, and Cécile Belleudy. Multi-level energy/poweraware design methodology for MPSoC. Journal of Parallel and Distributed Computing, 100(??):203–215, February 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516300041.

Ortin-Obon:2016:RCD

[OOSGVG^{+}16] Marta Ortín-Obón, Darío Suárez-Gracia, María Villarroya-Gaudó, Cruz Izu, and Víctor Viñals. Reactive circuits: Dynamic construction of circuits for reactive traffic in homogeneous CMPs. Journal of Parallel and Distributed Computing, 95(??):57–68, September 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300089.

Orhean:2018:NSA

[OPR18] Alexandru Iulian Orhean, Florin Pop, and Ioan Raicu. New scheduling approach using reinforcement learning for heterogeneous distributed systems. Journal of Parallel and Distributed Computing, 117(??):292–302, July 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301521.

Ochoa-Ruiz:2018:MFE

[ORWT^{+}18] Gilberto Ochoa-Ruiz, Pamela Wattebled, Maamar Touiza, Florent De Lamotte, El-Bay Bourennane, Samy Meftali, Jean-Luc Dekeyser, and Jean-Philippe Diguet. A modeling frontend for seamless design and generation of context-aware Dynamically Reconfigurable Systems-on-Chip. Journal of Parallel and Distributed Computing, 112 (part 1)(??):1–19, February 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302654.

Outrata:2019:PEP

[OT19] Jan Outrata and Martin Trnecka. Parallel exploration of partial solutions in Boolean matrix factorization. Journal of Parallel and Distributed Computing, 123(??):180–191, January 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518306968.

Oz:2012:TVP

[OTKT12] Isil Oz, Haluk Rahmi Topcuoğlu, Mahmut Kandemir, and Oguz Tosun. Thread vulnerability in parallel applications. Journal of Parallel and Distributed Computing, 72(10):1171–1185, October 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001207.

Ostovari:2014:SLR

[OWK14] Pouya Ostovari, Jie Wu, and Abdallah Khreishah. Symbollevel reliable broadcasting of sensitive data in error-prone wireless networks. Journal of Parallel and Distributed Computing, 74(7):2673–2685, July 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151400032X.

Oliker:2013:BPA

[OY13] Leonid Oliker and Katherine Yelick. Best paper awards: 26th International Parallel and Distributed Processing Symposium (IPDPS 2012). Journal of Parallel and Distributed Computing, 73(7):986, July 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000452.

Ozturk:2011:DLP

[Ozt11] Ozcan Ozturk. Data locality and parallelism optimization using a constraint-based approach. Journal of Parallel and Distributed Computing, 71(2):280–287, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Paudel:2015:HPT

[PA15] Jeeva Paudel and José Nelson Amaral. Hybrid parallel task placement in irregular applications. Journal of Parallel and Distributed Computing, 76(??):94–105, February 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001877.

Polig:2018:HCF

[PAG^{+}18] Raphael Polig, Kubilay Atasu, Heiner Giefers, Christoph Hagleitner, Laura Chiticariu, Frederick Reiss, Huaiyu Zhu, and Peter Hofstee. A hardware compilation framework for text analytics queries. Journal of Parallel and Distributed Computing, 111(??):260–272, January 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301727.

Pal:2014:PMH

[PARB14] Anirban Pal, Abhishek Agarwala, Soumyendu Raha, and Baidurya Bhattacharya. Performance metrics in a hybrid MPI–OpenMP based molecular dynamics simulation with short-range interactions. Journal of Parallel and Distributed Computing, 74(3):2203–2214, March 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002505.

Pauli:2015:IFT

[PAS15] Stefan Pauli, Peter Arbenz, and Christoph Schwab. Intrinsic fault tolerance of multilevel Monte Carlo methods. Journal of Parallel and Distributed Computing, 84(??):24–36, October 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515001239.

Pelt:2015:EAS

[PB15] Daniël M. Pelt and Rob H. Bisseling. An exact algorithm for sparse matrix bipartitioning. Journal of Parallel and Distributed Computing, 85(??):79–90, November 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001033.

Park:2019:AOP

[PB19] Jinsu Park and Woongki Baek. Analyzing and optimizing the performance and energy efficiency of transactional scientific applications on large-scale NUMA systems with HTM support. Journal of Parallel and Distributed Computing, 127(??): 1–17, May 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518304635.

Prasad:2017:KTT

[PBB^{+}17] Sushil K. Prasad, Ioana Banicescu, Martina Barnas, Domingo Giménez, and Andrew Lumsdaine. Keeping up with technology: Teaching parallel, distributed and high-performance computing. Journal of Parallel and Distributed Computing, 105(??):1–3, July 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731517300813.

Penmatsa:2011:GTS

[PC11] Satish Penmatsa and Anthony T. Chronopoulos. Gametheoretic static load balancing for distributed systems. Journal of Parallel and Distributed Computing, 71(4):537–555, April 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Paun:2016:FIH

[PCLP16] Andrei Paun, Clayton Chandler, Chokchai Box Leangsuksun, and Mihaela Paun. A failure index for HPC applications. Journal of Parallel and Distributed Computing, 93–94(??):146– 153, July 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516300156.

Panyala:2017:EPE

[PCMM^{+}17] Ajay Panyala, Daniel Chavarría-Miranda, Joseph B. Manzano, Antonino Tumeo, and Mahantesh Halappanavar. Exploring performance and energy tradeoffs for irregular applications: a case study on the Tilera many-core architecture. Journal of Parallel and Distributed Computing, 104(??):234– 251, June 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516300715.

Peng:2011:ISN

[PCX^{+}11] Miao Peng, Hui Chen, Yang Xiao, Suat Ozdemir, Athanasios V. Vasilakos, and Jie Wu. Impacts of sensor node distributions on coverage in sensor networks. Journal of Parallel and Distributed Computing, 71(12):1578–1591, December 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511000864. See retraction

[PCX^{+}14] for abuse of citation policy.

Peng:2014:RNI

[PCX^{+}14] Miao Peng, Hui Chen, Yang Xiao, Suat Ozdemir, Athanasios V. Vasilakos, and Jie Wu. Retraction notice to: “Impacts of sensor node distributions on coverage in sensor networks” [J. Parallel Distrib. Comput. (2011) 1578–1591]. Journal of Parallel and Distributed Computing, 74(5):2438, May 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514000203. See [PCX^{+}11].

Palchaudhuri:2019:DAV

[PD19] Ayan Palchaudhuri and Anindya Sundar Dhar. Design and automation of VLSI architectures for bidirectional scan based fault localization approach in FPGA fabric aware cellular automata topologies. Journal of Parallel and Distributed Computing, 130(??):110–125, August 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519302515.

Pinel:2013:SVL

[PDB13] Frédéric Pinel, Bernabé Dorronsoro, and Pascal Bouvry. Solving very large instances of the scheduling of independent tasks problem on the GPU. Journal of Parallel and Distributed Computing, 73(1):101–110, January 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000627.

Papakostas:2017:PPC

[PDP17] G. A. Papakostas, K. I. Diamantaras, and T. Papadimitriou. Parallel pattern classification utilizing GPU-based kernelized Slackmin algorithm. Journal of Parallel and Distributed Computing, 99(??):90–99, January 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151630106X.

Peng:2011:EVF

[Pen11]  Kun Peng.  Efficient VSS free of computational as-

sumption. Journal of Parallel and Distributed Computing, 71(12):1592–1597, December 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001444.

Petrica:2018:FOC

[Pet18] Lucian Petrica. FPGA optimized cellular automaton random number generator. Journal of Parallel and Distributed Computing, 111(??):251–259, January 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301892.

Petra:2019:MDQ

[Pet19] Cosmin G. Petra. A memory-distributed quasi-Newton solver for nonlinear programming problems with a small number of general constraints. Journal of Parallel and Distributed Computing, 133(??):337–348, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518307731.

Puente:2019:CCI

[PG19] Valentin Puente and José Ángel Gregorio. CLASSIC: a cortexinspired hardware accelerator. Journal of Parallel and Distributed Computing, 134(??):140–152, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518307706.

Prasad:2018:VEK

[PGKV18] Sushil K. Prasad, Sheikh Ghafoor, Christos Kaklamanis, and Ramachandran Vaidyanathan. VSI: Edu-2016 — keeping up with technology: Teaching parallel, distributed and highperformance computing. Journal of Parallel and Distributed Computing, 118 (part 1)(??):118–119, August 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0743731518302144.

Patterson:2012:SCM

[PGP^{+}12] Cameron Patterson, Jim Garside, Eustace Painkras, Steve Temple, Luis A. Plana, Javier Navaridas, Thomas Sharp, and Steve Furber. Scalable communications for a million-core neural processing architecture. Journal of Parallel and Distributed Computing, 72(11):1507–1520, November 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000287.

Pedrero:2017:ROS

[PGRP17] Manuel Pedrero, Eladio Gutierrez, Sergio Romero, and Oscar Plata. ReduxSTM: Optimizing STM designs for irregular applications. Journal of Parallel and Distributed Computing, 107(??):114–133, September 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301314.

Phatak:2016:NDA

[PH16] Dhananjay S. Phatak and Steven D. Houston. New distributed algorithms for fast sign detection in residue number systems (RNS). Journal of Parallel and Distributed Computing, 97(??):78–95, November 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300703.

Park:2018:AAA

[PH18] Jong Hyuk Park and Houcine Hassan. Advanced algorithms and applications for IoT cloud computing convergence. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):265–266, August 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518302727.

Phillips:2013:BPI

[Phi13] Cynthia Phillips. Best papers, IPDPS 2010. Journal of Parallel and Distributed Computing, 73(7):897, July 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000889.

Pennycook:2013:IPP

[PHW^{+}13] S. J. Pennycook, S. D. Hammond, S. A. Wright, J. A. Herdman, I. Miller, and S. A. Jarvis. An investigation of the performance portability of OpenCL. Journal of Parallel and Distributed Computing, 73(11):1439–1450, November 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512001669.

Piccialli:2018:TID

[PJ18] Francesco Piccialli and Jason J. Jung. Towards the Internet of Data: Applications, opportunities and future challenges. Journal of Parallel and Distributed Computing, 116(??):1–2, ???? 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151830162X.

Page:2010:MHD

[PKN10] Andrew J. Page, Thomas M. Keane, and Thomas J. Naughton. Multi-heuristic dynamic task allocation using genetic algorithms in a heterogeneous distributed system. Journal of Parallel and Distributed Computing, 70(7):758–766, July 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Penoff:2010:ETL

[PKW^{+}10] Brad Penoff, Humaira Kamal, Alan Wagner, Mike Tsai, Karol Mroz, and Janardhan Iyengar. Employing transport layer multi-railing in cluster networks. Journal of Parallel and Distributed Computing, 70(3):259–269, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Plantenga:2013:ISI

[Pla13] Todd Plantenga. Inexact subgraph isomorphism in MapReduce. Journal of Parallel and Distributed Computing, 73(2): 164–175, February 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731512002559.

Pang:2014:MSN

[PLD14] Jun Pang, Alvin R. Lebeck, and Christopher Dwyer. Modeling and simulation of a nanoscale optical computing system. Journal of Parallel and Distributed Computing, 74(6):2470– 2483, June 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513001329.

Pulido:2018:RVA

[PLK^{+}18] Jesus Pulido, Daniel Livescu, Kalin Kanov, Randal Burns, Curtis Canada, James Ahrens, and Bernd Hamann. Remote visual analysis of large turbulence databases at multiple scales. Journal of Parallel and Distributed Computing, 120(??):115–126, October 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518303927.

Pan:2018:VAF

[PLSM18] Zheng Pan, Shuai Liu, Arun Kumar Sangaiah, and Khan Muhammad. Visual attention feature (VAF): a novel strategy for visual tracking based on cloud platform in intelligent surveillance systems. Journal of Parallel and Distributed Computing, 120(??):182–194, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304532.

Park:2015:TSA

[PLY15] Junyoung Park, Sunggu Lee, and Sungjoo Yoo. Time slot assignment for convergecast in wireless sensor networks. Journal of Parallel and Distributed Computing, 83(??):70–82, September 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515000945.

Pascual:2011:OBM

[PMAL11] Jose A. Pascual, Jose Miguel-Alonso, and Jose A. Lozano. Optimization-based mapping framework for parallel applications. Journal of Parallel and Distributed Computing, 71(10):1377–1387, October 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001328.

Prasad:2018:DEZ

[PMCC18] N. Prasad, Priyajit Mukherjee, Santanu Chattopadhyay, and Indrajit Chakrabarti. Design and evaluation of ZMesh topology for on-chip interconnection networks. Journal of Parallel and Distributed Computing, 113(??):17–36, March 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302873.

Plentz:2011:PMD

[PMdO11] P. D. M. Plentz, C. Montez, and R. S. de Oliveira. AS prediction mechanism for distributed threads systems. Journal of Parallel and Distributed Computing, 71(10):1367–1376, October 2011. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511000888.

Pattanayak:2019:GMR

[PMHM19] Debasish Pattanayak, Kaushik Mondal, Ramesh H., and Partha Sarathi Mandal. Gathering of mobile robots with weak multiplicity detection in presence of crash-faults. Journal of Parallel and Distributed Computing, 123(??):145–155, January 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151830697X.

Perez-Miguel:2015:MAC

[PMMMA15] Carlos Pérez-Miguel, Alexander Mendiburu, and Jose Miguel-Alonso. Modeling the availability of Cassandra. Journal of Parallel and Distributed Computing, 86(??):29–44, December 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151500129X.

Park:2013:EHE

[PP13] Alfred J. Park and Kalyan S. Perumalla. Efficient heterogeneous execution on large multicore and accelerator platforms: Case study using a block tridiagonal solver. Journal of Parallel and Distributed Computing, 73(12):1578–1591, December 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513001500.

Pal:2014:RDR

[PPP14] Rajesh Kumar Pal, Kolin Paul, and Sanjiva Prasad. ReKonf: Dynamically reconfigurable multiCore architecture. Journal of Parallel and Distributed Computing, 74(11):3071–3086, November 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001014.

Pietracaprina:2015:SEP

[PPSV15] A. Pietracaprina, G. Pucci, F. Silvestri, and F. Vandin. Spaceefficient parallel algorithms for combinatorial search problems. Journal of Parallel and Distributed Computing, 76(??):58–65, February 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001701.

Paszynski:2010:PDS

[PPTV^{+}10] Maciej Paszyński, David Pardo, Carlos Torres-Verdín, Leszek Demkowicz, and Victor Calo. A parallel direct solver for the self-adaptive hp Finite Element Method. Journal of Parallel and Distributed Computing, 70(3):270–281, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pellegrini:2019:CSE

[PQ19] Alessandro Pellegrini and Francesco Quaglia. Cross-state events: a new approach to parallel discrete event simulation and its speculative runtime support. Journal of Parallel and Distributed Computing, 132(??):48–68, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303715.

Pal:2012:SLT

[PR12] Soumitra Pal and Abhiram Ranade. Scheduling light-trails on WDM rings. Journal of Parallel and Distributed Computing, 72(10):1226–1236, October 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001281.

Pichel:2013:SMV

[PR13] Juan C. Pichel and Francisco F. Rivera. Sparse matrixvector multiplication on the Single-Chip Cloud Computer many-core processor. Journal of Parallel and Distributed Computing, 73(12):1539–1550, December 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151300155X.

Prasanna:2016:MEC

[Pra16] Viktor K. Prasanna. Message from the Editor-in-Chief. Journal of Parallel and Distributed Computing, 87(??):iii, January 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151500204X.

Puthal:2019:SAL

[PRN^{+}19] Deepak Puthal, Rajiv Ranjan, Ashish Nanda, Priyadarsi Nanda, Prem Prakash Jayaraman, and Albert Y. Zomaya. Secure authentication and load balancing of distributed edge datacenters. Journal of Parallel and Distributed Computing, 124(??):60–69, February 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151830741X.

Prasanna:2014:IJS

[PRS14] Viktor K. Prasanna, Yves Robert, and Per Stenström. Introduction to the JPDC special issue on Perspectives on Parallel and Distributed Processing. Journal of Parallel and Distributed Computing, 74(7):2543, July 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000380.

Plimpton:2014:SDA

[PS14] Steven J. Plimpton and Tim Shead. Streaming data analytics via message passing with application to graph algorithms. Journal of Parallel and Distributed Computing, 74(8):2687–2698, August 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514000884.

Perez:2019:ATO

[PSB^{+}19] B. Pérez, E. Stafford, J. L. Bosque, R. Beivide, S. Mateo, X. Teruel, X. Martorell, and E. Ayguadé. Auto-tuned OpenCL kernel co-execution in OmpSs for heterogeneous systems. Journal of Parallel and Distributed Computing, 125(??):45–57, March 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518308189.

Peng:2016:BHS

[PSC^{+}16] Mengfei Peng, Wei Shi, Jean-Pierre Corriveau, Richard Pazzi, and Yang Wang. Black hole search in computer networks: State-of-the-art, challenges and future directions. Journal of Parallel and Distributed Computing, 88(??):1–15, February 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515001914.

Petit:2017:ROC

[PSGS17] Salvador Petit, Julio Sahuquillo, María E. Gómez, and Vicent Selfa. A research-oriented course on advanced multicore architecture: Contents and active learning methodologies. Journal of Parallel and Distributed Computing, 105(??): 63–72, July 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517300175.

Patt-Shamir:2012:DAC

[PSRS12] Boaz Patt-Shamir, Dror Rawitz, and Gabriel Scalosub. Distributed approximation of cellular coverage. Journal of Parallel and Distributed Computing, 72(3):402–408, March 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511002334.

Panigrahi:2019:DAM

[PST^{+}19] Chhabi Rani Panigrahi, Joy Lal Sarkar, Mayank Tiwary, Bibudhendu Pati, and Prasant Mohapatra. DATALET: an approach to manage big volume of data in cyber foraged environment. Journal of Parallel and Distributed Computing, 131(??):14–28, September 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519302680.

Peng:2013:SSM

[PTK^{+}13] Liu Peng, Guangming Tan, Rajiv K. Kalia, Aiichiro Nakano, Priya Vashishta, Dongrui Fan, Hao Zhang, and Fenglong Song. Scalability study of molecular dynamics simulation on Godson-T many-core architecture. Journal of Parallel and Distributed Computing, 73(11):1469–1482, November 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512001682.

Pena:2019:OCQ

[PTN^{+}19] David Peña, Andrei Tchernykh, Sergio Nesmachnow, Renzo Massobrio, Alexander Feoktistov, Igor Bychkov, Gleb Radchenko, Alexander Yu. Drozdov, and Sergey N. Garichev. Operating cost and quality of service optimization for multivehicle-type timetabling for urban bus systems. Journal of Parallel and Distributed Computing, 133(??):272–285, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518300297.

Panja:2019:HDC

[PV19] Rintu Panja and Sathish S. Vadhiyar. HyPar: a divideand-conquer model for hybrid CPU–GPU graph processing. Journal of Parallel and Distributed Computing, 132(??):8–20, October 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519304149.

Pandey:2018:ROC

[PVP18] Parul Pandey, Hariharasudhan Viswanathan, and Dario Pompili. Robust orchestration of concurrent application workflows in mobile device clouds. Journal of Parallel and Distributed Computing, 120(??):101–114, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303356.

Prades:2017:MTV

[PVRS17] Javier Prades, Blesson Varghese, Carlos Reaño, and Federico Silla. Multi-tenant virtual GPUs for optimising performance of a financial risk application. Journal of Parallel and Distributed Computing, 108(??):28–44, October 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300673.

Petagon:2016:EOA

[PW16] Roselin Petagon and Jeeraporn Werapun. Embedding the optimal all-to-all personalized exchange on multistage interconnection networks. Journal of Parallel and Distributed Computing, 88(??):16–30, February 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001902.

Petagon:2017:VVA

[PW17] Roselin Petagon and Jeeraporn Werapun. VA-DE: Valuable ATAPE with dynamic embedding and super-pipeline scheduling on partitionable multistage interconnection networks. Journal of Parallel and Distributed Computing, 102(??):1–15, April 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151630168X.

Park:2010:ISP

[PYP^{+}10] Jung-Wook Park, Hoon-Mo Yang, Gi-Ho Park, Shin-Dug Kim, and Charles C. Weems. An instruction-systolic programmable shader architecture for multi-threaded 3D graphics processing. Journal of Parallel and Distributed Computing, 70(11):1110– 1118, November 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Qiu:2017:SIS

[QGB^{+}17] Meikang Qiu, Saurabh Garg, Rujkumar Buyya, Bei Yu, and Shiyan Hu. Special issue on scalable cyber-physical systems. Journal of Parallel and Distributed Computing, 103(??):1–2, May 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300448.

Quislant:2017:ESB

[QGZP17] Ricardo Quislant, Eladio Gutierrez, Emilio L. Zapata, and Oscar Plata. Enhancing scalability in best-effort hardware transactional memory systems. Journal of Parallel and Distributed Computing, 104(??):73–87, June 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300084.

Quislant:2019:IHT

[QGZP19] Ricardo Quislant, Eladio Gutierrez, Emilio L. Zapata, and Oscar Plata. Improving hardware transactional memory parallelization of computational geometry algorithms using privatizing transactions. Journal of Parallel and Distributed Computing, 131(??):103–119, September 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518302247.

Rezvani:2011:SBM

[RA11] Mohammad Hossein Rezvani and Morteza Analoui. Strategic behavior modeling of multi-service overlay multicast networks based on auction mechanism design. Journal of Parallel and Distributed Computing, 71(8):1125–1141, August 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511000839.

Rho:2017:SPD

[RAN^{+}17] Jaeyong Rho, Takuya Azumi, Mayo Nakagawa, Kenya Sato, and Nobuhiko Nishio. Scheduling parallel and distributed processing for automotive data stream management system. Journal of Parallel and Distributed Computing, 109(??):286–300, November 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517302010.

Rao:2016:EPS

[Rao16] Dhananjai M. Rao. Efficient parallel simulation of spatiallyexplicit agent-based epidemiological models. Journal of Parallel and Distributed Computing, 93–94(??):102–119, July 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300107.

Rao:2012:AMH

[RB12] Jayanthi Rao and Subir Biswas. Analyzing multi-hop routing feasibility for sensor data harvesting using mobile sinks. Journal of Parallel and Distributed Computing, 72(6):764– 777, June 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512000615.

Reuther:2018:SSS

[RBA^{+}18] Albert Reuther, Chansup Byun, William Arcand, David Bestor, Bill Bergeron, Matthew Hubbell, Michael Jones, Peter Michaleas, Andrew Prout, Antonio Rosa, and Jeremy Kepner. Scalable system scheduling for HPC and big data. Journal of Parallel and Distributed Computing, 111(??):76–92, January 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517301983.

Russo:2017:MPG

[RBB17] Igor L. S. Russo, Heder S. Bernardino, and Helio J. C. Barbosa. A massively parallel grammatical evolution technique with OpenCL. Journal of Parallel and Distributed Computing, 109(??):333–349, November 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151730206X.

Rettkowski:2017:HSC

[RBG17] Jens Rettkowski, Andrew Boutros, and Diana Göhringer. HW/SW co-design of the HOG algorithm on a Xilinx Zynq SoC. Journal of Parallel and Distributed Computing, 109(??):50–62, November 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301569.

Rafique:2011:CAF

[RBN11] M. Mustafa Rafique, Ali R. Butt, and Dimitrios S. Nikolopoulos. A capabilities-aware framework for using computational accelerators in data-intensive computing. Journal of Parallel and Distributed Computing, 71(2):185–197, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rinke:2018:SAS

[RBOH^{+}18] Sebastian Rinke, Markus Butz-Ostendorf, Marc-André Hermanns, Mikaël Naveau, and Felix Wolf. A scalable algorithm for simulating the structural plasticity of the brain. Journal of Parallel and Distributed Computing, 120(??):251–266, October 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517303313.

Rezende:2011:IMM

[RBP^{+}11] Cristiano Rezende, Azzedine Boukerche, Richard W. Pazzi, Bruno P. S. Rocha, and Antonio A. F. Loureiro. The impact of mobility on Mobile Ad Hoc Networks through the perspective of complex networks. Journal of Parallel and Distributed Computing, 71(9):1189–1200, September 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731510002789.

Ramakrishnan:2011:DSW

[RCG^{+}11] Lavanya Ramakrishnan, Jeffrey S. Chase, Dennis Gannon, Daniel Nurmi, and Rich Wolski. Deadline-sensitive workflow orchestration without explicit resource control. Journal of Parallel and Distributed Computing, 71(3):343–353, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rapolu:2018:VAS

[RCG18] Naresh Rapolu, Srimat Chakradhar, and Ananth Grama. VAYU: Accelerating stream processing applications through dynamic network-aware topology re-optimization. Journal of Parallel and Distributed Computing, 111(??):13–23, January 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302125.

Rodrigues:2018:DME

[RDA18] Luiz A. Rodrigues, Elias P. Duarte, and Luciana Arantes. A distributed k-mutual exclusion algorithm based on autonomic spanning trees. Journal of Parallel and Distributed Computing, 115(??):41–55, May 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151830025X.

Rughetti:2017:MLB

[RDCQ17] Diego Rughetti, Pierangelo Di Sanzo, Bruno Ciciani, and Francesco Quaglia. Machine learning-based thread-parallelism regulation in software transactional memory. Journal of Parallel and Distributed Computing, 109(??):208–229, November 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517301909.

Rahman:2010:CPK

[REK10a] Sk. Md. Mizanur Rahman and Khalil El-Khatib. Corrigendum to “Private key agreement and secure communication for heterogeneous sensor networks” [J. Parallel Distrib. Comput. 70 (2010) 858–870]. Journal of Parallel and Distributed Computing, 70(11):1174, November 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). See [REK10b].

Rahman:2010:PKA

[REK10b] Sk. Md. Mizanur Rahman and Khalil El-Khatib. Private key agreement and secure communication for heterogeneous sensor networks. Journal of Parallel and Distributed Computing, 70 (8):858–870, August 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). See corrigendum [REK10a].

Ren:2011:ALP

[Ren11] Da Qi Ren. Algorithm level power efficiency optimization for CPU–GPU processing element in data intensive SIMD/SPMD computing. Journal of Parallel and Distributed Computing, 71 (2):245–253, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Razafindralambo:2017:SSP

[REZN17] Tahiry Razafindralambo, Milan Erdelj, Dimitrios Zorbas, and Enrico Natalizio. Spread and shrink: Point of interest discovery and coverage with mobile wireless sensors. Journal of Parallel and Distributed Computing, 102(??):16–27, April 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301472.

Regier:2019:CVU

[RFP^{+}19] Jeffrey Regier, Keno Fischer, Kiran Pamnany, Andreas Noack, Jarrett Revels, Maximilian Lama, Steve Howard, Ryan Giordano, David Schlegel, Jon McAuliffe, Rollin, Thomas, and Prabhat. Cataloging the visible universe through Bayesian inference in Julia at petascale. Journal of Parallel and Distributed Computing, 127(??):89–104, May 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304672.

Rajasekaran:2012:CEE

[RFS^{+}12] Sanguthevar Rajasekaran, Lance Fiondella, Dolly Sharma, Reda Ammar, and Nicholas Lownes. Communication and energy efficient routing protocols for single-hop radio networks. Journal of Parallel and Distributed Computing, 72(6):819–826, June 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151200069X.

Remis:2018:ESN

[RGAN18] Luis Remis, Maria Jesus Garzaran, Rafael Asenjo, and Angeles Navarro. Exploiting social network graph characteristics for efficient BFS on heterogeneous chips. Journal of Parallel and Distributed Computing, 120(??):282–294, October 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517303015.

Roy:2012:WAX

[RHH12] Amitabha Roy, Steven Hand, and Tim Harris. Weak atomicity for the x86 memory consistency model. Journal of Parallel and Distributed Computing, 72(10):1306–1317, October 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512001402.

Roh:2011:RAD

[RJKL11] Hyun-Gul Roh, Myeongjae Jeon, Jin-Soo Kim, and Joonwon Lee. Replicated abstract data types: Building blocks for collaborative applications. Journal of Parallel and Distributed Computing, 71(3):354–368, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

R:2018:EDP

[RK18] Elakkiya R. and Selvamani K. Enhanced dynamic programming approach for subunit modelling to handle segmentation and recognition ambiguities in sign language. Journal of Parallel and Distributed Computing, 117(??):246–255, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302137.

Rahmani:2016:SIE

[RLA^{+}16] Amir M. Rahmani, Pasi Liljeberg, Jose L. Ayala, Hannu Tenhunen, and Alexander V. Veidenbaum. Special issue on energy efficient multi-core and many-core systems, Part I. Journal of Parallel and Distributed Computing, 95(??):1–2, September 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516300272.

Rahmani:2017:SIE

[RLA^{+}17] Amir M. Rahmani, Pasi Liljeberg, Jose L. Ayala, Hannu Tenhunen, and Alexander V. Veidenbaum. Special issue on energy efficient multi-core and many-core systems, Part II. Journal of Parallel and Distributed Computing, 100(??):128–129, February 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301289.

Rajasegarar:2014:HCB

[RLP14] Sutharshan Rajasegarar, Christopher Leckie, and Marimuthu Palaniswami. Hyperspherical cluster based distributed anomaly detection in wireless sensor networks. Journal of Parallel and Distributed Computing, 74(1):1833–1847, January 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513002013.

Rachuri:2010:SER

[RM10] Kiran K. Rachuri and C. Siva Ram Murthy. On the scalability of expanding ring search for dense wireless sensor networks. Journal of Parallel and Distributed Computing, 70(9):917–929, September 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rachuri:2011:EEL

[RM11] Kiran K. Rachuri and C. Siva Ram Murthy. Energy efficient and low latency biased walk techniques for search in wireless sensor networks. Journal of Parallel and Distributed Computing, 71(3):512–522, March 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Reguly:2019:IRS

[RMGM19] I. Z. Reguly, G. R. Mudalige, M. B. Giles, and S. Maheswaran. Improving resilience of scientific software through a domain-specific approach. Journal of Parallel and Distributed Computing, 128(??):99–114, June 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519300917.

Rezaei:2017:DDN

[RMHR17] Arash Rezaei, Frank Mueller, Paul Hargrove, and Eric Roman. DINO: Divergent node cloning for sustained redundancy in HPC. Journal of Parallel and Distributed Computing, 109(??):350–362, November 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301995.

Requena:2014:EDP

[RMU14] Manuel Jesús Martín Requena, Pablo Moscato, and Manuel Ujaldón. Efficient data partitioning for the GPU computation of moment functions. Journal of Parallel and Distributed Computing, 74(1):1994–2004, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001342.

Rios:2018:EPM

[ROB^{+}18] Eyder Rios, Luiz Satoru Ochi, Cristina Boeres, Vitor N. Coelho, Igor M. Coelho, and Ricardo Farias. Exploring parallel multi-GPU local search strategies in a metaheuristic framework. Journal of Parallel and Distributed Computing, 111(??):39–55, January 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302009.

Rodrigo:2018:TUH

[RÖE^{+}18] Gonzalo P. Rodrigo, P.-O. Östberg, Erik Elmroth, Katie Antypas, Richard Gerber, and Lavanya Ramakrishnan. Towards understanding HPC users and systems: a NERSC case study. Journal of Parallel and Distributed Computing, 111(??):206–221, January 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302563.

Rodriguez:2019:PEA

[RPN19] Sebastián Rodríguez, Facundo Parodi, and Sergio Nesmachnow. Parallel evolutionary approaches for game playing and verification using Intel Xeon Phi. Journal of Parallel and Distributed Computing, 133(??):258–271, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151830491X.

Reano:2019:APP

[RPS19] Carlos Reaño, Javier Prades, and Federico Silla. Analyzing the performance/power tradeoff of the rCUDA middleware for future exascale systems. Journal of Parallel and Distributed Computing, 132(??):344–362, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303491.

Reano:2019:SIN

[RS19] Carlos Reaño and Federico Silla. On the support of internode P2P GPU memory copies in rCUDA. Journal of Parallel and Distributed Computing, 127(??):28–43, May 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300255.

Rolinger:2019:PCS

[RSK19] Thomas B. Rolinger, Tyler A. Simon, and Christopher D. Krieger. Performance considerations for scalable parallel tensor decomposition. Journal of Parallel and Distributed Computing, 129(??):83–98, July 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302897.

Ribeiro:2012:PDN

[RSL12] Pedro Ribeiro, Fernando Silva, and Luís Lopes. Parallel discovery of network motifs. Journal of Parallel and Distributed Computing, 72(2):144–154, February 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001729.

Ranjan:2019:NTT

[RSVW19] Rajiv Ranjan, Ellis Solaiman, Massimo Villari, and Paul Watson. A note on tools and techniques for end-to-end QoS monitoring in Internet of Things. Journal of Parallel and Distributed Computing, 132(??):160–163, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303740.

Ranjbari:2018:LAB

[RT18] Milad Ranjbari and Javad Akbari Torkestani. A learning automata-based algorithm for energy and SLA efficient consolidation of virtual machines in cloud data centers. Journal of Parallel and Distributed Computing, 113(??):55–62, March 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151730285X.

Rizvandi:2011:SOO

[RTZ11] Nikzad Babaii Rizvandi, Javid Taheri, and Albert Y. Zomaya. Some observations on optimal frequency selection in DVFSbased energy consumption minimization. Journal of Parallel and Distributed Computing, 71(8):1154–1164, August 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511000165.

Raghavan:2013:EAE

[RV13] Hari K. Raghavan and Sathish S. Vadhiyar. Efficient asynchronous executions of AMR computations and visualization on a GPU system. Journal of Parallel and Distributed Computing, 73(6):866–875, June 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000464.

R:2013:REM

[RWB^{+}13] Krish K. R., Guanying Wang, Puranjoy Bhattacharjee, Ali R. Butt, and Chris Gniady. On reducing energy management delays in disks. Journal of Parallel and Distributed Computing, 73(6):823–835, June 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000440.

Soliman:2011:FIP

[SA11] Mostafa I. Soliman and Ghada Y. Abozaid. FPGA implementation and performance evaluation of a high throughput crypto coprocessor. Journal of Parallel and Distributed Computing, 71(8):1075–1084, August 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000815.

Shamszaman:2019:ECC

[SA19] Zia Ush Shamszaman and Muhammad Intizar Ali. Enabling cognitive contributory societies using SIoT: QoS aware realtime virtual object management. Journal of Parallel and Distributed Computing, 123(??):61–68, January 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306518.

Soliman:2013:SMU

[SAJ13] Mostafa I. Soliman and Abdulmajid F. Al-Junaid. A shared matrix unit for a chip multi-core processor. Journal of Parallel and Distributed Computing, 73(8):1146–1156, August 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000488.

Sanchez-Artigas:2010:EPP

[SAL10] Marc Sánchez-Artigas and Pedro García López. Echo: a peerto-peer clustering framework for improving communication in DHTs. Journal of Parallel and Distributed Computing, 70(2): 126–143, February 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Samarah:2018:TAR

[SAR^{+}18] Samer Samarah, Mohammed Gh. Al Zamil, Majdi Rawashdeh, M. Shamim Hossain, Ghulam Muhammad, and Atif Alamri. Transferring activity recognition models in FOG computing architecture. Journal of Parallel and Distributed Computing, 122(??):122–130, December 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518305410.

Shamsi:2012:PSO

[SB12] Jawwad Shamsi and Monica Brockmeyer. Predictable service overlay networks: Predictability through adaptive monitoring and efficient overlay construction and management. Journal of Parallel and Distributed Computing, 72(1):70–82, January 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511001894.

Sharma:2015:LBD

[SB15] Gokarna Sharma and Costas Busch. A load balanced directory for distributed shared memory objects. Journal of Parallel and Distributed Computing, 78(??):6–24, April 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515000404.

Saule:2012:LBS

[SBÇ12a] Erik Saule, Erdeniz Ö. Bas, and Ümit V. Çatalyürek. Loadbalancing spatially located computations using rectangular partitions. Journal of Parallel and Distributed Computing, 72(10):1201–1214, October 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512001311.

Saule:2012:OSI

[SBÇ12b] Erik Saule, Doruk Bozdag, and Ümit V. Çatalyürek. Optimizing the stretch of independent tasks on a cluster: From sequential tasks to moldable tasks. Journal of Parallel and Distributed Computing, 72(4):489–503, April 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511002474.

Sulyok:2019:LOU

[SBRM19] András Attila Sulyok, Gábor Dániel Balogh, István Z. Reguly, and Gihan R. Mudalige. Locality optimized unstructured mesh algorithms on GPUs. Journal of Parallel and Distributed Computing, 134(??):50–64, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519301698.

Sun:2010:RAL

[SC10] Xian-He Sun and Yong Chen. Reevaluating Amdahl’s law in the multicore era. Journal of Parallel and Distributed Computing, 70(2):183–188, February 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sarkar:2010:POF

[SCG10] Arnab Sarkar, P. P. Chakrabarti, and Sujoy Ghose. Partition oriented frame based fair scheduler. Journal of Parallel and Distributed Computing, 70(7):707–718, July 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schulz:2013:ELS

[Sch13] Christian Schulz. Efficient local search on the GPUinvestigations on the vehicle routing problem. Journal of Parallel and Distributed Computing, 73(1):14–31, January 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000640.

Schreiber:2014:FBI

[Sch14] Robert Schreiber. A few bad ideas on the way to the triumph of parallel computing. Journal of Parallel and Distributed Computing, 74(7):2544–2547, July 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002177.

Su:2010:ESQ

[SCLL10] I-Fang Su, Yu-Chi Chung, Chiang Lee, and Yi-Ying Lin. Efficient skyline query processing in wireless sensor networks.

Journal of Parallel and Distributed Computing, 70(6):680–698, June 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Silber-Chaussumier:2013:GDT

[SCMH13] F. Silber-Chaussumier, A. Muller, and R. Habel. Generating data transfers for distributed GPU parallel programs. Journal of Parallel and Distributed Computing, 73(12):1649–1660, December 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513001603.

Shestak:2012:PRA

[SCMS12] Vladimir Shestak, Edwin K. P. Chong, Anthony A. Maciejewski, and Howard Jay Siegel. Probabilistic resource allocation in heterogeneous distributed systems with random failures. Journal of Parallel and Distributed Computing, 72(10):1186–1194, October 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512000688.

Salarian:2012:CWS

[SCN12] Hamidreza Salarian, Kwan-Wu Chin, and Fazel Naghdy. Coordination in wireless sensor-actuator networks: a survey. Journal of Parallel and Distributed Computing, 72(7):856–867, July 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000573.

Shen:2018:DAR

[SCW^{+}18] Bo Shen, Naveen Chilamkurti, Ru Wang, Xingshe Zhou, Shiwei Wang, and Wen Ji. Deadline-aware rate allocation for IoT services in data center network. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):296–306, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302666.

Singh:2017:NAA

[SDG17] Amreek Singh, Kusum Deep, and Pallavi Grover. A novel approach to accelerate calibration process of a k-nearest neighbours classifier using GPU. Journal of Parallel and Distributed Computing, 104(??):114–129, June 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300096.

Shriraman:2010:ITD

[SDS10] Arrvindh Shriraman, Sandhya Dwarkadas, and Michael L. Scott. Implementation tradeoffs in the design of flexible transactional memory support. Journal of Parallel and Distributed Computing, 70(10):1068–1084, October 2010. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Santos:2018:DSA

[SDS^{+}18] Ricardo Santos, Liana Duenha, Ana Caroline Silva, Matheus Sousa, Luiz Augusto Tedesco, João Carlos Melgarejo, Tony Santos, Rodolfo Azevedo, and Edward Moreno. Dark-silicon aware design space exploration. Journal of Parallel and Distributed Computing, 120(??):295–306, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303003.

Soliman:2015:SSM

[SE15] Mostafa I. Soliman and Elsayed A. Elsayed. Simple super-matrix processor: Implementation and performance evaluation. Journal of Parallel and Distributed Computing, 83(??):96–118, September 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515000994.

Sun:2017:SIS

[SFC17] Xian-He Sun, Marc Frincu, and Charalampos Chelmis. Special issue on scalable computing systems for big data applications. Journal of Parallel and Distributed Computing, 108(??):1–2, October 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517301776.

Su:2019:ULD

[SFHS19] Yi Su, Dan Feng, Yu Hua, and Zhan Shi. Understanding the latency distribution of cloud object storage systems. Journal of Parallel and Distributed Computing, 128(??):71–83, June 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518301175.

Seelam:2013:ESC

[SFT^{+}13] Seetharami Seelam, Liana Fong, Asser Tantawi, John Lewars, John Divirgilio, and Kevin Gildea. Extreme scale computing: Modeling the impact of system noise in multicore clustered systems. Journal of Parallel and Distributed Computing, 73(7):898–910, July 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000178.

Shi:2014:SBH

[SGAC14] Wei Shi, Joaquin Garcia-Alfaro, and Jean-Pierre Corriveau. Searching for a black hole in interconnected networks using mobile agents and tokens. Journal of Parallel and Distributed Computing, 74(1):1945–1958, January 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001640.

Szafaryn:2013:TPA

[SGdSS13] Lukasz G. Szafaryn, Todd Gamblin, Bronis R. de Supinski, and Kevin Skadron. Trellis: Portability across architectures with a high-level framework. Journal of Parallel and Distributed Computing, 73(10):1400–1413, October 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001275.

Sanchez-Gomez:2019:PMO

[SGVRP19] Jesus M. Sanchez-Gomez, Miguel A. Vega-Rodríguez, and Carlos J. Pérez. Parallelizing a multi-objective optimization approach for extractive multi-document text summarization. Journal of Parallel and Distributed Computing, 134(??):166–179, December 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519301583.

Sohanghpurwala:2017:HAS

[SHA17] Ali Asgar Sohanghpurwala, Mohamed W. Hassan, and Peter Athanas. Hardware accelerated SAT solvers — a survey. Journal of Parallel and Distributed Computing, 106(??):170–184, August 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301903.

Sun:2014:COA

[SHC14] Hongyang Sun, Wen-Jing Hsu, and Yangjie Cao. Competitive online adaptive scheduling for sets of parallel jobs with fairness and efficiency. Journal of Parallel and Distributed Computing, 74(3):2180–2192, March 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002372.

SeyedHassani:2019:BIP

[SHK19] Amirreza SeyedHassani, Mohammad Sayad Haghighi, and Ahmad Khonsari. Bayesian inference of private social network links using prior information and propagated data. Journal of Parallel and Distributed Computing, 125(??):72–80, March 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151830830X.

Shih:2013:TTL

[SHL^{+}13] Po-Chi Shih, Kuo-Chan Huang, Che-Rung Lee, I-Hsin Chung, and Yeh-Ching Chung. TLA: Temporal lookahead processor allocation method for heterogeneous multicluster systems. Journal of Parallel and Distributed Computing, 73(12):1661–1672, December 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513001561.

Soni:2019:PMC

[SHRM19] V. Soni, A. Hadjadj, O. Roussel, and G. Moebs. Parallel multi-core and multi-processor methods on point-value multiresolution algorithms for hyperbolic conservation laws. Journal of Parallel and Distributed Computing, 123(??):192–203, January 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518306981.

Sardar:2017:TPB

[SHSH17] Muhammad Usama Sardar, Osman Hasan, Muhammad Shafique, and Jörg Henkel. Theorem proving based Formal Verification of Distributed Dynamic Thermal Management schemes. Journal of Parallel and Distributed Computing, 100(??):157–171, February 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300764.

Su:2013:ADF

[SI13] Gong Su and Arun Iyengar. Avoiding disruptive failovers in transaction processing systems with multiple active nodes. Journal of Parallel and Distributed Computing, 73(5):630–640, May 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000087.

Siewe:2016:TMS

[Sie16] François Siewe. Towards the modelling of secure pervasive computing systems: a paradigm of Context-Aware Secure Action System. Journal of Parallel and Distributed Computing, 87(??):121–144, January 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001756.

Shin:2014:GSE

[SIY14] Seon-Ho Shin, Eun-Jin Im, and MyungKeun Yoon. A grand spread estimator using a graphics processing unit. Journal of Parallel and Distributed Computing, 74(2):2039–2047, February 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513002189.

Salehi:2012:QPA

[SJB12] Mohsen Amini Salehi, Bahman Javadi, and Rajkumar Buyya. QoS and preemption aware scheduling in federated and virtualized Grid computing environments. Journal of Parallel and Distributed Computing, 72(2):231–245, February 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511002073.

Surati:2019:BTP

[SJG19] Shivangi Surati, Devesh C. Jinwala, and Sanjay Garg. BMMItree: a peer-to-peer m-ary tree using 1 − m node splitting for an efficient multidimensional complex query search. Journal of Parallel and Distributed Computing, 125(??):1–17, March 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518307354.

Sun:2011:PPC

[SJS11] Yan Sun, Qiangfeng Jiang, and Mukesh Singhal. A Pre-Processed Cross Link Detection Protocol for geographic routing in mobile ad hoc and sensor networks under realistic environments with obstacles. Journal of Parallel and Distributed Computing, 71(7):1047–1054, July 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000025.

Santander-Jimenez:2019:CAG

[SJVRVVS19] Sergio Santander-Jiménez, Miguel A. Vega-Rodríguez, Jorge Vicente-Viola, and Leonel Sousa. Comparative assessment of GPGPU technologies to accelerate objective functions: a case study on parsimony. Journal of Parallel and Distributed Computing, 126(??):67–81, April 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304350.

Sun:2011:OCB

[SK11] Wen-Hung Sun and Chung-Ta King. ORN: a contentbased approach to improving supplier discovery in P2P VOD networks. Journal of Parallel and Distributed Computing, 71(12):1558–1569, December 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001614.

Seo:2015:AEC

[SKH15] Seog Chung Seo, Taehong Kim, and Seokhie Hong. Accelerating elliptic curve scalar multiplication over GF(2^{m}) on graphic hardwares. Journal of Parallel and Distributed Computing, 75(??):152–167, January 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001646.

Skinderowicz:2016:GBP

[Ski16] Rafal Skinderowicz. The GPU-based parallel Ant Colony System. Journal of Parallel and Distributed Computing, 98(??):48–60, December 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731516300284.

Shen:2014:DSL

[SKK14] Min Shen, Ajay D. Kshemkalyani, and Ashfaq Khokhar. Detecting stable locality-aware predicates. Journal of Parallel and Distributed Computing, 74(1):1971–1983, January 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513002050.

Song:2018:SDS

[SLG^{+}18] Shengli Song, Yishuai Lin, Bin Guo, Qiang Di, and Rong Lv. Scalable Distributed Semantic Network for knowledge management in cyber physical system. Journal of Parallel and Distributed Computing, 118 (part 1)(??):22–33, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S074373151730326X.

Sun:2019:CCO

[SLHS19] Hui Sun, Wei Liu, Jianzhong Huang, and Weisong Shi. Collaborative compaction optimization system using near-data processing for LSM-tree-based key-value stores. Journal of Parallel and Distributed Computing, 131(??):29–43, September 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518308645.

Spacey:2012:ICL

[SLKK12] Simon Spacey, Wayne Luk, Paul H. J. Kelly, and Daniel Kuhn. Improving communication latency with the writeonly architecture. Journal of Parallel and Distributed Computing, 72(12):1617–1627, December 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002146.

Spacey:2013:PPD

[SLKK13] Simon Spacey, Wayne Luk, Daniel Kuhn, and Paul H. J. Kelly. Parallel partitioning for distributed systems using sequential assignment. Journal of Parallel and Distributed Computing, 73(2):207–219, February 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002341.

Shen:2018:ESI

[SLL18] Xipeng Shen, Robert Lovas, and Xiaofei Liao. Editorial for the special issue on in-memory computing. Journal of Parallel and Distributed Computing, 120(??):322, October 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151830385X.

Sao:2019:CAA

[SLV19]  Piyush Sao, Xiaoye S. Li, and Richard Vuduc.  A

communication-avoiding 3D algorithm for sparse LU factorization on heterogeneous systems. Journal of Parallel and Distributed Computing, 131(??):218–234, September 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305197.

Shen:2010:SSC

[SLW10] Jingbo Shen, Jinlong Li, and Xufa Wang. SCDN: Stable Content Distribution Network based on demands. Journal of Parallel and Distributed Computing, 70(9):880–888, September 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sun:2019:SMC

[SLZ^{+}19] Daniel Sun, Guoqiang Li, Yuanyuan Zhang, Liming Zhu, and Raj Gaire. Statistically managing cloud operations for latency-tail-tolerance in IoT-enabled smart cities. Journal of Parallel and Distributed Computing, 127(??):184–195, May 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518300935.

Sousa:2010:AMP

[SMB10] Marcelo S. Sousa, Alba C. M. A. Melo, and Azzedine Boukerche. An adaptive multi-policy grid service for biological sequence comparison. Journal of Parallel and Distributed Computing, 70(2):160–172, February 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Solomonik:2014:MPT

[SMH^{+}14] Edgar Solomonik, Devin Matthews, Jeff R. Hammond, John F. Stanton, and James Demmel. A massively parallel tensor contraction framework for coupled-cluster computations. Journal of Parallel and Distributed Computing, 74(12):3176–3190, December 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151400104X.

Salimi:2014:TSU

[SMO14] Reza Salimi, Homayun Motameni, and Hesam Omranpour. Task scheduling using NSGA II with fuzzy adaptive operators for computational grids. Journal of Parallel and Distributed Computing, 74(5):2333–2350, May 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000161.

Shibata:2018:UDM

[SMO^{+}18] Masahiro Shibata, Toshiya Mega, Fukuhito Ooshita, Hirotsugu Kakugawa, and Toshimitsu Masuzawa. Uniform deployment of mobile agents in asynchronous rings. Journal of Parallel and Distributed Computing, 119(??):92–106, September 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518302156.

Sarma:2015:ERW

[SMP15] Atish Das Sarma, Anisur Rahaman Molla, and Gopal Pandurangan. Efficient random walk sampling in distributed networks. Journal of Parallel and Distributed Computing, 77(??):84–94, March 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515000039.

Shukla:2017:EDB

[SMP17] Shailendra Shukla, Rajiv Misra, and Animesh Prasad. Efficient disjoint boundary detection algorithm for surveillance capable WSNs. Journal of Parallel and Distributed Computing, 109(??):245–257, November 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301910.

Sanchez-Monedero:2011:BFB

[SMPMLVLS11] Javier Sanchez-Monedero, Javier Povedano-Molina, Jose M. Lopez-Vega, and Juan M. Lopez-Soler. Bloom filter-based discovery protocol for DDS middleware. Journal of Parallel and Distributed Computing, 71(10):1305–1317, October 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511000876.

Svendsen:2015:MMC

[SMT15] Michael Svendsen, Arko Provo Mukherjee, and Srikanta Tirthapura. Mining maximal cliques from a large graph using MapReduce: Tackling highly uneven subproblem sizes. Journal of Parallel and Distributed Computing, 79–80(??):104– 114, May 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001531.

Sahoo:2018:SBH

[SMW18] Prasan Kumar Sahoo, Suvendu Kumar Mohapatra, and Shih-Lin Wu. SLA based healthcare big data analysis and computing in cloud network. Journal of Parallel and Distributed Computing, 119(??):121–135, September 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518302454.

Shah:2012:ERT

[SNCP12] Sayed Chhattan Shah, Qurat-Ul-Ain Nizamani, Sajjad Hussain Chauhdary, and Myong-Soon Park. An effective and robust two-phase resource allocation scheme for interdependent tasks in mobile ad hoc computational Grids. Journal of Parallel and Distributed Computing, 72(12):1664–1679, December 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001736.

Salami:2016:PAP

[SNMB16] Bagher Salami, Hamid Noori, Farhad Mehdipour, and Mohammadreza Baharani. Physical-aware predictive dynamic thermal management of multi-core processors. Journal of Parallel and Distributed Computing, 95(??):42–56, September 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151600037X.

Soliman:2013:DIE

[Sol13] Mostafa I. Soliman. Design, implementation, and evaluation of a low-complexity vector-core for executing scalar/ vector instructions. Journal of Parallel and Distributed Computing, 73(6):836–850, June 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000282.

Shieh:2013:ETA

[SP13] Wann-Yun Shieh and Chin-Ching Pong. Energy and transition-aware runtime task scheduling for multicore processors. Journal of Parallel and Distributed Computing, 73(9):1225–1238, September 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513001007.

Saez:2017:TCF

[SPC^{+}17] Juan Carlos Saez, Adrian Pousa, Fernando Castro, Daniel Chaver, and Manuel Prieto-Matias. Towards completely fair scheduling on asymmetric single-ISA multicore processors. Journal of Parallel and Distributed Computing, 102(??):115– 131, April 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301873.

Seal:2013:RPC

[SPH13] Sudip K. Seal, Kalyan S. Perumalla, and Steven P. Hirshman. Revisiting parallel cyclic reduction and parallel prefixbased algorithms for block tridiagonal systems of equations. Journal of Parallel and Distributed Computing, 73(2):273–280, February 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512002535.

Sousa:2019:DFA

[SPPA19] Rafael Sousa, Marcio Pereira, Fernando Magno Quintão Pereira, and Guido Araujo. Data-flow analysis and optimization for data coherence in heterogeneous architectures. Journal of Parallel and Distributed Computing, 130(??):126–139, August 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519302631.

Saeed:2012:HPM

[SPRG^{+}12] Fahad Saeed, Alan Perez-Rathke, Jaroslaw Gwarnicki, Tanya Berger-Wolf, and Ashfaq Khokhar. A high performance multiple sequence alignment system for pyrosequencing reads from multiple reference genomes. Journal of Parallel and Distributed Computing, 72(1):83–93, January 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001584.

Song:2019:DSI

[SQQL19] Pengfei Song, Lei Qi, Xueming Qian, and Xiaoqiang Lu. Detection of ships in inland river using high-resolution optical satellite imagery based on mixture of deformable part models. Journal of Parallel and Distributed Computing, 132(??):1–7, October 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518307111.

Sudarsan:2016:CPP

[SR16] Rajesh Sudarsan and Calvin J. Ribbens. Combining performance and priority for scheduling resizable parallel applications. Journal of Parallel and Distributed Computing, 87(??):55–66, January 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001744.

Silva:2019:CUA

[SRB^{+}19] Samuel Henrique Silva, Paul Rad, Nicole Beebe, Kim-Kwang Raymond Choo, and Mahesh Umapathy. Cooperative unmanned aerial vehicles with privacy preserving deep vision for real-time object identification and tracking. Journal of Parallel and Distributed Computing, 131(??):147–160, September 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518308839.

Skyrme:2014:SSS

[SRI14] Alexandre Skyrme, Noemi Rodriguez, and Roberto Ierusalimschy. A survey of support for structured communication in concurrency control models. Journal of Parallel and Distributed Computing, 74(4):2266–2285, April 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002323.

Sundfeld:2018:PSD

[SRT^{+}18] Daniel Sundfeld, Caina Razzolini, George Teodoro, Azzedine Boukerche, and Alba Cristina Magalhaes Alves de Melo. PA-Star: a disk-assisted parallel A-Star strategy with localitysensitive hash for multiple sequence alignment. Journal of Parallel and Distributed Computing, 112 (part 2)(??):154–165, February 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517301508.

Shams:2011:OFD

[SS11] Ramtin Shams and Parastoo Sadeghi. On optimization of finite-difference time-domain (FDTD) computation on heterogeneous and GPU clusters. Journal of Parallel and Distributed Computing, 71(4):584–593, April 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sridharan:2017:DDC

[SS17] Aswinkumar Sridharan and André Seznec. Dynamic and discrete cache insertion policies for managing shared last level caches in large multicores. Journal of Parallel and Distributed Computing, 106(??):215–226, August 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300722.

Shukla:2018:MDS

[SS18] Anshu Shukla and Yogesh Simmhan. Model-driven scheduling for distributed stream processing systems. Journal of Parallel and Distributed Computing, 117(??):98–114, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300686.

Stivala:2010:LFP

[SSdlB^{+}10] Alex Stivala, Peter J. Stuckey, Maria Garcia de la Banda, Manuel Hermenegildo, and Anthony Wirth. Lock-free parallel dynamic programming. Journal of Parallel and Distributed Computing, 70(8):839–848, August 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Saez:2011:LWD

[SSFP11] Juan Carlos Saez, Daniel Shelepov, Alexandra Fedorova, and Manuel Prieto. Leveraging workload diversity through OS scheduling to maximize performance on single-ISA heterogeneous multicore systems. Journal of Parallel and Distributed Computing, 71(1):114–131, January 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Selfa:2018:ESM

[SSGG18] Vicent Selfa, Julio Sahuquillo, María E. Gómez, and Crispín Gómez. Efficient selective multicore prefetching under limited memory bandwidth. Journal of Parallel and Distributed Computing, 120(??):32–43, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303174.

Sundriyal:2013:ESS

[SSGZ13] Vaibhav Sundriyal, Masha Sosonkina, Alexander Gaenko, and Zhao Zhang. Energy saving strategies for parallel applications with point-to-point communication phases. Journal of Parallel and Distributed Computing, 73(8):1157–1169, August 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513000506.

Sariyuce:2015:RGC

[SSKÇ15] Ahmet Erdem Sariyüce, Erik Saule, Kamer Kaya, and Ümit V. Çatalyürek. Regularizing graph centrality computations. Journal of Parallel and Distributed Computing, 76(??):106–119, February 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001282.

Sagy:2011:TVA

[SSKS11] Guy Sagy, Izchak Sharfman, Daniel Keren, and Assaf Schuster. Top-k vectorial aggregation queries in a distributed environment. Journal of Parallel and Distributed Computing, 71(2): 302–315, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Salehi:2016:SBR

[SSM^{+}16] Mohsen Amini Salehi, Jay Smith, Anthony A. Maciejewski, Howard Jay Siegel, Edwin K. P. Chong, Jonathan Apodaca, Luis D. Briceño, Timothy Renner, Vladimir Shestak, Joshua Ladd, Andrew Sutton, David Janovy, Sudha Govindasamy, Amin Alqudah, Rinku Dewri, and Puneet Prakash. Stochastic-based robust dynamic resource allocation for independent tasks in a heterogeneous computing system. Journal of Parallel and Distributed Computing, 97(??):96–111, November 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516300739.

Stillwell:2010:RAA

[SSVC10] Mark Stillwell, David Schanzenbach, Frédéric Vivien, and Henri Casanova. Resource allocation algorithms for virtualized service hosting platforms. Journal of Parallel and Distributed Computing, 70(9):962–974, September 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shu:2014:SSS

[SSX14] Jiwu Shu, Zhirong Shen, and Wei Xue. Shield: a stackable secure storage system for file sharing in public storage. Journal of Parallel and Distributed Computing, 74(9):2872–2883, September 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001051.

Si:2010:OCA

[SSZ10] Weisheng Si, Selvadurai Selvakennedy, and Albert Y. Zomaya. An overview of channel assignment methods for multi-radio multi-channel wireless mesh networks. Journal of Parallel and Distributed Computing, 70(5):505–524, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shen:2012:ACD

[ST12] Zhenhui Shen and Srikanta Tirthapura. Approximate covering detection among content-based subscriptions using space filling curves. Journal of Parallel and Distributed Computing, 72(12):1591–1602, December 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002171.

Singh:2014:MPI

[ST14] Sneha Aman Singh and Srikanta Tirthapura. Monitoring persistent items in the union of distributed streams. Journal of Parallel and Distributed Computing, 74(11):3115–3127, November 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514001300.

Selvitopi:2012:RPU

[STA12] R. Oguz Selvitopi, Ata Turk, and Cevdet Aykanat. Replicated partitioning for undirected hypergraphs. Journal of Parallel and Distributed Computing, 72(4):547–563, April 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000056.

Stamoulis:2017:MBW

[Sta17] Georgios Stamoulis. The multi-budgeted and weighted bounded degree metric Steiner network problem. Journal of Parallel and Distributed Computing, 104(??):36–48, June 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301769.

Stewart:2017:SCH

[Ste17] Iain A. Stewart. Sufficient conditions for Hamiltonicity in multiswapped networks. Journal of Parallel and Distributed Computing, 101(??):17–26, March 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516301435.

Sinnen:2011:CAS

[STK11] Oliver Sinnen, Andrea To, and Manpreet Kaur. Contentionaware scheduling with task duplication. Journal of Parallel and Distributed Computing, 71(1):77–86, January 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shi:2012:TSN

[STKW12] Justin Y. Shi, Moussa Taifi, Abdallah Khreishah, and Jie Wu. Tuple switching network-when slower may be better. Journal of Parallel and Distributed Computing, 72(11):1521–1534, November 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512000263.

Si:2018:PGF

[STMZ18] Weisheng Si, Quincy Tse, Guoqiang Mao, and Albert Y. Zomaya. On the performance of greedy forwarding on Yao and Theta graphs. Journal of Parallel and Distributed Computing, 117(??):87–97, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518300832.

Strzodka:2012:DLO

[Str12] Robert Strzodka. Data layout optimization for multivalued containers in OpenCL. Journal of Parallel and Distributed Computing, 72(9):1073–1082, September 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511002115.

Spataro:2019:ESI

[STS19] William Spataro, Giuseppe A. Trunfio, and Georgios Ch. Sirakoulis. Editorial on the special issue on parallel computing in modelling and simulation. Journal of Parallel and Distributed Computing, 134(??):233–235, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519306173.

Sukhoroslov:2018:BWB

[Suk18] Oleg Sukhoroslov. Building web-based services for practical exercises in parallel and distributed computing. Journal of Parallel and Distributed Computing, 118 (part 1)(??):177–188, August 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0743731518301023.

Struharik:2018:SHA

[SV18] R. Struharik and B. Vukobratović. A system for hardware aided decision tree ensemble evolution. Journal of Parallel and Distributed Computing, 112 (part 1)(??):67–83, February 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302691.

Santambrogio:2019:R

[SV19] Marco D. Santambrogio and Ramachandran Vaidyanathan. RAW 2016. Journal of Parallel and Distributed Computing, 133(??):365–366, November 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519305635.

Saiedian:2012:CER

[SW12] Hossein Saiedian and Gabe Wishnie. A complex event routing infrastructure for distributed systems. Journal of Parallel and Distributed Computing, 72(3):450–461, March 2012.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511002218.

Sagonas:2018:CAA

[SW18] Konstantinos Sagonas and Kjell Winblad. A contention adapting approach to concurrent ordered sets. Journal of Parallel and Distributed Computing, 115(??):1–19, May 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303052.

Shoukourian:2017:AEC

[SWHB17] Hayk Shoukourian, Torsten Wilde, Herbert Huber, and Arndt Bode. Analysis of the efficiency characteristics of the first high-temperature direct liquid cooled petascale supercomputer and its cooling infrastructure. Journal of Parallel and Distributed Computing, 107(??):87–100, September 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301272.

Siwik:2019:FGP

[SWLP19] Leszek Siwik, Maciej Woźniak, Marcin Loś, and Maciej Paszyński. Fast and green parallel isogeometric analysis computations for multi-objective optimization of liquid fossil fuel reserve exploitation with minimal groundwater contamination. Journal of Parallel and Distributed Computing, 134(??):89–103, December 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518306890.

Song:2017:ERT

[SWLZ17] Chao Song, Jie Wu, Ming Liu, and Huanyang Zheng. Efficient routing through discretization of overlapped road segments in VANETs. Journal of Parallel and Distributed Computing, 102(??):57–70, April 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516301691.

Song:2017:PPF

[SWW^{+}17] Wei Song, Bing Wang, Qian Wang, Zhiyong Peng, Wenjing Lou, and Yihui Cui. A privacy-preserved full-text retrieval algorithm over encrypted data for cloud storage applications. Journal of Parallel and Distributed Computing, 99(??):14–27, January 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516300533.

Su:2016:CDW

[SZB16] Xing Su, Minjie Zhang, and Quan Bai. Coordination for dynamic weighted task allocation in disaster environments with time, space and communication constraints. Journal of Parallel and Distributed Computing, 97(??):47–56, November 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516300752.

Subrata:2010:CPA

[SZL10] Riky Subrata, Albert Y. Zomaya, and Bjorn Landfeldt. Cooperative power-aware scheduling in grid computing environments. Journal of Parallel and Distributed Computing, 70(2): 84–91, February 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shi:2013:REA

[SZMK13] Qi Shi, Ning Zhang, Madjid Merabti, and Kashif Kifayat. Resource-efficient authentic key establishment in heterogeneous wireless sensor networks. Journal of Parallel and Distributed Computing, 73(2):235–249, February 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002547.

Talbi:2019:UVP

[Tal19] El-Ghazali Talbi. A unified view of parallel multi-objective evolutionary algorithms. Journal of Parallel and Distributed Computing, 133(??):349–358, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151830279X.

Tamada:2018:MEP

[Tam18] Yoshinori Tamada. Memory efficient parallel algorithm for optimal DAG structure search using direct communication. Journal of Parallel and Distributed Computing, 119(??):27–35, September 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518302181.

Tariq:2019:MCD

[TAM^{+}19] Noshina Tariq, Muhammad Asim, Zakaria Maamar, M. Zubair Farooqi, Noura Faci, and Thar Baker. A mobile codedriven trust mechanism for detecting internal attacks in sensor node-powered IoT. Journal of Parallel and Distributed Computing, 134(??):198–206, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151930485X.

Tatrai:2011:PIB

[Tát11] Antal Tátrai. Parallel implementations of Brunotte’s algorithm. Journal of Parallel and Distributed Computing, 71(4): 565–572, April 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Taubenfeld:2016:FS

[Tau16] Gadi Taubenfeld. Fair synchronization. Journal of Parallel and Distributed Computing, 97(??):1–10, November 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300727.

Tang:2017:MSD

[TBG^{+}17] Qi Tang, Twan Basten, Marc Geilen, Sander Stuijk, and Ji-Bo Wei. Mapping of synchronous dataflow graphs on MP-SoCs based on parallelism enhancement. Journal of Parallel and Distributed Computing, 101(??):79–91, March 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151630171X.

Ting:2013:IGB

[TC13] I-Wei Ting and Yeim-Kuan Chang. Improved group-based cooperative caching scheme for mobile ad hoc networks. Journal of Parallel and Distributed Computing, 73(5):595–607, May 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512002948.

Tsai:2012:SFT

[TCHC12] Wen-Chung Tsai, Kuo-Chih Chu, Yu-Hen Hu, and Sao-Jie Chen. A scalable and fault-tolerant network routing scheme for many-core and multi-chip systems. Journal of Parallel and Distributed Computing, 72(11):1433–1441, November 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000275.

Tchernykh:2019:CCQ

[TCMB^{+}19] Andrei Tchernykh, Jorge M. Cortés-Mendoza, Igor Bychkov, Alexander Feoktistov, Loic Didelot, Pascal Bouvry, Gleb Radchenko, and Kirill Borodulin. Configurable cost-quality optimization of cloud-based VoIP. Journal of Parallel and Distributed Computing, 133(??):319–336, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304726.

Tu:2010:SCO

[TCS^{+}10] Shin-Chih Tu, Guey-Yun Chang, Jang-Ping Sheu, Wei Li, and Kun-Ying Hsieh. Scalable continuous object detection and tracking in sensor networks. Journal of Parallel and Distributed Computing, 70(3):212–224, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tuzov:2018:TSF

[TdAR18] Ilya Tuzov, David de Andrés, and Juan-Carlos Ruiz. Tuning synthesis flags to optimize implementation goals: Performance and robustness of the LEON3 processor as a case study. Journal of Parallel and Distributed Computing, 112 (part 1)(??):84–96, February 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302708.

Tang:2013:JSA

[TDBL13] Wei Tang, Narayan Desai, Daniel Buettner, and Zhiling Lan. Job scheduling with adjusted runtime estimates on production supercomputers. Journal of Parallel and Distributed Computing, 73(7):926–938, July 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000312.

Tripathy:2015:DTS

[TDP15] Binodini Tripathy, Smita Dash, and Sasmita Kumari Padhy. Dynamic task scheduling using a directed neural network. Journal of Parallel and Distributed Computing, 75(??):101–106, January 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001907.

Teng:2016:SCA

[Ten16] Yuan-Hsiang Teng. The spanning connectivity of the arrangement graphs. Journal of Parallel and Distributed Computing, 98(??):1–7, December 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516300909.

Terekhov:2016:HSP

[Ter16] Andrew V. Terekhov. A highly scalable parallel algorithm for solving Toeplitz tridiagonal systems of linear equations. Journal of Parallel and Distributed Computing, 87(??):102–108, January 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515001896.

Tarplee:2015:SLP

[TFMS15] Kyle M. Tarplee, Ryan Friese, Anthony A. Maciejewski, and Howard Jay Siegel. Scalable linear programming based resource allocation for makespan minimization in heterogeneous computing systems. Journal of Parallel and Distributed Computing, 84(??):76–86, October 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515001203.

Tumeo:2015:SIA

[TFV^{+}15] Antonino Tumeo, John Feo, Oreste Villa, Simone Secchi, and Timothy G. Mattson. Special issue on Architectures and Algorithms for Irregular Applications (AAIA)

— Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 76(??):1–2, February 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514002263.

Tumeo:2019:SIS

[TFV19] Antonino Tumeo, John Feo, and Oreste Villa. Special issue on: Systems for Learning, Inferencing, and Discovering (SLID). Journal of Parallel and Distributed Computing, 129(??):59– 60, July 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519302503.

Titos-Gil:2016:ASE

[TGPUC16] Rubén Titos-Gil, Oscar Palomar, Osman Unsal, and Adrian Cristal. Architectural support for efficient message passing on shared memory multi-cores. Journal of Parallel and Distributed Computing, 95(??):92–106, September 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516000204.

Talbi:2011:SIJ

[TH11] El-Ghazali Talbi and Geir Hasle. Special issue of the Journal of Parallel and Distributed Computing: Metaheuristics on GPU. Journal of Parallel and Distributed Computing, 71(4):621–622, April 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Talbi:2013:MG

[TH13] El-Ghazali Talbi and Geir Hasle. Metaheuristics on GPUs. Journal of Parallel and Distributed Computing, 73(1):1–3, January 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512002298.

Tran:2019:PPB

[TH19] Hong-Yen Tran and Jiankun Hu. Privacy-preserving big data analytics a comprehensive survey. Journal of Parallel and Distributed Computing, 134(??):207–218, December 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519300589.

Tang:2015:DDL

[THGY15] Jiqiang Tang, Hongyu Huang, Songtao Guo, and Yuanyuan Yang. Dellat: Delivery latency minimization in wireless sensor networks with mobile sink. Journal of Parallel and Distributed Computing, 83(??):133–142, September 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515000957.

Trahan:2010:RPM

[TJCB10] Jerry L. Trahan, Mingxian Jin, Wittaya Chantamas, and Johnnie W. Baker. Relating the power of the Multiple Associative Computing (MASC) model to that of reconfigurable bus-based models. Journal of Parallel and Distributed Computing, 70(5):458–466, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tuna:2017:SIS

[TKG^{+}17] Gurkan Tuna, Dimitrios G. Kogias, V. Cagri Gungor, Cengiz Gezer, Erhan Taskin, and Erman Ayday. A survey on information security threats and solutions for Machine to Machine (M2M) communications. Journal of Parallel and Distributed Computing, 109(??):142–154, November 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301867.

Tinati:2017:TET

[TKKH17] Melika Tinati, Roshanak Karimi, Somayyeh Koohi, and Shaahin Hessabi. Topology exploration of a thermally resilient wavelength-based ONoC. Journal of Parallel and Distributed Computing, 100(??):140–156, February 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300892.

Tian:2019:PPL

[TKR^{+}19] Yuan Tian, Mariya M. Kaleemullah, Mznah A. Rodhaan, Biao Song, Abdullah Al-Dhelaan, and Tinghuai Ma. A privacy preserving location service for cloud-of-things system. Journal of Parallel and Distributed Computing, 123(??):215–222, January 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151830649X.

Tziritas:2013:MRC

[TKX^{+}13] Nikos Tziritas, Samee Ullah Khan, Cheng-Zhong Xu, Thanasis Loukopoulos, and Spyros Lalis. On minimizing the resource consumption of cloud applications using process migrations. Journal of Parallel and Distributed Computing, 73(12):1690–1704, December 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513001585.

Tong:2018:EDM

[TLL^{+}18] Chao Tong, Jun Li, Chao Lang, Fanxin Kong, Jianwei Niu, and Joel J. P. C. Rodrigues. An efficient deep model for day-ahead electricity load forecasting with stacked denoising auto-encoders. Journal of Parallel and Distributed Computing, 117(??):267–273, July 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151730196X.

Tang:2010:LSD

[TLLL10] Xiaoyong Tang, Kenli Li, Guiping Liao, and Renfa Li. List scheduling with duplication for heterogeneous computing systems. Journal of Parallel and Distributed Computing, 70(4): 323–329, April 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tang:2010:RAS

[TLLV10] Xiaoyong Tang, Kenli Li, Renfa Li, and Bharadwaj Veeravalli. Reliability-aware scheduling strategy for heterogeneous distributed computing systems. Journal of Parallel and Distributed Computing, 70(9):941–952, September 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tang:2012:HRD

[TLQS12] Xiaoyong Tang, Kenli Li, Meikang Qiu, and Edwin H.- M. Sha. A hierarchical reliability-driven scheduling algorithm in grid systems. Journal of Parallel and Distributed Computing, 72(4):525–535, April 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511002346.

Tsai:2018:PMD

[TLW18] Chun-Wei Tsai, Shi-Jui Liu, and Yi-Chung Wang. A parallel metaheuristic data clustering framework for cloud. Journal of Parallel and Distributed Computing, 116(??):39–49, ???? 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302964.

Tan:2012:PAA

[TLY12] Wee Lum Tan, Wing Cheong Lau, and OnChing Yue. Performance analysis of an adaptive, energy-efficient MAC protocol for wireless sensor networks. Journal of Parallel and Distributed Computing, 72(4):504–514, April 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000330.

Torkestani:2010:CWA

[TM10] Javad Akbari Torkestani and Mohammad Reza Meybodi. Clustering the wireless ad hoc networks: a distributed learning automata approach. Journal of Parallel and Distributed Computing, 70(4):394–405, April 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Toce:2017:EHL

[TMK^{+}17] Andi Toce, Abbe Mowshowitz, Akira Kawaguchi, Paul Stone, Patrick Dantressangle, and Graham Bent. An efficient hypercube labeling schema for dynamic peer-to-peer networks. Journal of Parallel and Distributed Computing, 102(??):186– 198, April 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516302088.

Tao:2018:ACA

[TODQ18] Ming Tao, Kaoru Ota, Mianxiong Dong, and Zhuzhong Qian. AccessAuth: Capacity-aware security access authentication in federated-IoT-enabled V2G networks. Journal of Parallel and Distributed Computing, 118 (part 1)(??):107–117, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0743731517302587.

Turkyilmaz:2014:RBF

[TOR^{+}14] Ogun Turkyilmaz, Santhosh Onkaraiah, Marina Reyboz, Fabien Clermidy, Hraziia, Costin Anghel, Jean-Michel Portal, and Marc Bocquet. RRAM-based FPGA for “normally off, instantly on” applications. Journal of Parallel and Distributed Computing, 74(6):2441–2451, June 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001391.

Thebault:2018:AMC

[TP18] Loïc Thébault and Eric Petit. Asynchronous and multithreaded communications on irregular applications using vectorized divide and conquer approach. Journal of Parallel and Distributed Computing, 114(??):16–27, April 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303350.

Tu:2019:HPC

[TPJ^{+}19] Wanqing Tu, Florin Pop, Weijia Jia, Jie Wu, and Mauro Iacono. High-performance computing in edge computing networks. Journal of Parallel and Distributed Computing, 123 (??):230, January 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731518308086.

Tong:2018:FCM

[TPLY18] Zhou Tong, Scott Pakin, Michael Lang, and Xin Yuan. Fast classification of MPI applications using Lamport’s logical clocks. Journal of Parallel and Distributed Computing, 120(??):77–88, October 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151830340X.

Tiwary:2018:RTO

[TPS^{+}18] Mayank Tiwary, Deepak Puthal, Kshira Sagar Sahoo, Bibhudatta Sahoo, and Laurence T. Yang. Response time optimization for cloudlets in mobile edge computing. Journal of Parallel and Distributed Computing, 119(??):81–91, September 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518302430.

Tian:2016:LSP

[TR16] Yun Tian and Philip J. Rhodes. A location service for partial spatial replicas implementing an R-tree in a relational database. Journal of Parallel and Distributed Computing, 90–91(??):9–21, April 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731516000046.

Toharia:2012:SBD

[TRS^{+}12] Pablo Toharia, Oscar D. Robles, Ricardo Suárez, Jose Luis Bosque, and Luis Pastor. Shot boundary detection using Zernike moments in multi-GPU multi-CPU architectures. Journal of Parallel and Distributed Computing, 72(9):1127–1133, September 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511002103.

Trejo-Sanchez:2014:DAM

[TSFZ14] Joel Antonio Trejo-Sánchez and José Alberto Fernández-Zepeda. Distributed algorithm for the maximal 2-packing in geometric outerplanar graphs. Journal of Parallel and Distributed Computing, 74(3):2193–2202, March 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513002360.

Talia:2010:EDQ

[TT10] Domenico Talia and Paolo Trunfio. Enabling dynamic querying over distributed hash tables. Journal of Parallel and Distributed Computing, 70(12):1254–1265, December 2010. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Thomasian:2012:HRD

[TTH12] Alexander Thomasian, Yujie Tang, and Yang Hu. Hierarchical RAID: Design, performance, reliability, and recovery. Journal of Parallel and Distributed Computing, 72(12):1753–1769, December 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001633.

Turau:2012:ETD

[Tur12] Volker Turau. Efficient transformation of distance-2 selfstabilizing algorithms. Journal of Parallel and Distributed Computing, 72(4):603–612, April 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511002486.

Tripathi:2017:NCP

[TVT^{+}17] Rakesh Tripathi, S. Vignesh, Venkatesh Tamarapalli, Anthony T. Chronopoulos, and Hajar Siar. Non-cooperative power and latency aware load balancing in distributed data centers. Journal of Parallel and Distributed Computing, 107(??):76–86, September 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301284.

Thanakulwarapas:2015:OBS

[TW15] Tipraporn Thanakulwarapas and Jeeraporn Werapun. An optimized bitonic sorting strategy with midpoint-based dynamic communication. Journal of Parallel and Distributed Computing, 84(??):37–50, October 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731515000982.

Tu:2012:FBD

[TWQS12] Zhiliang Tu, Qiang Wang, Hairong Qi, and Yi Shen. Flocking based distributed self-deployment algorithms in mobile sensor networks. Journal of Parallel and Distributed Computing, 72(3):437–449, March 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511002322.

Tong:2014:PSD

[TXLL14] Zhao Tong, Zheng Xiao, Kenli Li, and Keqin Li. Proactive scheduling in distributed computing — a reinforcement learning approach. Journal of Parallel and Distributed Computing, 74(7):2662–2672, July 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151400063X.

Tong:2017:HAM

[TY17] Fei Tong and Zheng Yan. A hybrid approach of mobile malware detection in Android. Journal of Parallel and Distributed Computing, 103(??):22–31, May 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151630140X.

Torun:2016:FGC

[TYA16] Mustafa U. Torun, Onur Yilmaz, and Ali N. Akansu. FPGA, GPU, and CPU implementations of Jacobi algorithm for eigenanalysis. Journal of Parallel and Distributed Computing, 96(??):172–180, October 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300508.

Tu:2019:TRT

[TYD^{+}19] Weiping Tu, Yuhong Yang, Bo Du, Jiaxi Zheng, and Shuangxing Zhai. Towards a real-time production of immersive spatial audio of high individuality with an RBF neural network. Journal of Parallel and Distributed Computing, 131(??):120–129, September 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519300620.

Taheri:2011:FOL

[TZI11] Javid Taheri, Albert Y. Zomaya, and Mohsin Iftikhar. Fuzzy online location management in mobile computing environments. Journal of Parallel and Distributed Computing, 71(8):1142–1153, August 2011. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511000359.

Usui:2010:ALC

[UBES10] Takayuki Usui, Reimer Behrends, Jacob Evans, and Yannis Smaragdakis. Adaptive locks: Combining transactions and locks for efficient concurrency. Journal of Parallel and Distributed Computing, 70(10):1009–1023, October 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Uddin:2019:WSB

[Udd19] Md. Zia Uddin. A wearable sensor-based activity prediction system to facilitate edge computing in smart healthcare system. Journal of Parallel and Distributed Computing, 123(??):46–53, January 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518306270.

Utrera:2019:TPE

[UFF19] Gladys Utrera, Montse Farreras, and Jordi Fornes. Task packing: Efficient task scheduling in unbalanced parallel programs to maximize CPU utilization. Journal of Parallel and Distributed Computing, 134(??):37–49, December 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519305623.

Ulmer:2011:MPA

[UGG^{+}11] Craig Ulmer, Maya Gokhale, Brian Gallagher, Philip Top, and Tina Eliassi-Rad. Massively parallel acceleration of a document-similarity classifier to detect Web attacks. Journal of Parallel and Distributed Computing, 71(2):225–235, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Unlu:2017:BPA

[UM17] Eren Unlu and Christophe Moy. Bimodal packet aware scheduling for an OFDMA based on-chip RF interconnect. Journal of Parallel and Distributed Computing, 109(??):15–28, November 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517301533.

Umar:2018:ABP

[UMM^{+}18] Mariam Umar, Shirley V. Moore, Jeremy S. Meredith, Jeffrey S. Vetter, and Kirk W. Cameron. Aspen-based performance and energy modeling frameworks. Journal of Parallel and Distributed Computing, 120(??):222–236, October 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517303039.

Upadhyaya:2013:PAM

[Upa13] Sujatha R. Upadhyaya. Parallel approaches to machine learning — a comprehensive survey. Journal of Parallel and Distributed Computing, 73(3):284–292, March 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002705.

Wasi-ur-Rahman:2018:MAC

[uRIL^{+}18] Md. Wasi ur Rahman, Nusrat Sharmin Islam, Xiaoyi Lu, Dipti Shankar, and Dhabaleswar K. (DK) Panda. MR-Advisor: a comprehensive tuning, profiling, and prediction tool for MapReduce execution frameworks on HPC clusters. Journal of Parallel and Distributed Computing, 120(??):237–250, October 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517303027.

UlHassan:2019:DPR

[URK^{+}19] Muneeb Ul Hassan, Mubashir Husain Rehmani, Ramamohanarao Kotagiri, Jiekui Zhang, and Jinjun Chen. Differential privacy for renewable energy resources based smart metering. Journal of Parallel and Distributed Computing, 131(??):69–80, September 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518309201.

Vazquez:2019:PEF

[VAF19] Sergio Vázquez, Margarita Amor, and Basilio B. Fraguela. Portable and efficient FFT and DCT algorithms with the Heterogeneous Butterfly Processing Library. Journal of Parallel and Distributed Computing, 125(??):135–146, March 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151830858X.

Villar:2013:OOC

[VAS^{+}13] Juan A. Villar, Francisco J. Andújar, José L. Sánchez, Francisco J. Alfaro, José A. Gámez, and José Duato. Obtaining the optimal configuration of high-radix combined switches. Journal of Parallel and Distributed Computing, 73(9):1239–1250, September 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513000701.

Valdez-Balderas:2013:TAS

[VBDRC13] Daniel Valdez-Balderas, José M. Domínguez, Benedict D. Rogers, and Alejandro J. C. Crespo. Towards accelerating smoothed particle hydrodynamics simulations for free-surface flows on multi-GPU clusters. Journal of Parallel and Distributed Computing, 73(11):1483–1493, November 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001712.

Vinas:2013:EHP

[VBF13] Moisés Viñas, Zeki Bozkus, and Basilio B. Fraguela. Exploiting heterogeneous parallelism with the Heterogeneous Programming Library. Journal of Parallel and Distributed Computing, 73(12):1627–1638, December 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513001512.

V:2018:HAS

[VD18] Bibal Benifa J. V. and Dejey Dharma. HAS: Hybrid autoscaler for resource scaling in cloud environment. Journal of Parallel and Distributed Computing, 120(??):1–15, October 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518303022.

Veiga:2018:EME

[VETT18] Jorge Veiga, Roberto R. Expósito, Guillermo L. Taboada, and Juan Touriño. Enhancing in-memory efficiency for MapReduce-based data processing. Journal of Parallel and Distributed Computing, 120(??):323–338, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518302284.

Vinas:2017:HPM

[VFAD17] Moisés Viñas, Basilio B. Fraguela, Diego Andrade, and Ramón Doallo. High productivity multi-device exploitation with the Heterogeneous Programming Library. Journal of Parallel and Distributed Computing, 101(??):51–68, March 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301447.

Veloso:2018:SDA

[VLGV^{+}18] Bruno Veloso, Fátima Leal, Horacio González-Vélez, Benedita Malheiro, and Juan Carlos Burguillo. Scalable data analytics using crowdsourced repositories and streams. Journal of Parallel and Distributed Computing, 122(??):1–10, December 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518304544.

Vo:2014:MPU

[VLL^{+}14] Phuong Luu Vo, Tuan Anh Le, Sungwon Lee, Choong Seon Hong, Byeongsik Kim, and Hoyoung Song. Multi-path utility maximization and multi-path TCP design. Journal of Parallel and Distributed Computing, 74(1):1848–1857, January 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513001317.

Valery:2018:CCG

[VLW18] Olivier Valery, Pangfeng Liu, and Jan-Jan Wu. A collaborative CPU–GPU approach for principal component analysis on mobile heterogeneous platforms. Journal of Parallel and Distributed Computing, 120(??):44–61, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303411.

Vengerov:2010:ADA

[VMMB10] David Vengerov, Lykomidis Mastroleon, Declan Murphy, and Nick Bambos. Adaptive data-aware utility-based scheduling in resource-constrained systems. Journal of Parallel and Distributed Computing, 70(9):871–879, September 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Valls:2017:TFA

[VRGS17] Joan J. Valls, Alberto Ros, María E. Gómez, and Julio Sahuquillo. The Tag Filter Architecture: an energy-efficient cache and directory design. Journal of Parallel and Distributed Computing, 100(??):193–202, February 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300302.

Vupputuri:2010:UMD

[VRM10] Saamaja Vupputuri, Kiran K. Rachuri, and C. Siva Ram Murthy. Using mobile data collectors to improve network lifetime of wireless sensor networks with reliability constraints. Journal of Parallel and Distributed Computing, 70(7):767–778, July 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Venugopalan:2016:MLA

[VS16] Sarad Venugopalan and Oliver Sinnen. Memory limited algorithms for optimal task scheduling on parallel systems. Journal of Parallel and Distributed Computing, 92(??):35–49, May 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516000241.

Verma:2018:CCI

[VS18] Prabal Verma and Sandeep K. Sood. Cloud-centric IoT based disease diagnosis healthcare framework. Journal of Parallel and Distributed Computing, 116(??):27–38, ???? 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303301.

Wu:2013:LSA

[WBRT13] Di Wu, Lichun Bao, Amelia C. Regan, and Carolyn L. Talcott. Large-scale access scheduling in wireless mesh networks using social centrality. Journal of Parallel and Distributed Computing, 73(8):1049–1065, August 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000555.

Wang:2019:PSL

[WBS19] Zhu Wang, Jalil Boukhobza, and Zili Shao. Preserving SSD lifetime in deep learning applications with delta snapshots. Journal of Parallel and Distributed Computing, 133(??):63–76, November 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151830902X.

Wu:2018:ESS

[WCCH18] Libing Wu, Biwen Chen, Kim-Kwang Raymond Choo, and Debiao He. Efficient and secure searchable encryption protocol for cloud-based Internet of Things. Journal of Parallel and Distributed Computing, 111(??):152–161, January 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151730237X.

Wehe:2010:SPG

[WCEA10] André Wehe, Wen-Chieh Chang, Oliver Eulenstein, and Srinivas Aluru. A scalable parallelization of the gene duplication problem. Journal of Parallel and Distributed Computing, 70 (3):237–244, March 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2014:AIS

[WCF14] Y. Wang, S. D. Cotofana, and L. Fang. Analysis of the impact of spatial and temporal variations on the stability of SRAM arrays and the mitigation technique using independent-gate devices. Journal of Parallel and Distributed Computing, 74(6):2521–2529, June 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001354.

Wang:2017:PMC

[WCH^{+}17] Zhuo Wang, Qun Chen, Boyi Hou, Bo Suo, Zhanhuai Li, Wei Pan, and Zachary G. Ives. Parallelizing maximal clique and k-plex enumeration over graph data. Journal of Parallel and Distributed Computing, 106(??):79–91, August 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300837.

Wang:2013:CLO

[WCL^{+}13] Xingwei Wang, Hui Cheng, Keqin Li, Jie Li, and Jiajia Sun. A cross-layer optimization based integrated routing and grooming algorithm for green multi-granularity transport networks. Journal of Parallel and Distributed Computing, 73(6):807–822, June 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151300035X.

Wan:2017:OCD

[WCWO17] Lipeng Wan, Qing Cao, Feiyi Wang, and Sarp Oral. Optimizing checkpoint data placement with guaranteed burst buffer endurance in large-scale hierarchical storage systems. Journal of Parallel and Distributed Computing, 100(??):16–29, February 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301198.

Wen:2011:UDS

[WCXL11] Jigang Wen, Jiannong Cao, Kun Xie, and Renfa Li. User density sensitive P2P streaming in wireless mesh networks. Journal of Parallel and Distributed Computing, 71(4):573–583, April 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wittek:2013:ATM

[WD13] Peter Wittek and Sándor Darányi. Accelerating text mining workloads in a MapReduce-based distributed GPU environment. Journal of Parallel and Distributed Computing, 73(2):198–206, February 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731512002353.

Wisniewski:2018:BSB

[WD18] Mariusz Wiśniewski and Stanisław Deniziak. BMB synthesis of binary functions using symbolic functional decomposition for LUT-based FPGAs. Journal of Parallel and Distributed Computing, 120(??):16–22, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303149.

Wei:2018:PHP

[WDS^{+}18] Leyi Wei, Yijie Ding, Ran Su, Jijun Tang, and Quan Zou. Prediction of human protein subcellular localization using deep learning. Journal of Parallel and Distributed Computing, 117(??):212–217, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302393.

Wilkin:2013:MPA

[WE13] Gregory Aaron Wilkin and Patrick Eugster. Multicasting in the presence of aggregated deliveries. Journal of Parallel and Distributed Computing, 73(4):544–556, April 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002857.

Wang:2016:VDP

[WFLJ16] Xi Wang, Jianxi Fan, Cheng-Kuan Lin, and Xiaohua Jia. Vertex-disjoint paths in DCell networks. Journal of Parallel and Distributed Computing, 96(??):38–44, October 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300314.

Wang:2012:IST

[WFZJ12] Yan Wang, Jianxi Fan, Guodong Zhou, and Xiaohua Jia. Independent spanning trees on twisted cubes. Journal of Parallel and Distributed Computing, 72(1):58–69, January 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511001742.

Wu:2011:OEE

[WG11] Qishi Wu and Yi Gu. Optimizing end-to-end performance of data-intensive computing pipelines in heterogeneous network environments. Journal of Parallel and Distributed Computing, 71(2):254–265, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wlotzka:2017:EEM

[WH17] Martin Wlotzka and Vincent Heuveline. Energy-efficient multigrid smoothers and grid transfer operators on multicore and GPU clusters. Journal of Parallel and Distributed Computing, 100(??):181–192, February 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300363.

Wen:2018:EED

[WHC^{+}18] Shaojie Wen, Chuanhe Huang, Xi Chen, Jianhua Ma, Naixue Xiong, and Zongpeng Li. Energy-efficient and delay-aware distributed routing with cooperative transmission for Internet of Things. Journal of Parallel and Distributed Computing, 118 (part 1)(??):46–56, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0743731517302320.

Wu:2018:TCM

[WHS^{+}18] Yinfeng Wu, Yachao Hu, Yiwen Su, Ning Yu, and Renjian Feng. Topology control for minimizing interference with delay constraints in an ad hoc network. Journal of Parallel and Distributed Computing, 113(??):63–76, March 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302812.

Wang:2017:SSS

[WHW^{+}17] Jun Wang, Dan Huang, Huafeng Wu, Jiangling Yin, Xuhong Zhang, Xunchao Chen, and Ruijun Wang. SideIO: a Side I/O system framework for hybrid scientific workflow. Journal of Parallel and Distributed Computing, 108(??):45–58, October 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516300867.

Wu:2019:RDM

[WHW^{+}19] YouKe Wu, Haiyang Huang, Qun Wu, Anfeng Liu, and Tian Wang. A risk defense method based on microscopic state prediction with partial information observations in social networks. Journal of Parallel and Distributed Computing, 131(??):189–199, September 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305896.

Werapun:2012:EPC

[WIB12] Jeeraporn Werapun, Sarun Intakosum, and Veera Boonjing. An efficient parallel construction of optimal independent spanning trees on hypercubes. Journal of Parallel and Distributed Computing, 72(12):1713–1724, December 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001645.

Wassi:2018:FBS

[WIR^{+}18] G. Wassi, S. Iloga, O. Romain, B. Granado, and M. Tchuenté. FPGA-based simultaneous multichannel audio processor for musical genre indexing applications in broadcast band. Journal of Parallel and Distributed Computing, 119(??):146–161, September 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518300881.

Wei:2012:FAC

[WJ12] Zheng Wei and Joseph JaJa. A fast algorithm for constructing inverted files on heterogeneous platforms. Journal of Parallel and Distributed Computing, 72(5):728–738, May 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512000391.

Wu:2014:OFC

[WJ14] Jing Wu and Joseph JaJa. Optimized FFT computations on heterogeneous platforms with application to the Poisson equation. Journal of Parallel and Distributed Computing, 74(8):2745–2756, August 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514000653.

Wu:2010:AHS

[WL10] Hejun Wu and Qiong Luo. Adaptive holistic scheduling for query processing in sensor networks. Journal of Parallel and Distributed Computing, 70(6):657–670, June 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2011:IDG

[WL11] Yun Wang and Zhengdong Lun. Intrusion detection in a K-Gaussian distributed wireless sensor network. Journal of Parallel and Distributed Computing, 71(12):1598–1607, December 2011. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511001456.

Wang:2015:CSP

[WLCZ15] Chao Wang, Xi Li, Peng Chen, and Xuehai Zhou. A case study of parallel JPEG encoding on an FPGA. Journal of Parallel and Distributed Computing, 78(??):1–5, April 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001737.

Wang:2019:SML

[WLK^{+}19] Xianmin Wang, Jing Li, Xiaohui Kuang, Yu an Tan, and Jin Li. The security of machine learning in an adversarial setting: a survey. Journal of Parallel and Distributed Computing, 130(??):12–23, August 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518309183.

Wan:2016:NCA

[WLL16] Lanjun Wan, Kenli Li, and Keqin Li. A novel cooperative accelerated parallel two-list algorithm for solving the subsetsum problem on a hybrid CPU–GPU cluster. Journal of Parallel and Distributed Computing, 97(??):112–123, November 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516300880.

Wang:2016:ACS

[WLST16] Shinan Wang, Bing Luo, Weisong Shi, and Devesh Tiwari. Application configuration selection for energy-efficient execution on multicore systems. Journal of Parallel and Distributed Computing, 87(??):43–54, January 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001689.

Wei:2019:DPB

[WLYS19] Jianhao Wei, Yaping Lin, Xin Yao, and Voundi Koe Arthur Sandor. Differential privacy-based trajectory community recommendation in social network. Journal of Parallel and Distributed Computing, 133(??):136–148, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518309572.

Wu:2018:DDA

[WLZ^{+}18] Ziyan Wu, Zhihui Lu, Wei Zhang, Jie Wu, Shalin Huang, and Patrick C. K. Hung. A data-driven approach of performance evaluation for cache server groups in content delivery network. Journal of Parallel and Distributed Computing, 119(??):162–171, September 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518302697.

Wang:2018:HDA

[WMC^{+}18] Xuesong Wang, Yuting Ma, Yuhu Cheng, Liang Zou, and Joel J. P. C. Rodrigues. Heterogeneous domain adaptation network based on autoencoder. Journal of Parallel and Distributed Computing, 117(??):281–291, July 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517301922.

Wang:2012:PPL

[WMES12] Chao Wang, Frank Mueller, Christian Engelmann, and Stephen L. Scott. Proactive process-level live migration and back migration in HPC environments. Journal of Parallel and Distributed Computing, 72(2):254–267, February 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511002085.

Weyland:2013:MFS

[WMG13] Dennis Weyland, Roberto Montemanni, and Luca Maria Gambardella. A metaheuristic framework for stochastic combinatorial optimization problems based on GPGPU with a case study on the probabilistic traveling salesman problem with deadlines. Journal of Parallel and Distributed Computing, 73(1):74–85, January 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512001220.

Wang:2018:FBP

[WML^{+}18] Yu Wang, Weizhi Meng, Wenjuan Li, Jin Li, Wai-Xi Liu, and Yang Xiang. A fog-based privacy-preserving approach for distributed signature-based intrusion detection. Journal of Parallel and Distributed Computing, 122(??):26–35, December 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518305057.

Wang:2017:CQC

[WMY^{+}17] Hongbing Wang, Peisheng Ma, Qi Yu, Danrong Yang, Jiajie Li, and Huanhuan Fei. Combining quantitative constraints with qualitative preferences for effective nonfunctional properties-aware service composition. Journal of Parallel and Distributed Computing, 100(??):71–84, February 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731516301411.

Wolfson-Pou:2019:MAJ

[WPC19] Jordi Wolfson-Pou and Edmond Chow. Modeling the asynchronous Jacobi method without communication delays. Journal of Parallel and Distributed Computing, 128(??):84–98, June 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518304751.

Wang:2014:PAO

[WQL14] Peijian Wang, Yong Qi, and Xue Liu. Power-aware optimization for heterogeneous multi-tier clusters. Journal of Parallel and Distributed Computing, 74(1):2005–2015, January 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513001998.

Wei:2013:SSC

[WQZ^{+}13] Haitao Wei, Mingkang Qin, Weiwei Zhang, Junqing Yu, Dongrui Fan, and Guang R. Gao. StreamTMC: Stream compilation for tiled multi-core architectures. Journal of Parallel and Distributed Computing, 73(4):484–494, April 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002821.

Wang:2013:PDE

[WRW13] Hui Wang, Shahryar Rahnamayan, and Zhijian Wu. Parallel differential evolution with self-adapting control parameters and generalized opposition-based learning for solving high-dimensional optimization problems. Journal of Parallel and Distributed Computing, 73(1):62–73, January 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000639.

Wu:2011:OSP

[WSLC11] Jan-Jan Wu, Shu-Fan Shih, Pangfeng Liu, and Yi-Min Chung. Optimizing server placement in distributed systems in the presence of competition. Journal of Parallel and Distributed Computing, 71(1):62–76, January 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2019:ATA

[WSX^{+}19] Zhong Wang, Daniel Sun, Guangtao Xue, Shiyou Qian, Guoqiang Li, and Minglu Li. Ada-Things: an adaptive virtual machine monitoring and migration strategy for Internet of Things applications. Journal of Parallel and Distributed Computing, 132(??):164–176, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304404.

Wang:2016:MEM

[WTWZ16] Wenzhu Wang, Yusong Tan, Qingbo Wu, and Yaoxue Zhang. micMR: an efficient MapReduce framework for CPU–MIC heterogeneous architecture. Journal of Parallel and Distributed Computing, 93–94(??):120–131, July 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300132.

Wang:2018:IBQ

[WTY^{+}18] Hongbing Wang, Yong Tao, Qi Yu, Xin Lin, and Tianjing Hong. Incorporating both qualitative and quantitative preferences for service recommendation. Journal of Parallel and Distributed Computing, 114(??):46–69, April 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517303362.

Wu:2011:MAS

[Wu11] Zheng Da Wu. Modelling and analysis of strategies in the design of WSAN coordination systems. Journal of Parallel and Distributed Computing, 71(7):1055–1064, July 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000761.

Wang:2012:DMT

[WW12] Yunsheng Wang and Jie Wu. A dynamic multicast tree based routing scheme without replication in delay tolerant networks. Journal of Parallel and Distributed Computing, 72(3):424–436, March 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511002231.

Wang:2018:RDD

[WW18a] Ning Wang and Jie Wu. Rethink data dissemination in opportunistic mobile networks with mutually exclusive requirement. Journal of Parallel and Distributed Computing, 119(??):50–63, September 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518302193.

Wang:2018:DSW

[WW18b] Zhi Wang and Xizhao Wang. A deep stochastic weight assignment network and its application to chess playing. Journal of Parallel and Distributed Computing, 117(??):205–211, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302435.

Wu:2018:PBO

[WWA^{+}18] Huafeng Wu, Jun Wang, Raghavendra Rao Ananta, Vamsee Reddy Kommareddy, Rui Wang, and Prasant Mohapatra. Prediction based opportunistic routing for maritime search and rescue wireless sensor network. Journal of Parallel and Distributed Computing, 111(??):56–64, January 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302101.

Wang:2017:LAB

[WWW17a] Gang Wang, Zhiyue Wang, and Jie Wu. A local average broadcast gossip algorithm for fast global consensus over graphs. Journal of Parallel and Distributed Computing, 109(??):301–309, November 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517301594.

Wang:2017:NRM

[WWW17b] Jun Wang, Huafeng Wu, and Ruijun Wang. A new reliability model in replication-based big data storage systems. Journal of Parallel and Distributed Computing, 108(??):14–27, October 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517300497.

Wang:2018:PAB

[WWY^{+}18] Hongbing Wang, Lei Wang, Qi Yu, Zibin Zheng, and Zhengping Yang. A proactive approach based on online reliability prediction for adaptation of service-oriented systems. Journal of Parallel and Distributed Computing, 114(??):70–84, April 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517303374.

Wang:2019:CSS

[WXMZ19] Xu An Wang, Fatos Xhafa, Jianfeng Ma, and Zhiheng Zheng. Controlled secure social cloud data sharing based on a novel identity based proxy re-encryption plus scheme. Journal of Parallel and Distributed Computing, 130(??):153–165, August 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731519302345.

Wang:2018:TXN

[WXZ^{+}18] Ran Wang, Guangquan Xu, Xianjiao Zeng, Xiaohong Li, and Zhiyong Feng. TT-XSS: A novel taint tracking based dynamic detection framework for DOM cross-site scripting. Journal of Parallel and Distributed Computing, 118 (part 1)(??):100–106, August 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0743731517302186.

Wu:2013:EHP

[WYTX13] Qiang Wu, Canqun Yang, Tao Tang, and Liquan Xiao. Exploiting hierarchy parallelism for molecular dynamics on a petascale heterogeneous system. Journal of Parallel and Distributed Computing, 73(12):1592–1604, December 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001536.

Wang:2015:KBB

[WYW15] En Wang, Yongjian Yang, and Jie Wu. A knapsack-based buffer management strategy for delay-tolerant networks. Journal of Parallel and Distributed Computing, 86(??):1–15, December 2015. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515001264.

Wang:2013:RIA

[WZ13] Xiaonan Wang and Shan Zhong. Research on IPv6 address configuration for a VANET. Journal of Parallel and Distributed Computing, 73(6):757–766, June 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513000270.

Wang:2019:QPS

[WZH^{+}19] Shangguang Wang, Yali Zhao, Lin Huang, Jinliang Xu, and Ching-Hsien Hsu. QoS prediction for service recommendations in mobile edge computing. Journal of Parallel and Distributed Computing, 127(??):134–144, May 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151730268X.

Wu:2013:DAM

[WZQ^{+}13] Gang Wu, Huxing Zhang, Meikang Qiu, Zhong Ming, Jiayin Li, and Xiao Qin. A decentralized approach for mining event correlations in distributed system monitoring. Journal of Parallel and Distributed Computing, 73(3):330–340, March 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512002225.

Wang:2019:ESP

[WZX^{+}19] Shangguang Wang, Yali Zhao, Jinlinag Xu, Jie Yuan, and Ching-Hsien Hsu. Edge server placement in mobile edge computing. Journal of Parallel and Distributed Computing, 127(??):160–168, May 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518304398.

Wang:2019:FPL

[WZY^{+}19] Zhiqiang Wang, Ye Zhao, Huidan (Whitney) Yu, Chen Lin, and Alan P. Sawchuck. Fully parallelized Lattice Boltzmann scheme for fast extraction of biomedical geometry. Journal of Parallel and Distributed Computing, 128(??):126–136, June 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151930156X.

Wang:2017:DLW

[WZZ^{+}17] Jun Wang, Xuhong Zhang, Junyao Zhang, Jiangling Yin, Dezhi Han, Ruijun Wang, and Dan Huang. Deister: a lightweight autonomous block management in data-intensive file systems using deterministic declustering distribution. Journal of Parallel and Distributed Computing, 108(??):3–13, October 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516000265.

Xiong:2019:PRI

[XCC^{+}19] Mingfu Xiong, Dan Chen, Jun Chen, Jingying Chen, Benyun Shi, Chao Liang, and Ruimin Hu. Person re-identification with multiple similarity probabilities using deep metric learning for efficient smart security applications. Journal of Parallel and Distributed Computing, 132(??):230–241, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517303076.

Xu:2010:EEC

[XHZ^{+}10] Hongli Xu, Liusheng Huang, Yindong Zhang, He Huang, Shenglong Jiang, and Gang Liu. Energy-efficient cooperative data aggregation for wireless sensor networks. Journal of Parallel and Distributed Computing, 70(9):953–961, September 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:2016:FTV

[XHZZ16] Xirong Xu, Yazhen Huang, Peng Zhang, and Sijia Zhang. Fault-tolerant vertex-pancyclicity of locally twisted cubes LT Q_{n}. Journal of Parallel and Distributed Computing, 88 (??):57–62, February 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373151500194X.

Xiao:2011:PMU

[XL11] Xiang Xiao and Jaehwan John Lee. A parallel multi-unit resource deadlock detection algorithm with O(log_{2}(min(m, n))) overall run-time complexity. Journal of Parallel and Distributed Computing, 71(7):938–954, July 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151100013X.

Xiang:2018:AVD

[XLC^{+}18] Tao Xiang, Xiaoguo Li, Fei Chen, Yuanyuan Yang, and Shengyu Zhang. Achieving verifiable, dynamic and efficient auditing for outsourced database in cloud. Journal of Parallel and Distributed Computing, 112 (part 1)(??):97–107, February 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302800.

Xuan:2018:EOA

[XLH18] Yubo Xuan, Dayu Li, and Wei Han. Efficient optimization approach for fast GPU computation of Zernike moments. Journal of Parallel and Distributed Computing, 111(??):104–114, January 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517302204.

Xu:2013:DSS

[XLHT13] Yuming Xu, Kenli Li, Ligang He, and Tung Khac Truong. A DAG scheduling scheme on heterogeneous computing systems using double molecular structure-based chemical reaction optimization. Journal of Parallel and Distributed Computing, 73(9):1306–1322, September 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151300110X.

Xie:2015:HDE

[XLL15] Guoqi Xie, Renfa Li, and Keqin Li. Heterogeneity-driven end-to-end synchronized scheduling for precedence constrained tasks and messages on networked embedded systems. Journal of Parallel and Distributed Computing, 83(??):1–12, September 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515000684.

Xu:2019:MEC

[XLPL19] Hongzhi Xu, Renfa Li, Chen Pan, and Keqin Li. Minimizing energy consumption with reliability goal on heterogeneous embedded systems. Journal of Parallel and Distributed Computing, 127(??):44–57, May 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300243.

Xiao:2018:UCC

[XLW^{+}18] Zheng Xiao, Xiong Li, Le Wang, Qiuwei Yang, Jiayi Du, and Arun Kumar Sangaiah. Using convolution control block for Chinese sentiment analysis. Journal of Parallel and Distributed Computing, 116(??):18–26, ???? 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302940.

Xhafa:2017:SIH

[XMMD17] Fatos Xhafa, Constandinos X. Mavromoustakis, George Mastorakis, and Ciprian Dobre. Special issue on “High Performance and Parallelism for Large Data Sets”. Journal of Parallel and Distributed Computing, 110(??):1, December 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302538.

Xia:2010:PEI

[XP10] Yinglong Xia and Viktor K. Prasanna. Parallel exact inference on the Cell Broadband Engine processor. Journal of Parallel and Distributed Computing, 70(5):558–572, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:2012:UUR

[XRB12] Cheng-Zhong Xu, Jia Rao, and Xiangping Bu. URL: a unified reinforcement learning approach for autonomic cloud management. Journal of Parallel and Distributed Computing, 72(2):95–105, February 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511001924.

Xie:2011:URB

[XS11] Tao Xie and Yao Sun. Understanding the relationship between energy conservation and reliability in parallel disk arrays. Journal of Parallel and Distributed Computing, 71(2): 198–210, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:2018:RDA

[XSYG18] Guangwei Xu, Zhifeng Sun, Cairong Yan, and Yanglan Gan. A rapid detection algorithm of corrupted data in cloud storage. Journal of Parallel and Distributed Computing, 111(??):115–125, January 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302344.

Xu:2012:MTB

[XTN12] Meilian Xu, Parimala Thulasiraman, and Sima Noghanian. Microwave tomography for breast cancer detection on Cell broadband engine processors. Journal of Parallel and Distributed Computing, 72(9):1106–1116, September 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511002127.

Xie:2014:VCP

[XYZW14] Miao Xie, Qiusong Yang, Jian Zhai, and Qing Wang. A vertex centric parallel algorithm for linear temporal logic model checking in Pregel. Journal of Parallel and Distributed Computing, 74(11):3161–3174, November 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001312.

Yoo:2011:OTP

[YA11] Younghwan Yoo and Dharma P. Agrawal. Optimal transmission power with delay constraints in 2D and 3D MANETs. Journal of Parallel and Distributed Computing, 71(11):1484– 1496, November 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731511001006.

Yoo:2010:ISL

[YAA10] Younghwan Yoo, Sanghyun Ahn, and Dharma P. Agrawal. Impact of a simple load balancing approach and an incentivebased scheme on MANET performance. Journal of Parallel and Distributed Computing, 70(2):71–83, February 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yuan:2015:PCE

[YAK15] Hang Yuan, Ishfaq Ahmad, and C.-C. Jay Kuo. Performanceconstrained energy reduction in data centers for videosharing services. Journal of Parallel and Distributed Computing, 75(??):29–39, January 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514002019.

Yuan:2013:CAT

[YBM13] Man (Mike) Yuan, Johnnie W. Baker, and Will C. Meilander. Comparisons of air traffic control implementations on an associative processor with a MIMD and consequences for parallel computing. Journal of Parallel and Distributed Computing, 73(2):256–272, February 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151200127X.

Yan:2013:PPP

[YBX^{+}13] Weizhong Yan, Umang Brahmakshatriya, Ya Xue, Mark Gilder, and Bowden Wise. p-PIC: Parallel power iteration clustering for big data. Journal of Parallel and Distributed Computing, 73(3):352–359, March 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001487.

Yoon:2012:IDP

[YC12] MyungKeun Yoon and Shigang Chen. An incrementally deployable path address scheme. Journal of Parallel and Distributed Computing, 72(10):1215–1225, October 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512001098.

Yang:2010:DEP

[YCH^{+}10] X. Y. Yang, F. Cores, P. Hernández, A. Ripoll, and E. Luque. Designing an effective P2P system for a VoD system to exploit the multicast communication. Journal of Parallel and Distributed Computing, 70(12):1175–1192, December 2010. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yu:2018:VTP

[YDTZ18] Bin Yu, Zhenhua Duan, Cong Tian, and Nan Zhang. Verifying temporal properties of programs: A parallel approach. Journal of Parallel and Distributed Computing, 118 (part 1)(??):89– 99, August 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0743731517302575.

Yi:2018:CIC

[YDZ^{+}18] Lingzhi Yi, Xianjun Deng, Zenghui Zou, Dexin Ding, and Laurence T. Yang. Confident information coverage hole detection in sensor networks for uranium tailing monitoring. Journal of Parallel and Distributed Computing, 118 (part 1)(??):57– 66, August 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S074373151730093X.

You:2017:DIH

[YFBY17] Yang You, Haohuan Fu, David Bader, and Guangwen Yang. Designing and implementing a heuristic cross-architecture combination for graph traversal. Journal of Parallel and Distributed Computing, 108(??):95–105, October 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300375.

You:2015:SSV

[YFS^{+}15] Yang You, Haohuan Fu, Shuaiwen Leon Song, Amanda Randles, Darren Kerbyson, Andres Marquez, Guangwen Yang, and Adolfy Hoisie. Scaling support vector machines on modern HPC platforms. Journal of Parallel and Distributed Computing, 76(??):16–31, February 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001683.

Yao:2019:PPC

[YGWJ19] Zhongjiang Yao, Jingguo Ge, Yulei Wu, and Linjie Jian. A privacy preserved and credible network protocol. Journal of Parallel and Distributed Computing, 132(??):150–159, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518308323.

Yang:2010:LCM

[YGZ^{+}10] Jun Yang, Lan Gao, Youtao Zhang, Marek Chrobak, and Hsien-Hsin S. Lee. A low-cost memory remapping scheme for address bus protection. Journal of Parallel and Distributed Computing, 70(5):443–457, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ye:2018:ABS

[YHWY18a] Dayong Ye, Qiang He, Yanchun Wang, and Yun Yang. An agent-based service adaptation approach in distributed multitenant service-based systems. Journal of Parallel and Distributed Computing, 122(??):11–25, December 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304866.

Ye:2018:DTS

[YHWY18b] Dayong Ye, Qiang He, Yanchun Wang, and Yun Yang. Detection of transmissible service failure in distributed servicebased systems. Journal of Parallel and Distributed Computing, 119(??):36–49, September 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518301783.

Yum:2010:IAC

[YJKD10] Ki Hwan Yum, Yuho Jin, Eun Jung Kim, and Chita R. Das. Integration of admission, congestion, and peak power control in QoS-aware clusters. Journal of Parallel and Distributed Computing, 70(11):1087–1099, November 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yaseen:2016:LBW

[YJL16] Ashraf Yaseen, Hao Ji, and Yaohang Li. A load-balancing workload distribution scheme for three-body interaction computation on Graphics Processing Units (GPU). Journal of Parallel and Distributed Computing, 87(??):91–101, January 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515001884.

Yaseen:2012:AKB

[YL12] Ashraf Yaseen and Yaohang Li. Accelerating knowledgebased energy evaluation in protein structure modeling with Graphics Processing Units. Journal of Parallel and Distributed Computing, 72(2):297–307, February 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001948.

Yin:2015:GHD

[YLB^{+}15] Jiangtao Yin, Yong Liao, Mario Baldi, Lixin Gao, and Antonio Nucci. GOM-Hadoop: a distributed framework for efficient analytics on ordered datasets. Journal of Parallel and Distributed Computing, 83(??):58–69, September 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515000933.

Yang:2017:HCM

[YLL17] Wangdong Yang, Kenli Li, and Keqin Li. A hybrid computing method of SpMV on CPU–GPU heterogeneous computing systems. Journal of Parallel and Distributed Computing, 104(??): 49–60, June 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517300011.

Yu:2018:PEM

[YLZW18] Chunyan Yu, Qi Lian, Dong Zhang, and Chunming Wu. PAME: Evolutionary membrane computing for virtual network embedding. Journal of Parallel and Distributed Computing, 111(??):136–151, January 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302356.

Yuan:2014:SLB

[YMLP14] Xin Yuan, Santosh Mahapatra, Michael Lang, and Scott Pakin. Static load-balanced routing for slimmed fat-trees. Journal of Parallel and Distributed Computing, 74(5):2423– 2432, May 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731514000215.

Yildirim:2011:PWL

[YÖ11] Ahmet Artu Yildirim and Cem Özdogan. Parallel WaveCluster: a linear scaling parallel clustering algorithm implementation with application to very large datasets. Journal of Parallel and Distributed Computing, 71(7):955–962, July 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151100075X.

Yu:2016:ASR

[YPCW16] Licheng Yu, Yulong Pei, Tianzhou Chen, and Minghui Wu. Architecture supported register stash for GPGPU. Journal of Parallel and Distributed Computing, 89(??):25–36, March 2016. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151500221X.

Yan:2013:CPE

[YpGyLlC13] Jing Yan, Xin ping Guan, Xiao yuan Luo, and Cai lian Chen. A cooperative pursuit-evasion game in wireless sensor and actor networks. Journal of Parallel and Distributed Computing, 73(9):1267–1276, September 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001159.

Yu:2012:HHC

[YQTV12] Weikuan Yu, Xinyu Que, Vinod Tipparaju, and Jeffrey S. Vetter. HiCOO: Hierarchical cooperation for scalable communication in Global Address Space programming models on Cray XT systems. Journal of Parallel and Distributed Computing, 72(11):1481–1492, November 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000342.

Yu:2011:HDI

[YS11] Yanan Yu and Ashok Srinivasan. Hybrid dynamic iterations for the solution of initial value problems. Journal of Parallel and Distributed Computing, 71(11):1509–1517, November 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511000852.

Yin:2011:EAC

[YSS11] Bolian Yin, Hongchi Shi, and Yi Shang. An efficient algorithm for constructing a connected dominating set in mobile ad hoc. Journal of Parallel and Distributed Computing, 71(1):27–39, January 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yi:2019:ETP

[YTH^{+}19] Xun Yi, Zahir Tari, Feng Hao, Liqun Chen, Joseph K. Liu, Xuechao Yang, Kwok-Yan Lam, Ibrahim Khalil, and Albert Y. Zomaya. Efficient threshold password-authenticated secret sharing protocols for cloud computing. Journal of Parallel and Distributed Computing, 128(??):57–70, June 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303071.

Yang:2019:PPN

[YTZ19] Shumei Yang, Shaohua Tang, and Xiao Zhang. Privacypreserving k nearest neighbor query with authentication on road networks. Journal of Parallel and Distributed Computing, 134(??):25–36, December 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151930019X.

Ye:2013:PBI

[YWAT13] Jing Ye, Andrew M. Wallace, Abdallah Al Zain, and John Thompson. Parallel Bayesian inference of range and reflectance from LaDAR profiles. Journal of Parallel and Distributed Computing, 73(4):383–399, April 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512002845.

Yun:2015:IAW

[YWG15] Daqing Yun, Chase Qishi Wu, and Yi Gu. An integrated approach to workflow mapping and task scheduling for delay minimization in distributed environments. Journal of Parallel and Distributed Computing, 84(??):51–64, October 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515001227.

Yang:2018:MDM

[YWJ^{+}18] Jiachen Yang, Jiabao Wen, Bin Jiang, Zhihan Lv, and Arun Kumar Sangaiah. Marine depth mapping algorithm based on the edge computing in Internet of Things. Journal of Parallel and Distributed Computing, 114(??):95–103, April 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517303477.

Yu:2012:CME

[YWW12] Jiguo Yu, Nannan Wang, and Guanghui Wang. Constructing minimum extended weakly-connected dominating sets for clustering in ad hoc networks. Journal of Parallel and Distributed Computing, 72(1):35–47, January 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001377.

Yong:2018:IBI

[YXW^{+}18] Binbin Yong, Zijian Xu, Xin Wang, Libin Cheng, Xue Li, Xiang Wu, and Qingguo Zhou. IoT-based intelligent fitness system. Journal of Parallel and Distributed Computing, 118 (part 1)(??):14–21, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0743731517301570.

Yu:2013:DSA

[YXX13] Bo Yu, Cheng-Zhong Xu, and Bin Xiao. Detecting Sybil attacks in VANETs. Journal of Parallel and Distributed Computing, 73(6):746–756, June 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731513000191.

Yuan:2011:DMC

[YYLC11] Dong Yuan, Yun Yang, Xiao Liu, and Jinjun Chen. Ondemand minimum cost benchmarking for intermediate dataset storage in scientific cloud workflow systems. Journal of Parallel and Distributed Computing, 71(2):316–332, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yu:2019:TPP

[YYWZ19] Hui Yu, Jiahai Yang, Hui Wang, and Hui Zhang. Towards predictable performance via two-layer bandwidth allocation in cloud datacenter. Journal of Parallel and Distributed Computing, 126(??):34–47, April 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308694.

Yao:2019:TBR

[YZC^{+}19] Xin Yao, Yizhu Zou, Zhigang Chen, Ming Zhao, and Qin Liu. Topic-based rank search with verifiable social data outsourcing. Journal of Parallel and Distributed Computing, 134(??):1–12, December 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519300322.

Yin:2018:ADE

[YZG18] Jiangtao Yin, Yanfeng Zhang, and Lixin Gao. Accelerating distributed expectation–maximization algorithms with frequent updates. Journal of Parallel and Distributed Computing, 111(??):65–75, January 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302174.

Ye:2015:DDD

[YZS15] Dayong Ye, Minjie Zhang, and Danny Sutanto. Decentralised dispatch of distributed energy resources in smart grids via multi-agent coalition formation. Journal of Parallel and Distributed Computing, 83(??):30–43, September 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515000672.

Yu:2015:QMP

[YZW^{+}15] Li Yu, Zhou Zhou, Sean Wallace, Michael E. Papka, and Zhiling Lan. Quantitative modeling of power performance tradeoffs on extreme scale systems. Journal of Parallel and Distributed Computing, 84(??):1–14, October 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515001045.

Yuan:2011:SAM

[YZX11] Zhaohui Yuan, Yuping Zhang, and Chun Jason Xue. Sleepaware mode assignment in wireless embedded systems. Journal of Parallel and Distributed Computing, 71(7):1002–1010, July 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731510002388.

Zhai:2017:EEI

[ZAAB17] Xiaojun Zhai, Amine Ait Si Ali, Abbes Amira, and Faycal Bensaali. ECG encryption and identification based security solution on the Zynq SoC for connected health systems. Journal of Parallel and Distributed Computing, 106(??):143–152, August 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516301927.

Zarrin:2018:RDD

[ZAB18] Javad Zarrin, Rui L. Aguiar, and João Paulo Barraca. Resource discovery for distributed computing systems: a comprehensive survey. Journal of Parallel and Distributed Computing, 113(??):127–166, March 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517303088.

Zahavi:2012:FTR

[Zah12] Eitan Zahavi. Fat-tree routing and node ordering providing contention free traffic for MPI global collectives. Journal of Parallel and Distributed Computing, 72(11):1423–1432, November 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512000305.

Zhang:2011:TNT

[ZBR11] Zhenxia Zhang, Azzedine Boukerche, and Hussam Ramadan. TEASE: a novel Tunnel-based sEcure Authentication SchemE to support smooth handoff in IEEE 802.11 wireless networks. Journal of Parallel and Distributed Computing, 71(7):897–905, July 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731510002686.

Zhu:2017:DPC

[ZBW^{+}17] Yanhe Zhu, Dongyang Bie, Xiaolu Wang, Yu Zhang, Hongzhe Jin, and Jie Zhao. A distributed and parallel control mechanism for self-reconfiguration of modular robots using Lsystems and cellular automata. Journal of Parallel and Distributed Computing, 102(??):80–90, April 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301824.

Zoni:2017:BEF

[ZCF^{+}17] Davide Zoni, Andrea Canidio, William Fornaciari, Panayiotis Englezakis, Chrysostomos Nicopoulos, and Yiannakis Sazeides. BlackOut: Enabling fine-grained power gating of buffers in Network-on-Chip routers. Journal of Parallel and Distributed Computing, 104(??):130–145, June 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517300229.

Zhang:2012:EIS

[ZCMY12] Zhan Zhang, Shigang Chen, Zhen Mo, and MyungKeun Yoon. An efficient incentive scheme with a distributed authority infrastructure in peer-to-peer networks. Journal of Parallel and Distributed Computing, 72(12):1741–1752, December 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731512001906.

Zhang:2018:ABF

[ZCS^{+}18] Guangyan Zhang, Shuhan Cheng, Jiwu Shu, Qingda Hu, and Weimin Zheng. Accelerating breadth-first graph search on a single server by dynamic edge trimming. Journal of Parallel and Distributed Computing, 120(??):383–394, October 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517302617.

Zheng:2019:TFM

[ZCW19] Huanyang Zheng, Wei Chang, and Jie Wu. Traffic flow monitoring systems in smart cities: Coverage and distinguishability among vehicles. Journal of Parallel and Distributed Computing, 127(??):224–237, May 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151830488X.

Zhang:2018:LWL

[ZFT^{+}18] Zheng Zhang, Dan Feng, Zhipeng Tan, Laurence T. Yang, and Jiayang Zheng. A light-weight log-based hybrid storage system. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):307–315, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517303404.

Zaman:2013:CAB

[ZG13] Sharrukh Zaman and Daniel Grosu. Combinatorial auctionbased allocation of virtual machine instances in clouds. Journal of Parallel and Distributed Computing, 73(4):495–508, April 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512002870.

Zheng:2014:BBE

[ZGG^{+}14] Cong Zheng, Shuo Gu, Tong-Xiang Gu, Bing Yang, and Xing-Ping Liu. BiELL: a bisection ELLPACK-based storage format for optimizing SpMV on GPUs. Journal of Parallel and Distributed Computing, 74(7):2639–2647, July 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514000458.

Zedadra:2018:SIB

[ZGJ^{+}18] Ouarda Zedadra, Antonio Guerrieri, Nicolas Jouandeau, Giandomenico Spezzano, Hamid Seridi, and Giancarlo Fortino. Swarm intelligence-based algorithms within IoT-based systems: a review. Journal of Parallel and Distributed Computing, 122(??):173–187, December 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518305987.

Zhang:2019:OLO

[ZGW^{+}19] Feifei Zhang, Jidong Ge, Chifong Wong, Chuanyi Li, Xingguo Chen, Sheng Zhang, Bin Luo, He Zhang, and Victor Chang. Online learning offloading framework for heterogeneous mobile edge computing system. Journal of Parallel and Distributed Computing, 128(??):167–183, June 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519301492.

Zhang:2011:RLP

[Zha11] Nan Zhang. Resolving a L2-prefetch-caused parallel nonscaling on Intel Core microarchitecture. Journal of Parallel and Distributed Computing, 71(7):915–924, July 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511000736.

Zhu:2019:FET

[ZHG^{+}19] Guanghui Zhu, Qiu Hu, Rong Gu, Chunfeng Yuan, and Yihua Huang. ForestLayer: Efficient training of deep forests on distributed task-parallel platforms. Journal of Parallel and Distributed Computing, 132(??):113–126, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305392.

Zhang:2015:HAB

[ZHH15] Kai Zhang, Jiayu Hu, and Bei Hua. A holistic approach to build real-time stream processing system with GPU. Journal of Parallel and Distributed Computing, 83(??):44–57, September 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731515000921.

Zhu:2012:AEE

[ZHLQ12] Xiaomin Zhu, Chuan He, Kenli Li, and Xiao Qin. Adaptive energy-efficient scheduling for real-time tasks on DVSenabled heterogeneous clusters. Journal of Parallel and Distributed Computing, 72(6):751–763, June 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000706.

Zhu:2016:SCN

[ZHT16] Youwen Zhu, Zhiqiu Huang, and Tsuyoshi Takagi. Secure and controllable k-NN query over encrypted cloud data with key confidentiality. Journal of Parallel and Distributed Computing, 89(??):1–12, March 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731515002105.

Zhao:2019:GCS

[ZHW19] Shu-Li Zhao, Rong-Xia Hao, and Jie Wu. The generalized 3-connectivity of some regular networks. Journal of Parallel and Distributed Computing, 133(??):18–29, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518306725.

Zhang:2018:MSP

[ZKZF18] Xuechen Zhang, Ujjwal Khanal, Xinghui Zhao, and Stephen Ficklin. Making sense of performance in in-memory computing frameworks for scientific data analysis: a case study of the spark system. Journal of Parallel and Distributed Computing, 120(??):369–382, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302927.

Zhang:2012:EEM

[ZLCJ12] Liang Zhang, Wen Luo, Shigang Chen, and Ying Jian. End-to-end maxmin fairness in multihop wireless networks: Theory and protocol. Journal of Parallel and Distributed Computing, 72(3):462–474, March 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731511002292.

Zhou:2018:DNA

[ZLCZ18] Hang Zhou, Qing Li, Kim-Kwang Raymond Choo, and Hai Zhu. DADTA: a novel adaptive strategy for energy and performance efficient virtual machine consolidation. Journal of Parallel and Distributed Computing, 121(??):53–70, November 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518304520.

Zheng:2018:HVF

[ZLH^{+}18] Shengan Zheng, Hao Liu, Linpeng Huang, Yanyan Shen, and Yanmin Zhu. HMVFS: a versioning file system on DRAM/NVM hybrid memory. Journal of Parallel and Distributed Computing, 120(??):355–368, October 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151730299X.

Zhang:2019:PAS

[ZLJ^{+}19] Feng Zhang, Victor E. Lee, Ruoming Jin, Saurabh Garg, Kim-Kwang, Raymond Choo, Michele Maasberg, Lijun Dongab, and Chi Cheng. Privacy-aware smart city: a case study in collaborative filtering recommender systems. Journal of Parallel and Distributed Computing, 127(??):145–159, May 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517303465.

Zefreh:2019:TCP

[ZLKK19] Ebrahim Zarei Zefreh, Shahriar Lotfi, Leyli Mohammad Khanli, and Jaber Karimpour. Topology and computationalpower aware tile mapping of perfectly nested loops with dependencies on distributed systems. Journal of Parallel and Distributed Computing, 129(??):14–35, July 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519301753.

Zhao:2014:TEF

[ZLL14] Han Zhao, Xinxin Liu, and Xiaolin Li. Towards efficient and fair resource trading in community-based cloud computing. Journal of Parallel and Distributed Computing, 74(11):3087–3097, November 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001270.

Zhu:2014:OVS

[ZLMC14] Xiaojun Zhu, Qun Li, Weizhen Mao, and Guihai Chen. Online vector scheduling and generalized load balancing. Journal of Parallel and Distributed Computing, 74(4):2304–2309, April 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513002402.

Zlateski:2017:STC

[ZLS17] Aleksandar Zlateski, Kisuk Lee, and H. Sebastian Seung. Scalable training of 3D convolutional networks on multiand many-cores. Journal of Parallel and Distributed Computing, 106(??):195–204, August 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300746.

Zhao:2019:BBP

[ZLT^{+}19] Yanqi Zhao, Yiming Liu, Aikui Tian, Yong Yu, and Xiaojiang Du. Blockchain based privacy-preserving software updates with proof-of-delivery for Internet of Things. Journal of Parallel and Distributed Computing, 132(??):141–149, October 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151930098X.

Zhou:2012:EPD

[ZLWL12] Xu Zhou, Kai Lu, Xiaoping Wang, and Xu Li. Exploiting parallelism in deterministic shared memory multiprocessing. Journal of Parallel and Distributed Computing, 72(5):716– 727, May 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512000421.

Zhang:2018:MGS

[ZLWZ18] Xiao Zhang, Yinrun Lyu, Yanjun Wu, and Chen Zhao. MixHeter: a global scheduler for mixed workloads in heterogeneous environments. Journal of Parallel and Distributed Computing, 111(??):93–103, January 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302198.

Zhang:2019:CFV

[ZLZ^{+}19] Fei Zhang, Guangming Liu, Bo Zhao, Piotr Kasprzak, Xiaoming Fu, and Ramin Yahyapour. CBase: Fast virtual machine storage data migration with a new data center structure. Journal of Parallel and Distributed Computing, 124(??):14–26, February 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518307342.

Zhang:2011:DID

[ZMCP11] Yongpeng Zhang, Frank Mueller, Xiaohui Cui, and Thomas Potok. Data-intensive document clustering on graphics processing unit (GPU) clusters. Journal of Parallel and Distributed Computing, 71(2):211–224, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zeng:2016:RND

[ZMG^{+}16] Yanyan Zeng, K. Alex Mills, Shreyas Gokhale, Neeraj Mittal, S. Venkatesan, and R. Chandrasekaran. Robust neighbor discovery in multi-hop multi-channel heterogeneous wireless networks. Journal of Parallel and Distributed Computing, 92(??): 15–34, May 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516000162.

Zhang:2017:DIS

[ZMZJ17] Pei Zhang, Aaron Mills, Joseph Zambreno, and Phillip H. Jones. The design and integration of a software configurable and parallelized coprocessor architecture for LQR control. Journal of Parallel and Distributed Computing, 106(??):121– 131, August 2017. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731517300473.

Zhao:2014:DAC

[ZPK^{+}14] W. S. Zhao, J. M. Portal, W. Kang, M. Moreau, Y. Zhang, H. Aziza, J.-O. Klein, Z. H. Wang, D. Querlioz, D. Deleruyelle, M. Bocquet, D. Ravelosona, C. Muller, and C. Chappert. Design and analysis of crossbar architecture based on complementary resistive switching non-volatile memory cells. Journal of Parallel and Distributed Computing, 74(6):2484–2496, June 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731513001408.

Zhu:2011:OBF

[ZQMM11] Dakai Zhu, Xuan Qi, Daniel Mossé, and Rami Melhem. An optimal boundary fair scheduling algorithm for multiprocessor real-time systems. Journal of Parallel and Distributed Computing, 71(10):1411–1425, October 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001304.

Zhang:2014:PVS

[ZRN^{+}14] Jianfeng Zhang, Mostafizur Rahman, Pritish Narayanan, Santosh Khasanvis, and C. Andras Moritz. Parameter variation sensing and estimation in nanoscale fabrics. Journal of Parallel and Distributed Computing, 74(6):2504–2511, June 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151300141X.

Zheng:2013:SDS

[ZS13] Wei Zheng and Rizos Sakellariou. Stochastic DAG scheduling using a Monte Carlo approach. Journal of Parallel and Distributed Computing, 73(12):1673–1689, December 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731513001573.

Zsaki:2016:HAG

[Zsa16] Attila Michael Zsaki. Hardware-accelerated generation of 3D diffusion-limited aggregation structures. Journal of Parallel and Distributed Computing, 97(??):24–34, November 2016.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300740.

Zhong:2018:ESR

[ZSCX18] Hong Zhong, Lili Shao, Jie Cui, and Yan Xu. An efficient and secure recoverable data aggregation scheme for heterogeneous wireless sensor networks. Journal of Parallel and Distributed Computing, 111(??):1–12, January 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302083.

Zhang:2014:COS

[ZSW14] Tao Zhang, Wei Shu, and Min-You Wu. CUIRRE: an opensource library for load balancing and characterizing irregular applications on GPUs. Journal of Parallel and Distributed Computing, 74(10):2951–2966, October 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001269.

Zhang:2016:DAT

[ZTFK16] Yi-Fan Zhang, Yu-Chu Tian, Colin Fidge, and Wayne Kelly. Data-aware task scheduling for all-to-all comparison problems in heterogeneous distributed systems. Journal of Parallel and Distributed Computing, 93–94(??):87–101, July 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300144.

Zhu:2017:RSA

[ZTGL17] Xiaorui Zhu, Xianping Tao, Tao Gu, and Jian Lu. ReLog: a systematic approach for supporting efficient reprogramming in wireless sensor networks. Journal of Parallel and Distributed Computing, 102(??):132–148, April 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516301861.

Zheng:2012:UBP

[ZV12] Qin Zheng and Bharadwaj Veeravalli. Utilization-based pricing for power management and profit optimization in data centers. Journal of Parallel and Distributed Computing, 72(1):27– 34, January 2012. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731511001730.

Zeng:2014:OMR

[ZV14] Zeng Zeng and Bharadwaj Veeravalli. Optimal metadata replications and request balancing strategy on cloud data centers. Journal of Parallel and Distributed Computing, 74(10):2934–2940, October 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001129.

Zeng:2011:NSS

[ZVL11] Zeng Zeng, Bharadwaj Veeravalli, and Kenli Li. A novel serverside proxy caching strategy for large-scale multimedia applications. Journal of Parallel and Distributed Computing, 71 (4):525–536, April 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zeng:2015:SSA

[ZVL15] Lingfang Zeng, Bharadwaj Veeravalli, and Xiaorong Li. SABA: a security-aware and budget-aware workflow scheduling strategy in clouds. Journal of Parallel and Distributed Computing, 75(??):141–151, January 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731514001658.

Zhang:2011:BRE

[ZW11] Xiaoguang Zhang and Zheng Da Wu. The balance of routing energy consumption in wireless sensor networks. Journal of Parallel and Distributed Computing, 71(7):1024–1033, July 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731511000566.

Zhao:2013:BRH

[ZW13] Yaxiong Zhao and Jie Wu. Building a reliable and highperformance content-based publish/subscribe system. Journal of Parallel and Distributed Computing, 73(4):371–382, April 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S074373151200295X.

Zhao:2016:THP

[ZWQ^{+}16] Dongfang Zhao, Ke Wang, Kan Qiao, Tonglin Li, Iman Sadooghi, and Ioan Raicu. Toward high-performance key– value stores through GPU encoding and locality-aware encoding. Journal of Parallel and Distributed Computing, 96(??):27– 37, October 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731516300296.

Zheng:2017:MDS

[ZWW17] Huanyang Zheng, Ning Wang, and Jie Wu. Minimizing deep sea data collection delay with autonomous underwater vehicles. Journal of Parallel and Distributed Computing, 104(??):99–113, June 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517300126.

Zhou:2016:TNM

[ZWWX16] Jian Zhou, Jun Wang, Fei Wu, and Changsheng Xie. TEES: a novel multiple criteria optimization scheme for temperatureconstrained energy efficient storage. Journal of Parallel and Distributed Computing, 96(??):152–162, October 2016. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731516300405.

Zhang:2015:PAB

[ZWY^{+}15] Junyao Zhang, Qingdong Wang, Jiangling Yin, Jian Zhou, and Jun Wang. PERP: Attacking the balance among energy, performance and recovery in storage systems. Journal of Parallel and Distributed Computing, 78(??):65–77, April 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514002251.

Zhang:2014:ASP

[ZXB14] Yuanming Zhang, Gang Xiao, and Takanobu Baba. Accelerating sequential programs on commodity multi-core processors.

Journal of Parallel and Distributed Computing, 74(4):2257– 2265, April 2014. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731513002517.

Zheng:2018:IBS

[ZXGD18] Chunyuan Zheng, Chengyi Xia, Quantong Guo, and Matthias Dehmer. Interplay between SIR-based disease spreading and awareness diffusion on multiplex networks. Journal of Parallel and Distributed Computing, 115(??):20–28, May 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300078.

Zhou:2018:SPA

[ZXMR18] Xiaokang Zhou, Guangquan Xu, Jianhua Ma, and Ivan Ruchkin. Scalable platforms and advanced algorithms for IoT and cyber-enabled applications. Journal of Parallel and Distributed Computing, 118 (part 1)(??):1–4, August 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0743731518301527.

Zhang:2011:MAA

[ZXYO11] Yuping Zhang, Chun Jason Xue, Chengmo Yang, and Alex Orailoglu. Migration-aware adaptive MPSoC static schedules with dynamic reconfigurability. Journal of Parallel and Distributed Computing, 71(10):1400–1410, October 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511001353.

Zhao:2012:PSJ

[ZY12] Miao Zhao and Yuanyuan Yang. Packet scheduling with joint design of MIMO and network coding. Journal of Parallel and Distributed Computing, 72(3):376–388, March 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731511002309.

Zhan:2018:EOS

[ZZJ^{+}18] Jinyu Zhan, Xia Zhang, Wei Jiang, Yue Ma, and Ke Jiang. Energy optimization of security-sensitive mixed-criticality applications for distributed real-time systems. Journal of Parallel and Distributed Computing, 117(??):115–126, July 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300911.